A Complete Bibliography of Publications in the Journal of Mathematical Physics: 2015–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 March 2025 Version 1.33

Title word cross-reference

(0, ∞) [1670]. (1 + 1) [1516, 1384]. (1 ≤ n ≤ d) [1114]. (2 + 1) [9, 63, 1004, 1286]. (2, p) [1874]. (4 + 1) [1712]. (A_{1} + A′_{1})^{(1)} [376]. (A_{2} + A_{1})^{(1)} [376]. (p, q) [1494]. 1 [1890, 461, 1992, 2016, 1210, 103, 2194, 1558, 1759, 1532, 2124, 221, 979, 1373, 252, 843, 800, 2052]. 1 + 1 [166]. 1/2 [1776]. 16 [1070]. 2 [29, 1150, 1729, 798, 545, 1225, 1614, 1759, 1846, 1985, 2171, 762, 1105, 613, 2188, 1876, 1724, 2048, 1727]. 2 + 1 [1620, 1789]. 2k + 1 [820]. 2N [1177]. 2 × 2 [503]. 3 [588, 1337, 1030, 1992, 1140, 1031, 1995, 227, 614, 873, 937, 1162, 2279, 1984, 1225, 347, 2255, 1032, 1240, 1764, 1042, 70, 377, 824, 2158, 1733, 2037, 2006, 823, 1704, 351, 795]. 3 + 1 [627]. 33 [873]. 4 [2236, 1134, 1341, 695, 698]. 5 [1173, 1404, 1596, 2291]. 6 [1560]. [N] ⊗ [N] [1641, 1553]. ^{N} [921, 460, 341]. ^{N−1} [113]. _{1} [2131, 2196]. _{2} [2153, 2266, 2231, 2070]. _{3} [1764]. _{n} [2129]. _{q} [2057]. A [1332]. A_{2} [27]. AdS_{5} [663]. α [788, 368]. α − z [87]. B [538]. B(∞) [1288]. B(λ) [1288]. B = ⊗_{j∈Z}M_{d}^{(j)}(C) [229]. B_{2} [983]. B_{n}, l = 4 [1677]. Ū_{q}(sl_{2}) [1247]. β [248, 1350]. C^{∗} [601, 158]. C^{k} [1593]. C_{2} [194, 152]. C_{n}^{(1)} [845]. N = 2 [1351]. W_{5} [1764]. CP^{2} [590]. CP^{N−1} [590]. D [420, 438, 809, 2097, 1808, 1114, 291, 1460, 885]. D(2, 1; α) [684]. d = 3 [406, 1423]. ẍ + f(x)ẋ^{2} + g(x) = 0 [2323, 2322, 607, 13, 19]. δ [1410, 282, 1216, 549]. D ≤ 4 [1500]. Ḣ^{3/2} [579]. E [1314]. e_{6,−14} [1800]. e_{6,−26} [887]. E_{8} [1234, 1085, 1387]. ℓ^{2}(Z) [2039]. f [1825, 1813, 1923]. F (4) [1453]. G_{2} [983]. G_{2}^{(1)} [28]. Γ [2265, 1487]. GL(2, F_{q}) [613]. gl(m|n) [195, 474]. gl(N) [2001]. H(3) [1959]. H^{2} × S^{1} [1403]. H^{s} [1578]. Ĉ [1767]. P̂ [1767]. T̂ [1767]. I [2000]. K [1862, 1926, 2186, 705, 1126, 691]. κ [934]. kx^{−2} + ω^{2}x^{2} [1670]. l [970, 867]. L^{2} [1065, 1667, 307]. L^{2}(R) [1716]. L^{p} [1252, 535]. L_{∞} [1897]. L_{p} [2183]. Λ [1342]. λφ^{4} [1522]. LR^{n} [1608]. m [877]. m(r) = 1/(1 + r^{4}) [1759]. H^{3} [1034]. S [280]. S^{2} [1974]. S^{3} [1034]. Z(d) [280]. Z/pZ [2294]. O [2266]. }↕_{∈} [1972]. ⟨_{N} [rtimes]∫√(∈N , R)_{C} [2182]. ∫↕(∈, R)

[151]. μ [67, 1574]. N [1112, 768, 927, 1350, 2204, 265, 2152, 567, 932, 1087, 1608, 1055, 667, 118, 135, 1471, 1830]. n + 1 [1114]. n + 2 [667]. N = (4, 4) [659]. N = 2 [1015]. N = 2l [102]. N ≥ 2 [1061]. N ≤ 3 [2315]. O(1) [211]. O(3) [789, 1163]. O(N)^{3} [2140]. osp(n|2) [1802]. p [431, 1870, 596, 149, 1228, 305, 2274, 2259, 751]. p(x) [1325]. P_{IV} [1468]. P_{n} [1787]. P_{V} [1468]. P_{VI} [1318]. PGL(2, F_{q}) [613]. Φ [1255, 1996, 155]. Φ_{2}^{4} [247]. π [442]. P Z2n [776]. Q [1194, 107, 1412, 1364, 1926, 1058, 805, 651, 2042, 294, 1977, 213]. Q(N) [1334]. q = 1 [294]. R [73, 309, 542, 462, 53, 332, 688, 708, 1773, 118, 972]. R(p, q) [1970]. R^{2} [549]. R^{3} [947, 1955, 1322, 1709, 1216]. R^{6} [1353]. R^{N} [779, 266, 148, 149, 1668]. R_{1}^{3} [1971]. r ≥ 1 [328]. rmsl_{2} [753]. S [587, 600, 1832, 952, 1212]. S^{n} [108]. σ [461]. σ_{k} [1695]. sl(2) [848]. sl(2, R) [1129]. SL(3, C) [1628]. sl(3, R) [745]. sl_{N} [2179]. so(3, 2) [949]. so(3, R) [433]. so(3, R) [1129]. SO(N) [92]. Sp(1) [1245]. sp(4) [1403]. SP_{q}^{1|2} [1048]. spl(2, 1) [1386]. SU(1, 1) [1682, 1846, 1925]. SU(2) [1269, 1765, 1682, 1751]. SU(3) [1328]. SU(N) [537, 438, 568, 569]. SU(p, q) [1546]. su(r + 1) [1126]. SU_{2}(C) [1845, 1840]. ∑_{s} |p_{s}| + μ [382]. T [446]. τ [2282, 1779, 1889]. × [1976, 2222]. U(1) [469, 391, 1303]. U(1)^{2} [630]. U_{q}(sl_{2}) [1068]. U_{q}(G_{2}^{(1)}) [2267]. U_{q}(osp(1|2n)) [1175]. U_{q}(sl^{(M|N)})U_{q}(sl^{(M|N)})U_{q}(sl^{(M|N)}) [1799]. U_{q}(ĝl2) [753]. U_{q}(ˆ∫↕_{∈}) [2128]. U_{q}sl(2|1) [543]. U_{r,s}(osp(1|2n)) [542]. V [1772, 1777]. φ^{4} [49]. W [2083, 684, 1334, 30, 539]. W_{∞} [687]. W_{N} [1495]. ˆ∫↕_{∈} [2157]. X [1617, 1105]. Y (R)F^{2} [629]. Z [1105]. Z_{2} [548]. Z_{2} × Z_{2} [1351].

* [158]. *-derivations [158].

-adic [596]. -algebra [1830, 1334]. -algebras [158, 2265, 684, 30, 539]. -arm [1558]. -associated [2111]. -belts [2279]. -body [1596, 937, 1032, 135, 1471, 1112, 768]. -brane [2274]. -Camassa [1574]. -categorical [29]. -categories [446]. -class [1578]. -coherent [848]. -completions [601]. -conformal [867, 102]. -corrections [368]. -coupled [877]. -d [1890, 798, 227, 2016, 1210, 103, 2171, 979, 2052, 351]. -deformed [805, 2042, 1194, 1970]. -designs [1105]. -diagonal [1105]. -dimensional [1030, 1031, 1995, 1341, 2097, 63, 1808, 9, 1004, 1114, 291, 613, 2158, 1460, 2006, 1712, 885, 1286, 698, 809]. -discretization [1314]. -divergences [1813, 1923]. -embedding [1593]. -ensembles [1350]. -entropies [1255]. -exact-solvability [2097, 108, 1460]. -expansion [600, 1212]. -fold [1070, 2204]. -forms [705, 1225]. -function [248, 2282]. -functions [1779, 1889, 438, 1487]. -Gaussians [107]. -generalization [483]. -hypergeometric [1412]. -imaginary [155]. -interaction [1410]. -interactions [549]. -invariant [630, 391, 469, 2140]. -Kepler [1245, 211]. -Kirchhoff [1325, 1870]. -Laplacian [1494, 1874, 149, 1228, 305, 2259, 751]. -lattice [1765]. -level [1126, 118]. -local [545]. -matrices [1773, 118, 972, 708]. -matrix [73, 309, 542, 462, 332, 688, 952]. -models [461]. -modules [1802, 848]. -norm [1065]. -orthogonal [294]. -plane [1825]. -point [614, 1617]. -polynomials [1926]. -preserving [2000]. -pseudo-bosons [420]. -qubit [695]. -Racah [1977]. -Rényi [87]. -shell [1216]. -Sobolev [1996]. -spaces [2183]. -spacetime [934]. -spectral [1364]. -spectrum [587, 1252, 1832]. -sphere [588]. -stability [307]. -state [873]. -statistical [213]. -statistics [1058]. -step [691]. -super [970]. -supercritical [1667]. -supermanifolds [776]. -symmetries [2083]. -symplectic [1087, 1608]. -theory [1862, 1926, 2186]. -tori [1134]. -twistor [1055]. -vertex [873]. -Virasoro [651]. -wave [431]. -Yamabe [1695].

1/2 [2149].

2018 [2314].

3d [1887].

48 [841, 428]. 49 [119, 4, 8].

50 [1358]. 51 [63, 608]. 52 [427]. 53 [640]. 54 [1720, 426, 945, 2323, 916, 641, 2322, 1319, 607, 19]. 55 [145, 607, 301]. 56 [302, 1438, 1359, 1439, 842, 1130, 1946, 339]. 57 [915, 465, 1017, 1199, 977, 678, 2042, 1489, 642]. 58 [1160, 1389, 1320, 2319, 1947]. 59 [2043, 1980, 1719, 2327, 2328, 1982, 1569, 1840, 1898, 1839, 2087].

60 [2326, 2320, 1981, 2321].

= [170, 1865].

ABCD [435]. Abel [429]. abelian [1165, 1409, 1590, 382, 390, 2225, 1627, 1331, 562, 1303, 1647]. Ablowitz [2282, 52, 838, 1829, 1718]. Abraham [868, 632]. Absence [94, 2156, 1399, 549]. absolutely [94, 637, 1017, 1416]. absorbing [2309]. absorption [387]. AC [1196]. acceleration [174, 737, 447]. acceleration-dependent [737]. according [1623]. account [79]. Accumulation [1024, 657]. Achieving [725]. acknowledgment [1948]. acoustic [1696]. across [1027, 1965]. acting [1586]. action [1938, 268, 3, 599, 614, 1836, 1344, 2205, 1358, 1517, 4, 1113, 41, 1927]. action-angle [2205]. action-dependent [1517]. Actions [753, 1862, 1926, 413, 701, 703, 2150]. activation [2104]. adapted [527]. added [276, 1688]. Addendum [1359]. Addition [2095, 287]. additive [545, 1193, 1885, 408, 2273]. additivity [1836]. Adiabatic [759, 453, 2240, 2019]. adic [596]. Adjoint [715, 2199, 11, 1720, 907, 1917, 1297, 833, 2204]. adjointness [1075, 547, 380]. ADM [392]. admit [1343]. admits [1427]. AdS [933, 1002, 1764, 2231, 2070]. advantage [2137]. advection [303]. adversarial [2104]. affine [1364, 2131, 28, 692, 210, 1234, 564, 846, 1934, 1524, 2143, 845, 1085, 30, 539, 715]. affine-Virasoro [2131]. affinizations [1735]. AG [2153]. against [1749]. aggregate [1040]. AGT [1173, 988]. Aharonov [23, 476, 1291]. AKNS [1787]. algebra [2177, 167, 970, 1942, 600, 1173, 2179, 1606, 986, 1912, 328, 329, 888, 2131, 614, 2125, 194, 1041, 1939, 2324, 1847, 152, 28, 1865, 739, 853, 1970, 1736, 472, 236, 1170, 2234, 753, 1959, 2180, 651, 2182, 2038, 1334, 151, 32, 463, 924, 1766, 1830, 1067, 885, 2014, 711, 687, 2315, 601]. algebra-valued [463]. Algebraic [720, 2316, 2317, 1713, 335, 2248, 544, 1148, 1367, 1918, 2060, 894, 319, 2300, 656, 1678, 275, 115, 1927]. algebras [122, 1351, 1976, 1943, 1620, 73, 309, 714, 515, 394, 1636, 1443, 1659, 926, 1878, 1364, 889, 745, 1521, 683, 1660, 510, 1126, 462, 1773, 439, 799, 1619, 1618, 1958, 210, 26, 1813, 1923, 1288, 1581, 2196, 846, 112, 689, 158, 1879, 269, 932, 1147, 1817, 1547, 1336, 1739, 670, 2265, 1097, 886, 1963, 25, 686, 2143, 355, 684, 845, 1990, 1495, 30, 539, 137, 1172, 1411, 270, 1067, 421, 155, 446, 310, 153, 884, 2103]. algebro [178]. algebro-geometric [178]. Algebroid [1888]. algebroids [368, 1045]. algorithm [915, 836, 2279, 218, 1289, 1717, 1566, 591]. Algorithmic [1078, 1079]. Algorithms [438, 1998]. Alice [1767]. aligned [713, 1338]. All-order [962, 628]. Almost [497, 1257, 1587, 1391, 864, 1605, 2317, 1258, 2190, 2276, 1578, 1711]. almost-commutative [2317, 1258]. almost-symmetries [2190]. alone [396]. along [448]. alternating [2128]. ambient [1697]. ambiguity [34]. Ampère [1156]. amplifier [2270, 719]. amplitude [1671, 1003, 1413]. Analog [436, 1117]. analogs [2179]. analogue [357, 1263]. Analysis [1466, 761, 1841, 780, 182, 1209, 1529, 1425, 2116, 1703, 572, 796, 777, 1347, 2107, 258, 875, 737, 2206, 641, 1081, 1319, 911, 1572, 12, 1320, 1762, 163, 1604, 1014, 1211, 1273, 909, 880, 1090, 1213, 1437, 2246, 15, 642, 499, 1286]. Analytic [1222, 239, 280, 1161, 352, 616, 2162, 1454, 1142]. Analytical [201, 235, 580, 1149, 1072, 1501, 199, 377, 627]. Analyticity [2235, 1003, 2208]. Ancilla [165, 487]. Ancilla-approximable [165]. ancilla-free [487]. Anderson [1248, 1370, 875, 1086, 2113, 749]. angle [1534, 2081, 2205, 1511]. angled [68]. angles [1555]. angular [630, 1462, 2072]. anharmonic [2244]. Anisotropic [967, 110, 45, 935, 379, 832, 1872, 2028, 1674, 985, 410, 2036, 77]. annihilation [942]. annuli [1435]. annulus [1753]. Anomalous [486, 1526]. anomaly [404, 2068]. ansatz [53]. antenna [455]. Anti [1957, 568, 1592, 2159, 2035, 1363, 448]. Anti- [1957]. anti-de [568, 1592, 2159, 448]. Anticoherent [1143]. antilinear [871]. Antiquantization [1784]. Antisymmetric [564]. any [367, 221]. anyons [128]. Appearances [673]. Application [1703, 1815, 676, 227, 641, 201, 15, 642, 254, 969, 1462, 1742, 1636, 1926, 1039, 1702, 1699, 545, 1864, 140, 55, 1057, 501, 1042, 34, 476, 1122, 1043, 1565, 2085]. Applications [763, 837, 1655, 2116, 1280, 2281, 736, 2219, 1255, 707, 105, 1365, 1848, 901, 507, 1008, 158, 1256, 2218, 250, 1186, 693, 985, 1308, 1212, 1556, 1863, 723, 418, 527, 1487, 1805]. applied [1521, 1264, 718]. approach [1233, 1183, 1430, 186, 2298, 1267, 983, 198, 599, 1713, 1436, 1154, 741, 1595, 1071, 1761, 2203, 330, 331, 2312, 455, 429, 285, 1581, 1481, 1464, 1141, 1864, 902, 2218, 1192, 2195, 236, 123, 1455, 1740, 857, 1078, 1866, 506, 1050, 656, 2117, 1131, 778, 553, 1944, 2301, 966]. approaches [470, 1539, 868]. approximable [165]. approximants [43]. Approximate [1219, 1571, 608, 7, 1680]. approximately [1145]. Approximation [1998, 2245, 2158, 1417, 1825, 1058, 1293, 277, 1454, 1380, 1684, 1796, 1011, 395, 1216, 467, 1933, 243, 162, 520, 560, 723, 746, 774, 837, 718, 2115, 1238, 1907]. arbitrarily [1302, 527]. Arbitrary [605, 1306, 768, 1632, 393, 625, 1033, 127, 457, 1155, 2198, 1270]. Arbitrary-order [605]. Area [399, 475, 175, 2186, 1974]. area-weighted [175]. arising [987, 1134, 1707, 376, 1876]. arm [1558]. Arquès [1480]. arrays [1691]. arrival [1999]. artery [1274]. Askey [2179]. Aspects [970, 735, 2161, 600, 426, 174, 2325, 241, 1292, 14]. associated [1549, 1942, 874, 821, 873, 2111, 1386, 2129, 573, 1801, 886, 433, 65, 605, 539, 1100, 2306, 2315]. Associative [708]. Associativity [490]. asymmetric [414, 326, 552, 675, 1382, 1769]. asymmetrical [1754]. Asymptotic [624, 796, 1451, 2033, 987, 882, 1935, 2258, 1822, 1400, 1464, 158, 827, 2093, 1762, 1900, 1790, 89, 1233, 518, 1775, 777, 914, 227, 697, 611, 980, 2090, 1103, 1169, 1857, 578, 2311, 2149, 581, 1275, 1003, 679, 1764, 2304, 1486, 1936, 1322, 1826, 1324, 469]. asymptotically [1755, 392, 1448]. Asymptotics [1549, 1941, 98, 1410, 1372, 1672, 1769, 412, 69, 468, 2278, 728, 281, 1071, 2296, 282, 337, 1613, 175, 744, 260, 1903, 2289, 1986, 1551]. Atiyah [992, 2242]. atlas [1026]. atmosphere [2106]. atmosphere-ocean [2106]. atom [2107, 1580, 1584, 1809, 2244, 1072, 2329, 2243, 891, 1678]. Atomic [1684, 1150, 2144, 807, 1535, 201]. Atomic-pair [1684]. atoms [2015, 1319, 1368, 1141, 12, 1320, 1349]. attached [761]. attacks [1302, 1964]. attenuator [2270]. Attracting [1261]. attraction [1313]. attractive [304, 549, 752, 2114, 50]. attractor [1057, 373]. attractors [2164, 292, 492, 1569, 1515, 1228, 408, 2027, 374, 2002]. Audenaert [724]. automaton [1931]. automorphisms [2088]. autonomous [172, 1261, 865, 2203, 1124, 1838, 2201, 866, 2027, 373, 2002]. averaged [704]. averaging [269, 1758, 866]. Aw [1796]. axial [487]. Axiomatic [609]. axis [1188]. axisymmetric [1821].

B [1913]. Bach [1945, 1312]. backaction [618]. background [2160, 1541, 1264, 561, 559, 943]. background-controlled [943]. backgrounds [2316, 2317]. Bäcklund [63, 9, 1286]. Baker [591]. Bakry [567]. balance [1459, 2067]. Balian [2039]. ball [2207, 1114]. ballistic [927]. Banach [1838]. band [252, 1195, 2319]. Bang [2024]. Bannai [1880]. Bannai-Ito [1880]. Bardeen [523]. Bargmann [1317, 538, 1912, 2217, 215]. barotropic [1645, 1987]. barriers [616]. Barut [1912, 1126]. baryon [627]. based [587, 237, 1030, 1555, 477, 1218, 2207, 283, 1421, 429, 1532, 2182, 698, 1812]. bases [833, 941, 2098, 622, 845]. basic [450, 845]. basis [20, 145, 1710, 1412, 1737, 544, 2128]. Batalin [416, 1927]. Bateman [1503, 1687]. bath [716]. Baxter [708, 884]. Bayesian [804, 1812]. BC [2129]. BCS [1214]. be [861, 571, 678]. beam [2013, 209, 1699, 291, 1284, 900, 1843]. beams [1577]. Becchi [639]. Beddington [1467]. behavior [469, 226, 987, 748, 882, 1392, 611, 980, 2090, 81, 1612, 1984, 2258, 1822, 158, 581, 1275, 2093, 2124, 1003, 679, 223, 2201, 1486, 1936, 2106, 2122, 1324]. behaviors [1470, 1447, 1732]. behaviour [2307, 2149, 827, 47]. Belavin [689]. Bell [483, 1588, 1102, 131, 1104]. below [1915]. Beltrami [2006]. belts [2279]. Bénard [2276, 1264]. Bender [956]. bends [68]. Benjamin [1666, 1900]. Benney [1345]. Berestycki [2292, 1730]. Berestycki-Lions [2292]. Berezin [1479]. Berg [672]. Bergman [2217]. Bernoulli [1127, 586, 2283]. Bernstein [1658]. Berry [2268, 2062, 1020]. Berta [1420]. Besov [2091, 2090, 2176, 682]. Bessel [1935, 1975, 744, 905, 1372, 2214]. Bessel-function [1935]. Bessel-type [1975]. beta [58]. beta-ensembles [58]. Bethe [53, 1236, 1558]. between [2060, 1480, 157, 1849, 2192, 2244, 449, 2098, 691, 2231, 1866, 903, 1489, 1804, 135, 1375, 89, 2297]. Beyond [2234, 162, 728, 2178, 1787, 1138]. BFV [599]. BGG [1332]. BGK [2120, 790]. Bi [52, 702, 255, 188, 1237, 655, 299, 507, 433, 2056, 655]. bi-coherent [299]. bi-dimensional [702]. Bi-Hamiltonian [702, 255, 188, 1237, 433]. Bi-HKT [655]. Bi-integrable [52]. bi-linear [507]. bi-nonlocal [2056]. bialgebra [439, 711]. Bianchi [42, 1466, 935, 1259]. Bianchi-like [1466]. biased [2277]. bicoherent [907]. biconfluent [2327, 1600, 370]. bicovariant [1170]. bidifferential [424]. bifractional [1098]. Bifurcation [871, 46]. Bifurcations [937]. Big [2024]. bilinear [1015]. billiard [1883]. billiards [1686, 1413, 1805]. Bilocal [240]. bimetric [2248]. binary [2294, 1504, 1354]. Binding [2015, 1086, 1063]. binomial [580, 761, 213]. biochemical [1039]. Biorthogonal [1525, 839, 831, 941]. Biot [2257, 1951]. Bipartite [854, 1040, 2229]. birdtrack [1093]. Birkhoffian [407]. bisection [725]. bispectral [1412]. bistable [303]. Bivariate [1880, 2057, 1739]. BKP [2126]. black [170, 568, 569, 1592, 3, 1696, 2249, 2326, 132, 1209, 1305, 449, 1358, 1756, 4, 561, 2070, 629, 673, 754]. bliss [1505]. Bloch [957, 709, 801, 1937]. Bloch-bundles [709]. block [414, 518, 1679, 2113, 503]. block-modified [518]. Blocks [1382, 1769, 1777, 501]. blood [1274]. Blow [251, 2284, 2040, 2121, 171, 1185, 372, 2175, 579, 189, 2012, 529, 1275, 2224, 2010, 305, 2256, 1394, 1724, 1611]. Blow-up [251, 2284, 2040, 2121, 171, 1185, 372, 2175, 579, 189, 2012, 529, 2224, 2010, 305, 2256, 1394, 1724, 1611]. Blowup [1530, 389, 1440, 1667, 342, 803, 829, 752, 1693]. Bob [1767]. bodies [449, 1217, 358]. body [1239, 2269, 2281, 1683, 1992, 1112, 1998, 768, 1555, 1480, 1904, 133, 1596, 937, 2097, 81, 2208, 902, 2096, 1032, 1594, 443, 1373, 176, 135, 1471, 1460, 160, 2085]. Bogoliubov [945, 420, 17, 2184, 2134]. Bogolyubov [176]. Bogomolny [576]. Bogomolńyi [770, 460]. Bohm [23, 476, 1291]. Boiti [63, 9, 1286]. Boll [1124]. Boltzmann [301, 42, 251, 2260, 580, 1361, 2049, 827, 1259, 1893, 790, 21, 2158, 1208, 306]. Bona [1900]. Bondi [1174]. bonding [1742]. Bonnet [1318, 1652]. bootstrap [1897]. Borel [1118]. Born [176, 1417, 1684, 395, 746, 774, 837]. Bose [295, 59, 1451, 669, 2184, 756, 1352, 2134, 2096, 1185, 2133, 162, 443]. boson [360, 438, 830, 221, 2182, 949, 1810, 250]. Bosonic [1360, 1939, 2324, 251, 1528, 831, 1496, 1375]. bosonization [473, 1176, 435, 359]. bosons [420, 673, 1992, 274]. both [313]. bottoms [1662]. Bound [68, 1084, 1357, 77, 187, 1622, 396, 697, 1214, 1024, 552, 1410, 1749, 1101, 1305, 549, 1624, 617, 851, 657, 1921, 725, 658, 1043, 1565, 135, 553, 499, 765]. boundaries [566]. Boundary [295, 2009, 1577, 21, 301, 410, 272, 1390, 1774, 1401, 1030, 116, 106, 2072, 2288, 1904, 1470, 1530, 1214, 385, 2110, 2170, 1831, 2216, 1614, 534, 2058, 2227, 1597, 1027, 979, 1732, 264, 838, 2158, 1824, 1490, 1881, 1908, 531, 207, 1279, 362, 928, 1296, 1947, 1946, 1965, 527, 1715, 2036, 2222, 2122, 71, 2173, 2162]. bounded [508, 948, 147, 1493, 858, 1951, 823]. Boundedness [1058, 545, 1906]. Bounds [1461, 1419, 1164, 1744, 579, 962, 1886, 1633, 1548, 2208, 1104, 1616, 1186, 2088, 1465, 2132, 1758, 1301, 553, 298]. Boussinesq [2321, 2010, 96, 1940, 1989, 373, 2048]. box [2207, 1505, 306]. box-ball [2207]. BPS [1241]. bracket [270]. brackets [409, 416, 704, 253, 160]. Braid [924, 440]. braided [2279, 446, 1944, 156]. Braids [231]. branching [1289]. brane [403, 2225, 2274, 1560]. branes [933, 1637]. Brans [288]. Breaking [317, 345, 185, 2210, 1817, 364, 1767, 1845, 1840, 1631]. breather [760]. breathers [1787]. bridge [1918]. bridges [1266]. bright [264]. Brillouin [1438, 85, 718]. brine [597]. broadband [1164]. broken [943]. Brownian [54, 584, 2255]. BRST [1590, 1928]. BS [1080]. Bubbling [1328]. Buckley [643]. building [1679]. Bulk [494, 2186, 2167]. bump [844, 349]. Bunch [1965]. bundle [2189, 2267, 1260, 909]. bundles [709, 1835, 903, 1489, 531, 114]. Burger [1952]. Burgers [641, 642, 882, 1282, 1211, 1273, 644, 15]. Burghelea [1439, 498]. butterfly [1621]. BV [1659, 599]. BV-formalism [1659].

C [1298]. Cabibbo [626]. Calculation [1670, 259]. calculi [1048]. calculus [1012, 565, 1441, 424, 1118, 573, 1170]. Caldirola [1745]. Calogero [2153, 990, 1771, 1835, 2130, 2007]. calorons [1719, 1644]. Camassa [345, 914, 1873, 778, 1440, 908, 1574]. Campbell [591, 940]. can [1724]. Canonical [731, 482, 184, 589, 704, 717, 452, 890]. canonizations [943]. capacities [1301]. Capacity [1886, 2104, 505]. capture [2064]. Caputo [1572]. Carathéodory [1966]. carbon [546]. Cargo [2168]. Cargo-LeRoux [2168]. Carleman [415]. Carpentier [192]. carried [1627]. carriers [805, 2042]. Carroll [933]. Cartan [570, 1697, 1443, 818, 2274]. Cartesian [1456]. cascade [1387]. cascades [1706]. Case [304, 286, 202, 1030, 2317, 2294, 709, 1854, 675, 581, 221, 1445, 355, 1334, 808, 252, 843, 930, 751, 1469]. cases [376]. Casimir [1943, 2181]. Catalan [1931]. categorical [29, 903, 1489]. categories [2295, 1070, 1404, 271, 446, 1944]. Category [2266, 776]. Cattaneo [2328, 1433]. Cauchy [1017, 1640, 1676, 1910, 637, 817, 1209, 1568, 1834, 790, 2118, 2052, 67, 499, 2048]. Causal [284, 1645, 1383, 1535, 1930]. Causality [1932, 1258, 131, 2300]. cavitation [824]. cavities [927]. cavity [1520]. cell [445, 1603]. cellular [1931]. Center [1826, 43, 2269]. center-of-mass [2269]. centers [46]. Central [1116, 2133, 735, 245, 302, 1596, 1625, 1994, 1974, 1034, 83]. centrally [122, 650]. centrifugal [1293, 608, 7]. certain [572, 472, 457]. CGL [1282]. Chain [88, 475, 606, 1419, 1345, 1115, 1083, 1776, 2149, 1845, 1840, 2036]. chains [233, 2207, 700, 400, 1242, 2001, 605, 83]. Chandrasekhar [2121]. change [1421, 1507]. changed [1984]. changing [759, 1092, 1493, 2056, 2259, 2293]. channel [1302, 1964, 2021, 788, 1633, 455, 1082]. channels [1924, 597, 1815, 2104, 2299, 948, 513, 505, 1408, 961, 1817, 1547, 621, 1963, 1106, 556, 555, 1587, 1885, 1301, 1527, 2138, 1812]. chaos [575]. Chaotic [1470, 927, 1513, 244, 243, 952, 2078]. Chaplygin [2168, 1734, 920, 2115]. character [2289, 1332]. characterisations [1750]. Characteristic [1137, 56, 1260, 1166, 2220, 2069, 252]. characterization [1485, 1969, 1078, 1079, 1709]. characters [987, 2181, 883]. charge [2072, 805, 578, 989, 449, 2042, 1804, 1928, 2297]. charge-monopole [1804, 2297]. Charged [60, 852, 1109, 1356, 101, 1032, 154, 1994, 2070]. charges [1024, 1823, 667, 710]. chase [1344]. Chebyshev [259, 1159]. chemical [1742]. chemicals [1324]. chemoattractant [980, 308]. chemotactic [1901]. chemotaxis [645, 980, 2090, 843, 1724, 308, 1324, 1321]. chemotaxis-fluid [645]. chemotaxis-Navier [2090]. Chern [432, 268, 390, 129, 130, 146, 789, 1956, 1764, 1042, 2135, 1165, 698]. chi [40]. chi-squared [40]. choice [2141]. choosing [557]. Choquard [2175, 1870, 981, 2055, 793, 1731, 1911, 1725, 2044]. Choquard-type [2055]. circle [1926, 1848, 1585, 77, 1805, 2162]. circuit [901, 2271]. circuits [1544, 487, 1198, 1196]. circular [1778, 1883]. Circumspheres [1114]. CKP [1176, 1355]. class [1499, 2175, 2239, 1059, 918, 1203, 197, 1412, 202, 736, 773, 812, 245, 302, 266, 888, 1444, 1494, 1726, 1156, 181, 1865, 464, 1359, 1229, 1625, 1381, 1018, 1874, 1985, 1294, 1449, 872, 946, 1578, 1675, 2237, 34, 864, 1672, 1723, 1130, 2259, 1276, 351, 984]. classes [572, 5, 427, 1780, 1260]. Classical [426, 1302, 1198, 1952, 1759, 14, 47, 58, 539, 1529, 2269, 1949, 1924, 2179, 2137, 705, 1964, 1623, 1245, 105, 937, 1729, 2148, 1773, 281, 1847, 409, 1310, 136, 1009, 1984, 131, 2102, 505, 260, 1108, 1336, 1739, 786, 886, 315, 1933, 521, 1562, 2318, 979, 1191, 164, 118, 972, 30, 2256, 1208, 1727, 1262]. Classical-quantum [1302, 315]. classically [290]. Classification [1404, 889, 683, 2323, 1641, 2322, 689, 1524, 607, 1195, 2319, 13, 19, 2006, 1067, 986, 2023, 464, 1359, 832, 1356, 1206, 452, 1675, 1691, 1130, 885]. classifications [1878, 1734]. Clausen [286]. Clebsch [2206, 1289]. Clifford [316, 1689, 911, 1014]. Clifford-cyclotomic [316]. climate [530]. cloak [1571, 161]. cloaking [1164]. close [2060]. Closed [1277, 817, 549]. closure [1310, 942]. cluster [1938, 998, 250, 897, 2143]. clustering [1205]. Clusters [747, 1684]. co [1594]. co-existence [1594]. coagulation [946]. coagulation-fragmentation [946]. coalescing [1137, 1249]. coalgebras [446]. Coarse [1691]. Coarse-grained [1691]. coaxial [1411]. cocycle [206]. cocyles [884]. coding [2196]. coefficient [1168, 1203, 532, 914, 739, 1576, 1615, 754, 1526, 1276, 1901]. coefficients [1401, 226, 2241, 2260, 366, 1726, 378, 513, 1289, 163, 1967, 2118, 1900]. Coexistence [1220]. coherence [855, 1409]. Coherent [242, 734, 296, 1657, 811, 1014, 336, 615, 746, 774, 837, 2060, 276, 481, 2100, 550, 1126, 2212, 1792, 848, 1177, 299, 56, 2063, 911, 1846, 1546, 313, 1688, 154, 540, 2271, 1626, 1155, 2087]. Coherent-state [615]. coherently [66]. cohomogeneity [966]. cohomologies [705, 689]. Cohomology [1022, 1802, 1590, 1068, 153]. cohomotopy [403]. Colbeck [478]. collapse [1980, 314, 618, 1617, 1981]. Collection [2314]. collective [2307]. collinear [1596, 1032]. collisions [443]. collocation [674, 1281]. colored [1107]. Combinatorial [285, 1481, 1246, 845, 1069]. combined [2107, 2055]. combing [2279]. combustion [2286]. coming [1085]. Comment [841, 426, 1980, 915, 2323, 2325, 2327, 641, 63, 608, 2322, 1981, 1319, 1358, 1320, 640, 2042, 607, 1130, 642, 8, 2087, 339]. Comments [1358, 4]. communication [1924, 2137, 697]. Commutation [1799, 1194]. commutative [2317, 1258, 502, 1741, 519, 2234]. commuting [2127, 997, 1192, 203]. commuting-operator [997]. comodules [156]. Compact [1094, 1862, 983, 1592, 2123, 705, 683, 1314, 1289, 107, 1835, 500, 2318, 151, 1043, 1565, 1474, 2191, 310, 774]. compact-support [107]. compactified [1522]. compactly [80]. companion [2001]. Comparing [2078]. Comparison [1963, 1878, 470, 1395]. compass [1026]. compatible [1538]. competing [610]. complementarity [510]. Complete [1013, 1145, 994, 2310]. completely [508, 1699, 545, 1422, 521, 2229]. completeness [1861]. Completing [2025]. Completion [1829, 184]. completions [601]. Complex [1280, 496, 185, 1529, 777, 2034, 1341, 1641, 1553, 558, 35, 660, 2312, 2040, 1368, 1856, 313, 1311, 1082, 1290, 48, 721, 1487, 1469]. Complex-time [496]. complex-valued [1082]. complexes [1332]. Complexification [1781]. Complexiton [1283]. Complexity [2161, 1337, 1319, 12, 1320]. component [326, 440, 179, 1355, 923, 264, 644, 1019, 1135, 1049, 2036]. components [665, 800]. composed [665]. composite [1837, 225]. compositional [634, 1740]. compound [812]. compressible [116, 791, 1530, 748, 2276, 1729, 342, 103, 1492, 822, 262, 1955, 2089, 1875, 792, 147, 1891, 922, 2124, 1396, 1869, 2285, 524, 979, 1733, 2122, 1987, 1089, 1476]. computation [128]. computational [2216]. compute [915, 836]. Computing [2132, 453, 1289, 2329, 940]. concave [2092]. Concentration [1922, 920, 918, 1203, 224, 1464, 449, 1064, 1448, 2084, 70, 1202, 1201, 2044]. concept [131]. concepts [1045]. conceptual [2059, 2320]. concerning [1293]. concise [1104]. condensate [1185, 443]. condensates [756, 2133]. condensation [2295, 295, 59, 669]. condensed [758]. condition [914, 1904, 747, 260, 2227, 571, 678, 1452, 1967, 1446, 559, 2162]. conditional [90, 1227, 2102, 86, 1381, 1885, 1301]. conditioned [582]. Conditions [961, 897, 1390, 1966, 1690, 106, 1470, 2292, 1214, 2110, 1395, 534, 1254, 289, 1004, 1807, 1872, 1597, 1057, 1074, 264, 838, 21, 301, 2158, 1824, 410, 531, 1279, 362, 928, 1296, 1947, 1946, 527, 2173]. conductance [183]. conducting [822, 666, 358]. conduction [1149]. conductive [342]. conductivity [196, 546, 1987]. cone [445]. configuration [1642, 1288]. configurations [2161, 1596, 1034]. confined [1367, 1918, 2060, 1992, 1368]. confinement [388, 594, 1669]. confluent [1808, 1566, 2252]. Conformal [371, 1053, 1958, 2303, 1965, 122, 970, 650, 1943, 1231, 461, 1030, 889, 327, 799, 2, 428, 563, 1677, 1970, 269, 60, 1777, 102, 867, 501, 964, 2141, 2074, 1067, 153, 1335]. conformally [1652]. Congress [2314]. Congruence [1643]. congruences [1341]. conical [2019]. conifolds [1077]. conjecture [1306, 465, 516, 435, 520, 192]. conjectures [1816, 2290]. Connected [800, 1480, 1139, 903, 1489]. Connection [1849, 691, 2060, 157, 1344, 1777, 1243, 1751]. Connections [570, 1697, 1367, 1772, 384, 903, 1489]. Connes [465, 167, 157, 516, 1376]. Conormal [1441]. conservation [1483, 386, 470, 337, 832, 117, 125, 712, 1242, 1050, 2117, 2053]. Conservative [67, 1517, 526]. conserved [365, 1550]. considerations [2132, 746]. Consistency [764]. consistent [1918, 1531, 52, 1427]. Constant [113, 42, 569, 2282, 733, 1125, 1707, 1072, 1259, 1354, 2243, 273, 1226, 41]. Constantin [1491]. constants [469, 768]. constitutive [216]. Constrained [1304, 505, 1256, 1971]. Constraint [224, 1266, 1757, 2206, 1504, 1695]. constraints [2280, 1333, 459, 1993, 1414]. constructed [586]. Constructing [1945, 602, 2216, 140, 1387]. Construction [276, 1450, 365, 359, 1088, 1918, 2060, 1942, 1521, 988, 470, 477, 1134, 2203, 1991, 1002, 1578, 1050, 178]. constructions [1145]. consumption [980, 1901, 308]. Contact [901, 1683, 292, 1244, 524, 1291]. contain [167]. containing [1487]. content [2197]. Context [127, 950]. Context-invariant [127]. context-selection [950]. contextuality [1632]. continua [1181]. continuation [1362]. Continuity [670, 506, 1548, 1882, 44, 1228, 1301]. Continuous [2167, 1236, 296, 1648, 1011, 1574, 94, 694, 733, 2146, 1416, 2051, 317, 1856, 2264]. Continuous-time [1011]. continuum [1434, 456, 1137, 553]. Contraction [513, 136, 1213]. Contractions [1564]. Contractivity [1000]. Contragredient [16, 916]. contrast [1762, 2262]. contribution [441]. control [360, 56, 1577]. Controllability [1582, 1401, 1680]. controlled [943]. Controlling [222]. convection [1910, 429, 1264]. convective [2276]. Convergence [107, 1861, 2263, 955, 619, 497, 1497, 1083, 2079, 555, 308]. convergent [637, 1017, 51]. converse [813]. convertibility [1464]. Convex [506, 1786, 509, 2092]. Convolution [2228]. Convolutions [1479]. cooling [454]. Cooper [523]. coordinate [53]. coordinates [1456, 2269, 1153, 1833]. coprime [99]. Coprimeness [1007]. Coprimeness-preserving [1007]. corepresentations [2057]. corners [1637]. Corrected [247]. correction [1598, 1522, 1997]. Corrections [1340, 368, 91]. correlated [49, 107, 2061]. Correlation [738, 893, 1833, 585, 465, 516, 962, 282, 1776, 2149, 244, 243]. correlations [205, 177, 363, 1361, 214, 383, 1507, 952, 213, 1500]. correspondence [1173, 699, 1507, 1562]. corresponding [206]. corrugated [1082]. Coset [1002, 2157]. cosine [304, 1650]. Cosmic [935]. Cosmological [567, 42, 569, 1125, 817, 1209, 1305, 827, 1259, 1226, 41]. cosmology [1075, 767, 369, 1864, 967, 1510, 370, 2247]. cost [697, 1922, 2273]. costratification [1765]. Couette [1082]. Coulomb [1720, 1630, 11, 198, 2009, 1754, 1849, 2194, 1808, 1509, 2255, 1186, 693, 890, 808, 2240, 667, 2188]. Coulomb-like [2194, 808, 2240]. Coulomb-type [2009]. count [486, 1595, 1290]. counter [2018]. counter-factual [2018]. counterexamples [1807]. counting [618]. Coupled [909, 1819, 1323, 1953, 396, 2101, 2045, 631, 1162, 35, 422, 913, 1383, 998, 706, 150, 877, 346, 1204, 1076, 1025, 70, 629, 408, 190, 66, 834, 222, 348, 677, 627]. coupling [768, 2172, 219, 1071, 1040, 221, 1829, 389, 943, 2215]. couplings [52, 782]. Courant [368]. covariance [332]. Covariant [1048, 692, 1550, 409, 505, 1106, 632, 2074, 1527, 1378]. covering [5, 427]. covers [1780]. Coxeter [1803]. CP [113, 460]. creation [227, 942]. creation/annihilation [942]. creep [1674]. criteria [1218, 1647, 1556, 978]. criterion [200, 342, 189, 99, 535, 2010, 930, 2256, 1440]. Critical [1870, 828, 779, 17, 945, 1444, 2091, 756, 224, 1392, 2054, 1630, 1950, 1612, 2238, 1558, 787, 1064, 1018, 1523, 149, 921, 150, 648, 1915, 1445, 1820, 1877, 1202, 1839, 1673, 148, 751, 2044]. criticality [1647]. cross [1122]. cross-sectional [1122]. crossings [1195, 2319]. Crystallization [1173]. crystallographic [1803]. crystals [784, 1492, 331, 1288, 529, 2267]. cubic [2174, 679, 910, 1700, 778]. cubic-quintic [2174, 1700]. cubics [2306].

Cucker [1757, 2146, 1169]. Cummings [118, 1713, 2268, 719, 1680]. cumulants [825]. curl [1983, 985]. Current [2041, 240]. Currents [1810, 261, 1550]. Curvature [902, 1113, 59, 255, 1778, 1112, 113, 733, 249, 1260, 732, 273, 93, 1020]. curve [2178]. Curved [417, 500, 248, 2257, 1031, 2302, 937, 609, 1634, 659, 890, 1099, 1536, 1928, 135, 1471]. curves [549, 1751, 627]. Curving [1187]. cut [727, 1816]. cutoffs [1694]. cycle [1425, 1287]. cyclic [1171]. cyclotomic [316]. cylinder [1955, 995, 1820, 2141]. cylinders [303]. cylindrical [288, 1469].

D [1727, 170, 1890, 1992, 798, 382, 1532, 1240, 1560, 1173, 461, 1337, 1150, 2236, 227, 2016, 1210, 1729, 103, 1162, 2194, 1984, 1614, 1759, 347, 1846, 1985, 2124, 2255, 2171, 762, 221, 979, 1764, 1042, 70, 377, 824, 1373, 252, 843, 800, 1733, 2052, 2037, 2188, 1876, 2291, 1724, 823, 1704, 351, 795, 2048]. D-mappings [2236]. damped [841, 1516, 1, 22, 1846, 234, 1746, 1611, 1989]. damping [2013, 2091, 1985, 1869, 917, 1843, 525, 795]. Darboux [237, 863, 1386, 1643, 2155, 908, 834]. Dark [264, 1153, 693, 1379, 2025]. Dark-bright [264]. Darwin [786]. data [630, 914, 189, 2303, 792, 2264, 979, 1791, 2287, 1019, 823, 1574, 1491]. data-to-solution [1574]. Datta [724]. Davey [739, 780, 610]. Davies [1965]. day [1358]. DeAngelis [1467]. Decay [187, 177, 524, 1500, 2032, 795, 1476, 1709, 532, 2067, 1255, 1285, 1666, 661, 953, 922, 2287, 1440, 682]. decaying [200, 1705]. Decoding [814, 725]. decoherence [375, 2144]. decoherence-free [2144]. Decomposability [1814]. decomposable [194]. Decomposition [393, 204, 671, 445, 1458, 437, 897, 2077, 1181, 421, 216]. decompositions [462]. decoupling [1506]. decreased [2060]. defect [577]. defects [2294]. defined [861, 833, 2229, 1921]. definite [1059, 444, 984]. definiteness [2018]. Definition [969, 2060, 1222, 1572]. deflection [2081, 1511]. defocusing [264]. Deformation [847, 1795, 1367, 1818, 2157, 1619, 1618, 1958, 1187, 924, 270, 2162]. Deformations [1335, 2023, 33, 25, 173, 1830]. deformative [439]. Deformed [990, 236, 890, 1536, 1453, 2017, 483, 1173, 580, 1073, 1194, 245, 302, 805, 1410, 299, 1884, 1982, 1970, 2218, 380, 2042, 808, 1373, 1533]. deforming [2233]. degeneracy [1490]. degenerate [1271, 913, 1891, 2285, 1304, 1732, 1379, 2025, 1437, 2106]. degeneration [228, 1387]. deGennes [331]. degree [1595]. delay [1798, 2013, 1757, 927, 1857, 1569, 1515, 1228, 917, 244, 243]. delayed [1402]. delays [374]. delocalization [1968]. Delta [2053, 1907, 606, 1406, 974, 1458, 551, 2114, 904]. delta-function [606, 904]. delta-perturbed [551]. Delta-shocks [1907]. Demkov [953]. denominators [1934]. Dense [2196]. densities [240]. Density [1136, 1160, 894, 578, 989, 171, 162, 2048]. density-dependent [2048]. density-matrix-functional [989]. dependence [2264, 1019, 1491]. dependent [1168, 720, 36, 2254, 1580, 550, 1151, 994, 737, 636, 766, 299, 2110, 455, 2209, 576, 1299, 1348, 1759, 1682, 300, 1541, 1517, 1294, 1057, 159, 1240, 952, 1993, 34, 1414, 38, 1804, 1357, 1863, 485, 1906, 1901, 1973, 1155, 124, 2048, 1533, 2297]. depolarizing [854, 555]. depression [2212]. dequantization [810]. Derivation [1992, 1662, 2096, 2329, 1398, 277, 442, 1893]. derivations [545, 158, 885, 2014, 1335]. derivative [254, 1752, 1231, 969, 1463, 422, 1610, 974, 134, 1572, 102, 867, 291, 1445, 1511]. Derivatives [1197, 1059, 1462, 821, 869, 2218]. derive [1226]. derived [1430]. descent [1903]. describing [1142]. description [1367, 2060, 2167, 841, 1, 716, 1249, 1519, 1055]. descriptions [901, 1288]. designs [1105]. deSitter [2035]. Detailed [1459, 2067]. Detecting [314, 1507, 1300]. detection [722, 893]. detectors [240]. determinant [585, 282, 2129]. determinantal [1934]. determinants [98, 1972, 1435, 498, 1439]. Determination [2159, 1511]. determine [514]. determined [912]. Deterministic [1142]. Development [1916, 1122]. deviation [772]. deviations [875, 105, 107]. DFR [601]. DFR-algebra [601]. diagonal [73, 1310, 1105]. diagonalization [335, 949]. diagram [46]. Diagrammatics [1107, 2140]. diagrams [1638, 1480, 1921]. diameter [1305, 1043, 1565]. diamond [1851]. diatomic [395]. Dicke [288, 118]. Dickey [1895]. dielectric [1837, 441, 1526]. diffeomorphism [639]. diffeomorphisms [2000]. difference [142, 321, 419, 872, 418]. different [2327, 1600, 2229, 62, 527, 2306]. Differentiability [363]. Differential [709, 1244, 1966, 1403, 650, 1655, 796, 1599, 1362, 879, 1443, 915, 836, 745, 1048, 1347, 707, 188, 976, 741, 61, 95, 1472, 403, 1389, 1120, 2083, 181, 111, 1831, 1979, 419, 1834, 1170, 1685, 1221, 1326, 1142, 880, 1131, 1090, 1399, 457, 1766, 184, 1056, 491]. differential-difference [419]. diffraction [1285]. diffusion [1327, 2164, 226, 387, 645, 1000, 783, 2277, 429, 2170, 869, 1569, 1515, 2227, 2145, 458, 1011, 1221, 303, 2201, 143, 114, 530, 2032, 1467, 1727, 374, 2002]. diffusion-wave [143]. Diffusions [582, 378, 1052]. diffusive [2082, 1910, 646]. digital [1524]. dilations [2305]. dilaton [1113]. dilogarithm [294]. dilute [2184, 2134, 1349]. dilution [1464]. Dimension [1338, 1179, 820, 1683, 738, 1463, 625, 789, 1956, 206, 1783, 891, 938]. dimensional [970, 650, 2017, 1401, 2116, 1983, 1676, 268, 1337, 1030, 33, 426, 440, 889, 142, 69, 1664, 1031, 388, 633, 1470, 1737, 579, 791, 470, 1995, 1341, 288, 1457, 510, 1407, 1083, 2097, 185, 1619, 2134, 1749, 63, 1699, 2080, 1009, 1808, 1101, 1643, 505, 218, 602, 1184, 1809, 1481, 9, 112, 1007, 2313, 1141, 1624, 1372, 1004, 811, 225, 1072, 1501, 2098, 1114, 786, 55, 1193, 44, 1396, 405, 1003, 704, 14, 467, 291, 1954, 163, 1628, 1186, 2243, 964, 199, 273, 712, 2038, 1330, 1791, 1502, 1309, 851, 2064, 1919, 702, 613, 21, 301, 2158, 808]. dimensional [656, 1357, 909, 972, 809, 1460, 1908, 2006, 207, 1712, 1906, 212, 527, 1080, 885, 1135, 1175, 2122, 230, 1089, 1286, 1034, 698, 627, 77, 1805, 2162]. dimensionless [1705]. dimensions [286, 1620, 1306, 2281, 541, 2250, 1636, 1531, 1841, 1278, 1696, 980, 1225, 2089, 1625, 166, 1384, 1174, 872, 2004, 1993, 62, 1414, 623, 1500, 1080, 2198]. Dimer [1775, 826, 2251]. Ding [1173]. Dinucleon [1329]. dipolar [2284]. dipole [666, 762]. dipoles [1024, 657]. dipositronium [560]. Dirac [1720, 608, 1199, 606, 1398, 11, 820, 33, 1909, 1531, 396, 2009, 764, 2139, 1393, 393, 415, 992, 1630, 1424, 282, 896, 975, 744, 1363, 7, 1589, 1652, 168, 166, 2214, 1027, 1186, 1068, 1084, 451, 1357, 2155, 32, 1152]. Dirac- [282]. direct [218, 930]. directed [2004]. direction [447]. Directional [985]. directly [1819]. Dirichlet [1665, 2235, 713, 298, 2162]. Dirichlet-to-Neumann [713, 2235]. disappearance [1258]. discontinuous [1027]. Discrete [1241, 2233, 1285, 1630, 1650, 838, 312, 1461, 1110, 413, 2146, 333, 675, 2111, 2051, 1096, 183, 376, 1007, 99, 2313, 1653, 140, 706, 2151, 835, 700, 294, 1295, 1085, 1866, 84, 1563, 2065, 943, 1718, 1387, 1159, 1708, 2007, 2305]. discrete-time [1110, 2305]. discretely [2176]. discretisation [553]. discretization [179, 840, 1314, 1700, 2007]. discriminant [1780]. discrimination [961]. Discussing [2218]. discussion [1540]. disguise [1751]. disk [2162]. disk-lens [2162]. disorder [49, 584]. disordered [475, 1499, 548, 1498, 1496]. dispersal [1402]. Dispersion [1433, 2328, 486, 1349, 627]. Dispersionless [876, 1841]. Dispersive [2311, 1520]. dissipation [532, 1664, 1039, 878, 189, 1645, 1033, 1705, 1196]. Dissipative [729, 2289, 399, 82, 492, 906, 2258, 2010, 1519, 2029, 176, 372, 682]. distance [157, 2149, 1551, 135, 1375]. Distillation [812]. distinction [2192]. distinguishability [1997]. distributed [2013]. Distribution [1942, 2147, 480, 518, 1151, 1121, 1598, 801, 1128, 213, 1826]. distributional [2110, 2117]. Distributions [1714, 1786, 580, 1441, 761, 1158, 734, 107, 1413, 1432, 1138]. disturbance [2192]. divergence [2067]. divergences [1813, 1923]. DKP [2126, 1766]. DNLS [1795]. Does [1290]. domain [106, 547, 1543, 661, 147, 1951, 883, 1727]. domains [1529, 1036, 347, 2093, 1493, 2201, 823, 1469]. dominant [1486, 1936]. domination [297]. doomsday [935]. Doppler [2112]. Doss [1123]. Double [258, 1281, 435, 187, 1385, 1702, 1910, 2125, 674, 1458, 1501, 1871, 2114, 278, 2252]. Double-bosonization [435]. double-diffusive [1910]. double-oscillator [278]. double-well [187, 1501]. doubles [73]. Doubly [1456, 556]. Dresselhaus [38]. Dressing [531, 700]. Drinfeld [2125, 28, 689, 156]. driven [1966, 1347, 722, 794, 640, 176, 10]. Driving [1215]. drop [1313, 2119]. Droplet [1579, 1419]. dS [1078, 1079]. DtN [1060]. Dual [638, 2073, 1865, 1736, 2151]. dual- [1736]. Duality [1227, 2208, 326, 413, 699, 327, 1621, 1481, 2178, 686, 2181, 2246, 474]. dually [901]. due [449]. Duffin [1041, 1384, 2198]. Duffing [493, 1700]. Duistermaat [1926]. Dullin [2224, 372]. Dunkl [412, 2228]. dust [1703, 1109]. Dyck [175]. dynamic [1245, 2149, 758]. Dynamical [1506, 1447, 959, 451, 603, 553, 2086, 172, 570, 1697, 1599, 1110, 2254, 2200, 1261, 1255, 82, 1229, 1682, 1776, 2286, 640, 546, 708, 10, 100, 2078, 2305]. Dynamics [664, 1157, 2211, 787, 851, 2064, 1919, 1469, 43, 955, 735, 1798, 2261, 1178, 1992, 1140, 2233, 2288, 1497, 740, 707, 1421, 2139, 1310, 1645, 2030, 1137, 1584, 1103, 2051, 1002, 1249, 635, 2145, 2024, 125, 1253, 1680, 1326, 290, 485, 1020, 212, 1718, 1135]. dyonic [170, 568, 1592]. dyons [568, 1592, 638]. Dyson [905, 1069, 2152].

E-H [1844]. E8 [2111]. Eckart [151]. Eckhaus [1613]. Eden [495]. edge [1496]. Editor [1840]. Editorial [1200]. effect [990, 2071, 394, 2013, 2302, 1537, 85, 1438, 1857, 352, 1929, 670, 353, 2066, 421, 1291]. Effective [805, 716, 1348, 2042, 114, 1020, 1960, 720, 1938, 1151, 1298, 485, 785, 1927, 2297]. Effects [594, 443, 295, 1851, 1137, 919, 1876]. Ehrenfest [1138]. eigenfunction [282, 1413]. eigenfunction-correlation [282]. eigenfunctions [1884, 1982, 1954, 1074, 223]. eigenmodes [801]. Eigenstates [1584, 475, 994, 1249, 1536]. Eigenvalue [744, 69, 1598, 1265, 2110, 929, 910, 298]. Eigenvalues [414, 23, 459, 430, 1774, 312, 444, 871, 1602, 1137, 1281, 747, 1297, 1979, 995, 581, 1954, 551, 1290, 2156, 2165, 1296, 1947, 1021]. eigenvector [1968]. eigenvectors [2062]. eigenwaves [1349]. eikonal [1343]. Einstein [197, 295, 59, 42, 1639, 568, 569, 1592, 2123, 669, 599, 756, 327, 625, 1512, 2232, 1811, 2248, 1259, 1821, 319, 1185, 208, 2133, 1113, 443, 41]. Einstein-scalar [2123]. Eisenhart [1031]. Eisenstein [1702]. elastica [797]. elasticity [371]. elastostatics [1053]. electric [261, 2107, 2325, 241, 1292, 901, 1541]. electro [1929]. electro-gravitational [1929]. Electrodynamics [101, 1636, 174, 638, 406, 1509, 868, 560, 1694]. electromagnetic [1549, 1230, 784, 1161, 1520, 2076, 1356, 681, 289, 149, 1263, 358, 169, 216]. Electromagnetism [951]. electron [2016, 168, 441]. electrons [1651, 1500]. electrorheological [1490]. electrostatics [1122]. Elementary [1274]. elements [2127, 394, 544, 1603, 537, 2188, 474]. elephant [1116]. Ellipse [581]. Ellipsoidal [2120, 790]. ellipsoidal-BGK [790]. Elliptic [1710, 905, 1638, 286, 963, 876, 779, 1362, 1665, 612, 1054, 181, 1972, 1950, 2238, 2129, 1934, 18, 932, 921, 648, 2056, 1604, 2263, 1839, 1673, 1122, 1062, 339]. elliptical [1435]. embedded [1774, 1338, 995, 2156]. embedding [1005, 1593, 465, 516, 1376]. embezzlement [999]. Emergence [213, 788]. emerging [635]. Émery [567]. emission [1535]. EMRIs [1340]. endomorphisms [625, 319]. energetics [2075]. energies [46]. Energy [1233, 532, 1706, 952, 526, 453, 170, 1429, 74, 2250, 1426, 773, 24, 791, 1136, 1160, 224, 1071, 191, 2134, 2284, 1281, 1962, 2099, 81, 1092, 2208, 1396, 1873, 1915, 405, 2329, 163, 551, 632, 1357, 1863, 2155, 1723, 904]. Energy-dependent [952, 1357, 1863]. energy-momentum [632]. enhanced [1854]. enlarged [620]. Enneper [2202]. ensemble [258, 48]. ensembles [1255, 1996, 1350, 1169, 2079, 1308, 58, 1826]. entangled [2189, 2137, 313, 1452, 400, 695]. Entanglement [1499, 2197, 1026, 1817, 364, 621, 1961, 475, 957, 1419, 729, 999, 1218, 2270, 105, 1407, 2272, 1633, 507, 1300, 1634, 1464, 998, 1507, 1631, 1921, 1691, 623, 89, 2273]. Entanglement-breaking [1817, 364]. Entanglement-saving [621]. Entire [303]. Entropic [783, 1309, 1708, 624, 1721, 2191]. entropies [87, 1255, 88, 813, 1692, 1961]. Entropy [852, 696, 1319, 246, 12, 1320, 556, 2086, 2298, 1419, 580, 1146, 515, 724, 3, 772, 1421, 1548, 1634, 1831, 1810, 86, 1358, 1969, 4, 555, 1885, 506]. envelope [2099, 391]. enveloping [986]. Environment [375, 2277, 1747]. epidemic [646]. EPR [2018]. equation [1390, 1774, 1168, 876, 312, 1183, 1867, 2043, 187, 1325, 186, 779, 1327, 1203, 1401, 680, 2161, 42, 673, 57, 1483, 1890, 251, 1639, 304, 1772, 1182, 2260, 1793, 228, 1466, 1599, 1531, 1992, 806, 643, 532, 1370, 1559, 388, 2236, 245, 302, 1347, 2323, 633, 209, 914, 297, 144, 981, 882, 1251, 1457, 1841, 2223, 878, 1463, 2016, 1015, 1667, 1006, 1442, 179, 2327, 191, 775, 641, 803, 742, 2221, 63, 1865, 1699, 1013, 608, 1433, 2328, 1567, 1123, 344, 1265, 1124, 492, 1447, 1613, 1666, 2012, 307, 1808, 429, 381, 1882, 2170, 2050, 906, 1610, 2253]. equation [859, 529, 2049, 1952, 466, 2322, 1828, 7, 1184, 1395, 974, 9, 350, 1614, 22, 2040, 1007, 99, 256, 1600, 1141, 1132, 869, 2311, 1864, 1372, 968, 480, 347, 1501, 1189, 827, 419, 706, 1287, 1307, 1275, 1448, 1874, 1856, 2008, 149, 794, 458, 1128, 134, 265, 648, 1228, 1873, 1903, 2084, 493, 1384, 2224, 835, 1311, 771, 1011, 1883, 1834, 923, 1088, 917, 291, 163, 1628, 1578, 762, 1014, 361, 1366, 1445, 2289, 199, 273, 171, 1670, 1343, 607, 790, 1391, 2114, 1084, 1397, 1993, 264, 838, 1877, 2304, 2118, 1986, 234, 781, 1731, 2158, 1900, 1414, 1142, 1357, 2155]. equation [708, 1202, 1446, 96, 900, 349, 408, 1490, 143, 667, 176, 1551, 1062, 1563, 1911, 67, 13, 19, 2259, 1488, 1908, 1766, 15, 642, 778, 1208, 1712, 1843, 668, 1718, 306, 908, 938, 1049, 1611, 1715, 1201, 1725, 2293, 1574, 372, 804, 2044, 682, 1469, 1491, 343, 1091, 2297, 2198, 77, 2162]. equation-Case [304]. equations [1795, 1966, 1478, 650, 1655, 2175, 261, 1323, 1949, 345, 391, 469, 1398, 1207, 2123, 1909, 736, 1953, 116, 1561, 1607, 1362, 879, 1443, 396, 777, 1941, 386, 1698, 745, 2067, 488, 2082, 2169, 2073, 193, 266, 585, 1127, 1470, 1444, 1668, 1730, 2278, 2045, 579, 791, 1530, 1053, 748, 533, 424, 2280, 227, 1484, 707, 188, 976, 741, 764, 612, 1392, 61, 1210, 1729, 1393, 1278, 327, 95, 1269, 103, 2055, 1350, 35, 1661, 2030, 1389, 1120, 2083, 111, 1950, 2080, 210, 1571, 383, 1513, 333, 675, 1234, 2111, 1124, 1576, 1615, 422, 780, 1282, 913, 1570, 2307, 429]. equations [1149, 1081, 285, 352, 2258, 2089, 789, 1568, 2094, 832, 1008, 147, 1831, 376, 534, 2313, 416, 18, 126, 1662, 1069, 1272, 1624, 1979, 1206, 300, 1671, 1891, 1695, 793, 752, 1064, 1018, 1693, 1523, 1569, 1515, 2093, 1985, 2124, 2227, 2145, 921, 150, 1396, 1869, 1872, 2171, 535, 2142, 346, 2285, 1493, 1728, 524, 640, 858, 1821, 117, 2010, 397, 666, 1082, 353, 1756, 1204, 1232, 1190, 501, 1468, 679, 1240, 1994, 1211, 979, 1732, 1221, 2152, 305, 1838, 1273, 1316, 1242, 208, 377, 824, 1085, 1866, 1326, 1675, 864, 58, 1672, 21, 301, 460, 1734, 1556, 1863, 263, 303, 2240, 2201]. equations [2165, 880, 1131, 1090, 910, 10, 84, 2256, 644, 457, 647, 354, 418, 2252, 978, 1937, 2006, 1276, 272, 184, 148, 526, 1322, 1279, 1876, 121, 1387, 527, 1973, 1056, 222, 491, 920, 1167, 2032, 2115, 939, 2122, 883, 823, 1704, 351, 339, 341, 751, 984, 1016, 1907, 2027, 1989, 677, 2173, 2048, 1283, 374, 2002, 2315, 966]. equatorial [1825]. equidistant [1587]. equidistribution [1362]. equilibria [1933, 135]. equilibrium [196, 1083, 1707, 865, 2119, 214, 2096, 2079]. Equivalence [1480, 1576, 1615, 1276, 57, 2033, 726, 835, 1311, 903, 1489]. equivalent [434, 1572]. Equivariant [2265, 1862, 1239, 271, 1926]. ergodic [759, 455]. Ergodicity [701, 1248]. Ermakov [1610]. Erratum [2043, 145, 1720, 945, 916, 2324, 428, 1389, 1017, 2328, 1438, 119, 2320, 1199, 1359, 427, 1439, 842, 678, 1489, 301, 1839, 2321, 2319, 19, 1947, 1946, 607]. error [1997, 855, 1758, 553, 2273]. essential [731, 1832]. estimate [1168, 713, 760]. Estimates [1021, 251, 1362, 496, 1036, 1193, 828, 749, 2032, 682]. estimation [1962, 1103, 658, 804, 2085, 1827]. estimations [872]. Euclidean [1358, 4, 986, 1073, 3, 1518, 1114]. Euler [2321, 1555, 1530, 227, 502, 1210, 1707, 2069, 1865, 1875, 1891, 829, 752, 1693, 1893, 2264, 1869, 1821, 1834, 467, 797, 2240, 104, 1940, 1487, 2115, 1907, 2315]. Eulerian [1557]. Evaluation [1702, 1100, 544, 201, 2215]. even [543]. Eventually [1631]. Every [2137, 20, 145, 1480]. evidence [1812]. Evolution [1251, 1549, 1483, 1207, 1599, 1607, 777, 2101, 2100, 1698, 227, 1726, 2136, 464, 1359, 825, 620, 1838, 656, 1130, 1138, 1973, 124]. Evolutionary [2280]. evolutions [375, 1255, 1801]. Exact [1980, 245, 302, 1726, 316, 1584, 1884, 1982, 429, 466, 1617, 1981, 1249, 968, 1072, 1501, 1883, 1954, 199, 1994, 2004, 891, 1413, 1373, 1556, 892, 949, 1927, 2297, 2105, 1182, 1010, 386, 1053, 1841, 2097, 217, 85, 1438, 108, 441, 1675, 1734, 1700, 1460, 2247, 1286]. Exactly [954, 654, 1513, 1625, 166, 1099, 721]. examination [2244]. example [580, 1117, 613]. Examples [1642, 609, 966]. Exceptional [1006, 1453, 1738, 691, 138]. excess [989]. exchange [441]. Exchangeable [400]. Excitation [2212]. excitations [2197, 141]. excited [493, 154]. excitons [1101]. exclusion [326, 1382, 1769, 2283]. excursion [1974]. exhibiting [1682]. Existence [170, 2239, 261, 1327, 1203, 1902, 568, 1953, 2172, 1444, 1797, 791, 645, 1434, 2276, 224, 612, 611, 2055, 191, 1054, 1669, 422, 390, 1331, 782, 1064, 921, 844, 1228, 493, 467, 946, 1898, 1758, 2304, 1202, 2074, 135, 1322, 1056, 1135, 491, 1201, 1725, 341, 1609, 2044, 2011, 2027, 881, 1592, 200, 1065, 144, 552, 980, 189, 1101, 449, 1287, 794, 2285, 982, 1057, 1391, 1594, 900, 1394, 2046, 1167, 1324, 2176]. exotic [1456, 650]. expanded [687]. expanding [2093, 701]. expansion [600, 2241, 1938, 1775, 619, 484, 927, 697, 404, 379, 490, 1520, 1011, 717, 1212, 436, 595, 247, 2301]. expansions [1315, 2081, 1935, 1350, 1096, 1861, 1776, 250, 163, 1826]. expectations [1227]. expected [495]. experiments [1632, 1147]. Explaining [957]. Explicit [588, 337, 1363, 595, 658, 1134, 1352, 2030, 1764, 966]. explicitly [632]. explosion [794]. exponent [1444, 2054, 2238, 1558, 97, 648, 2304, 2240, 2044]. Exponential [453, 2013, 1255, 262, 1955, 39, 1902, 2164, 2067, 95, 660, 674, 1281, 217, 44, 150, 1428, 917, 1084, 1056]. exponential-type [1084]. exponentially [232]. exponentials [537]. exponents [17, 945, 1018, 1354]. expression [1352]. Extended [1273, 957, 122, 650, 1690, 2327, 641, 1865, 1600, 1662, 1625, 93, 15, 642]. Extending [936, 911]. extensibility [447]. Extension [256, 268, 1945, 1015, 1363, 1007, 2218, 433, 1965, 1080]. extensions [11, 1720, 237, 1917, 2063, 236, 691, 890, 1099, 892]. extensive [580, 2208]. exterior [1529, 883]. external [1920, 791, 766, 1400, 1733, 710, 2246, 1704]. extraction [1824]. Extremal [1649, 1132]. extreme [1408].

f [483]. Faber [797]. factor [1776, 758]. Factorial [700]. Factorial-type [700]. factorisation [1178]. factors [1590, 201]. factual [2018]. Faddeev [1768, 104]. faithful [1087]. false [987, 661]. Families [1099, 1654, 45, 182, 1192, 1678]. family [345, 2282, 989, 1052, 1079, 270, 2085, 1473]. famous [2329]. fan [225]. Fast [454, 1051, 22, 493, 2226]. fast-slow [2226]. Faster [1962]. fastly [200]. Fault [722]. featuring [768]. feedback [948, 973]. Felderhof [1604]. Feller [974]. Fermi [1160, 1580, 1136, 1352, 1071, 2020, 850, 904]. Fermion [2295, 359, 998, 1465, 1175]. fermion-phonon [359]. Fermionic [1115, 1070, 2130, 1480, 766, 1377, 2138]. fermions [196, 1677, 1783, 651, 1025, 274]. Ferroelectric [106]. ferromagnet [2132]. ferromagnetic [1558]. ferromagnets [456]. few [2208]. few-body [2208]. fewer [1962]. Feynman [1638, 592, 1480, 1407, 49, 1388, 2180, 361, 2245, 816, 1196]. fiber [821, 909, 114]. Fibonacci [1374, 1914]. fibration [1995]. Fick [1607]. fidelity [894, 1744]. field [1819, 1168, 1549, 714, 1639, 2123, 554, 1485, 193, 638, 826, 1251, 1144, 188, 1595, 2206, 2146, 1823, 805, 1939, 2324, 2325, 1677, 241, 1292, 409, 609, 382, 1854, 1331, 1225, 2210, 2211, 1520, 1072, 1532, 2255, 713, 946, 154, 2243, 964, 2182, 1264, 2042, 1263, 1042, 2152, 1329, 2300, 1560, 1824, 290, 1020, 2046, 343, 1262, 1478]. fields [852, 200, 1920, 731, 2107, 772, 733, 722, 2185, 901, 119, 1450, 564, 830, 919, 1333, 1356, 681, 289, 1541, 149, 1465, 1076, 169, 2287, 749, 2246, 6, 8, 2306]. Fifth [1456, 501]. Fifth-order [1456]. fight [1358]. filament [2221]. film [666]. filtering [722, 1567]. filtration [1272]. filtrations [620]. finding [366, 1616]. Fine [1591, 581]. Finetti [205, 1377, 1256, 1546]. Finite [142, 803, 1693, 1486, 1936, 170, 2175, 318, 889, 637, 1017, 1803, 1643, 112, 269, 2308, 1246, 1106, 684, 1334, 1919, 1527, 972, 1067, 1394, 1611, 1715, 2222, 1175, 230]. Finite-dimensional [142, 1643, 112, 972, 230]. Finite-part [1486, 1936]. Finsler [2302]. First [769, 735, 1082, 1088, 1349]. Fischer [437]. Fisher [730, 1111]. fit [40]. FitzHugh [866]. Five [160, 206]. Five-body [160]. fixed [399, 926, 46]. flag [357]. flames [1282]. flat [1593, 268, 1945, 392, 1553, 901, 2076, 1662, 1187, 785]. flatland [771]. flatness [274]. flavors [627]. Flexible [1256]. flights [62]. flocking [1757, 1169]. Floquet [1442]. flow [1010, 2067, 392, 783, 1269, 912, 1747, 528, 1892, 2308, 1796, 1082, 2088, 51, 1274, 1243, 410, 795]. flow/Min [727, 1816]. flows [1703, 1037, 1649, 342, 788, 1081, 1875, 1706, 922, 1925, 971, 977, 452, 2047, 104, 1987]. Fluctuation [1039, 772, 1627]. Fluctuation-dissipation [1039]. fluctuations [729, 93]. Fluid [1037, 1010, 1904, 645, 1707, 912, 262, 1955, 1400, 535, 571, 678, 2077, 1232, 208, 1490, 354]. fluid-particles [1400]. fluid-rigid [1904]. fluids [227, 662, 2276, 2090, 1645, 289, 311, 1733, 1476, 1907]. Flux [1796, 1340, 2117, 2115, 1907]. flux-approximation [2115]. Fluxons [348]. Fock [2139, 2217, 1656, 538, 1913, 1194, 484, 488, 1144, 614, 281, 2185, 334, 989, 1317, 870, 560]. Focusing [1842, 228, 1667, 2311, 838]. Fokas [1712, 1973, 1715]. Fokker [640, 1599, 2082, 1265, 58, 176, 10, 121]. fold [1070, 2204]. folding [198]. foliated [1755]. foliations [1654]. force [735, 1153, 791, 2147, 1208, 1704]. forced [1033]. forces [1400, 752, 868, 1733]. Forchheimer [2047]. forcing [2080, 134, 864, 1858]. forests [884]. form [109, 204, 887, 1800, 1598, 1776, 688, 632, 1326, 201, 263, 41]. formal [25]. Formalism [1863, 1918, 1415, 565, 1073, 515, 1230, 1659, 1031, 1015, 1512, 810, 203, 1427, 1510, 1295, 1971]. formation [597, 784, 1987]. forms [109, 1525, 705, 987, 1365, 1035, 1852, 865, 1225, 704, 1766, 64]. formula [27, 1926, 1480, 442, 660, 1741, 1363, 2129, 498, 1439, 2329, 138, 1366, 632, 546, 2118, 940, 658, 471, 527]. formulae [592, 1777]. formulas [1150, 1849, 287, 1336, 1739, 1303, 355, 605, 1566, 928]. Formulation [163, 2233, 1227, 641, 217, 973, 1882, 1344, 433, 15, 642, 1148, 1927]. formulations [783, 559, 354]. foundations [862]. Four [2097, 286, 664, 249, 1026, 1330, 2036]. Four-body [2097]. four-component [2036]. four-dimensional [1330]. four-qubit [1026]. fourfold [411]. Fourier [340, 1792, 828, 1487]. fourth [381, 921, 2056, 691, 501, 1672, 1611]. fourth-order [381, 921, 1672]. FQXL [778]. fractal [2166, 1196]. Fractional [1668, 1528, 378, 974, 1685, 1716, 418, 1966, 779, 1949, 254, 990, 969, 1462, 1742, 2172, 1870, 1665, 1494, 878, 1463, 1667, 1726, 2055, 2054, 1433, 2328, 492, 1514, 2258, 464, 1359, 350, 869, 1307, 1448, 1523, 2008, 1905, 2227, 2145, 458, 844, 1572, 2084, 2142, 1493, 1011, 353, 1186, 2289, 551, 1397, 234, 781, 1731, 1202, 1723, 408, 143, 880, 1130, 1131, 1090, 1839, 1673, 30, 2259, 121, 298, 499, 341, 804, 2027, 682]. Fractional-order [418]. Fracture [2047]. fragmentation [946]. frame [1292, 1629, 2001]. Framed [1637, 1463]. frames [1338, 1443, 2151, 2231, 2039]. framework [634, 461, 2316, 894, 385, 609, 1250]. Fredholm [742, 1832]. Free [1819, 2195, 714, 1590, 116, 773, 487, 1362, 233, 2016, 2144, 636, 2134, 849, 2020, 2208, 2248, 2058, 651, 979, 1732, 2122, 884]. freedom [1595]. freely [1705]. Frenkel [141]. Frenkel-like [141]. frequencies [209]. frequency [391, 1033, 455, 2113]. friction [636, 2030, 2240]. frictional [2013]. Friedlander [498, 1439]. Friedman [2076]. Friedmann [968, 1551, 1353]. Friedrichs [1280, 1181]. Frobenius [1513, 463]. frog [2310]. Fröhlich [547]. frustration [233, 849, 2058]. frustration-free [233, 849, 2058]. Fueter [2217]. fugacity [1352]. Full [1775, 938, 116, 2013, 979, 526]. fully [227]. function [248, 279, 1429, 606, 1710, 2213, 1775, 198, 1737, 740, 314, 2282, 157, 1278, 1935, 152, 379, 1077, 1281, 1161, 912, 1679, 55, 313, 544, 1063, 1295, 441, 613, 451, 2117, 557, 2037, 718, 904, 2215]. Functional [1598, 1972, 1435, 745, 1214, 2134, 641, 1118, 1836, 523, 989, 1206, 1922, 41, 15, 642, 1056, 491]. functionals [1485, 1896, 224, 2283]. functions [286, 1403, 1059, 1197, 2257, 33, 1179, 874, 861, 1406, 987, 592, 1364, 1698, 1164, 2197, 2219, 585, 1151, 477, 1154, 707, 259, 438, 1418, 78, 1935, 1482, 280, 54, 2203, 962, 180, 181, 1849, 1991, 56, 1235, 76, 1779, 896, 840, 183, 2126, 1934, 18, 681, 1780, 1776, 2149, 1372, 1957, 1629, 1777, 1027, 700, 1604, 1740, 551, 244, 243, 355, 1317, 213, 2001, 370, 1487, 1889, 339]. functors [1069, 950]. Fundamental [1421, 757, 458, 1164, 1123, 1336, 1025, 1615]. Fusing [2294]. fusion [2177, 27, 271, 715]. Future [1899]. fuzzy [318, 157].

g [2093]. g-Navier [2093]. Gabor [2151, 1213]. gain [2016]. Galilean [970, 1473]. Galilei [122, 650, 1976, 1943, 889, 102, 867]. game [1504]. games [997, 1376, 895, 522, 2022]. gap [453, 98, 849, 235]. gapless [233]. Gapped [233, 1783]. gaps [2058, 1769]. gas [1038, 2261, 2288, 2168, 636, 2134, 2286, 1557, 50, 1326, 162, 1734, 1349, 410, 904, 212, 920, 2115, 795]. gases [2184, 1352, 2284, 225, 2096]. gate [316]. gates [554, 39]. Gaudin [886, 195]. Gauge [699, 460, 1592, 1920, 962, 1765, 803, 1224, 1331, 1509, 489, 932, 862, 1507, 835, 1311, 1027, 2141, 1042, 2038, 1791, 1556, 1928, 1165, 1148]. gauge-invariant [962]. gauge/gravity [1507]. gauged [789, 1956, 1500]. Gauss [1652, 2141, 1751, 553]. Gaussian [640, 1499, 484, 2016, 204, 1721, 2270, 996, 2272, 317, 466, 86, 1546, 1253, 400, 1692, 10, 2138, 749, 1375, 240, 1586]. Gaussian-to-Gaussian [204]. Gaussians [107]. Gel’fand [1954, 1895, 1088]. Gell [660]. Gell-Mann [660]. General [590, 119, 839, 118, 476, 6, 8, 1690, 2316, 371, 388, 599, 1285, 764, 1968, 1071, 1579, 936, 1857, 1884, 1982, 1109, 1344, 1299, 1879, 1817, 1547, 1260, 1104, 2275, 1211, 546, 58, 808, 1908, 362, 928, 1296, 1947, 1946, 1100, 751, 447, 330, 1807, 746]. generalisation [590, 872]. generalisations [286]. Generalised [1561, 1029, 570, 1697, 1077, 135]. Generalization [1449, 483, 1231, 1224, 923, 1452, 930]. generalizations [1351, 1976, 90, 564, 472, 644, 1937, 155]. generalize [181]. Generalized [1656, 420, 1406, 1150, 1754, 1151, 1318, 283, 1126, 78, 2148, 649, 869, 899, 47, 613, 1414, 892, 1863, 1131, 949, 1233, 769, 1918, 1203, 572, 820, 450, 1941, 386, 643, 1745, 2100, 329, 1036, 193, 1870, 1444, 1730, 981, 761, 1463, 1125, 2030, 1054, 299, 1045, 1124, 2012, 1808, 1779, 1388, 1650, 833, 941, 158, 1220, 2311, 1206, 1064, 2028, 617, 571, 678, 2077, 357, 1616, 2066, 719, 1273, 1675, 903, 1489, 213, 540, 1566, 96, 1858, 2271, 75, 1486, 1936, 778, 834, 1159, 222, 1574, 1609, 2176, 1335, 1709, 881, 1473]. Generalizing [167, 345]. generated [1339, 1144, 1024, 2185, 1263, 1527]. Generating [1235, 355, 208, 2017, 259, 152, 1409, 1889]. Generation [154, 2188, 1364]. Generations [1110]. generator [1348]. Generators [338, 2143, 1367, 1918, 2060, 2095, 817]. Generic [1761, 915, 836, 138, 41].

Genuine [1218]. Genuine-multipartite [1218]. genus [1488, 687]. genus-two [1488]. geodesic [902]. Geodesics [1342, 1030]. Geometric [1595, 379, 975, 1199, 1001, 248, 1640, 709, 741, 2206, 1537, 292, 901, 1627, 1026, 2132, 1993, 1554, 178]. Geometrical [600, 902, 1473]. geometries [318, 1761, 1524]. Geometrization [289]. Geometry [238, 1771, 1833, 2136, 315, 1253, 1339, 1239, 2241, 796, 1636, 1364, 1341, 1154, 730, 157, 996, 445, 1932, 1103, 1741, 818, 489, 1244, 1807, 1693, 1608, 110, 2274, 506, 423, 1409]. Gerstenhaber [416]. ghost [2248]. ghost-free [2248]. GHZ [89]. Gibbs [580, 1434, 2207, 1532, 2096]. Ginibre [48]. Ginzburg [261, 2210, 1856, 408, 536, 2027, 1469]. Girardello [1912, 1126]. Girsanov [830]. Gisin [1567, 1221]. given [365]. giving [2118]. Glaisher [286]. glasses [497, 2211]. GLM [1049]. Global [43, 116, 1664, 2091, 1729, 103, 980, 2090, 1855, 1984, 2049, 347, 1287, 2124, 794, 2171, 2285, 979, 712, 530, 1208, 1394, 1611, 2046, 1167, 823, 1704, 432, 351, 1324, 1089, 2176, 1727, 2175, 2164, 645, 189, 492, 307, 1643, 789, 2120, 786, 1693, 1057, 361, 1391, 1049, 1989]. Gluing [1333, 498, 1439]. GNS [2067]. GNS-detailed [2067]. go [1055]. golden [850]. Goodbye [1868]. goodness [40]. Gordan [1289]. Gordon [1329, 1066, 187, 1442, 2253, 1648, 1578, 1391, 1791, 1937, 526, 527, 1353]. Gottwald [2224, 372]. governed [1282]. governing [353]. Gowdy [817]. Gowdy-symmetric [817]. GPU [1788]. Grad [1361]. graded [368, 1047]. Gradient [2067, 783, 2286, 1901, 2150]. grading [1564]. Gradings [887, 1800]. grained [1691]. Graph [1938, 699, 283, 1040, 2310, 2197]. graphene [1024, 805, 2042, 657, 2065]. graphs [1248, 486, 669, 2035, 366, 1407, 1071, 2296, 1579, 54, 76, 896, 2235, 1046, 1807, 850, 1139, 1428, 2180, 816]. Grassmann [1942, 1126, 1834]. Grassmannian [1409]. gravitating [227]. gravitation [417]. gravitational [3, 2302, 193, 1929, 1358, 4, 1342, 1028, 1378]. gravity [2199, 1640, 2075, 268, 570, 1697, 2073, 558, 1512, 1177, 1508, 1552, 168, 1004, 1695, 965, 1108, 862, 1507, 1304, 1076, 1764, 561, 93, 1753, 629, 746, 774, 837, 169, 1437, 1511]. Gray [2077]. Green [176, 2257, 740, 1278, 1161, 1662, 55, 1063, 551]. Gromov [1608]. Gross [2184, 2222, 1016]. Grothendieck [271]. Ground [1313, 680, 74, 1730, 2097, 2054, 2134, 1368, 793, 1523, 2008, 1493, 1428, 1877, 1694, 647, 1460, 453, 669, 2292, 1392, 1962, 2238, 1448, 1871, 1728, 1845, 1783, 1840, 904, 881, 1066]. ground-state [453]. Ground-states [1313]. Grounded [2003]. Group [832, 1356, 1211, 1675, 1862, 1339, 1012, 1459, 1703, 268, 863, 1941, 440, 1516, 413, 477, 1841, 641, 382, 1234, 1081, 1450, 2062, 1689, 528, 218, 602, 464, 1359, 1658, 1879, 1206, 1174, 1068, 436, 816, 1085, 701, 1130, 1527, 924, 537, 15, 642, 1387, 446, 1927]. group-invariant [1081]. groupoid [1098, 1276]. groupoids [1701]. groups [2127, 309, 1592, 340, 82, 1803, 1975, 1738, 1314, 1289, 689, 321, 1139, 1106, 1240, 2318, 613, 723, 774, 1247, 2191, 1708, 156]. growing [22]. growth [1902, 495, 2238, 1205, 919, 1064, 1523, 921, 150, 1872, 1877, 1202, 148, 1906]. GUE [1598]. Guldberg [745]. Gurevich [2052]. Gurevich-Zybin [2052]. gyrogroup [625, 319]. gyrokinetic [796]. gyrokinetic-waterbag [796].

Gyrokinetics [1758]. gyroscopic [665].

H [1844]. haarsch [1436]. Hadamard [1017, 283, 766, 637]. Hagedorn [1743]. hair [2249, 2326]. hairy [568, 569]. Half [1462, 1742, 770, 80, 694, 1083, 1613, 2312, 1096, 350, 534, 593, 2135, 1876, 362, 928, 1296, 1947, 1946]. half-line [694, 1083, 1613, 1096, 350, 534, 593, 928]. half-signature [2135]. half-space [2312]. Hall [1980, 1981, 990, 189, 1617, 353, 1709]. Hall-magnetohydrodynamics [189]. Hamilton [1230, 592, 1154, 1472, 1624, 2006, 1262]. Hamiltonian [841, 2031, 769, 255, 1599, 731, 1, 633, 737, 188, 733, 1237, 1035, 2136, 1442, 2108, 547, 1543, 1129, 1249, 1682, 140, 1193, 1925, 117, 433, 65, 2079, 221, 479, 1967, 1113, 702, 723, 290, 710, 1437, 668, 1155, 785, 966]. Hamiltonians [2319, 1506, 907, 1642, 1219, 873, 1360, 53, 1630, 2020, 833, 1589, 1192, 1294, 1685, 651, 656, 605, 760, 1195, 1960]. Hankel [585]. hard [1361, 2049]. Hardy [1623, 612, 415, 2238]. Harmonic [1852, 54, 1499, 242, 384, 841, 1516, 1, 1995, 2212, 1083, 1761, 1252, 2108, 1281, 1759, 1541, 1846, 1532, 2195, 691, 273, 47, 1746, 647, 924, 2188, 124, 1533]. harmonically [1135]. harmonics [588, 1742, 1601, 1122, 1974]. Harnack [1052]. Hartree [955, 1953, 488, 1497, 1144, 224, 2185, 1612, 989, 1728, 560, 1201]. Hasegawa [1664, 1447]. Hausdorff [566, 940, 591]. having [384, 107, 1979]. Hawking [2071, 1696]. heat [1401, 1941, 297, 1158, 342, 822, 1433, 2328, 1576, 1615, 1149, 1987]. heat-conducting [822]. heat-conductive [342]. heated [24]. heavy [875]. heavy-tailed [875]. Heckman [1926]. Heegaard [2037]. Hehl [216]. Hehl-Obukhov [216]. Heisenberg [1022, 177, 888, 379, 2234, 1959, 2132, 2039, 809]. Heisenberg-like [809]. helicity [1833, 1951]. Helmholtz [775, 2040, 2308, 55, 2162]. Henstock [361]. Hepp [17, 945]. Heralded [1633]. Herglotz [1164]. Hermite [2087, 954, 1914, 691, 1626]. Hermitian [1094, 2232, 1682, 425, 1743, 1431, 1919, 1888, 275, 949]. Hertz [1242]. heterostructures [1150]. Heun [2043, 2327, 1867, 1710, 202, 2278, 1808, 256, 1600, 2252, 370]. Heun-type [2043, 1867]. hidden [127]. hierarchies [1271, 1661, 1129, 1386, 2126, 1895, 1355, 1787]. hierarchy [1795, 876, 1176, 255, 1171, 52, 459, 433, 65, 797, 176, 463, 471, 2007, 178, 115]. Higgs [1887, 731, 1835, 2195, 2287, 1165, 1609]. Higgs-like [2195]. High [1425, 405, 677, 1071, 665, 1962, 1762, 163, 377, 2165, 527]. high-dimensional [527]. high-energy [1071, 163]. high-loss [665]. High-order [677]. high-precision [1962]. Higher [1231, 2179, 69, 488, 323, 1619, 492, 2010, 102, 1445, 2155, 1323, 1752, 2281, 1270, 1477, 288, 821, 1463, 393, 980, 1581, 1333, 1873, 405, 1003, 867, 872, 1340, 1764, 1791, 173, 389, 1906, 1511, 1021, 698, 1148]. Higher-derivative [1231, 102, 867, 1511]. Higher-dimensional [1619, 1906]. Higher-order [488, 1445, 2155, 821, 1873, 173]. highest [2014]. Highly [1204, 1794, 792, 1242, 2273]. Hilbert [842, 2095, 228, 599, 763, 1601, 2219, 1765, 671, 1512, 402, 127, 41, 778]. Hill [1864]. Hirota [2169, 1015, 834, 1283]. Hitchin [2073, 1780, 1043, 1565]. HKT [655, 1761]. Hodge [1780, 1894]. Hofstadter [1621, 1711]. Hohenberg [1542]. Hölder [44]. hole [569, 3, 132, 1358, 1756, 4, 561, 629]. holes [170, 568, 1592, 1696, 2249, 2326, 1209, 1305, 449, 754, 2070]. Holevo [1633, 725, 1301, 658]. Hollow [324]. Holm [345, 914, 1873, 2224, 778, 1440, 908, 1574, 372]. Holographic [561]. holomorphic [1741]. holomorphy [219]. holonomic [1795]. holonomy [903, 1489]. Holstein [1370]. Hom [72, 25, 446, 230]. Hom-structures [230]. Homogeneous [269, 1833, 1602, 135]. Homogenization [71, 2173, 1988, 2263]. Homology [1139, 2208]. homothetic [1553]. homotopic [2023]. homotopy [1659, 1991, 5, 427, 1681]. honeycomb [1398, 486, 1650]. honor [1200]. Hooke [1516, 1584]. Hopf [641, 642, 167, 2189, 72, 15, 1474, 446, 884]. Hopfield [2005]. Horava [1552]. horizon [1209]. horizons [817, 2070]. horizontal [189]. Horn [1197]. Horndeski [1550]. horospheric [811]. HS [67]. hsp [2182]. Hubbard [1451]. Hückel [1063]. Hulthén [2213, 1884, 1982]. Hunter [1446]. Hurwitz [1722, 1235, 944, 687]. Husimi [1138]. Hybrid [1284]. hydraulic [798]. hydrodynamic [1081]. hydrodynamics [2148]. hydrogen [2107, 1809, 2244, 1072, 2329, 2243, 1349, 1678]. hydrogen-like [2243, 1349]. hydrogenic [1319, 12, 1320, 1309, 809]. hyper [1261]. hyper-spherical [1261]. hyperbolic [1755, 1364, 761, 126, 681, 1204, 1357, 2053, 2006, 1394]. hyperbolic-parabolic [1204]. hyperbolicity [1029]. hyperboloidal [1122]. Hypercontractivity [508, 112, 401]. hyperdisk [1426]. hyperelastic [265]. hyperelastic-rod [265]. hyperelasticity [470]. hypergeometric [1638, 1197, 1412, 450, 1896, 1482, 2129]. hyperkähler [898]. hyperoctahedral [1658]. hyperplane [1220]. hyperspheres [2003]. hypersurface [2069]. hypersurfaces [649, 448]. hypoelliptic [1052]. hypothesis [728, 1705].

i.i.d [728]. i.i.d.-quantum [728]. ICMP [2314]. ideal [912, 2010, 1232, 1212]. ideals [986]. identical [1915]. identification [142]. identities [502, 1779, 2228]. identity [1975, 140]. Ignorance [1505]. II [469, 2161, 1697, 2317, 1742, 1270, 2282, 331, 1615, 130, 1923, 1026, 1314, 2196, 1828, 1368, 1108, 243, 220, 1431, 1342, 1090, 774, 1565, 677, 1469]. II. [1079, 1936]. III [1890, 2197, 1807, 598, 1200, 1131, 837]. III. [258]. illusion [1686]. ILW [1488]. Image [2109]. images [1028]. imaginary [155]. IMCF [1755]. imitative [826]. impedance [358]. implementation [830, 1717]. implicit [1472]. Improved [752, 1540]. impulsive [1342, 1056]. impurities [1157]. impurity [1101]. incidence [1402]. including [2200, 608, 7, 989]. inclusion [1918]. incommensurate [1150]. incompatibility [197, 2299, 317, 1547]. Incomplete [2214]. incompressible [470, 1891, 1556, 263, 1279, 1876, 1704, 351]. incorporating [1137]. indefinite [2008, 1919]. indefiniteness [398]. independence [559]. independent [1463]. index [992, 548, 1498, 2084]. indexed [220, 598, 942, 1977]. indices [2034, 364]. individual [1205]. Induced [367, 100, 1080, 375, 1410, 584, 499]. induction [1929]. inductions [435]. Inductive [2301]. inelastic [2049, 1208]. Inequalities [1660, 449, 503, 2298, 696, 1996, 1721, 1588, 415, 1102, 1001, 519, 1052, 1420, 1104, 1479, 1922, 1399]. inequality [508, 724, 2035, 86, 1396, 1885, 1760, 1043, 1565]. infective [646]. inference [506]. Infinite [2116, 799, 2307, 183, 1212, 970, 541, 510, 1661, 1699, 76, 505, 2186, 303, 623, 485, 499]. infinite-area [2186]. infinite-dimensional [510, 505]. Infinitely [1905, 190, 910]. infinitesimal [1332]. infinity [1949, 1214, 191, 2265, 1820, 2141]. inflationary [43, 1864]. inflow [1475]. influence [1782]. informal [1988]. Information [1103, 852, 2161, 814, 90, 1964, 1154, 517, 730, 510, 1721, 996, 1111, 1747, 455, 893, 1464, 246, 754, 1301]. information-spectrum [1464]. information-theoretic [2161]. informational [2021]. informationally [1145]. informative [365]. Inhomogeneous [1614, 1674, 2175, 268, 304, 1039, 387, 2174, 1392, 1310, 381, 905, 681, 171, 781, 1020, 222, 351]. Initial [2170, 2222, 630, 914, 2288, 1530, 2303, 344, 750, 792, 1614, 2264, 559, 1019, 1715, 823, 1491]. Initial-boundary [2170, 2222, 1530, 1614, 1715]. injectivity [1452]. Inomata [2023]. input [56, 961]. input-plant-output [56]. inspired [1544]. instabilities [902]. Instability [669, 981, 1793, 919, 2308, 1731, 1718]. instantaneous [946, 124]. instantons [775, 85, 1438, 2232]. instrument [699, 365]. insulators [2167, 548, 1498]. integer [2252, 1936, 1486]. Integrability [1457, 865, 99, 1562, 463, 1784, 2107, 188, 323, 1781, 1013, 2142, 574, 273, 58]. Integrable [2169, 179, 210, 923, 1937, 1603, 1795, 1412, 255, 863, 1443, 1156, 111, 2251, 1129, 52, 1386, 1007, 2313, 321, 1192, 706, 1311, 433, 2079, 1246, 1986, 1413, 1829, 843, 118, 972, 1770, 644, 141, 943, 1973, 2036, 1574]. Integral [413, 286, 628, 963, 1073, 1941, 2034, 2217, 1792, 35, 1661, 637, 1017, 692, 1584, 85, 1438, 1589, 334, 1540, 480, 1082, 872, 605, 1566, 253, 527, 883, 1827]. integral-difference [872]. integrals [1638, 735, 1772, 592, 1848, 49, 992, 637, 1017, 72, 1388, 1975, 2129, 857, 2245, 615, 1894, 1100]. integration [1714, 2016, 1343, 1556, 1486, 1936]. integrator [2030]. integro [796, 95]. integro-differential [796, 95]. intensity [1551]. inter [943]. inter-site [943]. Interacting [2024, 1770, 196, 1352, 1086, 716, 2325, 241, 618, 334, 1141, 593, 1651, 166, 1994, 2165, 1181, 2046]. Interaction [225, 396, 1904, 1497, 764, 1410, 1292, 352, 1400, 2255, 1216, 1074, 38, 2246, 1159]. interactions [1038, 1676, 2281, 1683, 1825, 2200, 649, 1169, 1101, 549, 2113, 1348, 1027, 1760, 264, 1274, 77, 1805]. Interbasis [1315, 2241]. interchange [773]. interest [1599]. interface [2167, 2294]. interfacial [1577]. interference [957]. interferometer [240]. interior [1416]. internal [1401, 1892, 1801, 1557]. International [2314]. Interplay [1757]. interpolants [1429]. interpolating [557]. interpretation [893, 1951]. Intersection [2204, 2019]. intertwiners [1925]. Intertwining [907, 2153, 2177, 802]. interval [2050, 1074, 1973, 1049, 1715, 2222]. intrinsic [570]. Introduction [1768, 2314, 2166, 1988]. invariance [1053, 803, 2108, 1224, 1240, 890, 1536, 1575]. Invariant [650, 1639, 2207, 2080, 2205, 1189, 134, 2106, 122, 630, 720, 391, 469, 1337, 2140, 962, 129, 130, 1679, 1081, 218, 602, 1380, 1682, 862, 1767, 127, 562, 1578, 229, 93, 2156, 880, 1090]. invariants [1521, 1737, 709, 1008, 1244, 2181, 626, 1637]. Inverse [1390, 80, 312, 1367, 2160, 1596, 1416, 929, 2262, 75, 1168, 1918, 2060, 1793, 426, 344, 419, 2227, 2214, 14, 688, 838, 451, 1016, 1842]. inverse-square [1793]. inversion [471]. inverted [1252]. Investigating [1866]. Investigation [574, 1141]. involving [1325, 680, 1665, 1444, 387, 2054, 648, 1834, 691, 2117, 1839, 1673, 2259, 1100]. Iohara [1173]. ion [1580, 719]. ionization [891]. ionized [2261, 1349]. Irreducible [1365, 1737, 357, 1087, 800, 1067]. irreducibly [1106, 1527]. Irregular [1777, 501]. irreversible [635]. isentropic [2168, 1530, 1210, 1278, 1707, 147, 2240, 2115]. Ising [1115, 1558, 2208, 1193]. Isochronicity [2064]. isochronous [768, 1642]. isoenergetic [1375]. isolating [1275]. isometric [2231]. isomonodromy [1784]. isomorphisms [1897]. isosceles [1933]. Isospectral [1155]. isothermal [1492, 1893, 939, 1907]. isotonic [702]. isotopies [1586]. Isotropic [946, 1833, 1737, 1995, 1571, 2195, 1705, 1805]. isotropy [1862]. issue [1768]. Itô [2309, 2083]. Iterated [1638, 628]. iteration [1233]. Ito [255, 1880]. IV [2097, 942]. Iwatsuka [1537]. Izergin [1604].

J [2043, 145, 1720, 841, 426, 1980, 915, 945, 302, 2323, 916, 1719, 1160, 465, 2324, 2325, 2327, 428, 2326, 641, 1017, 63, 608, 2328, 1438, 1982, 119, 2320, 1199, 2322, 1981, 1319, 427, 1439, 842, 1569, 1320, 977, 640, 678, 4, 1840, 1898, 2042, 607, 1489, 301, 1130, 1839, 2321, 2319, 19, 642, 8, 1947, 1946, 2087, 339]. J. [1389, 1359, 1358]. Jacobi [1655, 237, 1230, 987, 954, 257, 2302, 1154, 1935, 1472, 1350, 1779, 2275, 2187, 2006, 1826, 1262, 2182]. Jacobi-type [1655]. jamming [1302, 1964]. Janes [118]. Jaynes [719, 1680]. jet [225]. Jiang [1606]. JMP [1899]. Joint [1851, 1538, 127]. Jordan [514, 605, 2103]. Josephson [1810, 348]. Jost [1429, 2213]. Jumarie [1463]. jump [1039, 225]. jump-fan [225]. jumps [1347, 798, 866]. junctions [1810, 348].

Kac [2125, 739, 1097, 192, 325]. Kadomtsev [1203, 419, 706, 1895, 459, 463]. Kähler [1561, 655, 1744, 870, 966]. Kalau [207]. Kaluza [2316, 2317, 2025]. KAM [1436, 1699, 785]. Kanai [1745]. Kapovich [1925]. Kappeler [1439, 498]. Kappeler-gluing [1439, 498]. Kapustin [1818]. Kármán [2013]. Kastler [207]. Kato [717]. Katugampola [254]. Kaup [52, 1829, 1858]. Kawahara [1091]. Kazhikov [226]. KdV [1559, 2236, 1015, 344, 913, 1653, 1563, 834]. KdV-type [1559]. Keldysh [383]. Keller [843, 1906, 1724]. Kelvin [2308]. Kemmer [1041, 1384, 2198]. Kepler [1153, 1245, 934, 2098, 2028, 211, 199, 1502, 890, 540]. kernel [1941, 393, 1572]. Kernels [874, 946, 410]. Kerr [2040, 168, 1078, 1079, 169]. key [812]. Keyring [1455]. KGS [389]. Khintchine [592]. KIDs [1079]. Killing [320, 1478, 663, 2303, 60]. kind [769]. Kinematic [1076]. Kinematical [1620, 1618, 1240, 1619]. Kinetic [1349, 382]. Kirchhoff [1839, 1325, 2239, 186, 1065, 1870, 1444, 2292, 191, 2094, 1395, 1905, 921, 1493, 1673, 2011]. Kirchhoff-type [1444, 2292]. Kirkpatrick [2031]. Kirkwood [176]. kit [1126]. Kitaev [2140]. Kitaev-like [2140]. kite [963]. Klein [1791, 2316, 2317, 1442, 1578, 1391, 1329, 2025, 527, 1353, 1066]. KMS [2183, 993]. knots [1010]. knowledge [314]. known [1416, 1536]. Kobayashi [626]. Kochen [398]. Kohn [1542]. Kolmogorov [1705, 1675]. Kompaneets [1206]. Kontsevich [1388, 1894]. Korepin [1604]. Korteweg [1211, 1774, 1323, 469, 492, 2050, 1311, 1088, 1062, 1276]. Korteweg-de [1774, 2050]. KP [977, 2161, 1171, 1841, 971, 471, 2007]. KPI [742]. Kramers [85, 1438, 718]. Kratzer [1657]. Kratzer-type [1657]. Kraus [1459]. Krawtchouk [2057]. Krein [2112, 731, 1896]. Krein-like [1896]. Kubo [1150]. Kundu [1613]. Kunike [953]. Kupershmidt [1736]. Kuramoto [631, 1514, 2051, 97]. Kuznetsov [347]. KZ [272].

l [122]. l-oscillators [122]. Labyrinth [743]. Ladder [2063, 483]. Ladik [838, 1718]. Ladyzhenskaya [978]. Lagrange [2206]. Lagrangian [841, 1, 1995, 937, 1595, 385, 1744, 2275, 1519, 2029, 1042, 212]. Lagrangians [737, 1517, 1437]. Laguerre [1655, 954, 258, 1151, 281, 1350, 2311, 544, 110]. Laguerre- [1655]. Lakshmanan [1299, 278]. lambs [661]. Lamé [285]. laminated [1577, 1284]. Landau [261, 761, 2223, 2021, 330, 331, 307, 1679, 1741, 947, 2210, 1541, 1856, 408, 760, 31, 525, 536, 2027, 1469]. Landau-breather [760]. Lanford [1200]. Langlands [1621]. language [818]. Lannes [1491]. Laplace [1670, 2006]. Laplacian [2239, 2172, 694, 1665, 1494, 2296, 2054, 747, 2290, 1874, 149, 1228, 430, 305, 2259, 298, 751]. Laplacian-Spectral [2290]. Laplacians [1435, 183, 1186, 2156]. lapses [1859]. Large [748, 927, 1350, 1732, 2287, 74, 772, 791, 875, 105, 1407, 2284, 2080, 1033, 1984, 107, 18, 932, 2149, 2310, 1985, 979, 2152, 1309, 809, 2122, 823, 1704, 351, 339]. Large- [927, 932]. Large-time [748, 1984, 2149]. largest [1598, 919]. laser [24, 2016]. laser-heated [24]. Lattice [1038, 376, 1461, 2250, 496, 196, 2236, 731, 728, 1051, 214, 1765, 1558, 840, 1650, 1007, 55, 120, 1789, 1246, 221, 2088, 1468, 1465, 2038, 50, 141, 1603, 373, 115, 627]. lattices [1605, 1337, 1179, 418]. Laurent [574]. law [475, 411, 1607, 399, 470, 2081, 1968, 357, 1432, 2117, 2240]. Laws [2310, 1483, 386, 337, 832, 117, 125, 712, 1242, 1050, 2053, 1349]. Lax [63, 1318, 462, 1773, 9, 1316, 293, 2222]. layered [577]. layers [1537]. Leading [1899, 1107, 1598]. leads [1882]. Leaky [1520, 552, 2216]. Leaky-mode [1520]. learning [521]. Least [1723, 191, 1092, 365, 1717].

Lefever [1790]. Leffler [1128]. left [686]. left-right [686]. Legendre [1938, 1935, 292, 1011, 1100]. Legendre-polynomial [1011]. Legendrian [1763]. Leibniz [500, 153]. Lemaître [1353]. lemma [323]. length [1073, 1128, 1354, 1117, 808, 1373, 482, 1152, 1279]. lens [2162]. lensing [1028]. Leon [63, 9, 1286]. Leray [263]. LeRoux [2168]. Level [162, 37, 2281, 2160, 1126, 1679, 1294, 964, 118]. levels [761, 1136, 1160, 1741, 2329]. Leverett [643]. Levi [194]. Levin [95]. Levinson [1290]. Levitan [1088]. Lévy [640, 1966, 592, 1158, 1848, 337, 10, 2226]. Lewis [1846]. Lichnerowicz [2123]. Lie [2323, 641, 1389, 1359, 2322, 25, 607, 1130, 19, 642, 1620, 1862, 600, 1022, 1599, 1659, 1878, 745, 1521, 685, 1133, 340, 462, 439, 799, 152, 1619, 1618, 1958, 2203, 1120, 210, 1045, 1450, 1288, 1314, 1736, 464, 689, 1229, 2228, 269, 1206, 1336, 1739, 886, 686, 1170, 1449, 2180, 355, 2318, 684, 1334, 845, 656, 1131, 723, 774, 539, 13, 137, 1172, 1411, 15, 474, 1067, 155, 230, 711, 310, 275, 1335, 1286, 115, 698]. Lie-point [1389, 1120]. Lieb [328, 329, 1479, 2088, 651, 553]. Liénard [2323, 2322, 36, 607, 34, 13, 19]. Lifespan [1132]. Lifshitz [2223, 1552, 749]. lift [1653]. lifts [1237]. light [1030, 445, 2312, 577, 1511]. light-cone [445]. light-like [1030]. Lightlike [448]. like [779, 1030, 1466, 1107, 2140, 994, 1896, 330, 331, 2194, 344, 1892, 2195, 2243, 460, 808, 1900, 2240, 1349, 809, 958, 141, 1172, 1411, 1554]. Limit [826, 1612, 1308, 1425, 1146, 1370, 533, 1116, 1361, 1407, 1847, 948, 2089, 2130, 1776, 604, 2255, 946, 1578, 1366, 2152, 1391, 2133, 2053, 169, 1974, 1279, 1135, 920, 83]. Limitations [1264, 757]. Limiting [2201, 107]. limits [1164, 1421, 1210, 76, 1837, 1532, 2240]. Lindblad [1370, 2067]. line [80, 109, 461, 1778, 694, 1083, 85, 1438, 1613, 1096, 350, 534, 593, 1535, 362, 928, 1296, 1947, 1946]. Linear [1127, 1661, 1549, 1327, 680, 2260, 841, 705, 174, 1521, 1, 193, 258, 784, 997, 2190, 1773, 95, 671, 1576, 1615, 507, 429, 1209, 1288, 919, 2040, 1879, 1979, 782, 1448, 2142, 1762, 504, 1814, 1562, 1308, 50, 1967, 864, 1894, 1135, 275]. Linearisable [61]. linearity [644]. Linearizability [1207]. linearizable [1166, 1009]. Linearization [1896, 2143]. linearized [2199]. Linearly [1271, 683, 737, 346, 190]. lines [948]. linked [1938, 1058]. linking [1763, 2257]. Links [2329, 1443]. Lions [1730, 2292]. Liouville [558, 180, 674, 2094, 300, 1088, 1074, 1753, 2237, 2085]. Liouvillian [1756]. Lippmann [1141, 1883]. Lipschitz [1029]. liquid [1313, 1492, 2119, 331, 529, 441, 795]. liquid-gas [795]. liquids [822, 824]. Lissajous [236]. little [2062]. lived [2071]. lives [1358]. Lobachevsky [811]. Local [2031, 1178, 1968, 1492, 849, 2012, 789, 726, 1801, 1791, 232, 1165, 411, 1327, 1590, 1607, 1485, 257, 1035, 2099, 1103, 545, 1548, 1837, 131, 1287, 1705, 2088, 511, 223, 1016]. localisation [1585]. localised [1663, 1551]. localising [398]. Locality [105, 2300, 496, 2088]. Localization [1749, 1238, 1999, 1451, 875, 1436, 1086, 296, 2113, 749]. localizations [284, 1708]. localized [77, 1805]. localizer [2135]. locally [1591, 631, 232, 2074, 2191]. LOCC [1254, 1997]. loci [1780]. locking [631].

Loewner [1801]. Löf [2230]. log [379, 1760]. log-partition [379]. log-Sobolev [1760]. Logarithmic [1677, 714, 508, 1265, 2259]. logistic [1205, 1906, 1724]. Lohe [1822]. Lojasiewicz [97]. Long [1798, 914, 1613, 2149, 1903, 1986, 2065, 1313, 1038, 49, 2211, 2113, 2124, 1193, 348]. Long-distance [2149]. Long-range [2065, 1313, 49, 2113]. Long-time [1798, 914, 1613, 1903, 1986]. look [20, 145]. loop [286, 628, 1860, 440, 631, 49, 1350, 1522, 1177, 1508, 369, 528, 965, 1108, 1076, 1510, 247, 837, 2301]. Lorentz [404, 51, 436]. Lorentzian [427, 109, 1593, 1636, 1932, 528, 5, 1925, 2000]. Lorenz [1791, 1165]. loss [665]. lossless [665, 948]. Lotka [178]. Lounesto [2023]. Lovelock [1695]. Low [440, 1337, 468, 2016, 81, 163, 1791, 2070, 660, 2089, 2039]. low- [163]. Low-dimensional [440, 1337]. low-energy [81]. Lower [579, 1186, 1921, 1622, 1214, 135]. lowest [1679]. lowest-Landau-level [1679]. Löwner [876]. LSMR [1717]. Lucas [1374]. Ludwig [1768]. LUE [1598]. Lugiato [1790]. lump [1789]. lumps [322]. Lyapunov [2203, 1354].

M [320, 1637]. M-branes [1637]. M-theory- [320]. M5 [403, 2225, 1560]. M5-brane [403, 2225, 1560]. Macdonald [1477, 944, 1934]. Mach [2089]. macroscopic [1217]. Madelung [1882, 1575]. Magnetic [1573, 991, 1748, 2306, 852, 1168, 2116, 2107, 772, 1157, 1251, 1144, 1823, 805, 1939, 2324, 2325, 241, 1292, 747, 2185, 919, 260, 995, 1072, 1985, 713, 154, 2243, 2042, 2156, 349, 749, 2246, 1876, 1727, 343]. magnetized [42]. magneto [1476]. magneto-micropolar [1476]. magnetoelectric [2302]. magnetohydrodynamic [342, 103, 1396, 1732]. magnetohydrodynamics [1980, 189, 1617, 1981, 1475]. Mahony [1900]. Majid [435]. Mallows [1646]. managed [1285]. Manev [45]. manifold [59, 43, 1833, 2035, 1408, 498, 1439, 903, 1489, 2259]. manifolds [1591, 1755, 1005, 1593, 1983, 2123, 1258, 2072, 1945, 2045, 392, 368, 733, 1312, 292, 1513, 182, 528, 975, 1199, 1380, 659, 357, 2000, 870, 1243, 1399, 1888, 207, 1474, 2306, 2226]. Manin [1751]. Mann [660]. many [2269, 1992, 1998, 1480, 2096, 1905, 250, 1616, 358, 118, 176, 910, 1505, 190]. many-body [2269, 1992, 1998, 1480, 2096, 176]. many-box [1505]. many-mode [118]. map [814, 1719, 1644, 1060, 713, 2138, 1574]. Mapping [1184, 1804, 1415, 2297]. mappings [2236, 136]. Maps [1513, 1897, 1218, 507, 1850, 2313, 1422, 854, 504, 1814, 2088, 1631, 903, 1489, 800]. Marginally [109, 51]. Markov [338, 634, 1039, 1255, 2207, 2144, 1103, 1427, 586, 83]. Markovian [375, 716, 1859, 948]. Mars [1466]. Martin [2230]. Martin-Löf [2230]. martingales [1801]. Maskawa [626]. Mass [766, 630, 1401, 720, 2269, 226, 36, 1182, 2072, 392, 1151, 224, 1392, 1522, 1612, 1299, 1759, 1541, 159, 34, 38, 1804, 104, 485, 2291, 1155, 1201, 1826, 1533, 2297]. mass-resonance [2291]. masses [1240, 135]. Massive [17, 945, 638, 994, 417, 1363, 1055, 1687, 1437]. massless [1776, 2149, 101, 166, 2155, 1503]. master [906, 416, 1142]. Math [2043, 145, 1720, 841, 426, 1980, 915, 945, 302, 2323, 916, 1719, 1160, 465, 2324, 2325, 2327, 428, 2326, 641, 1389, 1017, 63, 608, 2328, 1438, 1982, 119, 2320, 1199, 1359, 2322, 1981, 1319, 427, 1439, 1358, 842, 1569, 1320, 977, 640, 678, 2329, 4, 1840, 1898, 2042, 607, 1489, 301, 1130, 1839, 2321, 2319, 19, 642, 8, 1947, 1946, 2087, 339]. Mathematical [2314, 1505, 1487, 660, 523, 857]. Mathews [1299, 278]. Mathieu [493, 285, 1711]. matrices [414, 411, 518, 2220, 444, 329, 142, 283, 1773, 1968, 465, 516, 53, 1661, 1454, 494, 581, 503, 1431, 2318, 2187, 595, 708, 252, 118, 972, 2026, 272]. Matrix [318, 1996, 1816, 139, 537, 474, 73, 309, 1851, 239, 542, 1255, 517, 927, 356, 1365, 462, 1482, 1545, 332, 335, 112, 989, 1652, 1684, 1693, 2142, 2178, 544, 688, 964, 244, 952, 1308, 1025, 626, 1316, 595, 39, 1635, 2188, 48, 362, 928, 1296, 1947, 1946, 2007]. matrix-valued [964]. matter [417, 693, 1379, 2025, 758]. Maupertuis [2302]. Max [727, 1816]. Max-flow [727, 1816]. Max-flow/Min-cut [727, 1816]. Maximal [1375, 1862, 1580, 989, 1594, 1923]. maximally [695, 1409]. Maximum [1548, 855, 1831, 506]. maximum-entropy [506]. Maxwell [2328, 432, 1066, 705, 396, 2280, 288, 1707, 1571, 1433, 666, 267, 1791, 1113, 2158, 910, 647, 354, 1937, 2246, 1163, 881]. Maxwell-dilaton [1113]. Maxwellian [2120]. Maxwellians [307]. McKinley [2023]. Mean [2182, 826, 2211, 1532, 2255, 946, 240]. Mean-field [2182, 826, 1532, 946]. meaning [1149]. measurability [1851]. measurable [1073]. measure [1632, 402, 842, 1646]. Measurement [197, 511, 1250, 1538, 856]. measurements [951, 1538, 1747, 2192, 1629, 127, 1298]. measures [582, 1477, 1145, 2207, 545, 369, 509, 1532, 2096, 44, 1309, 809, 421]. mechanical [1544, 821, 442, 605, 476, 819]. Mechanics [759, 957, 279, 254, 1605, 1223, 1231, 969, 1917, 1462, 363, 239, 284, 763, 477, 1134, 992, 653, 901, 2102, 335, 847, 1853, 1681, 991, 1748, 425, 102, 867, 693, 294, 1113, 293, 1298]. media [1037, 1436, 1837, 681, 982, 868, 1020]. medium [387, 2262, 577]. Meissner [985]. melonic [1854]. memory [1859, 1149, 1768, 787, 1827]. memoryless [812]. memristive [1198]. meromorphic [1363, 1990]. mesoscopic [2307]. message [1302, 2104]. metaplectic [128, 1586]. method [1233, 1325, 17, 945, 826, 685, 366, 2016, 1051, 1278, 2327, 674, 1281, 1643, 1040, 1395, 1831, 256, 1600, 829, 1458, 998, 1903, 2084, 163, 1264, 1670, 1343, 1326, 718, 531, 1387, 1973, 1715, 1725, 275, 804, 1283]. Methods [940, 227, 1886]. Metric [1751, 2302, 54, 500, 2275, 1226, 1919]. metrics [2232, 1379, 758]. Metrisability [1484]. Metzner [1174]. MHD [792, 1985, 2124, 1876, 823, 2176, 1709, 1727]. Mickelsson [206, 309]. Mickelsson-Rajeev [206]. microlocal [559]. micropolar [262, 1955, 535, 1733, 1476]. Microscopic [196, 1249, 2158]. MICZ [2098, 199, 1502, 540]. Mie [1463]. Mie-type [1463]. Miki [1173]. Mills [1887, 2287, 1609, 1590, 568, 569, 1592, 731, 297, 1269, 962, 1836, 26, 604, 1025, 613]. Millson [1925]. Mima [1664, 1447]. MIMO [1385]. Min-cut [727, 1816]. Minimal [1659, 402, 842, 1147, 745, 17, 945, 117, 808, 1373, 324, 482, 1152, 1724]. minimally [629]. Minimax [1325, 2298]. minimization [1179, 2159, 1612].

Minimizers [2191, 224]. Minimizing [2250]. Minimum [37, 1426, 1073, 1548, 654]. Minitwistors [1887]. Minkowski [428, 2, 563, 802, 868, 51, 2287]. Minkowskian [1591]. mirror [1119]. misleading [1117]. Mittag [1128]. Mittag-Leffler [1128]. Miura [344, 1085]. Mixed [2220, 1037, 730, 639, 1265, 1850, 1328, 1050, 252, 48, 2173, 1586]. Mixing [1346, 399]. mKdV [391, 797]. Mn [1298]. Mode [1188, 1520, 1562, 443, 909, 118, 949]. model [453, 1819, 238, 2116, 1248, 226, 990, 29, 2317, 106, 1775, 1622, 1451, 1107, 2140, 1477, 1370, 1664, 826, 2091, 2168, 470, 387, 1713, 2268, 676, 2139, 113, 590, 2146, 1492, 788, 2119, 1972, 379, 2185, 390, 1857, 1514, 1040, 1383, 1558, 1822, 2120, 1205, 2210, 107, 1400, 1956, 1810, 2205, 359, 2310, 827, 97, 646, 166, 1193, 120, 1796, 2286, 886, 770, 2178, 1597, 1063, 815, 719, 1025, 2047, 790, 891, 1560, 278, 843, 104, 476, 743, 325, 141, 1298, 530, 778, 1906, 2106, 1181, 1163, 2007, 795, 536, 1467, 1402]. modeling [1242]. modelling [127, 693]. Models [2150, 1313, 1038, 42, 1412, 461, 1912, 177, 2033, 1659, 2200, 1701, 873, 631, 1086, 655, 111, 382, 2251, 901, 56, 1198, 817, 2208, 905, 953, 489, 1328, 1455, 1246, 1604, 195, 1343, 967, 1647, 118, 2037, 1505, 1635, 1028, 949, 1473]. modes [2189, 200, 1713, 2268, 807, 2216, 452]. modification [1340]. modified [469, 518, 1945, 1679, 52, 1260, 1873, 2171, 1311, 923, 293, 1563, 1440, 908, 920, 2115]. modular [1070, 1679, 1539]. modular-invariant [1679]. Modulated [681, 1702, 55]. modulation [780]. Modulational [1718]. module [1660]. Modules [2014, 1802, 714, 986, 879, 888, 2131, 848, 72, 2265, 753, 1317, 845, 1990, 1067, 155, 885, 156]. Moduli [1010, 322, 2127, 445, 1780, 1186, 1751]. molecular [2030, 395]. molecules [1684, 1933]. moment [875, 2325, 241, 1292, 21, 301, 252, 2246]. momenta [630]. moments [2220, 633, 48]. momentum [36, 1462, 2072, 632, 654, 311]. Monge [1156]. monitored [1142]. monodromy [1995, 1035]. monogamy [1633]. monogenic [1317]. monoidal [1944]. monomer [826]. monomer-dimer [826]. monomial [1487]. monopole [1561, 2159, 578, 853, 1581, 1568, 1804, 2297]. Monopoles [775, 820, 991, 1748, 1751]. Monotone [859, 758, 1469]. Monotone-short [859]. Moody [739]. Mooney [470]. Mori [1227, 1827]. Morrey [2090]. Morse [608, 2250, 994, 7, 1141]. Morse-like [994]. Moscovici [167]. Moser [2153, 1771, 2007]. Moshinsky [160]. motility [1906]. Motion [132, 469, 820, 426, 822, 584, 2030, 1356, 1624, 1217, 14, 667, 1733, 447]. motions [54, 2108]. Motivations [2316]. moving [2325, 241, 868]. Moyal [1140, 1818, 931, 91, 1510, 2215]. MRW [1787]. Müller [2015, 989]. Multi [2139, 1086, 349, 1911, 66, 212, 791, 764, 179, 742, 913, 1103, 1331, 1689, 2063, 146, 2089, 793, 166, 844, 1355, 704, 923, 220, 598, 942, 1977, 595, 1718, 908, 1135]. Multi-bump [349, 844]. multi-component [179, 1355, 923, 1135]. multi-dimensional [791, 704]. multi-dimensions [2089]. multi-indexed [220, 598, 942, 1977]. multi-KdV [913]. multi-matrix [595]. multi-parameter [1103]. Multi-particle [1086]. Multi-peak [1911]. multi-qubit [1689]. multi-rogue [742, 1718]. multi-soliton [908]. Multi-speed [66]. multi-step [2063]. multi-string [1331]. Multi-symplectic [212]. Multi-time [2139, 764, 166]. multi-vortex [146]. multi-well [793]. multidimensional [1772, 828, 2118, 900]. multidimensions [61, 1781]. multifractal [337]. Multilocal [473]. multimodal [909]. multimode [2101, 484]. multiparameter [462]. multipartite [1218, 1407, 1588, 1257, 1104]. Multiple [1531, 266, 1494, 1874, 150, 84, 148, 343, 2254, 2200, 455, 1979, 622]. multiplets [659]. multiplication [1432]. Multiplicative [1234, 675, 794]. multiplicities [152]. Multiplicity [918, 1336, 2084, 1839, 1673, 2239, 1327, 2172, 1739, 1064, 921, 355, 1725, 341, 2044]. multipliers [257, 1850]. Multipole [2249, 2326, 216]. multiresonator [2101]. Multiscale [707, 1837, 777, 337]. Multistability [631]. multistate [1405]. multistep [1866]. Multisymplectic [1512, 1262]. multivariate [1420]. multivector [1635]. mutations [2143]. mutual [90, 455, 893, 1301]. Mutually [622, 20, 145].

N [170, 2147, 1865, 1873, 1841]. N=1 [1976]. Naghdi [1662]. Nagumo [866]. Nakajima [1227]. Nambu [683]. nanotubes [546]. nanowires [676]. Nash [519]. Navier [116, 579, 791, 748, 1729, 1278, 2090, 788, 1984, 2089, 147, 1891, 2093, 2171, 2285, 524, 1082, 979, 377, 824, 1556, 263, 978, 1279, 1167, 939, 2122, 1704, 351, 1089]. near [1924, 1137, 307, 1209, 2120, 2141, 1135]. near-linear [1135]. near-optimality [1924]. Necessary [1690, 1057, 1395, 1967]. negative [59, 569, 1580, 761, 512, 2109, 2293, 682]. negatively [2257]. Nehari [2084, 2259]. neighborhood [304]. neighborhoods [1195, 2319]. neighbourhood [865]. Nelson [1622, 1543]. nematic [1492, 331]. nets [1778]. network [2005, 973, 1857]. networks [1582, 631, 757, 948, 183]. Neumann [672, 515, 2034, 297, 1660, 1813, 1923, 2196, 1372, 2235, 1147, 127, 430, 713, 2193, 71]. neutral [1347, 2246, 491]. neutron [2121]. Neveu [1865, 2315]. Newell [52, 1829]. Newman [168, 169]. Newton [570, 1697, 1516, 2274, 1432]. Newtonian [767, 417, 1272, 2329, 370, 2247]. next [1107]. next-to-leading [1107]. Nichols [73]. Nijenhuis [1736]. Nikiforov [2327, 256, 1600]. nine [2098, 199, 1502]. nine-dimensional [2098, 199, 1502]. NLS [1795, 2174, 1787, 2291, 1842]. No [529, 2210, 123, 1055, 930]. no-go [1055]. no-signaling [123]. nodal [486, 191, 2008, 1877]. Noether [82, 1229, 407, 819]. Nohel [95]. noise [1966, 2104, 1058, 584, 2209, 1885, 408, 2106, 2226]. noiseless [2270]. noises [1127, 573, 794, 586]. noisy [513]. Non [1917, 174, 2131, 875, 1210, 2119, 2251, 1859, 948, 780, 1627, 833, 519, 1743, 2264, 2234, 1647, 1936, 1019, 2019, 1491, 1499, 1795, 1786, 59, 1549, 1327, 680, 172, 1605, 907, 384, 2112, 1599, 1607, 1623, 257, 245, 302, 914, 1530, 784, 662, 502, 1418, 1278, 1492, 1707, 716, 865, 2203, 417, 1803, 1124, 1741, 429, 512, 817, 1297, 549, 147, 1007, 1272, 1682, 425, 1625, 1684, 1517, 640, 2029, 1578, 762, 203, 2245, 1562, 1994, 1838, 1334, 1919, 701, 151, 629, 2201, 10, 190, 866, 1987, 1165, 2273, 275, 949, 2027, 373, 1469, 2002, 1805, 966, 196]. Non-Abelian [1627, 1165]. non-additive [2273]. Non-adiabatic [2019]. non-autonomous [172, 865, 2203, 1124, 1838, 2201, 866, 2027, 373, 2002]. non-barotropic [1987]. non-central [245, 302, 1625, 1994]. non-classical [1623]. non-closed [817, 549]. Non-commutative [519, 2234, 502, 1741]. non-commuting [203]. non-compact [151]. non-conservative [1517]. non-constant [1707]. non-convex [1786]. Non-criticality [1647]. non-crystallographic [1803]. non-cylindrical [1469]. non-dissipative [2029]. non-Gaussian [640, 1499, 10]. non-Hamiltonian [1599]. Non-Hermitian [1743, 1682, 425, 1919, 275, 949]. non-holonomic [1795]. Non-integer [1936]. Non-integrable [2251, 1007]. Non-interacting [196]. non-isentropic [1530, 1278, 1707, 147]. non-isothermal [1492]. non-isotropic [1805]. non-linear [1549, 1327, 680, 174, 784, 429, 1562]. non-local [1327, 1607, 257]. Non-Markovian [1859, 948, 716]. non-minimally [629]. non-negative [59, 512]. non-Newtonian [1272]. non-periodic [1605]. Non-perturbative [174, 875]. non-pure [762]. non-regular [1334]. Non-relativistic [1210, 417, 1684, 1578]. Non-self-adjoint [1917, 833, 907, 1297]. non-smooth [2245]. Non-spherical [2119]. non-standard [1418]. non-symmetric [190]. non-trivial [384]. Non-uniform [2264, 1019, 1491]. non-uniformly [701]. non-uniqueness [2112]. Non-viscous [780]. Non-weight [2131]. non-zero [914, 662]. nonassociative [1453]. nonautonomous [1229, 1016]. nonclassical [1918, 69, 374]. Noncommutative [1339, 1415, 2316, 2317, 1636, 1559, 477, 1134, 2183, 249, 1836, 2232, 2102, 847, 1498, 489, 1653, 991, 1748, 500, 1895, 1524, 1170, 460, 870, 253, 270]. noncommuting [1102]. nonconservative [819]. noncontextuality [1690]. Nondegeneracy [1789]. nondiagonal [644]. nonequilibrium [2288, 1424, 383, 292, 901, 31]. nonexistence [1444, 1305]. nonexpanding [1342]. nonholonomic [819]. nonhomogeneous [2171]. noninteracting [1770]. Nonlinear [915, 836, 2082, 1149, 350, 1190, 141, 306, 1390, 483, 1367, 1323, 779, 242, 1798, 2164, 1398, 1483, 1676, 1909, 1793, 228, 1953, 1992, 806, 1941, 1698, 745, 1521, 532, 257, 388, 209, 2172, 1470, 1797, 1668, 144, 387, 533, 645, 676, 227, 878, 2276, 61, 1667, 783, 871, 611, 1162, 803, 2080, 1013, 1669, 1570, 381, 2258, 1614, 534, 2040, 18, 126, 321, 2311, 1206, 782, 793, 1307, 2008, 2227, 134, 120, 877, 1872, 1903, 1728, 835, 640, 397, 467, 982, 1762, 291, 1898, 679, 1445, 2289, 815, 267, 1391, 70, 2114, 1397, 264, 1242, 47, 702]. nonlinear [1675, 1575, 34, 1986, 781, 1672, 21, 301, 2117, 892, 1202, 900, 349, 1490, 143, 880, 1090, 39, 10, 84, 1911, 457, 943, 1603, 66, 1712, 668, 527, 1049, 1611, 222, 339, 1467, 1402, 677, 2173]. nonlinearities [1730, 1285, 2055, 22, 1671, 787, 2304, 2092, 2165, 610]. nonlinearity [2086, 1870, 149, 1873, 2114, 234, 96, 910, 1839, 1673, 457, 2259, 778, 751]. Nonlocal [767, 1570, 835, 1734, 1390, 736, 2169, 1013, 1092, 2312, 895, 126, 1671, 1275, 1311, 2056, 522, 1986, 2032, 984, 1016, 1402, 2226]. nonlocality [105, 119, 123, 6, 8]. Nonlocally [499]. Nonperturbative [406, 2215, 910]. nonreflecting [1908]. Nonrelativistic [533, 1183, 1391, 32]. nonsmooth [1036]. nonstandard [1872]. nonstationary [21, 301]. nontrivial [1203, 264]. nonvanishing [383]. nonzero [1390, 2160, 264, 838]. Nordheim [251]. Nordström [2035]. Nordström-anti-deSitter [2035]. norm [508, 1065]. Normal [483, 204, 587, 865, 1118, 581, 514, 64]. normality [1103]. normalized [1884, 1982]. norms [955]. Note [302, 1719, 1160, 465, 1982, 1569, 977, 1840, 978, 1590, 1738, 22, 596, 1384, 223, 48]. Notes [1933]. notions [507]. Novel [768, 906, 1559, 333, 2244, 1572]. Novikov [345, 750, 1019]. nuclear [2116]. Null [1401, 1230, 1341, 817, 1930]. null-plane [1230]. number [1772, 1534, 1480, 2089, 895, 18, 126, 1295, 540, 2283, 1296, 1947, 339, 536]. number-phase [1295]. numbers [483, 1722, 1656, 2257, 1931, 631, 1235, 944, 2310, 377, 2135, 1028, 687]. numeric [891]. Numerical [1991, 2030, 503, 506, 1881]. NUT [1078, 1079, 853, 1226].

O [1956]. observable [365]. observables [705, 1115, 1102, 1369]. observation [1583]. observer [818]. obstructions [1042]. obtain [1599]. obtained [446]. obtaining [1192]. Obukhov [216]. occupation [582]. ocean [2106]. Octonic [354]. octonion [193, 638]. odd [1983, 1278, 416]. odd-dimensional [1983]. Off [2213]. Off-shell [2213]. on-resonance [1340]. One [947, 25, 808, 627, 2017, 1401, 2298, 1924, 1676, 1683, 33, 2254, 669, 1268, 388, 633, 1470, 1457, 1158, 1083, 1522, 1749, 2111, 1101, 218, 789, 1809, 1481, 2113, 1956, 1141, 1624, 1372, 1254, 1997, 1193, 44, 1139, 1954, 163, 1628, 1783, 851, 2064, 891, 21, 301, 909, 212, 1135, 2122, 1089, 966]. One-baryon [627]. one-connected [1139]. One-dimensional [808, 2017, 1401, 33, 388, 633, 1083, 1749, 1101, 218, 1809, 1481, 1141, 1624, 1372, 1193, 44, 1954, 163, 1628, 851, 2064, 21, 301, 909, 212, 1135, 1089]. one-frequency [2113]. one-loop [1522]. One-parameter [25]. one-particle [669]. one-shot [2298, 1924]. one-sided [1158]. One-species [947]. one-way [1254, 1997]. only [1642, 397, 1727]. Onsager [2179]. open [1249, 1990, 603, 1805]. open-string [1990]. Operational [1250, 893]. Operator [1254, 248, 1529, 680, 1983, 200, 796, 1145, 1559, 2009, 997, 78, 894, 2153, 1252, 404, 1886, 56, 455, 434, 896, 1297, 490, 1363, 1096, 509, 1682, 2235, 402, 842, 1817, 1547, 995, 44, 1963, 503, 1711, 522, 1068, 70, 223, 1839, 1673, 958, 471, 1951, 362, 928, 1296, 1947, 1946, 1159, 1353]. operator-algebraic [894]. operator-valued [1145]. operators [2177, 1403, 2199, 80, 1094, 1093, 1095, 483, 1367, 587, 1943, 1459, 1605, 11, 1720, 1913, 1405, 1461, 907, 1525, 1917, 2257, 33, 1599, 1364, 142, 1441, 1660, 740, 1136, 1160, 502, 393, 1792, 415, 663, 894, 962, 671, 672, 1118, 210, 2194, 2110, 2063, 744, 296, 839, 831, 2113, 260, 416, 1799, 269, 1807, 23, 1648, 2214, 1027, 872, 753, 1616, 1186, 2229, 380, 1832, 2182, 942, 1527, 2065, 1908, 1978, 1894, 2283, 2085]. opinions [2277]. Oppenheimer [859, 1417, 1684, 395, 746, 774, 837]. opposite [2277]. optical [982, 546, 75]. optics [1406, 1797, 1898, 1575]. optima [2031]. Optimal [705, 487, 1369, 922, 1446, 37, 1081, 218, 602, 961, 682]. optimality [1924, 1446]. optimization [2047]. optimizers [1721]. optomechanical [2101]. orbifolds [1959]. orbit [1071, 1765, 608, 840, 7, 38]. Orbital [265, 877, 2114, 1091, 1669]. orbits [937, 1803, 2204, 1032, 2028]. order [1456, 628, 122, 1323, 1425, 1207, 745, 488, 1480, 1347, 1737, 821, 728, 697, 962, 381, 1581, 802, 1831, 921, 1873, 2010, 2056, 397, 1340, 1445, 173, 1672, 1824, 605, 2155, 389, 1131, 457, 418, 2252, 1437, 1611, 677]. ordered [660, 217, 1299]. Ordering [1812, 483, 78, 1295, 34]. orders [963, 1107, 1486, 1936]. ordinary [745, 1831]. orientable [2279]. origin [563, 1741, 416, 1554]. Orlicz [918, 1674]. Ornstein [1266]. Orthogonal [1430, 2278, 281, 1431, 1867, 1655, 2043, 328, 1896, 726, 689, 1934, 2228, 691, 138, 700, 2181, 220, 294, 598, 942, 1977, 1308, 164, 1691, 2006]. orthogonal-orthogonal [2181]. orthosymplectic [309]. Oscar [1200]. oscillating [792, 1672]. oscillation [1713, 2268, 766]. oscillations [2029]. Oscillator [925, 1270, 1499, 1453, 242, 1425, 606, 1267, 36, 572, 841, 1516, 1745, 2100, 1, 1995, 2212, 1041, 1939, 2324, 1252, 2063, 1299, 2244, 1846, 2195, 236, 691, 273, 890, 1099, 702, 278, 1746, 32, 924, 1700, 1626, 1152, 2188, 2087, 1533]. oscillators [122, 2017, 954, 257, 454, 1083, 2108, 1281, 1532, 1914, 47, 34, 892, 124]. Oscillatory [391, 469, 1051, 1794, 1204]. osp [2228]. other [1153, 1961]. outcome [1817, 1547]. output [56, 1422, 83]. overcompleteness [615]. overdamped [1058]. overdamping [665]. overlaps [1741]. overview [1722]. Ozols [724].

packet [125, 1716, 485]. packets [1253, 851, 2064, 1138]. Painlevé [1385, 57, 1772, 2107, 585, 1484, 2282, 1006, 181, 333, 675, 2111, 1828, 376, 139, 1777, 691, 501, 1468, 1085, 1866, 1387, 1826]. Painlevé. [258]. pair [462, 63, 9, 1684, 1452, 1487, 2222]. pairings [2204]. pairs [16, 916, 1318, 839, 831, 941, 1651]. Pais [1267]. para [237, 1834, 1888]. para-Grassmann [1834]. para-Hermitian [1888]. para-Jacobi [237]. Parabolic [2098, 787, 1275, 1872, 1204, 2118, 2237, 843, 1901, 1394]. parabolic-parabolic [843, 1901]. paradox [1623, 2018]. Parameter [804, 542, 28, 2110, 1103, 1828, 846, 1864, 300, 1372, 25, 278, 658, 1486, 1936, 2293]. parameters [1197, 1206, 118]. Parametric [382, 721, 954, 1745, 2100, 719]. parametrization [1778]. Parasyuk [17, 945]. parity [1828]. part [637, 1017, 2128, 1486, 1936, 546]. Parthasarathy [133]. Partial [111, 1181, 885, 650, 2213, 106, 987, 1664, 707, 188, 61, 1669, 2142, 880, 1090, 930, 1056]. Partially [325, 2216]. Particle [820, 2231, 1172, 1411, 852, 426, 669, 174, 2197, 1434, 1848, 1086, 2307, 1002, 2209, 1356, 1759, 2195, 14, 154, 1055, 1994, 1971, 2246, 1020, 2032]. particle-field [1020]. Particle-like [1172, 1411]. particles [1999, 1703, 196, 994, 2147, 132, 1169, 1400, 593, 101, 1381, 2255, 166, 2024, 1139]. particular [377, 824, 843]. partite [507]. partition [1775, 2197, 379, 1077, 76, 2215, 2037]. partners [278]. passive [1164, 1706]. patch [2231, 1940, 2321]. Path [992, 1589, 1540, 253, 1073, 592, 1848, 1584, 217, 85, 1438, 1388, 1040, 1458, 1128, 860, 857, 2245, 615]. path-sums [217]. Pathological [577]. paths [175, 860]. Patlak [843]. Pauli [200, 487, 1994, 2297]. PBW [2128]. PDEs [122]. peak [1911]. peakons [1873]. peculiarity [377]. Pedal [1153]. Pempinelli [63, 9, 1286]. Penrose [2035]. Percival [1567, 1221]. Perfect [997, 999, 2077, 289, 571, 678, 208]. Perfect-fluid [2077, 571, 678, 208]. perfectly [358]. perimeter [495]. Periodic [1393, 147, 126, 2028, 45, 769, 1605, 1345, 209, 1436, 2207, 1442, 801, 2163, 2113, 1498, 1307, 1856, 2151, 134, 1032, 493, 713, 291, 1967, 864, 303, 96, 900, 1858, 577, 909, 1790, 84, 1733, 67, 1712, 1843, 1020, 306, 235, 536, 1727, 1491, 2085, 1091]. permittivity [2040, 2109]. Permutation [1959, 329, 816]. permutations [2229, 1646]. Perron [1513]. Persistence [1321, 2223]. persistent [1440]. perspective [1697]. perturb [232]. perturb-exponentially-locally [232]. Perturbation [98, 1864, 1267, 1480, 1444, 1297, 1916, 1535, 1756, 291, 717, 603, 668, 1725]. Perturbations [2183, 995, 164, 1890, 669, 1268, 2009, 1393, 1442, 1749, 282, 1685, 2263, 1838, 232]. perturbative [174, 875, 857, 2300]. Perturbed [585, 1281, 2244, 551, 1900]. Peter [436]. Petiau [1041, 1384, 2198]. Petviashvili [1203, 419, 706, 1895, 459, 463]. Petz [814, 1545]. Pfaff [876]. Pfaff-Toda [876]. Phan [2091]. Phan-Thein-Tanner [2091]. Phase [1860, 456, 1121, 1597, 1646, 1367, 1703, 228, 106, 2041, 1140, 592, 633, 1702, 2268, 631, 78, 801, 2062, 1001, 2210, 2205, 810, 55, 1871, 1925, 1295, 1881, 1635, 1238, 795, 819, 1586]. phase-locking [631]. phase-modulated [1702]. phase-space [633, 801]. phases [233, 2294, 1627, 1300, 1783, 1025, 800]. phenomena [1285, 665, 2012, 171, 372]. phenomenon [1203, 1530]. Phillips [293]. phonon [359]. phonons [124]. Photon [2062, 1688, 1223, 276, 722, 1853]. Photon-added [1688, 276]. photon-subtracted [1688]. photonic [120]. photons [618]. Phys [2043, 145, 1720, 841, 426, 1980, 915, 945, 302, 2323, 916, 1719, 1160, 465, 2324, 2325, 2327, 428, 2326, 641, 1017, 63, 608, 2328, 1438, 1982, 119, 2320, 1199, 2322, 1981, 1319, 427, 1439, 842, 1569, 1320, 977, 640, 678, 4, 1840, 1898, 2042, 607, 1489, 301, 1130, 1839, 2321, 2319, 19, 642, 8, 1947, 1946, 2087, 339]. Phys. [1389, 1359, 1358]. Physical [1149, 1892, 1655, 1525, 245, 302, 583, 1595, 2147, 839, 1616, 520, 418, 1487]. Physics [2314, 2329, 511, 758]. Pi [2329]. piecewise [620]. Pieri [1879]. Piron [367]. piston [1210]. Pitaevskii [2184, 2222, 1016]. pivotal [2295]. Planar [1282, 1902, 926, 902, 2028, 667, 1475]. Planck [640, 1599, 2082, 1265, 58, 176, 10, 121]. Plane [2076, 664, 1429, 1825, 2023, 1230, 777, 2200, 784, 1823, 1410, 660, 802, 452, 1911, 1876, 2215]. planetary [1594]. plant [56]. plasma [24]. plasmas [1707, 525]. Platonic [2296]. Plebański [1304]. plethora [1134]. PMT [1755]. Podolsky [197, 1811, 1230, 2139]. Poincaré [1339, 1926, 1996, 323, 330, 331, 1450, 1008, 1260, 2264, 2231]. Poincaré-like [330, 331]. Point [1442, 1401, 874, 1980, 926, 1833, 2323, 614, 1024, 1389, 1120, 801, 2322, 1184, 1617, 1981, 1229, 1934, 168, 2061, 1711, 607, 890, 162, 135, 13, 19, 1159]. pointer [1539]. points [399, 1114, 828, 972]. Poisson [947, 1902, 1913, 683, 1347, 468, 227, 1701, 611, 2054, 409, 1643, 489, 1891, 829, 1192, 752, 1893, 1905, 150, 844, 1869, 704, 467, 1877, 2092, 1723, 253, 270, 1167, 2011, 343]. Poisson-commuting [1192]. Pol [1425, 1470, 493]. polar [671, 1539]. polarization [264]. polarizations [1044]. polaron [74]. Pollard [1892]. Pollard-like [1892]. Poly [1701]. Poly-Poisson [1701]. poly-symplectic [1701]. Polyanalytic [215]. polygon [1686, 667]. Polygonal [1471]. polygons [1786]. polyhedra [1925]. polylogarithms [963]. polymer [2071]. Polynomial [78, 2126, 1069, 915, 836, 2200, 393, 2327, 1026, 1581, 1600, 1192, 1011, 1916, 691, 1308, 537]. polynomially [1182, 300]. Polynomials [2200, 1430, 1867, 1655, 2043, 237, 2220, 2057, 450, 2254, 1926, 2278, 1151, 356, 259, 1158, 502, 1896, 281, 944, 1779, 246, 92, 2228, 2311, 437, 1880, 825, 1914, 797, 138, 1246, 220, 294, 598, 942, 1431, 1977, 164, 252, 1159, 1100]. polytropic [171]. Ponderomotive [868]. porosity [1837]. porous [1798, 1037, 1837]. ports [1824]. Pöschl [237]. posed [347]. posedness [2175, 2091, 1910, 2174, 662, 1492, 2090, 1984, 2012, 792, 1662, 1287, 2171, 2224, 917, 1791, 121, 1165, 351, 1989]. position [720, 1151, 994, 2062, 1299, 1759, 1583, 159, 1240, 34, 654, 38, 1804, 485, 1155, 1533, 2297]. position-dependent [720, 1151, 994, 1759, 38, 1804, 1155, 1533, 2297]. Positive [136, 2238, 648, 2293, 1467, 2239, 1059, 918, 1902, 444, 2072, 1145, 2172, 1444, 1660, 1218, 509, 1422, 402, 842, 1259, 1448, 1874, 521, 1226, 46, 1202, 930, 1399, 2128, 148, 1322, 2044, 2011, 1812]. positive-operator [402, 842]. Positivity [504, 1752, 507]. positronium [560]. Potential [2099, 1422, 1233, 279, 779, 1949, 254, 237, 304, 1793, 2250, 2213, 426, 1782, 2219, 2107, 1668, 1497, 981, 579, 791, 1154, 1463, 612, 611, 1852, 191, 766, 1252, 759, 331, 608, 912, 1265, 1416, 1612, 1808, 2209, 466, 7, 1368, 1141, 1759, 1540, 1653, 300, 1458, 1501, 2008, 617, 14, 762, 1366, 679, 693, 1670, 2114, 50, 1357, 1208, 1842, 1806]. Potentials [202, 1915, 1456, 80, 187, 606, 245, 302, 1730, 2292, 2081, 875, 764, 1794, 1602, 2327, 35, 594, 1663, 2194, 1123, 1129, 2110, 1884, 1982, 947, 2049, 974, 1657, 1600, 1625, 793, 1648, 397, 2245, 1994, 1084, 1185, 1536, 1863, 190, 235, 721]. Potts [1040, 1383]. POVMs [1374]. Power [1196, 696, 724, 2081, 2055, 2192, 1975, 86, 300, 1885, 2114, 234, 2240, 1409, 682]. power-law [2081]. power-type [2055]. powers [1685, 504, 1814]. Prandtl [2256, 1876]. Prasad [576, 770, 460]. pre [2329, 2180]. pre-Lie [2180]. pre-Newtonian [2329]. precision [1962]. Preface [1768]. prefundamental [925, 1270]. premodular [1404]. preparation [653, 1962, 1811]. prepare [1632]. prepare-and-measure [1632]. prepotentials [898]. prescribed [1065, 209]. prescribing [447]. Prescription [557]. presence [1549, 261, 606, 2309, 35, 584]. preserving [1007, 2000, 2026]. Pressure [2115, 1111, 2286, 100, 1080, 920]. pressureless [227]. prevent [1724]. Prime [986]. primitive [2067]. principal [394, 637, 1017, 684, 903, 1489]. principle [1754, 772, 2206, 2166, 1395, 1517, 1510, 2191, 866]. principles [1097, 315, 1349]. prior [314]. privacy [855]. Private [510]. probabilistic [197]. probabilities [2102, 1381, 514]. Probability [1786, 2147, 1508, 107, 1220, 1294, 2234]. probe [1552]. problem [1168, 312, 45, 1239, 664, 1640, 57, 1676, 1683, 1426, 1112, 1555, 1480, 2288, 1754, 1904, 1870, 1665, 1910, 2168, 1530, 2081, 133, 2223, 1596, 2276, 937, 1210, 2097, 1054, 1092, 934, 1376, 750, 1568, 830, 1831, 1614, 902, 1695, 1541, 2098, 1060, 2227, 1717, 1871, 2286, 1032, 2028, 2242, 65, 2056, 1954, 521, 199, 979, 712, 790, 1502, 1594, 2118, 2158, 540, 808, 263, 46, 2262, 910, 39, 1940, 2321, 658, 75, 2052, 135, 1471, 67, 1460, 1715, 2222, 1475, 71, 2048, 2085]. problems [80, 2239, 918, 2269, 228, 1153, 1245, 1998, 768, 2292, 871, 385, 759, 180, 674, 1416, 1612, 2170, 929, 787, 2214, 211, 1674, 1074, 890, 2237, 451, 654, 1373, 476, 1839, 1673, 135, 410, 447]. Proca [1066]. procedure [600, 1378]. process [1367, 326, 597, 1782, 772, 509, 1427, 1117, 1382, 1769]. processes [1942, 412, 634, 773, 874, 495, 1039, 337, 905, 1934, 640, 2234, 946, 521, 10]. processing [763]. Prodi [535, 978]. product [109, 1365, 1047, 404, 1816, 182, 490, 2204, 332, 498, 1439, 670, 622, 1750, 2103]. production [772, 1810, 556, 2231]. products [368, 1236, 1102]. profile [2121, 1185]. profiles [344, 1322]. projected [1452]. projection [1094, 1521]. projections [328]. Projective [965, 1108, 274, 2127, 1244, 322, 732, 1304, 1076]. projector [766]. Prolate [161, 2098]. Proof [2181, 955, 1306, 27, 192]. propagating [982]. propagation [1398, 1666, 1882, 1743, 64]. Propagator [1458, 481, 277, 1670]. propagators [1792]. proper [745]. Properties [1367, 254, 2164, 450, 1485, 242, 1346, 276, 1150, 821, 1000, 2021, 214, 1161, 523, 906, 534, 596, 1892, 2248, 1915, 1003, 1246, 546, 703, 743, 421, 1574, 627]. property [766, 948, 99, 897, 574, 229, 856, 1321]. Propositional [2018]. Protected [554]. protocol [725]. proton [770]. provides [2137]. providing [745]. Pseudo [1071, 1774, 680, 420, 673, 533, 1341, 898, 466, 831, 1277, 1682, 2084, 732, 1201]. pseudo-bosonic [831]. pseudo-bosons [673]. pseudo-embedded [1774]. pseudo-Gaussian [466]. pseudo-hyperkähler [898]. pseudo-index [2084]. Pseudo-orbit [1071]. pseudo-projective [732]. pseudo-relativistic [680, 533, 1201]. pseudo-Riemannian [1341, 1277]. pseudoanalytic [681]. pseudoconvex [1930]. pseudodifferential [1441, 740]. pseudofermionic [550]. pseudoscalar [805, 2042]. Pseudospectra [425]. PT [1602, 1046, 721]. PT-symmetric [1602]. PT-symmetry [1046]. Publisher [465, 302, 1719, 1160, 1982, 1569, 977]. Pullback [374, 2002, 1228]. pulled [2004]. pulse [179, 2012, 1287, 923, 1316, 891, 1937]. Pure [1344, 2329, 1343, 484, 1588, 2061, 229, 762, 234]. purely [1312, 1501]. purification [479]. purifications [757]. purity [1422]. q [2057, 299]. q-deformed [299]. q-Krawtchouk [2057]. QCD [897, 627]. QED [1012, 2139, 1423]. QFT [2215]. qHV [127]. QRT [2236]. quadrant [1088]. Quadratic [853, 173, 572, 1676, 2323, 1136, 1160, 1360, 2322, 573, 607, 1760, 34, 656, 96, 13, 19, 531, 778, 2291, 2150]. quadrature [553]. quadrupole [2325, 241, 1292, 2246]. Qualitative [572, 534]. quantisation [274]. quantitative [1362, 1988, 1974]. quantity [1301]. Quantization [815, 1746, 2071, 592, 1642, 731, 692, 975, 1199, 1846, 1744, 810, 1479, 971, 977, 1187, 2193, 336, 561, 1503, 1687, 711]. quantized [36, 357]. Quantum [279, 1544, 1075, 1248, 1605, 1223, 1371, 1140, 1920, 401, 727, 2325, 2021, 759, 241, 609, 1585, 442, 584, 807, 1552, 944, 1813, 1923, 2102, 573, 1658, 334, 1299, 1141, 1119, 1624, 1997, 1547, 1046, 1355, 2329, 2066, 521, 1025, 1180, 890, 38, 1527, 2022, 290, 743, 1928, 40, 370, 586, 1409, 1456, 453, 957, 1339, 1453, 2017, 37, 1367, 2060, 1023, 2269, 254, 951, 1459, 338, 2086, 73, 1429, 1924, 309, 990, 1178, 1412, 969, 1917, 814, 729, 1146, 90, 1860, 554, 1815, 426, 1462, 1636, 177, 773, 165, 1851, 812, 1302, 2104, 1964, 669, 1992]. quantum [496, 2033, 399, 1659, 1998, 1406, 619, 954, 1745, 2100, 497, 1364, 363, 239, 724, 2299, 284, 2197, 763, 1048, 542, 1127, 1735, 1255, 1996, 477, 1134, 517, 366, 510, 128, 1126, 1421, 728, 697, 1000, 2016, 558, 1721, 2270, 105, 406, 88, 993, 2136, 1071, 1537, 2144, 992, 1579, 277, 35, 1310, 214, 653, 1749, 28, 722, 2284, 136, 1816, 1177, 56, 82, 948, 973, 1198, 1567, 1257, 1250, 1538, 76, 1508, 1103, 1369, 1747, 2192, 119, 369, 1376, 896, 512, 317, 513, 2209, 131, 505, 2196, 1509, 961, 830, 847, 726, 846, 689, 1249, 959, 1215, 1651]. quantum [1853, 690, 1681, 86, 635, 991, 1748, 425, 1254, 1217, 1583, 2235, 365, 1891, 965, 1108, 500, 850, 670, 514, 2096, 2024, 1104, 2285, 315, 1785, 236, 691, 14, 353, 713, 123, 1685, 753, 1845, 1106, 556, 1740, 1840, 2088, 950, 551, 1587, 2230, 717, 273, 294, 1562, 511, 1304, 1076, 967, 1885, 1068, 1221, 1510, 1099, 2300, 1753, 960, 1922, 34, 1413, 1142, 605, 892, 708, 1692, 1301, 1349, 476, 560, 1881, 39, 746, 774, 837, 293, 1247, 311, 161, 800, 1694, 1138, 1298, 2247, 2188, 272, 6, 8, 2283, 2191, 421, 2014, 1155, 1167]. quantum [1175, 1708, 1960, 1805, 83, 1812, 2305]. quantum-dynamical [2305]. quantum-group [1459]. quantum-mechanical [605, 476]. quantum-mechanics [1853]. quartic [2140, 1854, 2215]. Quasi [1897, 209, 1310, 1102, 1835, 1307, 1856, 1193, 291, 108, 2088, 1858, 1626, 2087, 1795, 1453, 1436, 281, 2097, 1442, 2113, 2265, 134, 127, 493, 617, 47, 1099, 96, 900, 135, 1460, 1712]. Quasi-additive [1193]. Quasi-Bell [1102]. quasi-bound [617]. quasi-classical [281]. Quasi-coherent [1626, 2087]. Quasi-compact [1835]. Quasi-diagonal [1310]. quasi-exactly [1099]. quasi-harmonic [47]. quasi-homogeneous [135]. quasi-integrable [1795]. Quasi-isomorphisms [1897]. quasi-local [2088]. Quasi-locality [2088]. quasi-nonassociative [1453]. Quasi-periodic [209, 1307, 1856, 291, 1858, 1436, 1442, 2113, 134, 493, 96, 900, 1712]. quasiclassical [1271]. quasicrystalline [676]. quasicrystals [2186]. Quasilinear [1166, 2165, 186, 918, 266, 1444, 1730, 612, 1054, 1950, 1064, 1018, 648, 1204, 1330, 2237, 1322, 1725, 2293, 2044, 881]. Quasiperiodic [2108, 1890, 1648, 178]. quasistatic [1164]. quaternion [335, 1979]. quaternion-valued [1979]. quaternionic [587, 1059, 671, 672, 1118, 335, 1717, 1832, 595]. quaternions [336]. qubit [316, 1102, 1257, 1300, 1689, 1026, 1142, 695, 658]. qubit-state [658]. qubits [1962]. qudits [696, 622]. queer [685, 1133]. questions [895]. quintic [2174, 1700]. quiver [1171].

r [210, 921, 341]. Racah [1023, 1912, 1977]. Radial [538, 1084, 540]. Radially [984]. radiating [858, 1190]. radiation [1696, 1343]. Radiative [1128, 24, 771, 1011]. radius [503]. Rainich [1004]. raised [2060]. Rajeev [2205, 206]. Randers [2302]. Random [517, 1389, 1120, 1788, 2145, 2318, 62, 408, 373, 453, 414, 411, 1690, 518, 2220, 94, 1116, 1794, 1588, 2147, 1859, 1383, 494, 2210, 107, 1380, 1114, 581, 1011, 1428, 1105, 828, 2230, 1308, 2187, 595, 1974, 749, 2078, 2027, 64, 244]. randomly [2277, 176]. range [1313, 1038, 1683, 49, 1101, 2113, 1193, 1881, 2065]. ranges [506]. Rank [435, 1862, 2179, 1268, 1404, 328, 799, 382, 652, 367, 2208, 2143, 355, 2152]. rank-3 [2152]. Rank-inductions [435]. rank-one [1268]. Ranken [2205]. ranks [1270]. rapid [1285]. rapidly [399]. rarefaction [1210]. rarefied [1557]. Rarita [1819]. Rascle [1796]. Rashba [38]. Rate [1497, 251, 772, 1024, 1205, 919, 1810, 1705, 2263, 657, 308]. rates [922, 524, 1961]. Rational [237, 768, 1698, 1281, 1686, 2063, 332, 236, 691, 700, 890, 2001, 892, 2014]. rationalities [434]. rationally [1625]. Rayleigh [919, 1264, 201]. reaction [1282, 429, 2170, 1569, 1515, 2145, 303, 2201, 2002]. reaction-advection-diffusion [303]. reaction-convection-diffusion [429]. reaction-diffusion [2170, 1569, 1515, 2145, 2201, 2002]. Real [709, 1297, 1681, 252, 109, 1188, 1917, 1341, 1641, 887, 1800, 1047, 948, 1368, 322, 537, 721, 2215]. realisation [1175]. realistic [363]. reality [1686]. Realization [1170, 745, 28, 2267, 1449, 521]. realizations [1521, 438, 686]. realized [1403, 905]. reciprocal [1575]. reconstruction [976, 1545]. Recovery [2260, 2138, 814, 1545]. rectangular [2290]. Recurrence [1581, 220, 598, 942, 1977, 79, 138]. recursion [1722, 1559, 210, 471]. Recursive [1882]. redistribution [697]. reduced [2236, 2185, 1513, 2204, 950, 1937]. Reducibility [1890]. reducible [1403]. Reducing [895]. Reduction [1521, 685, 976, 1803, 321, 1468, 476, 2161, 863, 1179, 1345, 879, 1555, 1643, 2047]. Reductions [1895, 972, 1437, 205, 24, 386, 633, 1256, 1330, 457, 834]. Referee [1948]. Refined [298]. reflecting [1220]. Reflection [1752, 2312, 914, 1046, 163]. reformulation [193]. refraction [1844]. regime [2184, 1776, 2149, 44]. regimes [2082, 1037]. regional [1327, 1665]. regions [1435, 1634, 430]. Regular [831, 629, 1248, 1429, 180, 132, 1844, 839, 941, 347, 1334, 667]. Regularity [821, 267, 1838, 1203, 1423, 492, 1666, 535, 2010, 1465, 1791, 1029, 978, 823].

Regularization [934, 248, 1702, 780]. regularized [987]. Regularizing [1388]. Reidemeister [1738]. Reinsch [591]. Reissner [2035]. rejuvenation [967]. related [1786, 863, 495, 1116, 210, 1045, 1234, 2307, 321, 313, 691, 1490, 531, 1067]. Relation [2192, 486, 1146, 2098, 1893, 1049]. relational [1701]. relations [624, 1194, 1534, 1972, 2251, 1433, 2328, 299, 802, 1799, 2181, 220, 598, 942, 1977, 1866, 1708]. relationship [1551]. Relative [1802, 72, 1634, 555, 960, 156, 1586, 87, 1146, 2069, 1969, 1933, 1692, 135]. Relativistic [740, 91, 1557, 2005, 680, 1999, 371, 284, 144, 533, 662, 1158, 2206, 1210, 1645, 417, 1952, 1509, 1684, 786, 2024, 1216, 1578, 215, 719, 1201, 2115, 525, 432, 1907, 178]. relativistically [166]. relativity [599, 936, 1109, 1344, 1260, 447]. relaxation [1514]. Remark [867]. remarkable [142]. Remarks [630, 1763, 1045, 1514, 1956, 51, 234, 781, 878, 1293, 26]. Remling [589]. Remote [1811]. Removal [34]. removed [646]. Renner [478]. Renormalizable [1854]. Renormalization [1269, 1703, 1941, 406, 528, 1927]. renormalized [1622, 2096]. Renyi [869, 87, 90, 2067, 88, 893, 813, 1692, 1961]. repeated [603]. Rephasing [626]. replica [2210]. Representation [1247, 1978, 957, 1403, 1233, 1183, 1430, 538, 255, 481, 231, 1261, 477, 2202, 28, 382, 333, 2102, 1736, 1481, 313, 1174, 613, 1049, 1175]. Representations [1689, 605, 1566, 543, 1351, 1976, 1860, 925, 1270, 1194, 440, 328, 329, 889, 142, 1521, 685, 1735, 1134, 280, 152, 742, 1236, 337, 1450, 26, 367, 1811, 1289, 1589, 1336, 1739, 1087, 1316, 870, 924, 474, 310]. representatives [1684]. representing [2141]. repulsive [669, 2114]. rescaled [1652]. Researchers [2329]. reservoirs [2047]. residual [639]. resistive [1198]. resistivity [103, 2124]. Resolvent [1096, 1363]. resonance [2116, 1745, 1340, 2291]. Resonances [1268, 1794, 1071, 1663, 635, 395, 1762]. Resonant [1825, 2061, 619, 1699, 1861]. resonant-state [619]. Resource [1632, 960]. respect [20, 145, 1197, 1372, 1527]. Response [2323, 1033, 1981, 1320, 642, 2312, 1837, 451]. Restoring [2111]. restricted [664, 2023, 92, 2085]. Restriction [1353]. restrictions [105]. result [628, 144]. results [2005, 1599, 1116, 1522, 1584, 1149, 1078, 1079, 2237, 891, 389, 2074, 1839, 1673, 2247, 235, 2162]. Resurgence [219]. Retraction [2326, 1898]. reversal [709, 548]. reverse [401]. revisited [815, 1889, 372]. Reynolds [377]. Rham [705]. ribbon [141]. Riccati [2142, 858]. Ricci [59, 1593, 1553, 1277, 899, 1243, 1043, 1565, 966]. Ricci-flat [1593]. Riemann [424, 228, 2168, 1121, 2286, 712, 2240, 778]. Riemannian [1339, 1239, 2123, 2112, 1945, 2045, 1341, 2136, 1277, 758, 114, 2306]. Riesenfeld [1846]. Riesz [969, 1426, 833, 941, 974, 421]. Rigged [1288, 763, 1601, 2219]. right [671, 686, 68]. right-angled [68]. rigid [1904, 734]. rigidity [936]. Rigorous [2005, 1398, 1893]. rings [271, 1573, 1247]. Rivlin [470]. RLC [1544]. RN [1839, 1494, 1950, 2094, 1673, 2002, 2145]. Robe [2085]. Robertson [678, 1551, 1353, 1125, 287, 571, 899, 2077, 1473]. Robin [747, 260, 2290]. Robinson [2088, 553]. robust [1749]. Robustness [1857]. rod [265]. rods [2233]. rogue [742, 1712, 1718, 677]. role [1058, 2144, 1300, 574]. roll [43]. Root [2242, 2254, 1934]. rooted [167, 884]. roots [651, 543, 537]. Rosen [608, 197, 1811, 7]. Rota [884]. rotating [756, 1292, 1821]. Rotation [1035, 1719, 1644, 1454]. rotational [1825]. rotationally [2074]. rotations [487, 356]. rotator [734]. rotopulsators [1471]. Rouet [639]. route [1616]. RPI [1755]. rule [1071, 850, 2193]. ruled [817]. rules [2177, 1093, 1938, 27, 88, 973, 1289, 553]. Rydberg [1319, 1320, 1368, 12].

S [239, 692]. S-matrix [239]. Sachdev [2140]. Sachs [1174]. saddle [779, 162]. saddle-like [779]. SAdS [1952]. Salpeter [1754]. same [20, 145]. sandpile [1647]. sandwich [1306]. sandwiched [2067]. Sanov [83]. Sasa [1903, 1715]. Sasakian [1312]. satisfied [1466]. satisfying [397]. Satsuma [1903, 1715, 834]. saturable [982]. Savart [2257, 1951]. saving [621]. Saxon [617]. Saxton [1446]. scalar [1549, 1075, 1752, 1640, 2123, 17, 945, 406, 249, 1236, 289, 1550, 2046, 2215]. scalar-tensor [1075, 1640]. Scale [1362, 862, 93, 1258, 707, 2108, 1610, 2253, 250, 866]. Scale-free [1362]. Scale-invariant [862, 93]. Scaled [1465, 412]. scales [407, 819]. scaling [258, 1385, 2251]. Scarf [1233, 1368, 1357]. scatterers [801, 1762]. Scattering [1405, 388, 1162, 617, 358, 2291, 362, 1390, 1429, 2160, 94, 914, 2081, 2174, 366, 759, 35, 344, 1416, 1141, 1060, 419, 1915, 1003, 244, 243, 838, 2064, 201, 1526, 577, 293, 2065, 75, 410, 1016, 1805, 1946]. scatterings [405]. scenario [994]. scheme [476]. Schlömilch [1702]. Schlömilch-type [1702]. Scholes [673]. Schramm [1801]. Schrieffer [523]. Schrödinger [302, 2327, 1905, 1947, 1946, 2011, 1390, 1168, 1351, 80, 312, 1529, 1323, 1327, 1902, 680, 1949, 1405, 1461, 1890, 197, 1676, 1182, 1793, 228, 1992, 488, 388, 245, 266, 1127, 1444, 1668, 1730, 2045, 144, 533, 1251, 1457, 878, 1463, 2016, 1392, 1667, 611, 1792, 439, 799, 1162, 2055, 2054, 191, 1252, 803, 2080, 136, 1013, 2194, 1669, 1567, 1123, 2110, 422, 1570, 2163, 381, 352, 2258, 1822, 466, 1184, 1096, 974, 260, 350, 1614, 534, 1600, 321, 2311, 1372, 1807, 480, 995, 1501, 472, 782, 1064, 1307, 1018, 1448, 1523, 2008, 44, 134, 150, 844, 877, 1648, 2084, 346, 835, 467]. Schrödinger [620, 163, 1628, 762, 361, 1366, 679, 1445, 2289, 199, 1240, 273, 1994, 1670, 70, 2114, 1397, 1993, 264, 1877, 1329, 864, 1986, 234, 781, 1672, 1414, 1863, 223, 1202, 2092, 1723, 349, 2165, 84, 2065, 1971, 2259, 1908, 943, 1603, 190, 148, 66, 1322, 362, 928, 1296, 938, 1049, 1159, 222, 2293, 785, 341, 804, 984, 677, 343, 881, 77]. Schrödinger-Poisson [343]. Schroedinger [806]. Schur [1850, 1481, 92, 700]. Schwarz [1865, 2315]. Schwarzian [1610]. Schwarzschild [2326, 2249, 1363, 1358, 1756]. Schwarzschild-anti-de [1363]. Schwinger [1438, 1819, 1605, 383, 85, 1141, 1069, 1883, 2152]. SdS [1952]. SDYM [1271]. search [1345, 1542]. Searching [1108]. Second [728, 122, 1176, 1798, 745, 1347, 697, 802, 1831, 1381, 215, 1432, 1824, 252, 457, 1437]. Second-order [728, 122, 745, 1824, 457]. Secrecy [2104]. Secret [1302, 2104, 812]. secret-key [812]. sectional [1122]. Sedeonic [1232]. seed [2143]. Segal [988, 1317]. Segel [843, 1906, 1724]. Segregated [346]. Segur [2282, 52, 1829]. Seiberg [1897, 2154]. Selection [2223, 950]. selective [1298]. Self [11, 1720, 326, 2073, 547, 380, 1075, 907, 1763, 1917, 1531, 396, 694, 227, 1144, 327, 2221, 2185, 52, 1297, 833, 2053, 1181, 2046, 2176]. Self-adjoint [11, 1720, 1297]. Self-adjointness [547, 380, 1075]. self-consistent [1531, 52]. Self-dual [2073]. Self-duality [326, 327]. self-generated [1144, 2185]. self-gravitating [227]. self-interacting [1181, 2046]. self-interaction [396]. self-linking [1763]. self-similar [694, 2221, 2053, 2176]. selfadjoint [80]. selfadjointness [731]. selfsimilar [2286]. semantics [1623]. Semi [1949, 260, 1529, 2112, 179, 82, 1108, 706, 1228, 2010, 708, 1563, 943, 2007, 941]. semi- [708]. Semi-classical [1949, 260, 1529, 1108]. semi-continuity [1228]. semi-discrete [706, 1563, 943, 2007]. semi-discretization [179]. semi-dissipative [2010]. semi-groups [82]. Semi-regular [941]. semi-Riemannian [2112]. semicircle [411]. Semiclassical [76, 693, 1138, 137, 751, 2116, 481, 1370, 1792, 277, 1744, 150, 125, 93, 851, 2064, 2246, 243]. semicontinuity [1569, 1515, 2027]. semidirect [2182]. semigroup [1549, 1000, 701, 703, 603]. semigroups [338, 1255, 2144, 2022, 586, 2283, 2305]. semilinear [2304, 647, 148]. semisimple [1592]. semitrivial [2163]. separability [2241, 1031, 566]. separable [402, 842, 2006]. Separated [2001, 806, 1634]. separating [1456]. Separation [1773, 597, 1597, 771, 1785, 1343, 1502]. sequence [1480, 2093]. Sequential [2103, 566, 1282, 670, 856, 1750]. Series [1867, 1655, 2043, 2081, 278, 1702, 807, 1975, 300, 1372, 828, 591]. Serrin [978, 535]. sesquilinear [1525]. set [20, 145, 994, 1538, 117, 695, 348, 115]. sets [1690, 1044, 316, 1114, 2151, 2229, 1820]. setting [1832]. settings [1544, 728]. several [259, 1482]. sextic [1501]. Shabat [1297, 300]. shadow [129, 130]. shallow [2308]. shallow-water [2308]. Shannon [455]. Shape [499, 720, 890, 1536]. shape-invariant [720]. shapes [798, 2119]. Shapovalov [688]. Sharma [133]. Sharp [2175, 394, 1534, 1052, 679]. shear [1081, 452]. shell [2213, 1027, 1216]. Shemesh [1815]. Sherrington [2031]. shift [2112, 1535]. Shifted [687]. shifts [1346, 1586]. shock [1844, 2311, 2053]. shocks [784, 1907]. short [2071, 1683, 179, 2012, 859, 2211, 1969, 923, 1316, 891, 1937]. short-lived [2071]. short-range [1683]. shot [2298, 1924]. should [861]. Shubin [1441]. SIC [1374]. SICS [2105, 1338]. sided [1158]. Sierpinski [1196]. sigma [1701, 113, 590, 655, 1679, 789, 1956, 815, 1163]. sigma-function [1679]. sign [2161, 2041, 1092, 1493, 2056, 2259, 2293, 984]. sign-changing [1092, 1493, 2056, 2259, 2293]. sign-definite [984]. signal [763, 1906]. signal-dependent [1906]. signaling [123]. signals [1321]. signature [2112, 1945, 2135]. Signatures [1881]. Signed [2056, 816]. silence [661]. similar [694, 2221, 2053, 2176]. Similarity [143, 583, 1855, 1875, 2244].

Simon [1466]. Simons [1165, 268, 390, 129, 130, 146, 789, 1956, 1764, 1042, 1163, 432, 698]. simple [955, 986, 326, 1516, 683, 888, 366, 2212, 1314, 269, 355, 891, 1382, 1769, 230]. simplicial [129, 130]. Simplification [1093]. Simplified [2105, 946]. Simplifying [591, 331]. simulability [1250]. simulation [1117, 723]. simulations [1584]. Simultaneous [1964, 1828, 726]. sinc [674, 1281, 1429]. sinc-collocation [1281]. Sinc-interpolants [1429]. Sine [2253, 187, 1782, 1650, 1937, 526]. sine-Gordon [187, 1937, 526]. Singer [992]. Single [1385, 250, 631, 316, 594, 722, 1853, 1287, 719]. single-cycle [1287]. single-loop [631]. single-qubit [316]. Single-user [1385]. single-well [594]. singlet [2095]. Singular [879, 258, 1872, 2187, 1495, 1417, 1325, 33, 239, 1668, 1497, 674, 344, 1272, 1572, 1685, 1366, 2263, 1332, 1185, 2262, 476, 2011]. Singular-unbounded [2187]. Singularities [1869, 1526, 486, 1980, 3, 956, 1617, 1981, 1358, 4, 1195, 2319]. Singularity [1243, 1987, 11, 1720, 496, 2024, 567]. Singularly [694, 1900]. sinks [1647]. SIR [646, 1402]. site [1348, 943]. site-dependent [1348]. Sitter [2324, 1079, 568, 1592, 2159, 1041, 1939, 859, 1363, 448, 1132, 32, 2287, 2006, 1965, 2046]. Six [106, 1477, 185, 1225, 1597, 1604, 207]. six-dimensional [207]. Six-vertex [106, 1597, 1604]. size [449]. Skew [1879, 146, 885]. skew-derivations [885]. skew-symmetric [146]. Skyrme [770, 104]. skyrmions [1241, 576]. slab [2040, 1526]. slip [1904, 1577, 1279]. slits [240]. Slow [2226, 43, 1514, 1578, 577]. slow-light [577]. slow-roll [43]. Slowly [759, 1864, 1821]. Smagulov [226]. Smale [1757, 2146, 1169]. Small [1588, 2029, 1258, 791, 189, 652, 1762, 358, 1247, 1486, 1936, 530, 1279]. Smoluchowski [2089, 1089]. Smooth [1408, 1403, 515, 1081, 1132, 2245, 616]. smoothing [352, 1378]. smoothness [1123, 1820]. Snyder [2324, 1041, 1939, 32]. SO* [1177]. Sobolev [918, 955, 508, 1444, 1996, 2238, 2258, 1568, 1018, 1674, 1760, 900, 491]. Sobolev-type [491]. soft [947]. softcore [1808]. sol [1833]. sol-manifold [1833]. solar [227]. Sole [192]. solenoid [1291]. solid [1703, 1972]. solid-on-solid [1972]. solids [2296]. Solitary [1272, 1203, 1282, 265, 877, 66, 1135, 222]. Soliton [835, 1311, 1889, 569, 1661, 1570, 2253, 126, 1767, 433, 65, 264, 1900, 843, 908, 1283, 966]. Soliton-generating [1889]. soliton-like [1900]. Solitons [1774, 63, 9, 391, 469, 1559, 1285, 1277, 2311, 982, 838, 1790, 1043, 1565, 1474, 966]. Solution [1183, 57, 2327, 35, 1265, 802, 1600, 2118, 1902, 1429, 304, 1182, 914, 791, 191, 1054, 1123, 1092, 2253, 2049, 1952, 466, 147, 256, 1856, 1985, 2008, 458, 648, 1789, 2224, 1311, 1883, 1088, 361, 2004, 377, 824, 1142, 1734, 1446, 843, 143, 1208, 1167, 1163, 1574, 432, 2011]. Solution-giving [2118]. Solutions [24, 1950, 1018, 45, 769, 1867, 1655, 2043, 2239, 261, 918, 1327, 680, 720, 1949, 1710, 1483, 251, 1639, 569, 228, 1345, 116, 1362, 1531, 806, 1980, 386, 532, 1065, 1664, 245, 302, 209, 2172, 266, 1444, 1730, 1494, 2292, 2045, 579, 1053, 645, 288, 1841, 2276, 590, 612, 2282, 1729, 1393, 1667, 1726, 611, 2202, 1006, 103, 1707, 2054, 2327, 803, 742, 2221, 1033, 608, 1615, 1666, 1984, 2238, 913, 1570, 1331, 429, 1081, 2258, 859, 218, 602, 262, 146, 789, 1828, 7, 1617, 1981, 534, 2040, 1600, 18, 126, 2248, 1132, 1272, 1372, 968, 347, 1072, 1501, 1189, 1328, 1259]. solutions [782, 786, 1064, 1307, 1275, 1693, 1874, 2124, 1905, 2264, 921, 120, 844, 1396, 1869, 1872, 1871, 2084, 346, 1767, 493, 2285, 1493, 835, 1916, 1821, 2056, 666, 1787, 1756, 291, 679, 199, 273, 1994, 979, 1732, 267, 712, 305, 1084, 264, 1877, 47, 208, 1675, 864, 1373, 1900, 278, 892, 2025, 629, 1202, 1723, 303, 2240, 96, 900, 1858, 349, 667, 910, 84, 1399, 2256, 1911, 647, 2252, 67, 1488, 530, 190, 148, 66, 526, 1322, 1843, 1718, 908, 616, 1611, 920, 1725, 2293, 1475, 1704, 339, 341, 1324, 1089, 984, 1609, 2044, 2176, 178, 677, 1286, 1709, 2048, 1283, 1727, 343, 881, 1091]. solvability [2097, 1299, 108, 1460]. Solvable [294, 1110, 2254, 1878, 954, 1698, 2200, 873, 53, 1513, 1625, 166, 1099, 654, 721]. Solve [1979]. Solving [1278, 1281, 1831]. Some [735, 1772, 878, 1293, 1292, 214, 1522, 26, 1381, 2074, 2247, 1323, 242, 1802, 1525, 1599, 592, 188, 1848, 149, 2171, 236, 1212, 880, 418, 1067, 2032]. Sommerfeld [576]. Sommerfield [770, 460]. Sommerfield-like [460]. sound [1798]. source [882, 832, 171, 1551, 1724, 222]. sources [812, 52, 666, 305, 1647, 710]. Space [1636, 286, 1403, 248, 295, 918, 109, 2095, 538, 1913, 1676, 1030, 969, 384, 705, 1073, 2112, 1140, 592, 328, 329, 1642, 231, 2233, 633, 1649, 614, 78, 1278, 1121, 281, 2097, 1696, 1125, 1765, 1886, 132, 445, 801, 1009, 2312, 2059, 2320, 818, 853, 2258, 2253, 1001, 789, 1568, 1952, 2120, 448, 1956, 334, 5, 427, 2205, 322, 1358, 402, 842, 811, 965, 810, 1873, 1925, 127, 1384, 313, 571, 678, 899, 2077, 1174, 753, 4, 503, 1187, 551, 1343, 1078, 1079, 522, 51, 1536, 654, 808, 1373, 143, 1090, 1551, 1928, 2252, 1460, 1635, 1930, 1394, 2176]. space [819, 1533, 1654]. space-fractional [551]. space-time [1568, 1358, 1384, 1174, 1187, 1343, 1078, 1079, 1090, 1928]. space-times [1030, 1125, 1952, 899, 2077, 1551, 1930]. spacelike [705, 448]. Spaces [860, 2127, 1656, 1942, 1802, 1593, 318, 2257, 1010, 1112, 413, 1031, 763, 1601, 2219, 2302, 2091, 1341, 1641, 2183, 1418, 2090, 566, 671, 901, 1508, 1568, 320, 1780, 1518, 322, 500, 1170, 1674, 273, 2318, 985, 1838, 890, 1099, 62, 531, 2006, 1353, 682, 2103]. Spacetime [818, 1339, 3, 288, 1855, 934, 1634, 1363, 168, 1132, 732, 91]. spacetimes [1553, 566, 609, 417, 287, 2024, 1820, 1029, 2074, 169, 2287, 567, 1473]. spacings [1136, 1160]. Sparse [1968, 494]. Spatial [575, 45, 468, 1820, 2141]. spatially [1182, 2076, 877]. Special [1823, 2314, 181, 1288, 1768, 972, 2252]. species [947, 1324, 1321]. Specification [703]. specified [2180]. Specker [398]. spectra [1917, 1630, 1706, 785]. Spectral [1529, 1606, 2296, 180, 300, 2058, 2178, 423, 587, 318, 33, 796, 1364, 1047, 1589, 596, 1780, 1372, 44, 2214, 65, 872, 2088, 2132, 451, 118, 743, 1751, 2135, 1654, 2290]. spectrally [1790]. Spectrum [1216, 2017, 1233, 587, 1461, 94, 694, 1442, 1252, 1265, 1464, 2235, 1046, 1856, 2186, 2061, 1616, 221, 1832, 1711, 1308, 559, 627]. spectrum-generating [2017]. speed [391, 1757, 756, 66]. sphere [588, 157, 931, 185, 1688, 2109, 236, 108, 1994, 526]. spheres [1361, 1109].

Spherical [2241, 1601, 588, 1742, 861, 1261, 2119, 1520, 1216, 1974, 765]. spherically [116, 396, 2009, 262]. spheroidal [2098, 2195, 161]. spin [1819, 1582, 1222, 1178, 1337, 1419, 729, 276, 1860, 773, 1477, 360, 861, 233, 497, 356, 105, 393, 1579, 608, 849, 1627, 7, 2211, 1481, 1333, 1776, 2149, 2058, 250, 313, 1845, 221, 1840, 1764, 2001, 38, 2036, 2222, 2007]. spin- [233]. spin-1 [2149, 2222]. spin-1/2 [2149]. spin-12 [1819]. spin-32 [1819]. spin-boson [360, 221]. spin-orbit [608, 7, 38]. spin-weighted [861]. SpinH [2154]. spinless [852, 1754]. spinor [2023, 1978, 1175]. spinor-plane [2023]. spinors [1478, 565, 384, 2112, 755, 663, 1852, 367, 320, 1244, 60, 423]. spins [1434, 456, 1835, 1783]. spiral [1282]. Split [193, 1749, 335, 229]. split-step [1749]. splitting [2037, 567]. Spontaneous [1535, 1845, 185, 1840]. Sq6 [2225]. square [1793, 426, 1861, 14, 273, 485, 1842]. square-integrability [273]. square-well [1861]. squared [40]. squares [1717, 310]. Squashed [623, 2003, 2270]. squeezed [2100]. Squeezing [1745]. Stability [1168, 1755, 1347, 1707, 2163, 1844, 1509, 646, 1873, 713, 1783, 1397, 1475, 261, 1188, 1267, 569, 1793, 1953, 2276, 95, 1669, 422, 307, 1857, 2020, 1081, 262, 1955, 1822, 1400, 265, 877, 917, 452, 1465, 2114, 1967, 2304, 1790, 306, 1056, 1135, 491, 1091]. Stabilization [308, 2013]. stabilizer [1257, 203]. stable [304, 2055, 1790, 2226]. Stäckel [2026]. standard [2317, 1418, 1567, 967, 1813]. standing [1793, 1953, 981, 533, 2055, 1669, 422, 2114, 1397, 1731]. Star [368]. Stark [1861]. stars [2121]. state [453, 29, 2137, 74, 165, 669, 396, 481, 619, 772, 1730, 2292, 873, 697, 2212, 2097, 1410, 1847, 2054, 53, 2134, 1962, 2238, 2192, 1811, 1861, 1368, 911, 1997, 2008, 120, 1871, 1493, 313, 229, 1783, 1014, 336, 615, 1221, 1084, 1427, 1877, 658, 647, 904, 1460, 1965, 881, 965]. States [20, 145, 1499, 1313, 1918, 2060, 187, 680, 242, 907, 1419, 276, 2100, 484, 550, 1434, 730, 2183, 2309, 283, 1126, 1365, 1392, 1024, 1588, 993, 1792, 552, 848, 996, 766, 1579, 653, 1749, 2272, 722, 1816, 1686, 1177, 299, 1257, 1741, 1548, 1539, 2063, 296, 961, 549, 726, 1657, 86, 1624, 811, 1108, 1846, 1744, 514, 1546, 793, 1448, 1523, 150, 1728, 617, 1688, 1428, 1845, 154, 1452, 1840, 68, 2230, 1304, 1076, 985, 400, 1496, 657, 1922, 540, 1357, 1692, 2271, 695, 746, 774, 837, 1694, 137, 1626, 1375, 89, 31, 235, 1155, 499, 765]. states [2087, 751, 1467, 1066, 1586, 77, 1812]. Static [2075, 441, 295, 170, 288, 1109, 1189, 208]. stationarity [1705]. Stationary [798, 304, 882, 2207, 1395, 1356, 273, 400, 1839, 1673, 31, 1475, 1238]. statistical [276, 1310, 901, 1513, 1147, 2079, 213]. Statistics [244, 243, 258, 157, 1058, 869, 1308, 1349, 83]. Stavskaya [1117]. Steady [104, 1036, 468, 772, 912, 120, 1427, 1876, 1467]. steady-state [772, 120, 1427]. steep [611, 1612, 349]. steepest [1903]. steerability [1455]. steering [197]. Stein [826]. Steklov [1021]. stellar [1190]. step [1774, 412, 1749, 344, 2063, 691, 2117]. step-like [344]. Stewartson [739, 780, 610]. sticky [2004]. Stieltjes [1486, 1936]. Stirling [483, 1656]. Stochastic [1966, 1477, 620, 1942, 1931, 2033, 1347, 454, 772, 1127, 133, 2223, 976, 741, 1389, 1120, 2083, 1198, 1567, 1447, 573, 334, 1348, 1569, 1515, 2093, 2145, 794, 2224, 521, 1988, 556, 693, 2318, 1221, 1142, 2201, 408, 2106, 1028, 1056, 866, 491, 2027, 373, 2226]. Stockwell [296]. Stokes [2089, 1891, 1167, 1089, 116, 1036, 579, 791, 748, 1729, 1278, 2090, 788, 1984, 147, 2093, 2171, 2285, 524, 1082, 979, 377, 824, 1556, 263, 978, 1279, 939, 2122, 1704, 351]. stopping [334]. Stora [639]. strained [1024]. strata [1765]. strategies [997, 1376]. Stratonovich [2309, 973]. Streater [2021]. stress [116]. strictly [2053]. String [699, 368, 1077, 1331, 459, 1990]. strings [1341, 1528, 1305, 1336]. Strip [2267]. Strong [813, 1731, 866, 2048, 116, 1664, 1904, 219, 103, 865, 917, 1732, 213, 939, 1704, 2011]. Strongly [675, 1992, 22, 627]. structural [520, 1603]. Structure [1361, 1106, 30, 628, 1819, 431, 1222, 255, 239, 2067, 914, 185, 639, 1538, 1408, 509, 2154, 2205, 514, 117, 65, 1628, 1087, 1557, 702, 758, 1505, 1172, 1411, 1163, 275]. structured [414]. structures [1398, 2250, 1237, 439, 1424, 1736, 1254, 1190, 1888, 230, 115]. Study [1795, 1667, 2144]. Stueckelberg [367]. stumbled [2329]. Sturm [180, 674, 300, 1088, 1074, 2085]. Sub [1724, 186, 2082, 1932, 1395, 2150]. sub- [1395]. sub-diffusive [2082]. Sub-logistic [1724]. sub-Lorentzian [1932]. sub-quadratic [2150]. sub-supersolution [186]. Subadditivity [1836]. subalgebra [2144]. subalgebras [745, 194, 1081, 1409]. Subcritical [787]. subdomains [1974]. subentropy [246]. subfactor [2125]. subinterval [1416]. subject [914]. subjected [1759]. submajorization [960]. submanifolds [109, 292, 448, 1518, 1142, 1353]. Submodels [387]. subspace [1578]. subspaces [1261, 1143, 1921]. substitutions [1354]. subtracted [1688]. subtraction [17, 945, 1212]. succinct [857]. sufficient [402, 842, 1147, 1057, 1967]. Sugawara [988]. sum [29, 265, 1873, 2182]. summation [1051]. sums [1702, 1051, 217, 55]. sunrise [286, 628]. Super [1386, 421, 2295, 970, 1351, 2104, 2082, 1119, 140, 711, 2236]. super-activation [2104]. super-bialgebra [711]. super-diffusive [2082]. super-discrete [140]. super-KdV [2236]. Super-QRT [2236]. superadditivity [1633]. superalgebra [542, 2228, 684, 1334, 474]. superalgebras [2017, 1022, 685, 1133, 663, 434, 1449, 539, 230, 1335]. Supercharacters [1133]. supercharge [182]. superconducting [261, 1500]. superconductivity [523]. superconductors [431, 1498, 985]. superconformal [1976, 1560]. superconvergent [1572]. supercritical [1667, 648, 305]. superfluids [31]. supergravity [170, 413]. superintegrability [1502, 173, 2026]. Superintegrable [1031, 159, 1191, 173, 1456, 1009, 853, 1581, 397, 236, 1180, 702]. Superlattices [2189]. superlinear [1393, 341]. supermanifolds [776, 1044, 1238]. supermatrix [489]. Supermembrane [604]. superoperators [204]. superoscillations [1251, 1215]. superpolynomials [1061]. Superposition [18, 339, 957]. Superscar [1686]. supersolution [186, 1395]. superspace [1714, 1048, 1761, 437]. superspaces [2, 428, 563]. Supersymmetric [2315, 16, 916, 113, 590, 1015, 655, 992, 2063, 1589, 1681, 489, 932, 278]. supersymmetrical [802]. supersymmetry [102, 1566]. supertorus [690]. supervaluationism [1371]. supnorm [263]. support [107]. supported [80, 649, 1410, 549]. sure [497]. surface [1778, 1662, 1132, 2195, 1971, 710]. surfaces [983, 1561, 1443, 1318, 113, 2202, 51, 324, 1978, 1960]. surgery [1303]. Survey [522]. susceptible [646]. susceptible-infective-removed [646]. Sutcliffe [2242]. Sutherland [990, 2153, 1835, 2130]. Sutter [1420]. Suzuki [723]. SYK [1107]. SYK-like [1107]. Sylvester [1062]. symbols [98, 575]. Symmetric [580, 1719, 1644, 2030, 1779, 246, 1032, 1604, 632, 1061, 45, 769, 1582, 2269, 2257, 116, 396, 1145, 413, 2009, 1737, 288, 709, 993, 1602, 817, 262, 146, 320, 1458, 617, 430, 1246, 2318, 1295, 252, 2074, 958, 531, 190, 721, 984]. Symmetries [2199, 741, 1895, 1326, 1678, 457, 2252, 735, 413, 257, 2323, 1219, 1553, 2190, 976, 1389, 1120, 2083, 1124, 1576, 1615, 818, 464, 1359, 2322, 1229, 2313, 1355, 607, 1734, 2074, 1130, 476, 13, 19, 721, 115]. symmetrized [1102]. symmetrizing [1957]. Symmetry [386, 1555, 663, 912, 2059, 2320, 690, 125, 943, 650, 411, 1459, 1245, 1841, 340, 1393, 871, 185, 639, 739, 1955, 2210, 1809, 548, 1119, 1801, 1682, 1046, 1767, 315, 867, 1845, 1840, 452, 1764, 1273, 1536, 1131, 1286]. Symmetry-broken [943]. symplectic [1415, 444, 1048, 1701, 1418, 194, 563, 1677, 689, 1244, 1087, 1608, 423, 212]. synchronization [1514, 2051, 2142]. synchronized [877, 346]. Synchronous [465, 516, 1504, 2022]. synthesis [37, 316, 2271]. system [1417, 1023, 1323, 1902, 2269, 2164, 2261, 1315, 1710, 255, 1065, 454, 468, 1385, 1494, 1910, 645, 997, 2207, 2148, 980, 1162, 2153, 716, 2147, 1669, 739, 1984, 2238, 2163, 947, 750, 853, 352, 218, 602, 146, 1875, 332, 1299, 2130, 1891, 829, 1259, 514, 782, 786, 706, 2264, 844, 877, 2286, 1294, 2275, 1082, 467, 2061, 712, 70, 1791, 1180, 1191, 1329, 21, 301, 2053, 2092, 389, 1858, 2321, 1062, 1399, 1503, 1687, 603, 2052, 943, 1603, 485, 190, 66, 526, 1019, 1901, 2291, 1440, 1394, 1724, 866, 834, 1155, 2222, 883, 432, 1165, 308, 1324, 1089, 2011, 2176, 373, 1286, 1709, 1321, 1727]. system [610, 881, 2226]. systematic [1387]. Systems [2142, 1499, 1456, 589, 1166, 37, 1367, 1918, 2060, 720, 1280, 172, 1178, 572, 2281, 983, 74, 863, 1860, 1599, 1110, 2254, 2160, 1771, 773, 360, 496, 1443, 399, 2101, 1036, 2200, 1818, 1031, 1261, 2172, 550, 583, 227, 821, 1434, 1126, 728, 105, 188, 1726, 1156, 611, 1472, 280, 1707, 2054, 665, 1579, 865, 2203, 214, 181, 722, 1033, 136, 849, 81, 1009, 2307, 1081, 1101, 317, 1643, 1026, 1581, 464, 1359, 1481, 792, 376, 1810, 1249, 1229, 1934, 139, 1215, 1835, 1504, 321, 1624, 300, 1517, 2058, 998, 140, 1893, 1905, 250, 346, 1767, 640]. systems [858, 2242, 397, 236, 691, 1519, 2029, 2079, 762, 2088, 551, 159, 2263, 244, 243, 952, 1562, 267, 1330, 1496, 1309, 1967, 1919, 173, 1575, 34, 451, 38, 1804, 1716, 1723, 880, 1130, 1131, 1090, 972, 176, 407, 1770, 1940, 10, 809, 100, 2032, 2078, 275, 949, 819, 1066, 2297].

table [1498]. tadpoles [715]. tail [578]. tailed [875]. tails [749]. Takiff [2266]. tame [1026]. Tangent [1237]. Tangherlini [2326, 2249]. Tanner [2091]. TASEP [1381]. tau [592, 2126, 1780]. tau-functions [2126]. tau-quantization [592]. Taub [853, 1226]. Tavis [1713, 2268]. Taylor [919]. technique [1941, 2235]. techniques [1179, 517, 807, 335]. teleportation [2196]. Teller [237]. temperature [2146, 2070]. temperatures [676]. tempered [458]. Temperley [328, 329, 651]. Temporal [124, 391, 222, 682]. temporally [905]. tension [1662]. Tensor [328, 329, 1075, 1640, 1466, 1107, 2140, 1833, 1737, 730, 757, 1854, 1312, 1627, 564, 732, 504, 1814, 632, 1076, 2152, 1550, 1554, 216]. tensorial [382]. tensors [367, 899]. term [1293, 403, 382, 608, 7, 2240, 104, 2011]. terms [963, 1867, 1655, 2043, 951, 1127, 2080, 99, 1249, 38, 1195, 2319, 1138, 2036, 222]. test [132]. testing [728, 40]. tetrahedral [863]. Tetrahedron [2313]. TFDW [1313]. Thein [2091]. their [1351, 1976, 1059, 1197, 276, 1406, 1642, 1735, 1701, 901, 1570, 1741, 1081, 236, 1253, 1308, 1122, 169, 155]. theorem [398, 589, 587, 1815, 2072, 772, 1116, 992, 660, 672, 1699, 2099, 82, 1658, 1498, 1377, 1542, 478, 1684, 1546, 1303, 1290, 1608, 1055, 754, 2039, 2133, 151, 407, 75, 1974, 207, 819]. theorems [738, 826, 1039, 759, 2094, 2290, 813, 752, 1308, 79, 567, 83]. theoretic [2161]. Theoretical [862, 1122]. theories [1075, 1752, 1640, 197, 1590, 1592, 554, 413, 2206, 1677, 1331, 1225, 989, 489, 1220, 932, 1695, 862, 405, 1003, 964, 1042, 960, 460, 1824, 1928, 1511, 1262, 1378]. Theory [1158, 1365, 1887, 1862, 1405, 1267, 1772, 568, 569, 1230, 2316, 2317, 1462, 1742, 1926, 17, 945, 2184, 517, 1144, 2016, 368, 1595, 1632, 49, 1269, 1472, 1619, 1618, 1958, 409, 609, 962, 1522, 1765, 330, 331, 382, 1854, 2099, 383, 2015, 1567, 780, 129, 130, 1837, 131, 367, 320, 839, 246, 1681, 2154, 2248, 1217, 1583, 480, 604, 2084, 2186, 1916, 2109, 1535, 562, 1174, 1088, 950, 2121, 717, 2182, 244, 511, 522, 1042, 1308, 2038, 2152, 2300, 247, 613, 891, 2070, 1560, 1550, 2025, 1349, 293, 290, 2065, 1195, 2319, 668, 362, 1946, 1163]. theory [785, 1609, 698, 1148, 216, 1262, 1806]. There [930]. thermal [2013, 2288, 214, 584, 530]. thermally [1263]. thermo [2164]. thermo-diffusion [2164]. Thermodynamic [583, 1776, 1427]. thermodynamical [703]. Thermodynamics [50, 1424, 292]. thermoelasticity [1798]. thermoelectricity [676]. thermostated [325]. theta [1179, 987, 280, 1934]. thin [1306, 666, 2201, 1960]. Third [397, 57, 1207, 1737, 756, 457]. third- [457]. Third-order [397]. Thomas [1580]. Thorpe [1043, 1565]. Three [1823, 434, 1871, 1330, 1460, 1034, 1750, 1239, 664, 268, 2281, 1683, 2250, 440, 1664, 1555, 1737, 579, 1713, 2268, 53, 81, 1129, 1300, 352, 2313, 902, 1625, 811, 998, 786, 1396, 1174, 1186, 1562, 1993, 1594, 264, 443, 1243, 930, 1603, 2085, 627]. Three-body [1460, 1239, 2281, 1683, 1555, 81, 902, 1594, 443, 2085]. three-by-three [930]. three-component [264]. Three-dimensional [1034, 1664, 1737, 579, 811, 1396, 1186]. three-manifolds [1243]. three-mode [1562]. three-qubit [1300]. three-state [53]. Threshold [1285, 1392, 2175, 1137, 849, 1275]. Tietz [1540]. Tight [1301, 1338, 1086, 1063]. tight-binding [1086]. tiles [575]. tilings [1786, 1788]. tilted [2019]. Time [1999, 2254, 2101, 2100, 550, 1463, 636, 660, 299, 1169, 576, 656, 1843, 2007, 286, 248, 852, 1168, 37, 2175, 1798, 1890, 1110, 1636, 496, 777, 914, 1580, 748, 1757, 927, 709, 764, 2139, 1726, 766, 803, 132, 378, 217, 1433, 2328, 1613, 1984, 1857, 2051, 455, 2209, 1610, 464, 1359, 1568, 548, 1682, 2149, 1358, 1541, 1693, 2124, 2227, 166, 1903, 1294, 825, 1384, 1011, 571, 678, 1174, 1057, 4, 1187, 244, 243, 1343, 1732, 1078, 1079, 1993, 1986, 1414, 303, 143, 880, 1130, 1090, 407, 1733, 647, 1928, 958, 1138, 485, 1394, 1973, 616, 866, 1611, 124, 2122, 819, 1654]. time [2305]. Time-asymptotic [1169]. time-delay [1757, 927, 1857]. Time-dependent [2254, 550, 299, 576, 1168, 1580, 766, 2209, 1682, 1057, 1993, 1414, 485, 1973, 124]. Time-evolution [2100, 825]. time-fractional [1433, 2328, 2227, 1011]. time-harmonic [647]. time-like [958]. Time-ordered [660, 217]. time-reversal [709]. time-varying [852, 378]. timelike [705, 1132]. times [295, 1030, 1125, 1952, 334, 899, 2077, 1551, 1930]. timescales [2211]. Timoshenko [2164, 1284]. Toda [876, 983, 2148, 1007, 99, 1789, 115]. Toeplitz [98]. together [1074]. tolerant [722]. Tolman [859]. Tomamichel [1420]. tomography [959, 2271]. Topological [1077, 1225, 1042, 1496, 1722, 2167, 461, 568, 1592, 554, 2294, 631, 1111, 390, 1331, 146, 548, 1498, 2186, 595, 613, 1637, 2037, 100, 1080]. topologically [1437]. topology [1591, 1932, 1507, 860]. Topos [950]. tori [1134, 558, 2080, 1314, 2205, 500, 134, 1594, 938]. toroidal [796, 753, 155]. torsion [1738]. torsional [1674]. Torus [2202, 810, 249, 1447, 1679, 1741, 1119, 1328, 1355, 1138, 885, 2014, 785]. total [72, 311]. towers [1338]. toy [1298]. Trace [928, 2085, 955, 1388, 260, 1420, 1711, 1375]. tracer [1706]. traffic [1796]. trajectories [570, 1697, 1140, 1823, 1071, 211, 240, 447]. transcendent [691]. Transcendental [383]. transcendents [1385]. transfer [272]. transform [1390, 2160, 2034, 78, 2203, 344, 296, 830, 419, 1486, 1936, 1487]. Transformation [2026, 63, 1184, 9, 89, 908, 1016]. transformations [1098, 420, 863, 165, 371, 755, 912, 1386, 2142, 890, 1085, 2155, 482, 834, 1286]. transforms [541, 340, 2217, 1650, 911, 215, 1014, 1253]. Transition [1095, 597, 1508, 514, 1294, 1340, 1860, 733, 2051, 2162]. transitions [2041, 456, 1881, 2019]. transitive [701]. Translation [229]. translationally [1337, 2156]. translations [1001]. Transmission [1060, 312, 1302, 2104, 1964, 948, 2312, 1526, 910]. transmutations [681]. transparent [2216]. Transport [2209, 24, 1150, 1128, 771, 1011]. transportation [1922]. transpose [930]. transversal [1312]. transverse [177, 1115]. trapped [109, 2284, 1141, 719, 51, 1135]. Traveling [1483, 913, 1205, 1242, 1402, 646, 1789, 1091]. Travelling [806, 429]. Travelling-wave [806]. treatment [1759, 857]. tree [1346, 596]. tree-shifts [1346]. trees [167, 1860]. tri [507, 52]. tri-integrable [52]. tri-partite [507]. triangular [141, 348]. triangulations [1383]. triconfluent [202]. tridiagonal [1233, 1183, 1430]. trigonometric [237, 2157]. triple [545, 332]. triples [318, 1606, 2009, 1047, 1654]. trivial [384, 1257]. tropical [530]. Trotter [1366, 723]. Trudi [1779]. truncated [2063]. Tsallis [1058, 1961]. Tunneling [661, 2061, 616]. Turaev [562, 1303]. Turaev-Virelizier [1303]. turbulence [1833, 1705, 946]. turbulent [1703, 788, 1706]. twist [2001]. Twisted [1862, 714, 1735, 652, 870, 1339, 384, 1926, 888, 992, 182, 846]. twisting [1553]. twistor [1344, 1055]. twistors [1852, 1344]. Two [412, 1145, 49, 28, 1129, 528, 846, 593, 55, 2243, 964, 2038, 278, 644, 1805, 2162, 286, 628, 650, 37, 1703, 2281, 326, 228, 2250, 426, 2160, 1531, 1878, 1833, 542, 575, 2045, 470, 1553, 676, 133, 188, 1696, 1707, 2327, 2134, 1137, 1699, 2080, 1009, 1169, 602, 1828, 1184, 1007, 1600, 1141, 1835, 1206, 1625, 225, 659, 1072, 1501, 166, 2024, 1294, 2178, 14, 467, 68, 273, 355, 1074, 2004, 712, 1536, 443, 1373, 1414, 656, 1357, 46, 658, 1488, 1908, 1019, 1028, 240, 866, 1049, 795, 949, 1324, 1321, 77]. two-body [2281, 133, 1373]. Two-component [644, 326, 1019, 1049]. Two-dimensional [55, 2243, 964, 2038, 1805, 2162, 650, 470, 2134, 2080, 1009, 602, 1184, 1007, 225, 1072, 712, 656, 1357]. two-fluid [1707]. two-interval [1074]. two-level [2160]. Two-loop [49, 528, 286, 628]. two-manifolds [659]. two-matrix [2178]. two-mode [443, 949]. Two-parameter [28, 846, 278, 1828, 658]. two-phase [1703, 228, 795]. two-point [1833]. two-species [1324, 1321]. Two-step [412]. two-time-scale [866]. type [1867, 1655, 2043, 2239, 73, 11, 1720, 538, 1913, 42, 1676, 27, 1590, 1953, 1271, 592, 954, 1745, 1065, 1559, 2009, 935, 2323, 1870, 1444, 1702, 2292, 2131, 2035, 1641, 1553, 340, 1463, 1156, 1968, 415, 1537, 799, 2055, 1677, 1765, 28, 1803, 1576, 1615, 1975, 1844, 2196, 2322, 2129, 1657, 321, 1328, 1693, 921, 1648, 535, 1170, 753, 700, 1604, 607, 1332, 305, 1084, 2237, 34, 278, 354, 13, 19, 978, 207, 1876, 474, 2014, 491, 2150, 343, 1066]. type-I [42]. types [1835, 886, 2143, 1977]. Typical [2272]. Tyutin [639].

U [2232]. Uhlenbeck [1267, 1266]. ultra [295]. ultra-static [295]. ultradiscrete [1828]. Ultrarelativistic [765]. Ultraviolet [1423]. unavoidable [1747]. unbiased [20, 145, 622]. unbounded [587, 1506, 1890, 1434, 456, 347, 250, 380, 2187, 2201, 603]. unboundedness [1058]. uncertainties [851, 2064]. Uncertainty [1097, 1510, 624, 1964, 1534, 1754, 2166, 299, 1309, 654, 809, 2191, 1708]. Unchanged [447]. Unconventional [17, 945]. Unentangled [1629]. unification [1379]. Unified [275, 1512, 163, 79, 1016, 178]. Uniform [2051, 175, 1454, 475, 1251, 1939, 2324, 2264, 154, 1301, 1019, 1491]. Uniform-in-time [2051]. uniformly [1112, 701]. unique [1362]. Uniqueness [1538, 369, 146, 1728, 2112, 1897, 1904, 791, 1434, 1392, 2145, 467, 946, 1646, 491, 939]. uniserial [889]. unit [1757, 1114, 1603, 526]. unit-speed [1757]. unitaries [316, 1105]. Unitary [1105, 2271, 2305, 2127, 1012, 1605, 2220, 258, 1850, 1975, 1289, 726, 967, 474, 1826]. unity [651, 543]. Universal [128, 81, 2070, 1684, 754, 856, 1100]. Universality [2251, 252, 494]. universe [2076, 2046]. universes [935]. unstable [290].

Unsteady [1082, 1556]. upon [2329]. UPPE [2216]. Upper [1305, 1569, 1515, 697, 1228, 1104, 1043, 1565, 2027]. use [1406, 960, 64]. user [1385]. Using [1599, 1233, 1183, 1941, 1945, 633, 2327, 217, 1388, 1450, 1811, 2196, 2216, 1600, 1458, 813, 1916, 1088, 452, 1326, 710, 1978, 804]. Uvarov [2327, 256, 1600].

V [487, 1977, 1826]. Vacuum [1275, 1641, 1729, 1189, 979, 1820, 2048]. vacuums [1907]. Validity [1671]. value [398, 2288, 1530, 385, 637, 1017, 750, 2170, 1831, 1614, 2158, 559, 1715, 2222, 71]. valued [1145, 509, 1979, 402, 842, 2151, 1082, 964, 463]. values [1774, 637, 1017, 1539, 615, 1486, 1936]. Vanishing [531, 1279, 453, 1949, 226, 191, 2240, 920]. Variable [739, 1117, 1203, 1401, 806, 532, 1730, 2292, 1726, 1576, 1615, 2312, 317, 127, 1834, 2304, 2118, 1900, 1276]. variable-coefficient [532]. variable-exponent [2304]. Variable-length [1117]. Variables [1502, 1690, 259, 1126, 1773, 1482, 2059, 2320, 1828, 107, 2205, 771, 1785, 1343, 2001]. Variance [828]. variant [398]. variants [1174]. variation [183]. Variational [1518, 983, 2206, 385, 529, 2218, 140, 315, 1758, 1951, 184, 526]. variations [565, 23, 536]. variety [1597]. Various [507, 1539]. varying [852, 1182, 1302, 1782, 1112, 378, 1864, 222]. Vector [1364, 2045, 2151, 588, 733, 1418, 901, 564, 659, 346]. Vector-valued [2151]. vectors [1525, 1236, 1033, 1495]. Velocity [287, 1833, 1985, 351]. Verma [155]. version [835]. versus [420, 925, 1270, 1545]. vertex [2177, 106, 1477, 1364, 1735, 873, 1071, 111, 28, 434, 1799, 1807, 2265, 1597, 753, 1959, 1246, 1604, 1990, 247, 2301]. vesicles [2004]. vessel [1088]. Vessiot [745]. VI [1006]. via [453, 814, 705, 228, 1659, 328, 2159, 2073, 826, 1151, 875, 976, 1619, 1618, 1958, 404, 1886, 1141, 1060, 1507, 1903, 2084, 1186, 2143, 838, 725, 38, 1894, 553, 1715, 1725, 2297]. view [570]. Vilkovisky [416, 1927]. violation [1104, 897]. violations [1588]. Virasoro [1173, 268, 888, 2131, 614, 439, 799, 739, 1970, 472, 651, 1894, 1067, 711]. Virelizier [1303]. Viro [562]. virtual [2281]. viscoelastic [532, 922, 794]. viscosities [1984]. viscosity [226, 919, 2285, 1732, 2053, 1279]. viscous [882, 662, 342, 822, 780, 262, 524, 795, 2048]. Vishik [1388]. visualized [1151]. Vlasov [304, 633, 468, 1580, 748, 947, 786, 668, 432]. Vlasov/compressible [748]. VOA [2157]. void [1358]. voids [3, 4]. Volkoff [859]. Volterra [178]. volume [996, 1762]. Vortex [431, 353, 1940, 2321, 1066, 1010, 2221, 146, 982, 536]. vortices [664, 1797, 1157, 390, 1898, 324]. vorticity [662]. Vries [1774, 1323, 469, 492, 2050, 1311, 1088, 1211, 1062, 1276]. vs [2175, 2309, 1058].

W [89]. Waals [769, 2107, 225]. walk [1116, 1749, 1011]. Walker [678, 1125, 287, 571, 899, 2077, 1551, 1353, 1473]. walks [1920, 1859]. wall [106, 2159]. Wallis [442]. walls [661, 2004]. Walsh [1480]. Walze [207].

Wang [575]. Wannier [1991]. warped [498, 1439]. Wasserstein [1649]. water [1825, 798, 63, 9, 1892, 2308]. water-waves [798]. waterbag [796]. Wave [485, 431, 1183, 345, 1398, 1483, 736, 198, 806, 532, 1470, 314, 1210, 1849, 913, 429, 352, 529, 832, 2076, 22, 1272, 2311, 1957, 794, 265, 946, 917, 125, 1253, 70, 305, 1397, 1242, 851, 2064, 2304, 2001, 1274, 2053, 1716, 143, 1551, 1138, 66, 1718, 1181, 1611, 677, 1091]. wave-breaking [345]. Wavefront [1044]. wavefronts [110]. Wavefunction [618]. wavefunctions [1679, 512, 1604]. waveguide [713]. waveguides [807, 68, 1790, 1960]. wavelets [1213]. Wavepackets [1020, 1743]. waves [1203, 1825, 1793, 1953, 2213, 981, 533, 784, 2055, 798, 742, 63, 1669, 422, 1282, 2163, 1209, 1205, 9, 1892, 225, 646, 265, 877, 524, 1204, 2114, 1242, 358, 1731, 1342, 1712, 1135, 222, 1402, 64]. way [1070, 1254, 1997, 1226]. Weak [1904, 227, 2079, 1074, 70, 305, 939, 261, 532, 219, 791, 1837, 2124, 2151, 1396, 535, 2285, 267, 615, 67, 1843, 1167, 785, 2215]. weak- [535]. Weak-strong [1904, 939]. weakly [1352, 1086, 1410, 492, 618, 549, 1349, 372]. Weber [2034]. webs [1166, 2006]. Wegner [760, 749]. Weierstrass [2202, 1679]. Weight [840, 1739, 986, 888, 2131, 152, 2014]. Weight-lattice [840]. Weighted [1722, 1036, 2283, 861, 1235, 175]. Weil [1403]. welcome [1868]. Well [2174, 792, 2224, 917, 121, 2175, 187, 2091, 1910, 662, 611, 1492, 2090, 594, 1984, 2012, 1861, 1662, 1501, 793, 1287, 2171, 1791, 485, 1165, 499, 765, 351, 1989]. Well-posedness [2174, 792, 2224, 121, 2175, 2091, 1910, 662, 1492, 2090, 1984, 2012, 1662, 1287, 2171, 1791, 1165, 351, 1989]. wells [1612, 349]. Wentzel [1438, 85, 718]. Wess [403]. Weyl [2177, 1012, 1005, 1640, 33, 481, 78, 327, 1847, 639, 672, 1234, 840, 1481, 810, 899, 357, 686, 1449, 1014, 91, 436, 2039, 1295, 1085, 1527, 1554]. Weyl-diffeomorphism [639]. Weyl-geometric [1640]. Weyl-like [1554]. Weyl-orbit [840]. where [580, 1416]. which [667]. while [2329]. whiskered [1594]. white [2209]. white-noise [2209]. Whitham [117]. Whittaker [874]. whole [2120]. Wick [825]. widely [1634]. Widom [2282]. width [2308]. Wigner [541, 1151, 477, 1418, 1968, 2062, 1454, 810, 1253, 1295, 151, 718]. Wigner-type [1968]. Wilson [1023, 2179]. winding [631]. Winfree [1857]. winning [1376]. Winter [238]. wire [1651]. wires [552]. wiretap [1302]. within [869]. without [279, 2091, 314, 803, 855, 2124, 2145, 1572, 530, 2291, 1291]. Witnessing [2299]. Witt [1830]. Witten [1897, 1818, 403, 2154, 1894]. WKB [467]. Wonderland [738]. Wong [1380]. Woods [617]. World [85, 1438, 461, 1026]. World-line [85, 1438, 461]. wormholes [2003]. wrong [2041]. Wronskian [742]. Wu [956].

X [1980, 1981]. X-point [1980, 1981]. XOR [1376]. XXZ [1419, 1579, 1776, 2149, 1604]. XXZ-type [1604]. XY [475].

Yamabe [392, 1695]. Yamaguti [25]. Yang [568, 569, 1592, 731, 1887, 1590, 297, 1269, 962, 1836, 26, 604, 1025, 613, 708, 2287, 1609]. Yangian [332]. Yangians [652]. YdKN [1124]. Ye [2140]. Yetter [156]. yielding [1548]. Young [1094, 1921, 240]. Yukawa [389, 1806]. Yvon [176].

Z2 [1976]. Zakai [1380]. Zakharov [255, 386, 1297, 300, 347]. Zeno [2066]. Zernike [1315, 1710, 2241, 2219, 1180, 1191]. Zero [255, 1663, 200, 914, 662, 103, 1101, 168, 855, 2155, 169, 235, 1080, 1987, 2273, 2162]. zero-dimensional [1080]. zero-error [855, 2273]. zero-gap [235]. zero-gravity [168, 169]. Zeros [1698, 1429, 450, 915, 836, 2200, 1121, 164]. Zeta [896, 498, 1439, 613]. zeta-determinants [498, 1439]. Zimmermann [945, 17]. zippers [94]. ZND [2286]. Zumino [403]. Zwanzig [1227, 1827]. Zybin [2052].

References

Chandrasekar:2007:LHD

[1] V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. On the Lagrangian and Hamiltonian description of the damped linear harmonic oscillator. Journal of Mathematical Physics, 48(3):032701, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/
p032701_s1.

Fioresi:2007:MCS

[2] R. Fioresi, M. A. Lledó, and V. S. Varadarajan. The Minkowski and conformal superspaces. Journal of Mathematical Physics, 48(11):113505, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113505_s1. See erratum [428].

Castro:2008:EGA

[3] Carlos Castro. The Euclidean gravitational action as black hole entropy, singularities, and spacetime voids. Journal of Mathematical Physics, 49(4):042501, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p042501_s1. See comments [4].

Mitra:2009:CEG

[4] Abhas Mitra. Comments on “The Euclidean gravitational action as black hole entropy, singularities, and space–time voids”, [J. Math. Phys. 49, 042501 (2008)]. Journal of Mathematical Physics, 50(4):042502, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042502_s1. See [3, 1358].

Kim:2011:LCS

[5] Do-Hyung Kim. Lorentzian covering space and homotopy classes. Journal of Mathematical Physics, 52(12):122501, December 2011. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i12/

p122501_s1. See erratum [427].

Wang:2008:GNQ

[6] Hai-Jun Wang. General nonlocality in quantum fields. Journal of Mathematical Physics, 49(3):033513, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i3/p033513_s1. See comment [8]

and erratum [119].

Ikhdair:2010:ASD

[7] Sameer M. Ikhdair. Approximate solutions of the Dirac equation for the Rosen–Morse potential including the spin-orbit centrifugal term. Journal of Mathematical Physics, 51(2):023525, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023525_s1. See

comment [608].

Wang:2010:CGN

[8] Hai-Jun Wang. Comment on “General nonlocality in quantum fields”, [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 51(5): 054102, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p054102_s1. See [6].

Jiang:2010:SBT

[9] Yan Jiang, Bo Tian, Wen-Jun Liu, Min Li, Pan Wang, and Kun Sun. Solitons, Bäcklund transformation, and Lax pair for the (2 + 1)-dimensional Boiti–Leon–Pempinelli equation for the water waves. Journal of Mathematical Physics, 51(9):093519, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i9/p093519_s1. See [63].

Sun:2012:FPE

[10] Xu Sun and Jinqiao Duan. Fokker–Planck equations for nonlinear dynamical systems driven by non-Gaussian Lévy processes. Journal of Mathematical Physics, 53(7):072701, July 2012. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p072701_s1. See comment

[640].

Arrizabalaga:2013:SAE

[11] Naiara Arrizabalaga, Javier Duoandikoetxea, and Luis Vega. Self-adjoint extensions of Dirac operators with Coulomb type singularity. Journal of Mathematical Physics, 54(4):041504, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p041504_s1. See erratum

[1720].

Lopez-Rosa:2013:ECA

[12] S. López-Rosa, I. V. Toranzo, P. Sánchez-Moreno, and J. S. Dehesa. Entropy and complexity analysis of hydrogenic Rydberg atoms. Journal of Mathematical Physics, 54(5):052109, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1319, 1320].

Tiwari:2013:CLP

[13] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0. Journal of Mathematical Physics, 54(5): 053506, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [19], comments [607, 2322], and response [2323].

Martinez-y-Romero:2013:TDM

[14] R. P. Martínez y Romero, H. N. Núñez-Yépez, and A. L. Salas-Brito. The two dimensional motion of a particle in an inverse square potential: Classical and quantum aspects. Journal of Mathematical Physics, 54(5): 053509, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [426].

Waclawczyk:2013:AEL

[15] Marta Waclawczyk and Martin Oberlack. Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation. Journal of Mathematical Physics, 54(7):072901, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [641] and response [642].

Chuah:2013:CSP

[16] Meng-Kiat Chuah. Contragredient supersymmetric pairs. Journal of Mathematical Physics, 54(9):091702, September 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [916].

Carvalho:2013:UMS

[17] Paulo R. S. Carvalho and Marcelo M. Leite. Unconventional minimal subtraction and Bogoliubov–Parasyuk–Hepp–Zimmermann method: Massive scalar theory and critical exponents. Journal of Mathematical Physics, 54(9):093301, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [945].

Khare:2014:SEF

[18] Avinash Khare and Avadh Saxena. Superposition of elliptic functions as solutions for a large number of nonlinear equations. Journal of Mathematical Physics, 55(3):032701, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [339].

Tiwari:2014:ECL

[19] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Erratum: “Classification of Lie point symmetries for quadratic Liénard type equation ẍ+f(x)ẋ^{2}+g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 55(5):059901, May 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13, 607].

Amburg:2014:SLS

[20] Ilya Amburg, Roshan Sharma, Daniel M. Sussman, and William K. Wootters. States that “look the same” with respect to every basis in a mutually unbiased set. Journal of Mathematical Physics, 55(12):122206, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [145].

Sakabekov:2014:BCO

[21] Auzhan Sakabekov and Yerkanat Auzhani. Boundary conditions for the one-dimensional nonlinear nonstationary Boltzmann’s moment system equations. Journal of Mathematical Physics, 55(12):123507, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [301].

Kalantarov:2015:NSD

[22] Varga Kalantarov and Sergey Zelik. A note on a strongly damped wave equation with fast growing nonlinearities. Journal of Mathematical Physics, 56(1):011501, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lena:2015:EVA

[23] Corentin Léna. Eigenvalues variations for Aharonov–Bohm operators. Journal of Mathematical Physics, 56(1):011502, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Broadbridge:2015:SRR

[24] P. Broadbridge and N. M. Ivanova. Solutions and reductions for radiative energy transport in laser-heated plasma. Journal of Mathematical Physics, 56(1):011503, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2015:OPF

[25] Yao Ma, Liangyun Chen, and Jie Lin. One-parameter formal deformations of Hom–Lie–Yamaguti algebras. Journal of Mathematical Physics, 56(1):011701, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herscovich:2015:SRR

[26] Estanislao Herscovich. Some remarks on representations of Yang–Mills algebras. Journal of Mathematical Physics, 56(1):011702, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barker:2015:NPF

[27] Amy Barker, David Swinarski, Lauren Vogelstein, and John Wu. A new proof of a formula for the type A_{2} fusion rules. Journal of Mathematical Physics, 56(1):011703, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2015:TPQ

[28] Yun Gao, Naihong Hu, and Honglian Zhang. Two-parameter quantum affine algebra of type G_{2}^{(1)}, Drinfeld realization and vertex representation. Journal of Mathematical Physics, 56(1):011704, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baratin:2015:CSS

[29] Aristide Baratin and Laurent Freidel. A 2-categorical state sum model. Journal of Mathematical Physics, 56(1):011705, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suh:2015:SCA

[30] Uhi Rinn Suh. Structure of classical affine and classical affine fractional W -algebras. Journal of Mathematical Physics, 56(1):011706, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wreszinski:2015:LSN

[31] Walter F. Wreszinski. Landau superfluids as nonequilibrium stationary states. Journal of Mathematical Physics, 56(1):011901, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stetsko:2015:DON

[32] M. M. Stetsko. Dirac oscillator and nonrelativistic Snyder–de Sitter algebra. Journal of Mathematical Physics, 56(1):012101, January 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigl:2015:SDS

[33] Alexander Beigl, Jonathan Eckhardt, Aleksey Kostenko, and Gerald Teschl. On spectral deformations and singular Weyl functions for onedimensional Dirac operators. Journal of Mathematical Physics, 56(1): 012102, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruby:2015:ROA

[34] V. Chithiika Ruby, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. Removal of ordering ambiguity for a class of position dependent mass quantum systems with an application to the quadratic Liénard type nonlinear oscillators. Journal of Mathematical Physics, 56(1):012103, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Franz:2015:SCI

[35] Jan Franz. Solution of coupled integral equations for quantum scattering in the presence of complex potentials. Journal of Mathematical Physics, 56(1):012104, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagchi:2015:QLO

[36] B. Bagchi, A. Ghose Choudhury, and Partha Guha. On quantized Liénard oscillator and momentum dependent mass. Journal of Mathematical Physics, 56(1):012105, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albertini:2015:MTO

[37] Francesca Albertini and Domenico D’Alessandro. Minimum time optimal synthesis for two level quantum systems. Journal of Mathematical Physics, 56(1):012106, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:2015:QSP

[38] Alexandre G. M. Schmidt, L. Portugal, and Anderson L. de Jesus. Quantum systems with position-dependent mass and spin-orbit interaction via Rashba and Dresselhaus terms. Journal of Mathematical Physics, 56(1): 012107, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steeb:2015:EMN

[39] Willi-Hans Steeb and Yorick Hardy. Exponential of a matrix, a nonlinear problem, and quantum gates. Journal of Mathematical Physics, 56(1): 012201, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Temme:2015:QCS

[40] Kristan Temme and Frank Verstraete. Quantum chi-squared and goodness of fit testing. Journal of Mathematical Physics, 56(1):012202, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tolksdorf:2015:EHA

[41] Jürgen Tolksdorf. The Einstein–Hilbert action with cosmological constant as a functional of generic form. Journal of Mathematical Physics, 56(1):012301, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ayissi:2015:BTM

[42] Raoul Domingo Ayissi and Norbert Noutchegueme. Bianchi type-I magnetized cosmological models for the Einstein–Boltzmann equation with the cosmological constant. Journal of Mathematical Physics, 56(1):

012501, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alho:2015:GDI

[43] Artur Alho and Claes Uggla. Global dynamics and inflationary center manifold and slow-roll approximants. Journal of Mathematical Physics, 56(1):012502, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:HCS

[44] Wencai Liu and Xiaoping Yuan. Hölder continuity of the spectral measures for one-dimensional Schrödinger operator in exponential regime. Journal of Mathematical Physics, 56(1):012701, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alberti:2015:NFS

[45] Angelo Alberti and Claudio Vidal. New families of symmetric periodic solutions of the spatial anisotropic Manev problem. Journal of Mathematical Physics, 56(1):012901, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seri:2015:PTF

[46] Marcello Seri. The problem of two fixed centers: bifurcation diagram for positive energies. Journal of Mathematical Physics, 56(1):012902, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:2015:GNO

[47] C. Quesne. Generalized nonlinear oscillators with quasi-harmonic behaviour: Classical solutions. Journal of Mathematical Physics, 56(1): 012903, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walters:2015:NMM

[48] Meg Walters and Shannon Starr. A note on mixed matrix moments for the complex Ginibre ensemble. Journal of Mathematical Physics, 56 (1):013301, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dudka:2015:TLF

[49] M. Dudka. Two-loop Feynman integrals for φ^{4} theory with long-range correlated disorder. Journal of Mathematical Physics, 56(1):013302, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pirjol:2015:TLG

[50] Dan Pirjol and Carlos Schat. Thermodynamics of a lattice gas with linear attractive potential. Journal of Mathematical Physics, 56(1):013303, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:2015:RCF

[51] Bennett Palmer. Remarks on a convergent flow for marginally trapped surfaces in Lorentz–Minkowski space. Journal of Mathematical Physics, 56(1):013501, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2015:BIC

[52] Baiying He, Liangyun Chen, and Yan Cao. Bi-integrable couplings and tri-integrable couplings of the modified Ablowitz–Kaup–Newell–Segur hierarchy with self-consistent sources. Journal of Mathematical Physics, 56 (1):013502, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fonseca:2015:MTS

[53] T. Fonseca, L. Frappat, and E. Ragoucy. R matrices of three-state Hamiltonians solvable by coordinate Bethe ansatz. Journal of Mathematical Physics, 56(1):013503, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fitzsimmons:2015:HFB

[54] Patrick J. Fitzsimmons and Kristin E. Kuter. Harmonic functions of Brownian motions on metric graphs. Journal of Mathematical Physics, 56(1):013504, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Linton:2015:TDP

[55] C. M. Linton. Two-dimensional, phase modulated lattice sums with application to the Helmholtz Green’s function. Journal of Mathematical Physics, 56(1):013505, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2015:COF

[56] John E. Gough. Characteristic operator functions for quantum inputplant-output models and coherent control. Journal of Mathematical Physics, 56(1):013506, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagderina:2015:SEP

[57] Yulia Yu. Bagderina and Nikolai N. Tarkhanov. Solution of the equivalence problem for the third Painlevé equation. Journal of Mathematical Physics, 56(1):013507, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rumanov:2015:CIB

[58] Igor Rumanov. Classical integrability for beta-ensembles and general Fokker–Planck equations. Journal of Mathematical Physics, 56(1): 013508, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akant:2015:BEC

[59] Levent Akant, Emine Ertugrul, Ferzan Tapramaz, and O. Teoman Turgut. Bose–Einstein condensation on a manifold with non-negative Ricci curvature. Journal of Mathematical Physics, 56(1):013509, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lischewski:2015:CCK

[60] Andree Lischewski. Charged conformal Killing spinors. Journal of Mathematical Physics, 56(1):013510, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakos:2015:LNP

[61] M. Dimakos and A. S. Fokas. Linearisable nonlinear partial differential equations in multidimensions. Journal of Mathematical Physics, 56(1): 013511, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reimberg:2015:RFT

[62] Paulo H. F. Reimberg and L. Raul Abramo. Random flights through spaces of different dimensions. Journal of Mathematical Physics, 56(1): 013512, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2015:CSB

[63] Xin-Yi Gao. Comment on “Solitons, Bäcklund transformation, and Lax pair for the (2 + 1)-dimensional Boiti–Leon–Pempinelli equation for the water waves” [J. Math. Phys. 51, 093519 (2010)]. Journal of Mathematical Physics, 56(1):014101, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [9].

deSuzzoni:2015:UNF

[64] Anne-Sophie de Suzzoni. On the use of normal forms in the propagation of random waves. Journal of Mathematical Physics, 56(2):021501, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manukure:2015:SHA

[65] Solomon Manukure and Wen-Xiu Ma. A soliton hierarchy associated with a new spectral problem and its Hamiltonian structure. Journal of Mathematical Physics, 56(2):021502, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2015:MSS

[66] Zhong Wang and Shangbin Cui. Multi-speed solitary wave solutions for a coherently coupled nonlinear Schrödinger system. Journal of Mathematical Physics, 56(2):021503, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tiglay:2015:CWS

[67] Feride Tiglay. Conservative weak solutions of the periodic Cauchy problem for μHS equation. Journal of Mathematical Physics, 56(2):021504, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nazarov:2015:BSW

[68] S. A. Nazarov, K. Ruotsalainen, and P. Uusitalo. Bound states of waveguides with two right-angled bends. Journal of Mathematical Physics, 56 (2):021505, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Camus:2015:HDN

[69] Brice Camus and Nils Rautenberg. Higher dimensional nonclassical eigenvalue asymptotics. Journal of Mathematical Physics, 56(2):021506, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pastor:2015:WCW

[70] Ademir Pastor. Weak concentration and wave operator for a 3D coupled nonlinear Schrödinger system. Journal of Mathematical Physics, 56(2):

021507, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2015:HBV

[71] Jie Zhao. Homogenization of the boundary value for the Neumann problem. Journal of Mathematical Physics, 56(2):021508, February 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2015:RHH

[72] Shuangjian Guo, Xiaohui Zhang, and Shengxiang Wang. Relative Hom– Hopf modules and total integrals. Journal of Mathematical Physics, 56 (2):021701, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Angiono:2015:MQD

[73] Iván Angiono and Hiroyuki Yamane. The R-matrix of quantum doubles of Nichols algebras of diagonal type. Journal of Mathematical Physics, 56 (2):021702, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benguria:2015:GSE

[74] Rafael D. Benguria, Rupert L. Frank, and Elliott H. Lieb. Ground state energy of large polaron systems. Journal of Mathematical Physics, 56 (2):021901, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takayanagi:2015:ISP

[75] Kazuo Takayanagi and Mariko Oishi. Inverse scattering problem and generalized optical theorem. Journal of Mathematical Physics, 56(2): 022101, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guneysu:2015:SLQ

[76] Batu Güneysu. Semiclassical limits of quantum partition functions on infinite graphs. Journal of Mathematical Physics, 56(2):022102, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dePrunele:2015:BST

[77] E. de Prunelé. Bound states for two dimensional Schrödinger equation with anisotropic interactions localized on a circle. Journal of Mathematical Physics, 56(2):022103, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Domingo:2015:GWT

[78] Herbert B. Domingo and Eric A. Galapon. Generalized Weyl transform for operator ordering: Polynomial functions in phase space. Journal of Mathematical Physics, 56(2):022104, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wallace:2015:RTU

[79] David Wallace. Recurrence theorems: A unified account. Journal of Mathematical Physics, 56(2):022105, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2015:IPS

[80] Tuncay Aktosun, Paul Sacks, and Mehmet Unlu. Inverse problems for selfadjoint Schrödinger operators on the half line with compactly supported potentials. Journal of Mathematical Physics, 56(2):022106, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gridnev:2015:ULE

[81] Dmitry K. Gridnev. Universal low-energy behavior in three-body systems. Journal of Mathematical Physics, 56(2):022107, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2015:NTD

[82] John E. Gough, Tudor S. Ratiu, and Oleg G. Smolyanov. Noether’s theorem for dissipative quantum dynamical semi-groups. Journal of Mathematical Physics, 56(2):022108, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanHorssen:2015:SCL

[83] Merlijn van Horssen and Madalin Guta. Sanov and central limit theorems for output statistics of quantum Markov chains. Journal of Mathematical Physics, 56(2):022109, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2015:MSD

[84] Jijiang Sun and Shiwang Ma. Multiple solutions for discrete periodic nonlinear Schrödinger equations. Journal of Mathematical Physics, 56 (2):022110, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordon:2015:WLI

[85] James Gordon and Gordon W. Semenoff. World-line instantons and the Schwinger effect as a Wentzel–Kramers–Brillouin exact path integral. Journal of Mathematical Physics, 56(2):022111, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1438].

Koenig:2015:CEP

[86] Robert Koenig. The conditional entropy power inequality for Gaussian quantum states. Journal of Mathematical Physics, 56(2):022201, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Audenaert:2015:RRE

[87] Koenraad M. R. Audenaert and Nilanjana Datta. α − z-Rényi relative entropies. Journal of Mathematical Physics, 56(2):022202, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dupuis:2015:CRQ

[88] Frédéric Dupuis. Chain rules for quantum Rényi entropies. Journal of Mathematical Physics, 56(2):022203, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vrana:2015:AET

[89] Péter Vrana and Matthias Christandl. Asymptotic entanglement transformation between w and GHZ states. Journal of Mathematical Physics, 56(2):022204, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berta:2015:RGC

[90] Mario Berta, Kaushik P. Seshadreesan, and Mark M. Wilde. Rényi generalizations of the conditional quantum mutual information. Journal of Mathematical Physics, 56(2):022205, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Much:2015:RCM

[91] A. Much. Relativistic corrections to the Moyal–Weyl spacetime. Journal of Mathematical Physics, 56(2):022301, February 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kemp:2015:RSP

[92] Garreth Kemp. SO(N) restricted Schur polynomials. Journal of Mathematical Physics, 56(2):022302, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pinamonti:2015:SIC

[93] Nicola Pinamonti and Daniel Siemssen. Scale-invariant curvature fluctuations from an extended semiclassical gravity. Journal of Mathematical Physics, 56(2):022303, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boumaza:2015:AAC

[94] Hakim Boumaza and Laurent Marin. Absence of absolutely continuous spectrum for random scattering zippers. Journal of Mathematical Physics, 56(2):022701, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dung:2015:ESL

[95] Nguyen Tien Dung. On exponential stability of linear Levin–Nohel integro-differential equations. Journal of Mathematical Physics, 56(2): 022702, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2015:QPS

[96] Yanling Shi, Junxiang Xu, and Xindong Xu. On quasi-periodic solutions for generalized Boussinesq equation with quadratic nonlinearity. Journal of Mathematical Physics, 56(2):022703, February 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2015:LEK

[97] Zhuchun Li, Xiaoping Xue, and Daren Yu. On the Lojasiewicz exponent of Kuramoto model. Journal of Mathematical Physics, 56(2):022704, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Charlier:2015:ATD

[98] Christophe Charlier and Tom Claeys. Asymptotics for Toeplitz determinants: Perturbation of symbols with a gap. Journal of Mathematical Physics, 56(2):022705, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanki:2015:ICT

[99] Masataka Kanki, Jun Mada, and Tetsuji Tokihiro. Integrability criterion in terms of coprime property for the discrete Toda equation. Journal of Mathematical Physics, 56(2):022706, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xing:2015:ITP

[100] Zhitao Xing and Ercai Chen. Induced topological pressure for topological dynamical systems. Journal of Mathematical Physics, 56(2):022707, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lechner:2015:EMC

[101] Kurt Lechner. Electrodynamics of massless charged particles. Journal of Mathematical Physics, 56(2):022901, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masterov:2015:HDM

[102] Ivan Masterov. Higher-derivative mechanics with N = 2l-conformal Galilei supersymmetry. Journal of Mathematical Physics, 56(2):022902, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2015:GSS

[103] Jishan Fan and Yuxi Hu. Global strong solutions to the 1-D compressible magnetohydrodynamic equations with zero resistivity. Journal of Mathematical Physics, 56(2):023101, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Slobodeanu:2015:SEF

[104] Radu Slobodeanu. Steady Euler flows and the Faddeev–Skyrme model with mass term. Journal of Mathematical Physics, 56(2):023102, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeRoeck:2015:LNC

[105] W. De Roeck, C. Maes, K. Netocný, and M. Schütz. Locality and nonlocality of classical restrictions of quantum spin systems with applications to quantum large deviations and entanglement. Journal of Mathematical Physics, 56(2):023301, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bleher:2015:SVM

[106] Pavel Bleher and Karl Liechty. Six-vertex model with partial domain wall boundary conditions: Ferroelectric phase. Journal of Mathematical Physics, 56(2):023302, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jauregui:2015:CPL

[107] Max Jauregui and Constantino Tsallis. Convergence of the probability of large deviations in a model of correlated random variables having compact-support Q-Gaussians as limiting distributions. Journal of Mathematical Physics, 56(2):023303, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miller:2015:QES

[108] Willard Miller, Jr. and Alexander V. Turbiner. (quasi)-exact-solvability on the sphere S^{n}. Journal of Mathematical Physics, 56(2):023501, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anciaux:2015:MTS

[109] Henri Anciaux and Yamile Godoy. Marginally trapped submanifolds in Lorentzian space forms and in the Lorentzian product of a space form by the real line. Journal of Mathematical Physics, 56(2):023502, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:2015:AWL

[110] Bennett Palmer. Anisotropic wavefronts and Laguerre geometry. Journal of Mathematical Physics, 56(2):023503, February 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galleas:2015:PDE

[111] W. Galleas. Partial differential equations from integrable vertex models. Journal of Mathematical Physics, 56(2):023504, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Junge:2015:HFD

[112] Marius Junge, Carlos Palazuelos, Javier Parcet, and Mathilde Perrin. Hypercontractivity in finite-dimensional matrix algebras. Journal of Mathematical Physics, 56(2):023505, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delisle:2015:CCS

[113] L. Delisle, V. Hussin, I. Yurdusen, and W. J. Zakrzewski. Constant curvature surfaces of the supersymmetric CP^{N−1} sigma model. Journal of Mathematical Physics, 56(2):023506, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valdes:2015:EDR

[114] Carlos Valero Valdes. Effective diffusion on Riemannian fiber bundles. Journal of Mathematical Physics, 56(2):023507, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2015:NSS

[115] Xiao ying Zhu, Da jun Zhang, and Zong cheng Li. New set of symmetries and Lie algebraic structures of the Toda lattice hierarchy. Journal of Mathematical Physics, 56(2):023508, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bian:2015:GSS

[116] Dongfen Bian, Boling Guo, and Jingjun Zhang. Global strong spherically symmetric solutions to the full compressible Navier–Stokes equations with stress free boundary. Journal of Mathematical Physics, 56 (2):023509, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maltsev:2015:MSC

[117] A. Ya. Maltsev. On the minimal set of conservation laws and the Hamiltonian structure of the Whitham equations. Journal of Mathematical Physics, 56(2):023510, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2015:GIL

[118] T. Skrypnyk. General integrable n-level, many-mode Janes–Cummings– Dicke models and classical r-matrices with spectral parameters. Journal of Mathematical Physics, 56(2):023511, February 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2015:EGN

[119] Lei Han and Hai-Jun Wang. Erratum: “General nonlocality in quantum fields” [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 56(2):029901, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [6].

Liu:2015:SSS

[120] Chungen Liu and Qiang Ren. On the steady-state solutions of a nonlinear photonic lattice model. Journal of Mathematical Physics, 56(3):031501, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2015:WPF

[121] Jinlong Wei and Rongrong Tian. Well-posedness for the fractional Fokker–Planck equations. Journal of Mathematical Physics, 56(3): 031502, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2015:OSO

[122] N. Aizawa, Z. Kuznetsova, and F. Toppan. l-oscillators from second-order invariant PDEs of the centrally extended conformal Galilei algebras. Journal of Mathematical Physics, 56(3):031701, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendez:2015:NSA

[123] J. M. Méndez and Jesús Urías. On the no-signaling approach to quantum nonlocality. Journal of Mathematical Physics, 56(3):032101, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2015:TEI

[124] Zhenshan Yang. Temporal evolution of instantaneous phonons in timedependent harmonic oscillators. Journal of Mathematical Physics, 56 (3):032102, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohsawa:2015:SCL

[125] Tomoki Ohsawa. Symmetry and conservation laws in semiclassical wave packet dynamics. Journal of Mathematical Physics, 56(3):032103, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:2015:PHS

[126] Avinash Khare and Avadh Saxena. Periodic and hyperbolic soliton solutions of a number of nonlocal nonlinear equations. Journal of Mathematical Physics, 56(3):032104, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loubenets:2015:CIQ

[127] Elena R. Loubenets. Context-invariant quasi hidden variable (qHV) modelling of all joint von Neumann measurements for an arbitrary Hilbert space. Journal of Mathematical Physics, 56(3):032201, March 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2015:UQC

[128] Shawn X. Cui and Zhenghan Wang. Universal quantum computation with metaplectic anyons. Journal of Mathematical Physics, 56(3):032202, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hahn:2015:SCSa

[129] Atle Hahn. From simplicial Chern–Simons theory to the shadow invariant I. Journal of Mathematical Physics, 56(3):032301, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hahn:2015:SCSb

[130] Atle Hahn. From simplicial Chern–Simons theory to the shadow invariant II. Journal of Mathematical Physics, 56(3):032302, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hofer-Szabo:2015:CBL

[131] Gábor Hofer-Szabó and Péter Vecsernyés. On the concept of Bell’s local causality in local classical and quantum theory. Journal of Mathematical Physics, 56(3):032303, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia:2015:MTP

[132] Alberto García, Eva Hackmann, Jutta Kunz, Claus Lämmerzahl, and Alfredo Macías. Motion of test particles in a regular black hole space–time. Journal of Mathematical Physics, 56(3):032501, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cresson:2015:SPS

[133] J. Cresson, F. Pierret, and B. Puig. The Sharma–Parthasarathy stochastic two-body problem. Journal of Mathematical Physics, 56(3):032701, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:ITD

[134] Jie Liu and Jianguo Si. Invariant tori for a derivative nonlinear Schrödinger equation with quasi-periodic forcing. Journal of Mathematical Physics, 56(3):032702, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tibboel:2015:ELB

[135] Pieter Tibboel. Existence of a lower bound for the distance between point masses of relative equilibria for generalised quasi-homogeneous nbody problems and the curved n-body problem. Journal of Mathematical Physics, 56(3):032901, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgiou:2015:PCM

[136] Tryphon T. Georgiou and Michele Pavon. Positive contraction mappings for classical and quantum Schrödinger systems. Journal of Mathematical Physics, 56(3):033301, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsobanjan:2015:SSL

[137] Artur Tsobanjan. Semiclassical states on Lie algebras. Journal of Mathematical Physics, 56(3):033501, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miki:2015:NRF

[138] Hiroshi Miki and Satoshi Tsujimoto. A new recurrence formula for generic exceptional orthogonal polynomials. Journal of Mathematical Physics, 56(3):033502, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawakami:2015:MPS

[139] Hiroshi Kawakami. Matrix Painlevé systems. Journal of Mathematical Physics, 56(3):033503, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2015:SDV

[140] Xin-Yue Li, Qiu-Lan Zhao, Yu-Xia Li, and Huan-He Dong. A superdiscrete variational identity and its application for constructing superdiscrete Hamiltonian systems. Journal of Mathematical Physics, 56(3): 033504, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vakhnenko:2015:NIM

[141] Oleksiy O. Vakhnenko. Nonlinear integrable model of Frenkel-like excitations on a ribbon of triangular lattice. Journal of Mathematical Physics, 56(3):033505, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2015:FDR

[142] Francesco Calogero. Finite-dimensional representations of difference operators and the identification of remarkable matrices. Journal of Mathematical Physics, 56(3):033506, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

SilvaCosta:2015:SSF

[143] F. Silva Costa, J. A. P. F. Marão, J. C. Alves Soares, and E. Capelas de Oliveira. Similarity solution to fractional nonlinear space–time diffusionwave equation. Journal of Mathematical Physics, 56(3):033507, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2015:ERR

[144] Yongkuan Cheng and Jun Yang. The existence result for a relativistic nonlinear Schrödinger equation. Journal of Mathematical Physics, 56 (3):033508, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amburg:2015:ESL

[145] Ilya Amburg, Roshan Sharma, Daniel M. Sussman, and William K. Wootters. Erratum: “States that ‘look the same’ with respect to every basis in a mutually unbiased set” [J. Math. Phys. 55, 122206 (2014)]. Journal of Mathematical Physics, 56(3):039901, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [20].

Huang:2015:UTM

[146] Hsin-Yuan Huang, Youngae Lee, and Chang-Shou Lin. Uniqueness of topological multi-vortex solutions for a skew-symmetric Chern–Simons system. Journal of Mathematical Physics, 56(4):041501, April 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2015:PSN

[147] Chunhua Jin. Periodic solution for a non-isentropic compressible Navier– Stokes equations in a bounded domain. Journal of Mathematical Physics, 56(4):041502, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2015:MPS

[148] Jun Wang, Dianchen Lu, Junxiang Xu, and Fubao Zhang. Multiple positive solutions for semilinear Schrödinger equations with critical growth in R^{N}. Journal of Mathematical Physics, 56(4):041503, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2015:SLE

[149] Sihua Liang and Jihui Zhang. On some p-Laplacian equation with electromagnetic fields and critical nonlinearity in R^{N}. Journal of Mathematical Physics, 56(4):041504, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:MSS

[150] Zhisu Liu, Shangjiang Guo, and Yanqin Fang. Multiple semiclassical states for coupled Schrödinger–Poisson equations with critical exponential growth. Journal of Mathematical Physics, 56(4):041505, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sellaroli:2015:WET

[151] Giuseppe Sellaroli. Wigner–Eckart theorem for the non-compact algebra ∫↕(∈, R). Journal of Mathematical Physics, 56(4):041701, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez-Nunez:2015:GFW

[152] José Fernández-Núñez, Wifredo García-Fuertes, and Askold M. Perelomov. On the generating function of weight multiplicities for the representations of the Lie algebra C_{2}. Journal of Mathematical Physics, 56 (4):041702, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:CLC

[153] Jiao Zhang. On the cohomology of Leibniz conformal algebras. Journal of Mathematical Physics, 56(4):041703, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mojaveri:2015:GEC

[154] B. Mojaveri and A. Dehghani. Generation of excited coherent states for a charged particle in a uniform magnetic field. Journal of Mathematical Physics, 56(4):041704, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2015:IVM

[155] Chengkang Xu, Shaobin Tan, and Qing Wang. ϕ-imaginary Verma modules and their generalizations for the toroidal Lie algebras. Journal of Mathematical Physics, 56(4):041705, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2015:RYD

[156] Haixing Zhu. Relative Yetter–Drinfeld modules and comodules over braided groups. Journal of Mathematical Physics, 56(4):041706, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Devi:2015:CDF

[157] Yendrembam Chaoba Devi, Shivraj Prajapat, Aritra K. Mukhopadhyay, Biswajit Chakraborty, and Frederik G. Scholtz. Connes distance function on fuzzy sphere and the connection between geometry and statistics. Journal of Mathematical Physics, 56(4):041707, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2015:ABG

[158] Hark-Mahn Kim and Ick-Soon Chang. Asymptotic behavior of generalized *-derivations on C^{∗}-algebras with applications. Journal of Mathematical Physics, 56(4):041708, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikitin:2015:SSP

[159] A. G. Nikitin and T. M. Zasadko. Superintegrable systems with position dependent mass. Journal of Mathematical Physics, 56(4):042101, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2015:FBM

[160] Shuyuan Xiao, Xueli Mu, Zhixuan Deng, and Hong Chen. Five-body Moshinsky brackets. Journal of Mathematical Physics, 56(4):042102, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Syue:2015:PSQ

[161] Cheng-De Syue and De-Hone Lin. Prolate spheroidal quantum cloak. Journal of Mathematical Physics, 56(4):042103, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ribeiro:2015:LDB

[162] A. D. Ribeiro. Level density of a Bose gas: Beyond the saddle point approximation. Journal of Mathematical Physics, 56(4):042104, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyazawa:2015:FUM

[163] Toru Miyazawa. Formulation of a unified method for low- and highenergy expansions in the analysis of reflection coefficients for onedimensional Schrödinger equation. Journal of Mathematical Physics, 56 (4):042105, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sasaki:2015:PAZ

[164] Ryu Sasaki. Perturbations around the zeros of classical orthogonal polynomials. Journal of Mathematical Physics, 56(4):042106, April 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blass:2015:AAQ

[165] Andreas Blass and Yuri Gurevich. Ancilla-approximable quantum state transformations. Journal of Mathematical Physics, 56(4):042201, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lienert:2015:RIE

[166] Matthias Lienert. A relativistically interacting exactly solvable multitime model for two massless Dirac particles in 1 + 1 dimensions. Journal of Mathematical Physics, 56(4):042301, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agarwala:2015:GCM

[167] Susama Agarwala and Colleen Delaney. Generalizing the Connes– Moscovici Hopf algebra to contain all rooted trees. Journal of Mathematical Physics, 56(4):042302, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiessling:2015:DPE

[168] M. K.-H. Kiessling and A. S. Tahvildar-Zadeh. The Dirac point electron in zero-gravity Kerr–Newman spacetime. Journal of Mathematical Physics, 56(4):042303, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tahvildar-Zadeh:2015:ZGL

[169] A. Shadi Tahvildar-Zadeh. On a zero-gravity limit of the Kerr–Newman spacetimes and their electromagnetic fields. Journal of Mathematical Physics, 56(4):042501, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akbar:2015:ESD

[170] Fiki T. Akbar and Bobby E. Gunara. Existence of static dyonic black holes in 4 d n = 1 supergravity with finite energy. Journal of Mathematical Physics, 56(4):042502, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novruzov:2015:BPP

[171] Emil Novruzov. Blow-up phenomena for polytropic equation with inhomogeneous density and source. Journal of Mathematical Physics, 56 (4):042701, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anzaldo-Meneses:2015:NAD

[172] A. Anzaldo-Meneses. On non-autonomous dynamical systems. Journal of Mathematical Physics, 56(4):042702, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2015:SDS

[173] Manuel F. Rañada. Superintegrable deformations of superintegrable systems: Quadratic superintegrability and higher-order superintegrability. Journal of Mathematical Physics, 56(4):042703, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burton:2015:NPA

[174] David A. Burton, Stephen P. Flood, and Haibao Wen. Non-perturbative aspects of particle acceleration in non-linear electrodynamics. Journal of Mathematical Physics, 56(4):042901, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haug:2015:UAA

[175] Nils Haug and Thomas Prellberg. Uniform asymptotics of area-weighted Dyck paths. Journal of Mathematical Physics, 56(4):043301, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sliusarenko:2015:BBG

[176] O. Yu. Sliusarenko, A. V. Chechkin, and Yu. V. Slyusarenko. The Bogolyubov–Born–Green–Kirkwood–Yvon hierarchy and Fokker–Planck equation for many-body dissipative randomly driven systems. Journal of Mathematical Physics, 56(4):043302, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjornberg:2015:DTC

[177] Jakob E. Björnberg and Daniel Ueltschi. Decay of transverse correlations in quantum Heisenberg models. Journal of Mathematical Physics, 56(4): 043303, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2015:UCA

[178] Peng Zhao and Engui Fan. A unified construction for the algebrogeometric quasiperiodic solutions of the Lotka–Volterra and relativistic Lotka–Volterra hierarchy. Journal of Mathematical Physics, 56(4): 043501, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:2015:ISD

[179] Bao-Feng Feng, Ken ichi Maruno, and Yasuhiro Ohta. Integrable semidiscretization of a multi-component short pulse equation. Journal of Mathematical Physics, 56(4):043502, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fucci:2015:SFR

[180] Guglielmo Fucci, Curtis Graham, and Klaus Kirsten. Spectral functions for regular Sturm–Liouville problems. Journal of Mathematical Physics, 56(4):043503, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaffet:2015:SCP

[181] B. Gaffet. A special class of Painlevé differential systems that generalize the elliptic functions. Journal of Mathematical Physics, 56(4):043504, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harju:2015:ATS

[182] Antti J. Harju. Analysis of twisted supercharge families on product manifolds. Journal of Mathematical Physics, 56(4):043505, April 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:2015:INV

[183] Palle Jørgensen and Feng Tian. Infinite networks and variation of conductance functions in discrete Laplacians. Journal of Mathematical Physics, 56(4):043506, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Voicu:2015:CVC

[184] Nicoleta Voicu and Demeter Krupka. Canonical variational completion of differential equations. Journal of Mathematical Physics, 56(4):043507, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Etesi:2015:CSS

[185] Gábor Etesi. Complex structure on the six dimensional sphere from a spontaneous symmetry breaking. Journal of Mathematical Physics, 56 (4):043508, April 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2015:SSA

[186] Claudianor O. Alves and Francisco Julio S. A. Corrêa. A subsupersolution approach for a quasilinear Kirchhoff equation. Journal of Mathematical Physics, 56(5):051501, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ali:2015:DBS

[187] Amir Ali, Hadi Susanto, and Jonathan A. D. Wattis. Decay of bound states in a sine-Gordon equation with double-well potentials. Journal of Mathematical Physics, 56(5):051502, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeSole:2015:ISB

[188] Alberto De Sole, Victor G. Kac, and Refik Turhan. On integrability of some bi-Hamiltonian two field systems of partial differential equations. Journal of Mathematical Physics, 56(5):051503, May 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Fei:2015:BCS

[189] Minggang Fei and Zhaoyin Xiang. On the blow-up criterion and small data global existence for the Hall-magnetohydrodynamics with horizontal dissipation. Journal of Mathematical Physics, 56(5):051504, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2015:IMS

[190] Chunhua Wang and Jing Yang. Infinitely many solutions to a linearly coupled Schrödinger system with non-symmetric potentials. Journal of Mathematical Physics, 56(5):051505, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueiredo:2015:ELE

[191] G. M. Figueiredo and J. R. Santos Júnior. Existence of a least energy nodal solution for a Schrödinger–Kirchhoff equation with potential vanishing at infinity. Journal of Mathematical Physics, 56(5):051506, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stubis:2015:PCC

[192] Keaton Stubis. A proof of the conjecture by Carpentier–De Sole–Kac. Journal of Mathematical Physics, 56(5):051701, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chanyal:2015:SOR

[193] B. C. Chanyal. Split octonion reformulation of generalized linear gravitational field equations. Journal of Mathematical Physics, 56(5):051702, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Douglas:2015:LDS

[194] Andrew Douglas and Joe Repka. Levi decomposable subalgebras of the symplectic algebra C_{2}. Journal of Mathematical Physics, 56(5):051703, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mukhin:2015:GM

[195] Evgeny Mukhin, Benoît Vicedo, and Charles Young. Gaudin models for gl(m|n). Journal of Mathematical Physics, 56(5):051704, May 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Bru:2015:MCL

[196] J.-B. Bru, W. de Siqueira Pedra, and C. Hertling. Microscopic conductivity of lattice fermions at equilibrium. I. Non-interacting particles. Journal of Mathematical Physics, 56(5):051901, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banik:2015:MIS

[197] Manik Banik. Measurement incompatibility and Schrödinger–Einstein– Podolsky–Rosen steering in a class of probabilistic theories. Journal of Mathematical Physics, 56(5):052101, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blokhintsev:2015:NAF

[198] L. D. Blokhintsev, A. S. Kadyrov, A. M. Mukhamedzhanov, and D. A. Savin. New approach to folding with the Coulomb wave function. Journal of Mathematical Physics, 56(5):052102, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2015:EAS

[199] Thanh-Son Nguyen, Dai-Nam Le, Tuan-Quoc N. Thoi, and Van-Hoang Le. Exact analytical solutions of the Schrödinger equation for the ninedimensional MICZ–Kepler problem. Journal of Mathematical Physics, 56(5):052103, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benguria:2015:CEZ

[200] R. D. Benguria and H. Van Den Bosch. A criterion for the existence of zero modes for the Pauli operator with fastly decaying fields. Journal of Mathematical Physics, 56(5):052104, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Safari:2015:AEA

[201] L. Safari, J. P. Santos, P. Amaro, K. Jänkälä, and F. Fratini. Analytical evaluation of atomic form factors: Application to Rayleigh scattering. Journal of Mathematical Physics, 56(5):052105, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Batic:2015:PHC

[202] D. Batic, D. Mills-Howell, and M. Nowakowski. Potentials of the Heun class: The triconfluent case. Journal of Mathematical Physics, 56(5): 052106, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ni:2015:NCS

[203] Xiaotong Ni, Oliver Buerschaper, and Maarten Van den Nest. A noncommuting stabilizer formalism. Journal of Mathematical Physics, 56(5): 052201, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DePalma:2015:NFD

[204] Giacomo De Palma, Andrea Mari, Vittorio Giovannetti, and Alexander S. Holevo. Normal form decomposition for Gaussian-to-Gaussian superoperators. Journal of Mathematical Physics, 56(5):052202, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnon-Friedman:2015:FRC

[205] Rotem Arnon-Friedman and Renato Renner. de Finetti reductions for correlations. Journal of Mathematical Physics, 56(5):052203, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lipponen:2015:MRC

[206] Henri Lipponen. Mickelsson-Rajeev cocycle corresponding to dimension five. Journal of Mathematical Physics, 56(5):052301, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2015:KKW

[207] Jian Wang and Yong Wang. The Kastler–Kalau–Walze type theorem for six-dimensional manifolds with boundary. Journal of Mathematical Physics, 56(5):052501, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quevedo:2015:GSP

[208] Hernando Quevedo and Saken Toktarbay. Generating static perfect-fluid solutions of Einstein’s equations. Journal of Mathematical Physics, 56 (5):052502, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2015:QPS

[209] Jing Chang, Yixian Gao, and Yong Li. Quasi-periodic solutions of nonlinear beam equation with prescribed frequencies. Journal of Mathematical Physics, 56(5):052701, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gerdjikov:2015:IER

[210] V. S. Gerdjikov, D. M. Mladenov, A. A. Stefanov, and S. K. Varbev. Integrable equations and recursion operators related to the affine Lie algebras a r ( 1 ). Journal of Mathematical Physics, 56(5):052702, May 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meng:2015:TKP

[211] Guowu Meng. On the trajectories of O(1)-Kepler problems. Journal of Mathematical Physics, 56(5):052901, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Webb:2015:MSL

[212] G. M. Webb. Multi-symplectic, Lagrangian, one-dimensional gas dynamics. Journal of Mathematical Physics, 56(5):053101, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruiz:2015:ESF

[213] G. Ruiz and C. Tsallis. Emergence of q-statistical functions in a generalized binomial distribution with strong correlations. Journal of Mathematical Physics, 56(5):053301, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frohlich:2015:SPC

[214] Jürg Fröhlich and Daniel Ueltschi. Some properties of correlations of quantum lattice systems in thermal equilibrium. Journal of Mathematical Physics, 56(5):053302, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mouayn:2015:PRS

[215] Zouhaïr Mouayn. Polyanalytic relativistic second Bargmann transforms. Journal of Mathematical Physics, 56(5):053501, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLange:2015:MTH

[216] O. L. de Lange and R. E. Raab. Multipole theory and the hehl-Obukhov decomposition of the electromagnetic constitutive tensor. Journal of Mathematical Physics, 56(5):053502, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giscard:2015:EFT

[217] P.-L. Giscard, K. Lui, S. J. Thwaite, and D. Jaksch. An exact formulation of the time-ordered exponential using path-sums. Journal of Mathematical Physics, 56(5):053503, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2015:DAO

[218] Xiaorui Hu, Yuqi Li, and Yong Chen. A direct algorithm of onedimensional optimal system for the group invariant solutions. Journal of Mathematical Physics, 56(5):053504, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cherman:2015:RHW

[219] Aleksey Cherman, Peter Koroteev, and Mithat Ünsal. Resurgence and holomorphy: From weak to strong coupling. Journal of Mathematical Physics, 56(5):053505, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2015:RRM

[220] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials. II. Journal of Mathematical Physics, 56(5):053506, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muminov:2015:SLS

[221] M. Muminov, H. Neidhardt, and T. Rasulov. On the spectrum of the lattice spin-boson Hamiltonian for any coupling: 1D case. Journal of Mathematical Physics, 56(5):053507, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2015:CTS

[222] Yunqing Yang, Zhenya Yan, and Dumitru Mihalache. Controlling temporal solitary waves in the generalized inhomogeneous coupled nonlinear Schrödinger equations with varying source terms. Journal of Mathematical Physics, 56(5):053508, May 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seo:2015:NLB

[223] Ihyeok Seo. A note on local behavior of eigenfunctions of the Schrödinger operator. Journal of Mathematical Physics, 56(6):061501, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deng:2015:CMM

[224] Yinbin Deng, Lu Lu, and Wei Shuai. Constraint minimizers of mass critical Hartree energy functionals: Existence and mass concentration.

Journal of Mathematical Physics, 56(6):061503, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lai:2015:IJF

[225] Geng Lai. Interaction of jump-fan composite waves in a two-dimensional jet for van der Waals gases. Journal of Mathematical Physics, 56(6): 061504, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Araruna:2015:BKS

[226] F. D. Araruna, P. Braz e Silva, R. R. Carvalho, and M. A. Rojas-Medar. On the behavior of Kazhikov–Smagulov mass diffusion model for vanishing diffusion and viscosity coefficients. Journal of Mathematical Physics, 56(6):061505, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colombeau:2015:WAM

[227] M. Colombeau. Weak asymptotic methods for 3-D self-gravitating pressureless fluids. application to the creation and evolution of solar systems from the fully nonlinear Euler–Poisson equations. Journal of Mathematical Physics, 56(6):061506, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertola:2015:DTP

[228] Marco Bertola and Pietro Giavedoni. A degeneration of two-phase solutions of the focusing nonlinear Schrödinger equation via Riemann–Hilbert problems. Journal of Mathematical Physics, 56(6):061507, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohari:2015:TIP

[229] Anilesh Mohari. Translation invariant pure state on B = ⊗_{j∈Z}M_{d}^{(j)}(C) and its split property. Journal of Mathematical Physics, 56(6):061701, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2015:HSF

[230] Jixia Yuan and Wende Liu. Hom-structures on finite-dimensional simple Lie superalgebras. Journal of Mathematical Physics, 56(6):061702, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartin:2015:BRS

[231] Daniel Cartin. Braids as a representation space of SU(5). Journal of Mathematical Physics, 56(6):061703, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roeck:2015:LPP

[232] W. De Roeck and M. Schütz. Local perturbations perturb-exponentially-locally. Journal of Mathematical Physics, 56(6):061901, June 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bravyi:2015:GGP

[233] Sergey Bravyi and David Gosset. Gapped and gapless phases of frustration-free spin- 1 2 chains. Journal of Mathematical Physics, 56 (6):061902, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2015:RDF

[234] Tarek Saanouni. Remarks on damped fractional Schrödinger equation with pure power nonlinearity. Journal of Mathematical Physics, 56(6): 061502, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xie:2015:ARZ

[235] Qiongtao Xie and Shiguang Rong. Analytical results of zero-gap states in periodic potentials. Journal of Mathematical Physics, 56(6):062101, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marquette:2015:DOA

[236] Ian Marquette and Christiane Quesne. Deformed oscillator algebra approach of some quantum superintegrable Lissajous systems on the sphere and of their rational extensions. Journal of Mathematical Physics, 56(6): 062102, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagchi:2015:RET

[237] B. Bagchi, Y. Grandati, and C. Quesne. Rational extensions of the trigonometric Darboux–Pöschl–Teller potential based on para-Jacobi polynomials. Journal of Mathematical Physics, 56(6):062103, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aglietti:2015:GWM

[238] U. G. Aglietti and P. M. Santini. Geometry of Winter model. Journal of Mathematical Physics, 56(6):062104, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Camblong:2015:ASM

[239] Horacio E. Camblong, Luis N. Epele, Huner Fanchiotti, and Carlos A. García Canal. Analytic structure of the s-matrix for singular quantum mechanics. Journal of Mathematical Physics, 56(6):062105, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Withers:2015:BCD

[240] L. P. Withers, Jr. and F. A. Narducci. Bilocal current densities and mean trajectories in a Young interferometer with two Gaussian slits and two detectors. Journal of Mathematical Physics, 56(6):062106, June 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fonseca:2015:QAM

[241] I. C. Fonseca and K. Bakke. Quantum aspects of a moving magnetic quadrupole moment interacting with an electric field. Journal of Mathematical Physics, 56(6):062107, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2325].

Amir:2015:CSN

[242] Naila Amir and Shahid Iqbal. Coherent states for nonlinear harmonic oscillator and some of its properties. Journal of Mathematical Physics, 56(6):062108, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novaes:2015:STDb

[243] Marcel Novaes. Statistics of time delay and scattering correlation functions in chaotic systems. II. Semiclassical approximation. Journal of Mathematical Physics, 56(6):062109, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novaes:2015:STDa

[244] Marcel Novaes. Statistics of time delay and scattering correlation functions in chaotic systems. I. Random matrix theory. Journal of Mathematical Physics, 56(6):062110, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chabab:2015:ESD

[245] M. Chabab, A. El Batoul, and M. Oulne. Exact solutions of deformed Schrödinger equation with a class of non-central physical potentials. Journal of Mathematical Physics, 56(6):062111, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [302].

Jozsa:2015:SPI

[246] Richard Jozsa and Graeme Mitchison. Symmetric polynomials in information theory: Entropy and subentropy. Journal of Mathematical Physics, 56(6):062201, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rivasseau:2015:CLV

[247] Vincent Rivasseau and Zhituo Wang. Corrected loop vertex expansion for Φ_{2}^{4} theory. Journal of Mathematical Physics, 56(6):062301, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agarwala:2015:GFC

[248] Susama Agarwala. The geometric β-function in curved space–time under operator regularization. Journal of Mathematical Physics, 56(6):062302, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fathizadeh:2015:SCN

[249] Farzad Fathizadeh. On the scalar curvature for the noncommutative four torus. Journal of Mathematical Physics, 56(6):062303, June 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lohmann:2015:SSC

[250] Martin Lohmann. Single scale cluster expansions with applications to many Boson and unbounded spin systems. Journal of Mathematical Physics, 56(6):063301, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bandyopadhyay:2015:BRE

[251] J. Bandyopadhyay and J. J. L. Velázquez. Blow-up rate estimates for the solutions of the bosonic Boltzmann–Nordheim equation. Journal of Mathematical Physics, 56(6):063302, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shcherbina:2015:USM

[252] Tatyana Shcherbina. Universality of the second mixed moment of the characteristic polynomials of the 1D band matrices: Real symmetric case. Journal of Mathematical Physics, 56(6):063303, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valtancoli:2015:PIN

[253] P. Valtancoli. Path integral and noncommutative Poisson brackets. Journal of Mathematical Physics, 56(6):063501, June 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2015:PKF

[254] Douglas R. Anderson and Darin J. Ulness. Properties of the Katugampola fractional derivative with potential application in quantum mechanics. Journal of Mathematical Physics, 56(6):063502, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baxter:2015:ZCR

[255] Mathew Baxter, S. Roy Choudhury, and Robert A. Van Gorder. Zero curvature representation, bi-Hamiltonian structure, and an integrable hierarchy for the Zakharov–Ito system. Journal of Mathematical Physics, 56(6):063503, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karayer:2015:ENU

[256] H. Karayer, D. Demirhan, and F. Büyükkiliç. Extension of Nikiforov– Uvarov method for the solution of Heun equation. Journal of Mathematical Physics, 56(6):063504, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carinena:2015:JMN

[257] J. F. Cariñena, J. de Lucas, and M. F. Rañada. Jacobi multipliers, nonlocal symmetries, and nonlinear oscillators. Journal of Mathematical Physics, 56(6):063505, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:SLS

[258] Min Chen and Yang Chen. Singular linear statistics of the Laguerre unitary ensemble and Painlevé. III. double scaling analysis. Journal of Mathematical Physics, 56(6):063506, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Damaskinsky:2015:CGF

[259] E. V. Damaskinsky, P. P. Kulish, and M. A. Sokolov. On calculation of generating functions of Chebyshev polynomials in several variables. Journal of Mathematical Physics, 56(6):063507, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kachmar:2015:SCT

[260] Ayman Kachmar and Marwa Nasrallah. Semi-classical trace asymptotics for magnetic Schrödinger operators with Robin condition. Journal of Mathematical Physics, 56(7):071501, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almog:2015:ESS

[261] Yaniv Almog, Leonid Berlyand, Dmitry Golovaty, and Itai Shafrir. Existence and stability of superconducting solutions for the Ginzburg– Landau equations in the presence of weak electric currents. Journal of Mathematical Physics, 56(7):071502, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2015:ESS

[262] Lan Huang and Ruxu Lian. Exponential stability of spherically symmetric solutions for compressible viscous micropolar fluid. Journal of Mathematical Physics, 56(7):071503, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schutz:2015:SFL

[263] Lineia Schütz, Janaína P. Zingano, and Paulo R. Zingano. On the supnorm form of Leray’s problem for the incompressible Navier–Stokes equations. Journal of Mathematical Physics, 56(7):071504, July 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prinari:2015:DBS

[264] Barbara Prinari, Federica Vitale, and Gino Biondini. Dark-bright soliton solutions with nontrivial polarization interactions for the threecomponent defocusing nonlinear Schrödinger equation with nonzero boundary conditions. Journal of Mathematical Physics, 56(7):071505, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:OSS

[265] Xingxing Liu. Orbital stability of the sum of N solitary waves for the hyperelastic-rod wave equation. Journal of Mathematical Physics, 56(7):

071506, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:MSC

[266] Caisheng Chen. Multiple solutions for a class of quasilinear Schrödinger equations in R^{N}. Journal of Mathematical Physics, 56(7):071507, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2015:RWS

[267] Xing-Bin Pan. Regularity of weak solutions to nonlinear Maxwell systems. Journal of Mathematical Physics, 56(7):071508, July 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barnich:2015:CSA

[268] Glenn Barnich, Gastón Giribet, and Mauricio Leston. Chern–Simons action for inhomogeneous Virasoro group as extension of three dimensional flat gravity. Journal of Mathematical Physics, 56(7):071701, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kolesnikov:2015:HAO

[269] Pavel Kolesnikov. Homogeneous averaging operators on simple finite conformal Lie algebras. Journal of Mathematical Physics, 56(7):071702, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2015:NPB

[270] Zhaoting Wei. The noncommutative Poisson bracket and the deformation of the family algebras. Journal of Mathematical Physics, 56(7):071703, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burciu:2015:GRE

[271] Sebastian Burciu. On the Grothendieck rings of equivariant fusion categories. Journal of Mathematical Physics, 56(7):071704, July 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vlaar:2015:BTM

[272] Bart Vlaar. Boundary transfer matrices and boundary quantum KZ equations. Journal of Mathematical Physics, 56(7):071705, July 2015.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Noguera:2015:SIS

[273] Norman Noguera and Krzysztof Rózga. On square-integrability of solutions of the stationary Schrödinger equation for the quantum harmonic oscillator in two dimensional constant curvature spaces. Journal of Mathematical Physics, 56(7):072101, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2015:PFQ

[274] Siye Wu. Projective flatness in the quantisation of bosons and fermions. Journal of Mathematical Physics, 56(7):072102, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:UAM

[275] Hong-Biao Zhang, Guang-Yuan Jiang, and Gang-Cheng Wang. Unified algebraic method to non-Hermitian systems with Lie algebraic linear structure. Journal of Mathematical Physics, 56(7):072103, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berrada:2015:CPA

[276] K. Berrada. Construction of photon-added spin coherent states and their statistical properties. Journal of Mathematical Physics, 56(7):072104, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fischer:2015:DSA

[277] Stefan G. Fischer and Andreas Buchleitner. On the derivation of the semiclassical approximation to the quantum propagator. Journal of Mathematical Physics, 56(7):072105, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2015:TPD

[278] Axel Schulze-Halberg and Jie Wang. Two-parameter double-oscillator model of Mathews–Lakshmanan type: Series solutions and supersymmetric partners. Journal of Mathematical Physics, 56(7):072106, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2015:QMP

[279] A. D. Alhaidari and M. E. H. Ismail. Quantum mechanics without potential function. Journal of Mathematical Physics, 56(7):072107, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Evangelides:2015:ART

[280] P. Evangelides, C. Lei, and A. Vourdas. Analytic representations with theta functions for systems on Z(d) and on S. Journal of Mathematical Physics, 56(7):072108, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Englis:2015:OPL

[281] Miroslav Englis and S. Twareque Ali. Orthogonal polynomials, Laguerre Fock space, and quasi-classical asymptotics. Journal of Mathematical Physics, 56(7):072109, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gebert:2015:AEC

[282] Martin Gebert. The asymptotics of an eigenfunction-correlation determinant for Dirac-δ perturbations. Journal of Mathematical Physics, 56(7): 072110, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2015:GGS

[283] Shawn X. Cui, Nengkun Yu, and Bei Zeng. Generalized graph states based on Hadamard matrices. Journal of Mathematical Physics, 56(7): 072201, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castrigiano:2015:CLR

[284] Domenico P. L. Castrigiano and Andreas D. Leiseifer. Causal localizations in relativistic quantum mechanics. Journal of Mathematical Physics, 56(7):072301, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2015:CAM

[285] Wei He. Combinatorial approach to Mathieu and Lamé equations. Journal of Mathematical Physics, 56(7):072302, July 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2015:TLS

[286] Luise Adams, Christian Bogner, and Stefan Weinzierl. The two-loop sunrise integral around four space–time dimensions and generalisations of the Clausen and Glaisher functions towards the elliptic case. Journal of Mathematical Physics, 56(7):072303, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klein:2015:VAF

[287] David Klein and Jake Reschke. Velocity addition formulas in Robertson– Walker spacetimes. Journal of Mathematical Physics, 56(7):072501, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciftci:2015:BDM

[288] Dilek K. Çiftci and Özgür Delice. Brans–Dicke–Maxwell solutions for higher dimensional static cylindrical symmetric spacetime. Journal of Mathematical Physics, 56(7):072502, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krongos:2015:GCP

[289] D. S. Krongos and C. G. Torre. Geometrization conditions for perfect fluids, scalar fields, and electromagnetic fields. Journal of Mathematical Physics, 56(7):072503, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strauss:2015:QFT

[290] Y. Strauss, L. P. Horwitz, J. Levitan, and A. Yahalom. Quantum field theory of classically unstable Hamiltonian dynamics. Journal of Mathematical Physics, 56(7):072701, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mi:2015:QPS

[291] Lufang Mi and Hongzi Cong. Quasi-periodic solutions for d-dimensional beam equation with derivative nonlinear perturbation. Journal of Mathematical Physics, 56(7):072702, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goto:2015:LSC

[292] Shin itiro Goto. Legendre submanifolds in contact manifolds as attractors and geometric nonequilibrium thermodynamics. Journal of Mathematical Physics, 56(7):073301, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strauss:2015:MLP

[293] Y. Strauss. A modified Lax–Phillips scattering theory for quantum mechanics. Journal of Mathematical Physics, 56(7):073501, July 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2015:SDQ

[294] Satoru Odake and Ryu Sasaki. Solvable discrete quantum mechanics: q-orthogonal polynomials with q = 1 and quantum dilogarithm. Journal of Mathematical Physics, 56(7):073502, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akant:2015:BEB

[295] Levent Akant, Emine Ertugrul, Yusuf Gül, and O. Teoman Turgut. Boundary effects on Bose–Einstein condensation in ultra-static space– times. Journal of Mathematical Physics, 56(7):073503, July 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hutnikova:2015:CST

[296] Mária Hutníková and Anna Misková. Continuous Stockwell transform: Coherent states and localization operators. Journal of Mathematical Physics, 56(7):073504, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Charalambous:2015:NDY

[297] Nelia Charalambous and Leonard Gross. Neumann domination for the Yang–Mills heat equation. Journal of Mathematical Physics, 56(7): 073505, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yolcu:2015:REB

[298] Selma Yildirim Yolcu and Türkay Yolcu. Refined eigenvalue bounds on the Dirichlet fractional Laplacian. Journal of Mathematical Physics, 56(7):073506, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gouba:2015:TDQ

[299] Laure Gouba. Time-dependent q-deformed bi-coherent states for generalized uncertainty relations. Journal of Mathematical Physics, 56(7): 073507, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kravchenko:2015:SPP

[300] Vladislav V. Kravchenko, Sergii M. Torba, and Ulises Velasco-García. Spectral parameter power series for Sturm–Liouville equations with a potential polynomially dependent on the spectral parameter and Zakharov– Shabat systems. Journal of Mathematical Physics, 56(7):073508, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakabekov:2015:EBC

[301] A. Sakabekov and Y. Auzhani. Erratum: “Boundary conditions for the one-dimensional nonlinear nonstationary Boltzmann’s moment system equations” [J. Math. Phys. 55, 123507 (2014)]. Journal of Mathematical Physics, 56(7):079901, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [21].

Chabab:2015:PNE

[302] M. Chabab, A. El Batoul, and M. Oulne. Publisher’s note: “Exact solutions of deformed Schrödinger equation with a class of non-central physical potentials” [J. Math. Phys. 56, 062111 (2015)]. Journal of Mathematical Physics, 56(7):079902, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Corrections of typographical errors in three equations; see [245].

Sheng:2015:EST

[303] Wei-Jie Sheng and Jia-Bing Wang. Entire solutions of time periodic bistable reaction-advection-diffusion equations in infinite cylinders. Journal of Mathematical Physics, 56(8):081501, August 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barre:2015:NIS

[304] Julien Barré and Yoshiyuki Y. Yamaguchi. On the neighborhood of an inhomogeneous stable stationary solution of the Vlasov equation-case of an attractive cosine potential. Journal of Mathematical Physics, 56(8): 081502, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pei:2015:WSB

[305] Pei Pei, Mohammad A. Rammaha, and Daniel Toundykov. Weak solutions and blow-up for wave equations of p-Laplacian type with supercritical sources. Journal of Mathematical Physics, 56(8):081503, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2015:NSB

[306] Kung-Chien Wu. Nonlinear stability of the Boltzmann equation in a periodic box. Journal of Mathematical Physics, 56(8):081504, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2015:SLE

[307] Seung-Yeal Ha and Qinghua Xiao. L^{2}-stability of the Landau equation near global Maxwellians. Journal of Mathematical Physics, 56(8):081505, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:SCR

[308] Qingshan Zhang and Yuxiang Li. Stabilization and convergence rate in a chemotaxis system with consumption of chemoattractant. Journal of Mathematical Physics, 56(8):081506, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashton:2015:MMA

[309] Thomas Ashton and Andrey Mudrov. R-matrix and Mickelsson algebras for orthosymplectic quantum groups. Journal of Mathematical Physics, 56(8):081701, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeier:2015:SRC

[310] Robert Zeier and Zoltán Zimborás. On squares of representations of compact Lie algebras. Journal of Mathematical Physics, 56(8):081702, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suto:2015:TMQ

[311] András Sütő. The total momentum of quantum fluids. Journal of Mathematical Physics, 56(8):081901, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2015:IPT

[312] Tuncay Aktosun and Vassilis G. Papanicolaou. Inverse problem with transmission eigenvalues for the discrete Schrödinger equation. Journal of Mathematical Physics, 56(8):082101, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lv:2015:NCF

[313] Cui hong Lv and Hong yi Fan. New complex function space related to both entangled state representation and spin coherent state. Journal of Mathematical Physics, 56(8):082102, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cowan:2015:DWF

[314] Charles Wesley Cowan and Roderich Tumulka. Detecting wave function collapse without prior knowledge. Journal of Mathematical Physics, 56 (8):082103, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luz:2015:GSQ

[315] Esther Bonet Luz and Cesare Tronci. Geometry and symmetry of quantum and classical-quantum variational principles. Journal of Mathematical Physics, 56(8):082104, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forest:2015:ESS

[316] Simon Forest, David Gosset, Vadym Kliuchnikov, and David McKinnon. Exact synthesis of single-qubit unitaries over Clifford-cyclotomic gate sets. Journal of Mathematical Physics, 56(8):082201, August 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heinosaari:2015:BGI

[317] Teiko Heinosaari, Jukka Kiukas, and Jussi Schultz. Breaking Gaussian incompatibility on continuous variable quantum systems. Journal of Mathematical Physics, 56(8):082202, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barrett:2015:MGF

[318] John W. Barrett. Matrix geometries and fuzzy spaces as finite spectral triples. Journal of Mathematical Physics, 56(8):082301, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:2015:AEE

[319] Lajos Molnár and Dániel Virosztek. On algebraic endomorphisms of the Einstein gyrogroup. Journal of Mathematical Physics, 56(8):082302, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hustler:2015:MTK

[320] Noel Hustler and Andree Lischewski. (m-theory-)killing spinors on symmetric spaces. Journal of Mathematical Physics, 56(8):082303, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konstantinou-Rizos:2015:RGR

[321] S. Konstantinou-Rizos, A. V. Mikhailov, and P. Xenitidis. Reduction groups and related integrable difference systems of nonlinear Schrödinger type. Journal of Mathematical Physics, 56(8):082701, August 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krusch:2015:MSL

[322] Steffen Krusch and Abera A. Muhamed. Moduli spaces of lumps on real projective space. Journal of Mathematical Physics, 56(8):082901, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Demessie:2015:HPL

[323] G. A. Demessie and C. Sämann. Higher Poincaré lemma and integrability. Journal of Mathematical Physics, 56(8):082902, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Traizet:2015:HVM

[324] Martin Traizet. Hollow vortices and minimal surfaces. Journal of Mathematical Physics, 56(8):083101, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tossounian:2015:PTK

[325] Hagop Tossounian and Ranjini Vaidyanathan. Partially thermostated Kac model. Journal of Mathematical Physics, 56(8):083301, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belitsky:2015:SDT

[326] V. Belitsky and G. M. Schütz. Self-duality for the two-component asymmetric simple exclusion process. Journal of Mathematical Physics, 56 (8):083302, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunajski:2015:EWC

[327] M. Dunajski, E. V. Ferapontov, and B. Kruglikov. On the Einstein–Weyl and conformal self-duality equations. Journal of Mathematical Physics, 56(8):083501, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsko:2015:TSRa

[328] Andrei Bytsko. Tensor space representations of Temperley–Lieb algebra via orthogonal projections of rank r ≥ 1. Journal of Mathematical Physics, 56(8):083502, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsko:2015:TSRb

[329] Andrei Bytsko. Tensor space representations of Temperley–Lieb algebra and generalized permutation matrices. Journal of Mathematical Physics, 56(8):083503, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2015:PLAa

[330] Giuseppe Gaeta. Poincaré-like approach to Landau theory. I. General theory. Journal of Mathematical Physics, 56(8):083504, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2015:PLAb

[331] Giuseppe Gaeta. Poincaré-like approach to Landau theory. II. Simplifying the Landau–deGennes potential for nematic liquid crystals. Journal of Mathematical Physics, 56(8):083505, August 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2015:YCT

[332] Xiao-Yu Jia, Shao-Kui Yao, Ke Wu, and Wei-Zhong Zhao. On Yangian covariance of the triple product system with the rational R-matrix. Journal of Mathematical Physics, 56(8):083506, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grammaticos:2015:NRD

[333] B. Grammaticos and A. Ramani. On a novel representation of discrete Painlevé equations. Journal of Mathematical Physics, 56(8):083507, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kang:2015:QST

[334] Yuanbao Kang. Quantum stopping times stochastic integral in the interacting Fock space. Journal of Mathematical Physics, 56(8):083508, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2015:ATD

[335] Tongsong Jiang, Ziwu Jiang, and Zhaozhong Zhang. Algebraic techniques for diagonalization of a split quaternion matrix in split quaternionic mechanics. Journal of Mathematical Physics, 56(8):083509, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muraleetharan:2015:CSQ

[336] B. Muraleetharan and K. Thirulogasanthar. Coherent state quantization of quaternions. Journal of Mathematical Physics, 56(8):083510, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorska:2015:ERM

[337] K. Górska and W. A. Woyczyński. Explicit representations for multiscale Lévy processes and asymptotics of multifractal conservation laws. Journal of Mathematical Physics, 56(8):083511, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Androulakis:2015:GQM

[338] George Androulakis and Matthew Ziemke. Generators of quantum Markov semigroups. Journal of Mathematical Physics, 56(8):083512, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:CSE

[339] Yi Zhang and Ji-Bin Li. Comment on “Superposition of elliptic functions as solutions for a large number of nonlinear equations” [J. Math. Phys. 56, 032104 (2015)]. Journal of Mathematical Physics, 56(8):084101, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [18].

Craddock:2015:FTT

[340] Mark Craddock. Fourier type transforms on Lie symmetry groups. Journal of Mathematical Physics, 56(9):091501, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:EMS

[341] Hui Zhang, Junxiang Xu, and Fubao Zhang. Existence and multiplicity of solutions for superlinear fractional Schrödinger equations in R^{N}. Journal of Mathematical Physics, 56(9):091502, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Du:2015:BCV

[342] Lili Du and Yongfu Wang. A blowup criterion for viscous, compressible, and heat-conductive magnetohydrodynamic flows. Journal of Mathematical Physics, 56(9):091503, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2015:MSS

[343] Anqiang Zhu and Xiaomei Sun. Multiple solutions for Schrödinger-Poisson type equation with magnetic field. Journal of Mathematical Physics, 56(9):091504, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grudsky:2015:IST

[344] Sergei Grudsky, Christian Remling, and Alexei Rybkin. The inverse scattering transform for the KdV equation with step-like singular Miura initial profiles. Journal of Mathematical Physics, 56(9):091505, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2015:FWB

[345] Stephen C. Anco, Priscila Leal da Silva, and Igor Leite Freire. A family of wave-breaking equations generalizing the Camassa–Holm and Novikov equations. Journal of Mathematical Physics, 56(9):091506, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Long:2015:SSV

[346] Wei Long and Qingfang Wang. Segregated and synchronized vector solutions to linearly coupled systems of Schrödinger equations. Journal of Mathematical Physics, 56(9):091507, September 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Larkin:2015:GRS

[347] N. A. Larkin. Global regular solutions for the 3D Zakharov–Kuznetsov equation posed on unbounded domains. Journal of Mathematical Physics, 56(9):091508, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yukon:2015:FTS

[348] Stanford P. Yukon and Boris A. Malomed. Fluxons in a triangular set of coupled long Josephson junctions. Journal of Mathematical Physics, 56 (9):091509, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirai:2015:MBS

[349] Shin ichi Shirai. Multi-bump solutions to a nonlinear Schrödinger equation with steep magnetic wells. Journal of Mathematical Physics, 56(9): 091510, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaikina:2015:NFS

[350] Elena I. Kaikina. Nonlinear fractional Schrödinger equation on a halfline. Journal of Mathematical Physics, 56(9):091511, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhai:2015:GWP

[351] Cuili Zhai and Ting Zhang. Global well-posedness to the 3-D incompressible inhomogeneous Navier–Stokes equations with a class of large velocity. Journal of Mathematical Physics, 56(9):091512, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoshino:2015:ASE

[352] Gaku Hoshino and Tohru Ozawa. Analytic smoothing effect for a system of Schrödinger equations with three wave interaction. Journal of Mathematical Physics, 56(9):091513, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Medina:2015:VEG

[353] Luciano Medina. Vortex equations governing the fractional quantum Hall effect. Journal of Mathematical Physics, 56(9):091514, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanisli:2015:OFM

[354] Murat Tanisli, Süleyman Demir, and Neslihan Sahin. Octonic formulations of Maxwell type fluid equations. Journal of Mathematical Physics, 56(9):091701, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2015:GFM

[355] José Fernández Núñez, Wifredo García Fuertes, and Askold M. Perelomov. Generating functions and multiplicity formulas: The case of rank two simple Lie algebras. Journal of Mathematical Physics, 56(9):091702, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Curtright:2015:MRS

[356] Thomas L. Curtright. More on rotations as spin matrix polynomials. Journal of Mathematical Physics, 56(9):091703, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matassa:2015:AWL

[357] Marco Matassa. An analogue of Weyl’s law for quantized irreducible generalized flag manifolds. Journal of Mathematical Physics, 56(9):091704, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramm:2015:SEW

[358] A. G. Ramm. Scattering of electromagnetic waves by many small perfectly conducting or impedance bodies. Journal of Mathematical Physics, 56(9):091901, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:2015:CBF

[359] Edwin Langmann and Per Moosavi. Construction by bosonization of a fermion-phonon model. Journal of Mathematical Physics, 56(9):091902, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boscain:2015:CSB

[360] Ugo Boscain, Paolo Mason, Gianluca Panati, and Mario Sigalotti. On the control of spin-boson systems. Journal of Mathematical Physics, 56(9): 092101, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nathanson:2015:GSS

[361] Ekaterina S. Nathanson and Palle E. T. Jørgensen. A global solution to the Schrödinger equation: From Henstock to Feynman. Journal of Mathematical Physics, 56(9):092102, September 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weder:2015:STM

[362] Ricardo Weder. Scattering theory for the matrix Schrödinger operator on the half line with general boundary conditions. Journal of Mathematical Physics, 56(9):092103, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1946].

Cabrera:2015:DCR

[363] Alejandro Cabrera, Edson de Faria, Enrique Pujals, and Charles Tresser. Differentiability of correlations in realistic quantum mechanics. Journal of Mathematical Physics, 56(9):092104, September 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lami:2015:EBI

[364] L. Lami and V. Giovannetti. Entanglement-breaking indices. Journal of Mathematical Physics, 56(9):092201, September 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuramochi:2015:CLI

[365] Yui Kuramochi. Construction of the least informative observable conserved by a given quantum instrument. Journal of Mathematical Physics, 56(9):092202, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cottrell:2015:SMF

[366] Seth S. Cottrell. A simple method for finding the scattering coefficients of quantum graphs. Journal of Mathematical Physics, 56(9):092203, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horwitz:2015:IRT

[367] Lawrence P. Horwitz and Meir Zeilig-Hess. Induced representations of tensors and spinors of any rank in the Stueckelberg–Horwitz–Piron theory. Journal of Mathematical Physics, 56(9):092301, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deser:2015:SPG

[368] Andreas Deser. Star products on graded manifolds and α′-corrections to Courant algebroids from string theory. Journal of Mathematical Physics, 56(9):092302, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hanusch:2015:UML

[369] Maximilian Hanusch. Uniqueness of measures in loop quantum cosmology. Journal of Mathematical Physics, 56(9):092303, September 2015.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vieira:2015:QNC

[370] H. S. Vieira and V. B. Bezerra. Quantum Newtonian cosmology and the biconfluent Heun functions. Journal of Mathematical Physics, 56(9): 092501, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brito:2015:CTG

[371] Irene Brito. Conformal transformations in general relativistic elasticity. Journal of Mathematical Physics, 56(9):092502, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2015:BPW

[372] Zaiyun Zhang, Jianhua Huang, and Mingbao Sun. Blow-up phenomena for the weakly dissipative Dullin–Gottwald–Holm equation revisited. Journal of Mathematical Physics, 56(9):092701, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2015:RAN

[373] Min Zhao and Shengfan Zhou. Random attractor of non-autonomous stochastic Boussinesq lattice system. Journal of Mathematical Physics, 56(9):092702, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2015:PAN

[374] Kaixuan Zhu and Chunyou Sun. Pullback attractors for nonclassical diffusion equations with delays. Journal of Mathematical Physics, 56(9): 092703, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbone:2015:EID

[375] Raffaella Carbone, Emanuela Sasso, and Veronica Umanità. Environment induced decoherence for Markovian evolutions. Journal of Mathematical Physics, 56(9):092704, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joshi:2015:LEA

[376] Nalini Joshi, Nobutaka Nakazono, and Yang Shi. Lattice equations arising from discrete Painlevé systems. I. (A_{2}+A_{1})^{(1)} and (A_{1}+A′_{1})^{(1)} cases.

Journal of Mathematical Physics, 56(9):092705, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rabinowitch:2015:PAS

[377] Alexander S. Rabinowitch. On a particular analytical solution to the 3D Navier–Stokes equations and its peculiarity for high Reynolds numbers. Journal of Mathematical Physics, 56(9):093101, September 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garra:2015:FDT

[378] Roberto Garra, Enzo Orsingher, and Federico Polito. Fractional diffusions with time-varying coefficients. Journal of Mathematical Physics, 56 (9):093301, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gandolfo:2015:GEL

[379] Daniel Gandolfo, Suren Poghosyan, and Jean Ruiz. Geometric expansion of the log-partition function of the anisotropic Heisenberg model. Journal of Mathematical Physics, 56(9):093302, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Much:2015:SAD

[380] Albert Much. Self-adjointness of deformed unbounded operators. Journal of Mathematical Physics, 56(9):093501, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayashi:2015:IFO

[381] Nakao Hayashi and Pavel I. Naumkin. On the inhomogeneous fourthorder nonlinear Schrödinger equation. Journal of Mathematical Physics, 56(9):093502, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geloun:2015:PRR

[382] Joseph Ben Geloun and Reiko Toriumi. Parametric representation of rank d tensorial group field theory: Abelian models with kinetic term ∑_{s} |p_{s}| + μ. Journal of Mathematical Physics, 56(9):093503, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giraldi:2015:TES

[383] Filippo Giraldi. Transcendental equations in the Schwinger–Keldysh nonequilibrium theory and nonvanishing correlations. Journal of Mathematical Physics, 56(9):093504, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bei:2015:SCH

[384] Francesco Bei and Nils Waterstraat. On the space of connections having non-trivial twisted harmonic spinors. Journal of Mathematical Physics, 56(9):093505, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2015:LVF

[385] Alexander Figotin and Guillermo Reyes. Lagrangian variational framework for boundary value problems. Journal of Mathematical Physics, 56 (9):093506, September 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buhe:2015:SRE

[386] Eerdun Buhe and George W. Bluman. Symmetry reductions, exact solutions, and conservation laws of the generalized Zakharov equations. Journal of Mathematical Physics, 56(10):101501, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chirkunov:2015:SMN

[387] Yu. A. Chirkunov. Submodels of model of nonlinear diffusion in the inhomogeneous medium involving absorption. Journal of Mathematical Physics, 56(10):101502, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carles:2015:SNS

[388] Rémi Carles and Clément Gallo. Scattering for the nonlinear Schrödinger equation with a general one-dimensional confinement. Journal of Mathematical Physics, 56(10):101503, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2015:BRK

[389] Qi-Hong Shi, Wan-Tong Li, and Shu Wang. Blowup results for the KGS system with higher order Yukawa coupling. Journal of Mathematical Physics, 56(10):101504, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2015:ETV

[390] Boling Guo and Fangfang Li. Existence of topological vortices in an Abelian Chern–Simons model. Journal of Mathematical Physics, 56(10):

101505, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2015:OSI

[391] Stephen C. Anco, Abdus Sattar Mia, and Mark R. Willoughby. Oscillatory solitons of U(1)-invariant mKdV equations. I. Envelope speed and temporal frequency. Journal of Mathematical Physics, 56(10):101506, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2015:YFA

[392] Liang Cheng and Anqiang Zhu. Yamabe flow and ADM mass on asymptotically flat manifolds. Journal of Mathematical Physics, 56(10):101507, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eelbode:2015:DPK

[393] D. Eelbode, T. Raeymaekers, and J. Van der Jeugt. Decomposition of the polynomial kernel of arbitrary higher spin Dirac operators. Journal of Mathematical Physics, 56(10):101701, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Binczak:2015:SPE

[394] G. Bińczak and J. Kaleta. Sharp and principal elements in effect algebras. Journal of Mathematical Physics, 56(10):102101, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez:2015:BOA

[395] André Martinez and Vania Sordoni. On the Born–Oppenheimer approximation of diatomic molecular resonances. Journal of Mathematical Physics, 56(10):102102, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bradford:2015:SSB

[396] R. A. W. Bradford. A spherically symmetric bound state of the coupled Maxwell–Dirac equations with self-interaction alone. Journal of Mathematical Physics, 56(10):102103, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marchesiello:2015:TOS

[397] A. Marchesiello, S. Post, and L. Snobl. Third-order superintegrable systems with potentials satisfying only nonlinear equations. Journal of Mathematical Physics, 56(10):102104, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abbott:2015:VKS

[398] Alastair A. Abbott, Cristian S. Calude, and Karl Svozil. A variant of the Kochen–Specker theorem localising value indefiniteness. Journal of Mathematical Physics, 56(10):102201, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandao:2015:ALF

[399] Fernando G. S. L. Brandão, Toby S. Cubitt, Angelo Lucia, Spyridon Michalakis, and David Perez-Garcia. Area law for fixed points of rapidly mixing dissipative quantum systems. Journal of Mathematical Physics, 56(10):102202, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parthasarathy:2015:ESE

[400] K. R. Parthasarathy and Ritabrata Sengupta. Exchangeable, stationary, and entangled chains of Gaussian states. Journal of Mathematical Physics, 56(10):102203, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cubitt:2015:QRH

[401] Toby Cubitt, Michael Kastoryano, Ashley Montanaro, and Kristan Temme. Quantum reverse hypercontractivity. Journal of Mathematical Physics, 56(10):102204, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuramochi:2015:MSP

[402] Yui Kuramochi. Minimal sufficient positive-operator valued measure on a separable Hilbert space. Journal of Mathematical Physics, 56(10):102205, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [842].

Fiorenza:2015:WZW

[403] D. Fiorenza, H. Sati, and U. Schreiber. The Wess–Zumino–Witten term of the M5-brane and differential cohomotopy. Journal of Mathematical Physics, 56(10):102301, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fredenhagen:2015:LAO

[404] Stefan Fredenhagen, Jens Hoppe, and Mariusz Hynek. The Lorentz anomaly via operator product expansion. Journal of Mathematical Physics, 56(10):102302, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maharana:2015:HES

[405] Jnanadeva Maharana. High energy scatterings in higher dimensional theories. Journal of Mathematical Physics, 56(10):102303, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimock:2015:NRS

[406] J. Dimock. Nonperturbative renormalization of scalar quantum electrodynamics in d = 3. Journal of Mathematical Physics, 56(10):102304, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2015:NTB

[407] Chuan-Jing Song and Yi Zhang. Noether theorem for Birkhoffian systems on time scales. Journal of Mathematical Physics, 56(10):102701, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shu:2015:RAS

[408] Ji Shu, Ping Li, Jia Zhang, and Ou Liao. Random attractors for the stochastic coupled fractional Ginzburg–Landau equation with additive noise. Journal of Mathematical Physics, 56(10):102702, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forger:2015:CPB

[409] Michael Forger and Mário O. Salles. On covariant Poisson brackets in classical field theory. Journal of Mathematical Physics, 56(10):102901, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

To:2015:BCG

[410] Quy-Dong To, Van-Huyen Vu, Guy Lauriat, and Céline Léonard. Boundary conditions for gas flow problems from anisotropic scattering kernels. Journal of Mathematical Physics, 56(10):103101, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alt:2015:LSL

[411] Johannes Alt. The local semicircle law for random matrices with a fourfold symmetry. Journal of Mathematical Physics, 56(10):103301, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andraus:2015:TSA

[412] Sergio Andraus and Seiji Miyashita. Two-step asymptotics of scaled Dunkl processes. Journal of Mathematical Physics, 56(10):103302, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbone:2015:IGA

[413] Lisa Carbone, Scott H. Murray, and Hisham Sati. Integral group actions on symmetric spaces and discrete duality symmetries of supergravity theories. Journal of Mathematical Physics, 56(10):103501, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aljadeff:2015:EBS

[414] Johnatan Aljadeff, David Renfrew, and Merav Stern. Eigenvalues of block structured asymmetric random matrices. Journal of Mathematical Physics, 56(10):103502, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Enblom:2015:HCT

[415] Alexandra Enblom. Hardy–Carleman type inequalities for Dirac operators. Journal of Mathematical Physics, 56(10):103503, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaufmann:2015:OOG

[416] Ralph M. Kaufmann, Benjamin C. Ward, and J. Javier Zúñiga. The odd origin of Gerstenhaber brackets, Batalin–Vilkovisky operators, and master equations. Journal of Mathematical Physics, 56(10):103504, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geracie:2015:CNR

[417] Michael Geracie, Kartik Prabhu, and Matthew M. Roberts. Curved non-relativistic spacetimes, Newtonian gravitation and massive matter. Journal of Mathematical Physics, 56(10):103505, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tarasov:2015:FOD

[418] Vasily E. Tarasov. Fractional-order difference equations for physical lattices and some applications. Journal of Mathematical Physics, 56(10): 103506, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2015:IST

[419] Qi Li, Shu hua Su, Ming hui Li, and Jian bing Zhang. The inverse scattering transform for the differential-difference Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 56(10):103507, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2015:GBT

[420] F. Bagarello and A. Fring. Generalized Bogoliubov transformations versus D-pseudo-bosons. Journal of Mathematical Physics, 56(10):103508, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xie:2015:SQM

[421] Yongjian Xie, Aili Yang, and Fang Ren. Super quantum measures on effect algebras with the Riesz decomposition properties. Journal of Mathematical Physics, 56(10):103509, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2015:ESS

[422] Boling Guo, Daiwen Huang, and Zhiqiang Wang. Existence and stability of standing waves for coupled derivative Schrödinger equations. Journal of Mathematical Physics, 56(10):103510, October 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vassilevich:2015:SGS

[423] Dmitri Vassilevich. Spectral geometry of symplectic spinors. Journal of Mathematical Physics, 56(10):103511, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chvartatskyi:2015:REB

[424] O. Chvartatskyi, F. Müller-Hoissen, and N. Stoilov. “Riemann equations” in bidifferential calculus. Journal of Mathematical Physics, 56 (10):103512, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krejcirik:2015:PNH

[425] D. Krejcirík, P. Siegl, M. Tater, and J. Viola. Pseudospectra in non-Hermitian quantum mechanics. Journal of Mathematical Physics, 56 (10):103513, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bietenholz:2015:CTD

[426] Wolfgang Bietenholz, Chryssomalis Chryssomalakos, and Marcelo Salgado. Comment on “The two dimensional motion of a particle in an inverse square potential: Classical and quantum aspects” [J. Math. Phys. 54, 053509 (2013)]. Journal of Mathematical Physics, 56(10):104101, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [14].

Kim:2015:ELC

[427] Do-Hyung Kim. Erratum: “Lorentzian covering space and homotopy classes” [j. math. phys. 52, 122501 (2011)]. Journal of Mathematical Physics, 56(10):109901, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5].

Fioresi:2015:EMC

[428] R. Fioresi, M. A. Lledó, and V. S. Varadarajan. Erratum: “The Minkowski and conformal superspaces” [j. math. phys. 48, 113505 (2007)]. Journal of Mathematical Physics, 56(10):109902, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2].

Harko:2015:ETW

[429] T. Harko and M. K. Mak. Exact travelling wave solutions of nonlinear reaction-convection-diffusion equations — an Abel equation based approach. Journal of Mathematical Physics, 56(11):111501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marrocos:2015:ENL

[430] Marcus A. M. Marrocos and Antônio L. Pereira. Eigenvalues of the Neumann Laplacian in symmetric regions. Journal of Mathematical Physics, 56(11):111502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alama:2015:VSW

[431] Stan Alama, Lia Bronsard, and Xavier Lamy. Vortex structure in pwave superconductors. Journal of Mathematical Physics, 56(11):111503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2015:GSR

[432] Jianjun Yuan. Global solution of the relativistic Vlasov–Maxwell–Chern– Simons system. Journal of Mathematical Physics, 56(11):111504, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manukure:2015:IES

[433] Solomon Manukure and Wen-Xiu Ma. An integrable extension of a soliton hierarchy associated with so(3, R) and its bi-Hamiltonian formulation. Journal of Mathematical Physics, 56(11):111505, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2015:TER

[434] Jianzhi Han and Chunrui Ai. Three equivalent rationalities of vertex operator superalgebras. Journal of Mathematical Physics, 56(11):111701, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2015:DBM

[435] Hongmei Hu and Naihong Hu. Double-bosonization and Majid’s conjecture, (I): Rank-inductions of ABCD. Journal of Mathematical Physics, 56(11):111702, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Perlov:2015:APW

[436] Leonid Perlov. Analog of the Peter–Weyl expansion for Lorentz group. Journal of Mathematical Physics, 56(11):111703, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lavicka:2015:FDP

[437] Roman Lávicka and Dalibor Smíd. Fischer decomposition for polynomials on superspace. Journal of Mathematical Physics, 56(11):111704, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dhand:2015:ABR

[438] Ish Dhand, Barry C. Sanders, and Hubert de Guise. Algorithms for SU(n) boson realizations and D-functions. Journal of Mathematical Physics, 56(11):111705, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fa:2015:LBS

[439] Huanxia Fa, Junbo Li, and Yelong Zheng. Lie bialgebra structures on the deformative Schrödinger–Virasoro algebras. Journal of Mathematical Physics, 56(11):111706, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruillard:2015:LDR

[440] Paul Bruillard, Liang Chang, Seung-Moon Hong, Julia Yael Plavnik, Eric C. Rowell, and Michael Yuan Sun. Low-dimensional representations of the three component loop braid group. Journal of Mathematical Physics, 56(11):111707, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qian:2015:SDF

[441] Zhixin Qian. Static dielectric function with exact exchange contribution in the electron liquid. Journal of Mathematical Physics, 56(11):111901, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedmann:2015:QMD

[442] Tamar Friedmann and C. R. Hagen. Quantum mechanical derivation of the Wallis formula for π. Journal of Mathematical Physics, 56(11): 112101, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabin:2015:ETB

[443] Carlos Sabín, Pablo Barberis-Blostein, Cristopher Hernández, Robert B. Mann, and Ivette Fuentes. Effects of three-body collisions in a two-mode Bose–Einstein condensate. Journal of Mathematical Physics, 56(11): 112102, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhatia:2015:SEP

[444] Rajendra Bhatia and Tanvi Jain. On symplectic eigenvalues of positive definite matrices. Journal of Mathematical Physics, 56(11):112201, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garner:2015:GLC

[445] David Garner and Sanjaye Ramgoolam. The geometry of the light-cone cell decomposition of moduli space. Journal of Mathematical Physics, 56(11):112301, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

You:2015:HHG

[446] Miman You, Nan Zhou, and Shuanhong Wang. Hom–Hopf group coalgebras and braided T -categories obtained from Hom–Hopf algebras. Journal of Mathematical Physics, 56(11):112302, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delaFuente:2015:UDM

[447] Daniel de la Fuente, Alfonso Romero, and Pedro J. Torres. Unchanged direction motion in general relativity: The problems of prescribing acceleration and the extensibility of trajectories. Journal of Mathematical Physics, 56(11):112501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Izumiya:2015:LHA

[448] Shyuichi Izumiya. Lightlike hypersurfaces along spacelike submanifolds in anti-de Sitter space. Journal of Mathematical Physics, 56(11):112502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khuri:2015:IBS

[449] Marcus A. Khuri. Inequalities between size and charge for bodies and the existence of black holes due to concentration of charge. Journal of Mathematical Physics, 56(11):112503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bihun:2015:PZG

[450] Oksana Bihun and Francesco Calogero. Properties of the zeros of generalized basic hypergeometric polynomials. Journal of Mathematical Physics, 56(11):112701, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakhnovich:2015:DSD

[451] A. L. Sakhnovich. Dynamical and spectral Dirac systems: response function and inverse problems. Journal of Mathematical Physics, 56(11):

112702, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nold:2015:NSM

[452] Andreas Nold, Martin Oberlack, and Alexei F. Cheviakov. On new stability modes of plane canonical shear flows using symmetry classification. Journal of Mathematical Physics, 56(11):113101, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adame:2015:EVG

[453] J. Adame and S. Warzel. Exponential vanishing of the ground-state gap of the quantum random energy model via adiabatic quantum computing. Journal of Mathematical Physics, 56(11):113301, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:FCS

[454] Yongxin Chen, Tryphon T. Georgiou, and Michele Pavon. Fast cooling for a system of stochastic oscillators. Journal of Mathematical Physics, 56(11):113302, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hachem:2015:SMI

[455] Walid Hachem, Aris Moustakas, and Leonid A. Pastur. The Shannon’s mutual information of a multiple antenna time and frequency dependent channel: An ergodic operator approach. Journal of Mathematical Physics, 56(11):113501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daletskii:2015:PTC

[456] Alexei Daletskii, Yuri Kondratiev, and Yuri Kozitsky. Phase transitions in continuum ferromagnets with unbounded spins. Journal of Mathematical Physics, 56(11):113502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tamizhmani:2015:SRO

[457] K. M. Tamizhmani, K. Krishnakumar, and P. G. L. Leach. Symmetries and reductions of order for certain nonlinear third- and second-order differential equations with arbitrary nonlinearity. Journal of Mathematical Physics, 56(11):113503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liemert:2015:FST

[458] André Liemert and Alwin Kienle. Fundamental solution of the tempered fractional diffusion equation. Journal of Mathematical Physics, 56(11): 113504, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:ESC

[459] Shaowei Liu. Eigenvalues of the string constraints for the Kadomtsev– Petviashvili hierarchy. Journal of Mathematical Physics, 56(11):113505, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sako:2015:GTN

[460] Akifumi Sako, Toshiya Suzuki, and Hiroshi Umetsu. Gauge theories on noncommutative CP^{N} and Bogomolńyi–Prasad–Sommerfield-like equations. Journal of Mathematical Physics, 56(11):113506, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baulieu:2015:WLF

[461] L. Baulieu, N. L. Holanda, and F. Toppan. A world-line framework for 1D topological conformal σ-models. Journal of Mathematical Physics, 56(11):113507, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrogowska:2015:MLP

[462] Alina Dobrogowska. R-matrix, Lax pair, and multiparameter decompositions of Lie algebras. Journal of Mathematical Physics, 56(11):113508, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strachan:2015:IFA

[463] Ian A. B. Strachan and Dafeng Zuo. Integrability of the Frobenius algebra-valued Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 56(11):113509, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2015:LSG

[464] Qing Huang and Shoufeng Shen. Lie symmetries and group classification of a class of time fractional evolution systems. Journal of Mathematical Physics, 56(12):123504, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1130] and erratum [1359].

Dykema:2015:PNS

[465] Kenneth J. Dykema and Vern Paulsen. Publisher’s note: “Synchronous correlation matrices and Connes embedding conjecture” [J. Math. Phys. 57, 015214 (2016)]. Journal of Mathematical Physics, 56(12):129901, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Correction to footnoted author in [516].

Iacob:2015:ESS

[466] Felix Iacob and Marina Lute. Exact solution to the Schrödinger’s equation with pseudo-Gaussian potential. Journal of Mathematical Physics, 56(12):121501, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masaki:2015:EUT

[467] Satoshi Masaki and Takayoshi Ogawa. Existence and uniqueness of two dimensional Euler–Poisson system and WKB approximation to the nonlinear Schrödinger–Poisson system. Journal of Mathematical Physics, 56(12):121502, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:VPS

[468] Zili Chen. The Vlasov–Poisson system with low steady spatial asymptotics. Journal of Mathematical Physics, 56(12):121503, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2015:OSU

[469] Stephen C. Anco, Abdus Sattar Mia, and Mark R. Willoughby. Oscillatory solitons of U(1)-invariant modified Korteweg–de Vries equations. II. Asymptotic behavior and constants of motion. Journal of Mathematical Physics, 56(12):121504, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheviakov:2015:CCL

[470] A. F. Cheviakov and S. St.Jean. A comparison of conservation law construction approaches for the two-dimensional incompressible Mooney– Rivlin hyperelasticity model. Journal of Mathematical Physics, 56(12): 121505, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tu:2015:IFR

[471] Guizhang Tu. An inversion formula for recursion operator of KP hierarchy. Journal of Mathematical Physics, 56(12):121506, December 2015.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2015:CGS

[472] Haisheng Li and Jiancai Sun. On certain generalizations of the Schrödinger–Virasoro algebra. Journal of Mathematical Physics, 56(12): 121701, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anguelova:2015:MB

[473] Iana I. Anguelova. Multilocal bosonization. Journal of Mathematical Physics, 56(12):121702, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Werry:2015:MED

[474] Jason L. Werry, Mark D. Gould, and Phillip S. Isaac. Matrix elements and duality for type 2 unitary representations of the Lie superalgebra gl(m|n). Journal of Mathematical Physics, 56(12):121703, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdul-Rahman:2015:UAL

[475] Houssam Abdul-Rahman and Günter Stolz. A uniform area law for the entanglement of eigenstates in the disordered XY chain. Journal of Mathematical Physics, 56(12):121901, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:2015:RSS

[476] A. G. Smirnov. Reduction by symmetries in singular quantummechanical problems: General scheme and application to Aharonov– Bohm model. Journal of Mathematical Physics, 56(12):122101, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chowdhury:2015:WFN

[477] S. Hasibul Hassan Chowdhury and S. Twareque Ali. Wigner functions for noncommutative quantum mechanics: a group representation based construction. Journal of Mathematical Physics, 56(12):122102, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landsman:2015:CRT

[478] Klaas Landsman. On the Colbeck–Renner theorem. Journal of Mathematical Physics, 56(12):122103, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Orsucci:2015:HP

[479] Davide Orsucci, Daniel Burgarth, Paolo Facchi, Hiromichi Nakazato, Saverio Pascazio, Kazuya Yuasa, and Vittorio Giovannetti. Hamiltonian purification. Journal of Mathematical Physics, 56(12):122104, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lange:2015:DTS

[480] Rutger-Jan Lange. Distribution theory for Schrödinger’s integral equation. Journal of Mathematical Physics, 56(12):122105, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braun:2015:SCS

[481] Carol Braun, Feifei Li, Anupam Garg, and Michael Stone. The semiclassical coherent state propagator in the Weyl representation. Journal of Mathematical Physics, 56(12):122106, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valtancoli:2015:CTM

[482] P. Valtancoli. Canonical transformations and minimal length. Journal of Mathematical Physics, 56(12):122107, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aleixo:2015:NON

[483] A. N. F. Aleixo and A. B. Balantekin. Normal ordering for nonlinear deformed ladder operators and the f -generalization of the Stirling and Bell numbers. Journal of Mathematical Physics, 56(12):122108, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cariolaro:2015:FEM

[484] Gianfranco Cariolaro and Gianfranco Pierobon. Fock expansion of multimode pure Gaussian states. Journal of Mathematical Physics, 56(12): 122109, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vubangsi:2015:WPD

[485] M. Vubangsi, M. Tchoffo, L. C. Fai, and Yu. M. Pisma’k. Wave packet dynamics for a system with position and time-dependent effective mass in an infinite square well. Journal of Mathematical Physics, 56(12):122110, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Band:2015:ANC

[486] Ram Band, Gregory Berkolaiko, and Tracy Weyand. Anomalous nodal count and singularities in the dispersion relation of honeycomb graphs. Journal of Mathematical Physics, 56(12):122111, December 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blass:2015:OAF

[487] Andreas Blass, Alex Bocharov, and Yuri Gurevich. Optimal ancilla-free Pauli + V circuits for axial rotations. Journal of Mathematical Physics, 56(12):122201, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carles:2015:HOS

[488] Rémi Carles, Wolfgang Lucha, and Emmanuel Moulay. Higher-order Schrödinger and Hartree–Fock equations. Journal of Mathematical Physics, 56(12):122301, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimcik:2015:SMP

[489] Ctirad Klimcík. On supermatrix models, Poisson geometry, and noncommutative supersymmetric gauge theories. Journal of Mathematical Physics, 56(12):122302, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holland:2015:AOP

[490] Jan Holland and Stefan Hollands. Associativity of the operator product expansion. Journal of Mathematical Physics, 56(12):122303, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2015:EUS

[491] Xuetao Yang and Quanxin Zhu. Existence, uniqueness, and stability of stochastic neutral functional differential equations of Sobolev-type. Journal of Mathematical Physics, 56(12):122701, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2015:HRG

[492] Yantao Guo, Ming Wang, and Yanbin Tang. Higher regularity of global attractors of a weakly dissipative fractional Korteweg de Vries equation. Journal of Mathematical Physics, 56(12):122702, December 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2015:EQP

[493] Lin Lu and Xuemei Li. Existence of quasi-periodic solutions of fast excited van der Pol–Mathieu–Duffing equation. Journal of Mathematical Physics, 56(12):122703, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2015:BUS

[494] Jiaoyang Huang, Benjamin Landon, and Horng-Tzer Yau. Bulk universality of sparse random matrices. Journal of Mathematical Physics, 56(12):123301, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouch:2015:EPE

[495] Gabriel Bouch. The expected perimeter in Eden and related growth processes. Journal of Mathematical Physics, 56(12):123302, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouch:2015:CTS

[496] Gabriel Bouch. Complex-time singularity and locality estimates for quantum lattice systems. Journal of Mathematical Physics, 56(12):123303, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buzinski:2015:ASC

[497] David Buzinski and Elizabeth Meckes. Almost sure convergence in quantum spin glasses. Journal of Mathematical Physics, 56(12):123304, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirsten:2015:BFK

[498] Klaus Kirsten and Yoonweon Lee. The Burghelea–Friedlander–Kappelergluing formula for zeta-determinants on a warped product manifold and a product manifold. Journal of Mathematical Physics, 56(12):123501, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1439].

Zaba:2015:NIF

[499] Mariusz Zaba and Piotr Garbaczewski. Nonlocally induced (fractional) bound states: Shape analysis in the infinite Cauchy well. Journal of Mathematical Physics, 56(12):123502, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Latremoliere:2015:CNT

[500] Frédéric Latrémolière. Curved noncommutative tori as Leibniz quantum compact metric spaces. Journal of Mathematical Physics, 56(12):123503, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nagoya:2015:ICB

[501] Hajime Nagoya. Irregular conformal blocks, with an application to the fifth and fourth Painlevé equations. Journal of Mathematical Physics, 56(12):123505, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeAngelis:2015:EPI

[502] Valerio De Angelis and Christophe Vignat. Euler polynomials and identities for non-commutative operators. Journal of Mathematical Physics, 56(12):123506, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moslehian:2016:IOS

[503] Mohammad Sal Moslehian and Mostafa Sattari. Inequalities for operator space numerical radius of 2 × 2 block matrices. Journal of Mathematical Physics, 57(1):015201, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Hermes:2016:PLM

[504] Alexander Müller-Hermes, David Reeb, and Michael M. Wolf. Positivity of linear maps under tensor powers. Journal of Mathematical Physics, 57 (1):015202, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holevo:2016:CCC

[505] A. S. Holevo. On the constrained classical capacity of infinite-dimensional covariant quantum channels. Journal of Mathematical Physics, 57(1): 015203, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rodman:2016:CME

[506] Leiba Rodman, Ilya M. Spitkovsky, Arleta Szkola, and Stephan Weis. Continuity of the maximum-entropy inference: Convex geometry and numerical ranges approach. Journal of Mathematical Physics, 57(1):015204, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2016:VNP

[507] Kyung Hoon Han and Seung-Hyeok Kye. Various notions of positivity for bi-linear maps and applications to tri-partite entanglement. Journal of Mathematical Physics, 57(1):015205, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigi:2016:HLS

[508] Salman Beigi and Christopher King. Hypercontractivity and the logarithmic Sobolev inequality for the completely bounded norm. Journal of Mathematical Physics, 57(1):015206, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jencova:2016:CSP

[509] Anna Jencová. On the convex structure of process positive operator valued measures. Journal of Mathematical Physics, 57(1):015207, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crann:2016:PAQ

[510] Jason Crann, David W. Kribs, Rupert H. Levene, and Ivan G. Todorov. Private algebras in quantum information and infinite-dimensional complementarity. Journal of Mathematical Physics, 57(1):015208, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okamura:2016:MTL

[511] Kazuya Okamura and Masanao Ozawa. Measurement theory in local quantum physics. Journal of Mathematical Physics, 57(1):015209, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hastings:2016:HQN

[512] M. B. Hastings. How quantum are non-negative wavefunctions? Journal of Mathematical Physics, 57(1):015210, January 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiai:2016:CCN

[513] Fumio Hiai and Mary Beth Ruskai. Contraction coefficients for noisy quantum channels. Journal of Mathematical Physics, 57(1):015211, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leung:2016:TPN

[514] Chi-Wai Leung, Chi-Keung Ng, and Ngai-Ching Wong. Transition probabilities of normal states determine the Jordan structure of a quantum system. Journal of Mathematical Physics, 57(1):015212, January 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berta:2016:SEF

[515] Mario Berta, Fabian Furrer, and Volkher B. Scholz. The smooth entropy formalism for von Neumann algebras. Journal of Mathematical Physics, 57(1):015213, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dykema:2016:SCM

[516] Kenneth J. Dykema and Vern Paulsen. Synchronous correlation matrices and Connes’ embedding conjecture. Journal of Mathematical Physics, 57 (1):015214, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Collins:2016:RMT

[517] Benoît Collins and Ion Nechita. Random matrix techniques in quantum information theory. Journal of Mathematical Physics, 57(1):015215, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arizmendi:2016:ADB

[518] Octavio Arizmendi, Ion Nechita, and Carlos Vargas. On the asymptotic distribution of block-modified random matrices. Journal of Mathematical Physics, 57(1):015216, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kastoryano:2016:NCN

[519] Michael Kastoryano and Kristan Temme. Non-commutative Nash inequalities. Journal of Mathematical Physics, 57(1):015217, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shultz:2016:SPA

[520] Fred Shultz. The structural physical approximation conjecture. Journal of Mathematical Physics, 57(1):015218, January 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Monras:2016:QLC

[521] Alex Monràs and Andreas Winter. Quantum learning of classical stochastic processes: The completely positive realization problem. Journal of Mathematical Physics, 57(1):015219, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palazuelos:2016:SNG

[522] Carlos Palazuelos and Thomas Vidick. Survey on nonlocal games and operator space theory. Journal of Mathematical Physics, 57(1):015220, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hainzl:2016:BCS

[523] C. Hainzl and R. Seiringer. The Bardeen–Cooper–Schrieffer functional of superconductivity and its mathematical properties. Journal of Mathematical Physics, 57(2):021101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2016:DRV

[524] Shixiang Ma and Jing Wang. Decay rates to viscous contact waves for the compressible Navier–Stokes equations. Journal of Mathematical Physics, 57(2):021501, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Young:2016:LDR

[525] Brent Young. Landau damping in relativistic plasmas. Journal of Mathematical Physics, 57(2):021502, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2016:ECS

[526] Qin Wang and Kyungwoo Song. Energy conservative solutions to the system of full variational sine-Gordon equations in a unit sphere. Journal of Mathematical Physics, 57(2):021503, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2016:IFA

[527] Xinyuan Wu and Changying Liu. An integral formula adapted to different boundary conditions for arbitrarily high-dimensional nonlinear Klein–Gordon equations with its applications. Journal of Mathematical Physics, 57(2):021504, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hesamifard:2016:TLR

[528] F. Hesamifard and M. M. Rezaii. Two-loop renormalization group flow for Lorentzian manifolds. Journal of Mathematical Physics, 57(2):021505, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2016:NBV

[529] Jingchi Huang and Yuxi Zheng. No blow-up to a variational wave equation in liquid crystals. Journal of Mathematical Physics, 57(2):021506, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wan:2016:GSS

[530] Renhui Wan. Global small solutions to a tropical climate model without thermal diffusion. Journal of Mathematical Physics, 57(2):021507, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valchev:2016:DMQ

[531] T. I. Valchev. Dressing method and quadratic bundles related to symmetric spaces. vanishing boundary conditions. Journal of Mathematical Physics, 57(2):021508, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2016:EDS

[532] Xiaomin Cao. Energy decay of solutions for a variable-coefficient viscoelastic wave equation with a weak nonlinear dissipation. Journal of Mathematical Physics, 57(2):021509, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2016:NLS

[533] Woocheol Choi and Jinmyoung Seok. Nonrelativistic limit of standing waves for pseudo-relativistic nonlinear Schrödinger equations. Journal of Mathematical Physics, 57(2):021510, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalantarov:2016:QPS

[534] Varga K. Kalantarov and Türker Özsari. Qualitative properties of solutions for nonlinear Schrödinger equations with nonlinear boundary conditions on the half-line. Journal of Mathematical Physics, 57(2):021511, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loayza:2016:WPS

[535] Miguel Loayza and Marko A. Rojas-Medar. A weak-L^{p} Prodi–Serrin type regularity criterion for the micropolar fluid equations. Journal of Mathematical Physics, 57(2):021512, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:VVN

[536] Peng Zhang. On the variations of the vortex number in a periodic Ginzburg–Landau model. Journal of Mathematical Physics, 57(2): 021513, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanKortryk:2016:MEG

[537] T. S. Van Kortryk. Matrix exponentials, SU(N) group elements, and real polynomial roots. Journal of Mathematical Physics, 57(2):021701, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Asai:2016:RBR

[538] Nobuhiro Asai, Marek Bozejko, and Takahiro Hasebe. Radial Bargmann representation for the Fock space of type B. Journal of Mathematical Physics, 57(2):021702, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suh:2016:CAA

[539] Uhi Rinn Suh. Classical affine W -algebras associated to Lie superalgebras. Journal of Mathematical Physics, 57(2):021703, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salazar-Ramirez:2016:NRC

[540] M. Salazar-Ramírez, D. Ojeda-Guillén, and R. D. Mota. The number radial coherent states for the generalized MICZ–Kepler problem. Journal of Mathematical Physics, 57(2):021704, February 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beltita:2016:WTI

[541] Ingrid Beltita, Daniel Beltita, and Marius Mantoiu. On Wigner transforms in infinite dimensions. Journal of Mathematical Physics, 57(2): 021705, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:MTP

[542] Jialei Chen and Junli Liu. R-matrix for the two parameter quantum superalgebra U_{r,s}(osp(1|2n)). Journal of Mathematical Physics, 57(2): 021706, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Semikhatov:2016:RER

[543] A. M. Semikhatov and I. Yu. Tipunin. Representations of U_{q}sl(2|1) at even roots of unity. Journal of Mathematical Physics, 57(2):021707, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McCoy:2016:AEM

[544] A. E. McCoy and M. A. Caprio. Algebraic evaluation of matrix elements in the Laguerre function basis. Journal of Mathematical Physics, 57(2): 021708, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamhalter:2016:BCA

[545] Jan Hamhalter, Karimbergen Kudaybergenov, Antonio M. Peralta, and Bernard Russo. Boundedness of completely additive measures with application to 2-local triple derivations. Journal of Mathematical Physics, 57 (2):021709, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rasmussen:2016:OPC

[546] Morten Grud Rasmussen, Benjamin Ricaud, and Baptiste Savoie. On the optical properties of carbon nanotubes. Part I. A general formula for the dynamical optical conductivity. Journal of Mathematical Physics, 57 (2):021901, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Griesemer:2016:SAD

[547] M. Griesemer and A. Wünsch. Self-adjointness and domain of the Fröhlich Hamiltonian. Journal of Mathematical Physics, 57(2):021902, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katsura:2016:IDT

[548] Hosho Katsura and Tohru Koma. The Z_{2} index of disordered topological insulators with time reversal symmetry. Journal of Mathematical Physics, 57(2):021903, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jex:2016:ABS

[549] Michal Jex and Vladimir Lotoreichik. On absence of bound states for weakly attractive δ′-interactions supported on non-closed curves in R^{2}. Journal of Mathematical Physics, 57(2):022101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cherbal:2016:TDP

[550] O. Cherbal and M. Maamache. Time-dependent pseudofermionic systems and coherent states. Journal of Mathematical Physics, 57(2):022102, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nayga:2016:GFE

[551] M. M. Nayga and J. P. Esguerra. Green’s functions and energy eigenvalues for delta-perturbed space-fractional quantum systems. Journal of Mathematical Physics, 57(2):022103, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2016:EBS

[552] Pavel Exner and Semjon Vugalter. On the existence of bound states in asymmetric leaky wires. Journal of Mathematical Physics, 57(2):022104, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woods:2016:DEB

[553] M. P. Woods and M. B. Plenio. Dynamical error bounds for continuum discretisation via Gauss quadrature rules — a Lieb–Robinson bound approach. Journal of Mathematical Physics, 57(2):022105, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beverland:2016:PGT

[554] Michael E. Beverland, Oliver Buerschaper, Robert Koenig, Fernando Pastawski, John Preskill, and Sumit Sijher. Protected gates for topological quantum field theories. Journal of Mathematical Physics, 57(2): 022201, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Hermes:2016:REC

[555] Alexander Müller-Hermes, Daniel Stilck França, and Michael M. Wolf. Relative entropy convergence for depolarizing channels. Journal of Mathematical Physics, 57(2):022202, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Hermes:2016:EPD

[556] Alexander Müller-Hermes, Daniel Stilck França, and Michael M. Wolf. Entropy production of doubly stochastic quantum channels. Journal of Mathematical Physics, 57(2):022203, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takimi:2016:PCI

[557] Tomohisa Takimi. Prescription for choosing an interpolating function. Journal of Mathematical Physics, 57(2):022301, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

David:2016:LQG

[558] François David, Rémi Rhodes, and Vincent Vargas. Liouville quantum gravity on complex tori. Journal of Mathematical Physics, 57(2):022302, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stottmeister:2016:MSC

[559] Alexander Stottmeister and Thomas Thiemann. The microlocal spectrum condition, initial value formulations, and background independence. Journal of Mathematical Physics, 57(2):022303, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sok:2016:PDH

[560] Jérémy Sok. The positronium and the dipositronium in a Hartree–Fock approximation of quantum electrodynamics. Journal of Mathematical Physics, 57(2):022304, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:2016:HQG

[561] I. Y. Park. Holographic quantization of gravity in a black hole background. Journal of Mathematical Physics, 57(2):022305, February 2016.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mathieu:2016:ABT

[562] P. Mathieu and F. Thuillier. Abelian BF theory and Turaev–Viro invariant. Journal of Mathematical Physics, 57(2):022306, February 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fioresi:2016:SOC

[563] R. Fioresi and E. Latini. The symplectic origin of conformal and Minkowski superspaces. Journal of Mathematical Physics, 57(2):022307, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Houri:2016:ATG

[564] Tsuyoshi Houri, Yoshiyuki Morisawa, and Kentaro Tomoda. Antisymmetric tensor generalizations of affine vector fields. Journal of Mathematical Physics, 57(2):022501, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Backdahl:2016:FCV

[565] Thomas Bäckdahl and Juan A. Valiente Kroon. A formalism for the calculus of variations with spinors. Journal of Mathematical Physics, 57 (2):022502, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flores:2016:HSB

[566] J. L. Flores, J. Herrera, and M. Sánchez. Hausdorff separability of the boundaries for spacetimes and sequential spaces. Journal of Mathematical Physics, 57(2):022503, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woolgar:2016:CST

[567] Eric Woolgar and William Wylie. Cosmological singularity theorems and splitting theorems for N–Bakry–Émery spacetimes. Journal of Mathematical Physics, 57(2):022504, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baxter:2016:ETH

[568] J. Erik Baxter. Existence of topological hairy dyons and dyonic black holes in anti-de Sitter su(N) Einstein–Yang–Mills theory. Journal of Mathematical Physics, 57(2):022505, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baxter:2016:SSH

[569] J. Erik Baxter and Elizabeth Winstanley. On the stability of soliton and hairy black hole solutions of su(N) Einstein–Yang–Mills theory with a negative cosmological constant. Journal of Mathematical Physics, 57 (2):022506, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bekaert:2016:CDT

[570] Xavier Bekaert and Kevin Morand. Connections and dynamical trajectories in generalised Newton–Cartan gravity I. An intrinsic view. Journal of Mathematical Physics, 57(2):022507, February 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantica:2016:CPF

[571] Carlo Alberto Mantica, Luca Guido Molinari, and Uday Chand De. A condition for a perfect-fluid space–time to be a generalized Robertson– Walker space–time. Journal of Mathematical Physics, 57(2):022508, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [678].

Bagchi:2016:QAC

[572] Bijan Bagchi, Samiran Ghosh, Barnali Pal, and Swarup Poria. Qualitative analysis of certain generalized classes of quadratic oscillator systems. Journal of Mathematical Physics, 57(2):022701, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ji:2016:QSC

[573] Un Cig Ji and Kalyan B. Sinha. Quantum stochastic calculus associated with quadratic quantum noises. Journal of Mathematical Physics, 57 (2):022702, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mase:2016:IRL

[574] Takafumi Mase. Investigation into the role of the Laurent property in integrability. Journal of Mathematical Physics, 57(2):022703, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:SCW

[575] Jin-Yu Chen, Yu-Jie Chen, Wen-Guei Hu, and Song-Sun Lin. Spatial chaos of Wang tiles with two symbols. Journal of Mathematical Physics, 57(2):022704, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:2016:TDB

[576] Theodora Ioannidou and Árpád Lukács. Time-dependent Bogomolny– Prasad–Sommerfeld skyrmions. Journal of Mathematical Physics, 57(2): 022901, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shipman:2016:PSD

[577] Stephen P. Shipman and Aaron T. Welters. Pathological scattering by a defect in a slow-light periodic layered medium. Journal of Mathematical Physics, 57(2):022902, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harland:2016:CDA

[578] Derek Harland and Daniel Nogradi. On the charge density and asymptotic tail of a monopole. Journal of Mathematical Physics, 57(2):022903, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheskidov:2016:LBP

[579] Alexey Cheskidov and Karen Zaya. Lower bounds of potential blowup solutions of the three-dimensional Navier–Stokes equations in Ḣ^{3/2}. Journal of Mathematical Physics, 57(2):023101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:2016:SDB

[580] H. Bergeron, E. M. F. Curado, J. P. Gazeau, and Ligia M. C. S. Rodrigues. Symmetric deformed binomial distributions: An analytical example where the Boltzmann–Gibbs entropy is not extensive. Journal of Mathematical Physics, 57(2):023301, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2016:FAB

[581] Seung-Yeop Lee and Roman Riser. Fine asymptotic behavior for eigenvalues of random normal matrices: Ellipse case. Journal of Mathematical Physics, 57(2):023302, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Angeletti:2016:DCO

[582] Florian Angeletti and Hugo Touchette. Diffusions conditioned on occupation measures. Journal of Mathematical Physics, 57(2):023303, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciccariello:2016:TSP

[583] Salvino Ciccariello. Thermodynamic similarity of physical systems. Journal of Mathematical Physics, 57(2):023304, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frohlich:2016:QBM

[584] Jürg Fröhlich and Jeffrey Schenker. Quantum Brownian motion induced by thermal noise in the presence of disorder. Journal of Mathematical Physics, 57(2):023305, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:PHD

[585] Min Chen, Yang Chen, and Engui Fan. Perturbed Hankel determinant, correlation functions and Painlevé equations. Journal of Mathematical Physics, 57(2):023501, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2016:QMS

[586] Caishi Wang and Jinshu Chen. Quantum Markov semigroups constructed from quantum Bernoulli noises. Journal of Mathematical Physics, 57(2): 023502, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alpay:2016:STQ

[587] Daniel Alpay, Fabrizio Colombo, and David P. Kimsey. The spectral theorem for quaternionic unbounded normal operators based on the Sspectrum. Journal of Mathematical Physics, 57(2):023503, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Achour:2016:EVS

[588] J. Ben Achour, E. Huguet, J. Queva, and J. Renaud. Explicit vector spherical harmonics on the 3-sphere. Journal of Mathematical Physics, 57 (2):023504, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Acharya:2016:RTC

[589] Keshav Raj Acharya. Remling’s theorem on canonical systems. Journal of Mathematical Physics, 57(2):023505, February 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delisle:2016:GSS

[590] L. Delisle, V. Hussin, and W. J. Zakrzewski. General solutions of the supersymmetric CP^{2} sigma model and its generalisation to CP^{N−1}. Journal of Mathematical Physics, 57(2):023506, February 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Van-Brunt:2016:SRA

[591] Alexander Van-Brunt and Matt Visser. Simplifying the Reinsch algorithm for the Baker–Campbell–Hausdorff series. Journal of Mathematical Physics, 57(2):023507, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Butko:2016:FFP

[592] Yana A. Butko, Martin Grothaus, and Oleg G. Smolyanov. Feynman formulae and phase space Feynman path integrals for tau-quantization of some Lévy–Khintchine type Hamilton functions. Journal of Mathematical Physics, 57(2):023508, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerner:2016:TIP

[593] Joachim Kerner and Tobias Mühlenbruch. Two interacting particles on the half-line. Journal of Mathematical Physics, 57(2):023509, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gannot:2016:ECS

[594] Oran Gannot. Effects of confinement for single-well potentials. Journal of Mathematical Physics, 57(2):023510, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Redelmeier:2016:EMM

[595] C. E. I. Redelmeier. Explicit multi-matrix topological expansion for quaternionic random matrices. Journal of Mathematical Physics, 57(2): 023511, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:2016:NSP

[596] Slawomir Klimek, Sumedha Rathnayake, and Kaoru Sakai. A note on spectral properties of the p-adic tree. Journal of Mathematical Physics, 57(2):023512, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berti:2016:TSP

[597] Alessia Berti, Ivana Bochicchio, and Mauro Fabrizio. Transition and separation process in brine channels formation. Journal of Mathematical Physics, 57(2):023513, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2016:RRMa

[598] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials. III. Journal of Mathematical Physics, 57(2):023514, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cattaneo:2016:BBA

[599] Alberto S. Cattaneo and Michele Schiavina. BV–BFV approach to general relativity: Einstein–Hilbert action. Journal of Mathematical Physics, 57(2):023515, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Artebani:2016:GAL

[600] M. Artebani, R. Caroca, M. C. Ipinza, D. M. Peñafiel, and P. Salgado. Geometrical aspects of the Lie algebra S-expansion procedure. Journal of Mathematical Physics, 57(2):023516, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forger:2016:CDA

[601] Michael Forger and Daniel V. Paulino. C^{∗}-completions and the DFRalgebra. Journal of Mathematical Physics, 57(2):023517, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2016:CTD

[602] Xiaorui Hu, Yuqi Li, and Yong Chen. Constructing two-dimensional optimal system of the group invariant solutions. Journal of Mathematical Physics, 57(2):023518, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tamura:2016:DSU

[603] Hiroshi Tamura and Valentin A. Zagrebnov. Dynamical semigroup for unbounded repeated perturbation of an open system. Journal of Mathematical Physics, 57(2):023519, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lechtenfeld:2016:SLY

[604] Olaf Lechtenfeld and Alexander D. Popov. Supermembrane limit of Yang–Mills theory. Journal of Mathematical Physics, 57(2):023520, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2016:AOJ

[605] Axel Schulze-Halberg. Arbitrary-order Jordan chains associated with quantum-mechanical Hamiltonians: Representations and integral formulas. Journal of Mathematical Physics, 57(2):023521, February 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aouadi:2016:DOP

[606] O. Aouadi, Y. Chargui, and M. S. Fayache. The Dirac oscillator in the presence of a chain of delta-function potentials. Journal of Mathematical Physics, 57(2):023522, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paliathanasis:2016:CCL

[607] A. Paliathanasis and P. G. L. Leach. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)] and its erratum [j. math. phys. 55, 059901 (2014)]. Journal of Mathematical Physics, 57(2):024101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13, 19].

Ghoumaid:2016:CAS

[608] A. Ghoumaid, F. Benamira, and L. Guechi. Comment on “Approximate solutions of the Dirac equation for the Rosen–Morse potential including the spin-orbit centrifugal term” [j. math. phys. 51, 023525 (2010)]. Journal of Mathematical Physics, 57(2):024102, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [7].

Fredenhagen:2016:QFT

[609] Klaus Fredenhagen and Kasia Rejzner. Quantum field theory on curved spacetimes: Axiomatic framework and examples. Journal of Mathematical Physics, 57(3):031101, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2016:DSS

[610] Shihui Zhu. On the Davey–Stewartson system with competing nonlinearities. Journal of Mathematical Physics, 57(3):031501, March 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Du:2016:EAB

[611] Miao Du, Lixin Tian, Jun Wang, and Fubao Zhang. Existence and asymptotic behavior of solutions for nonlinear Schrödinger–Poisson systems with steep potential well. Journal of Mathematical Physics, 57 (3):031502, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deng:2016:ESQ

[612] Yinbin Deng, Yuxia Guo, and Jiaquan Liu. Existence of solutions for quasilinear elliptic equations with Hardy potential. Journal of Mathematical Physics, 57(3):031503, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roche:2016:GZF

[613] Ph. Roche. Generalized zeta function representation of groups and 2dimensional topological Yang–Mills theory: The example of GL(2, F_{q}) and PGL(2, F_{q}). Journal of Mathematical Physics, 57(3):031701, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cox:2016:PVA

[614] Ben Cox, Elizabeth Jurisich, and Renato A. Martins. The 3-point Virasoro algebra and its action on a Fock space. Journal of Mathematical Physics, 57(3):031702, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parisio:2016:CSO

[615] Fernando Parisio. Coherent-state overcompleteness, path integrals, and weak values. Journal of Mathematical Physics, 57(3):032101, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2016:AST

[616] Zhi Xiao and Hai Huang. Analytic solutions of tunneling time through smooth barriers. Journal of Mathematical Physics, 57(3):032102, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lutfuoglu:2016:SBQ

[617] B. C. Lütfüoglu, F. Akdeniz, and O. Bayrak. Scattering, bound, and quasi-bound states of the generalized symmetric Woods–Saxon potential. Journal of Mathematical Physics, 57(3):032103, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harrell:2016:WCT

[618] L. E. Harrell. Wavefunction collapse through backaction of counting weakly interacting photons. Journal of Mathematical Physics, 57(3): 032104, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brown:2016:CQR

[619] J. M. Brown, P. Jakobsen, A. Bahl, J. V. Moloney, and M. Kolesik. On the convergence of quantum resonant-state expansion. Journal of Mathematical Physics, 57(3):032105, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Menguturk:2016:SSE

[620] Levent Ali Mengütürk. Stochastic Schrödinger evolution over piecewise enlarged filtrations. Journal of Mathematical Physics, 57(3):032106, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lami:2016:ESC

[621] L. Lami and V. Giovannetti. Entanglement-saving channels. Journal of Mathematical Physics, 57(3):032201, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McNulty:2016:MUP

[622] Daniel McNulty, Bogdan Pammer, and Stefan Weigert. Mutually unbiased product bases for multiple qudits. Journal of Mathematical Physics, 57(3):032202, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirokov:2016:SEI

[623] M. E. Shirokov. Squashed entanglement in infinite dimensions. Journal of Mathematical Physics, 57(3):032203, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adamczak:2016:AEU

[624] Radosław Adamczak, Rafał Latała, Zbigniew Puchała, and Karol Życzkowski. Asymptotic entropic uncertainty relations. Journal of Mathematical Physics, 57(3):032204, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frenkel:2016:EEG

[625] Péter E. Frenkel. On endomorphisms of the Einstein gyrogroup in arbitrary dimension. Journal of Mathematical Physics, 57(3):032301, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

PerezR:2016:RIC

[626] H. Pérez R., P. Kielanowski, and S. R. Juárez W. Rephasing invariants of the Cabibbo–Kobayashi–Maskawa matrix. Journal of Mathematical Physics, 57(3):032302, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daVeiga:2016:OBS

[627] Paulo A. Faria da Veiga, Michael O’Carroll, and José C. Valencia Alvites. One-baryon spectrum and analytical properties of one-baryon dispersion curves in 3 + 1 dimensional strongly coupled lattice QCD with three flavors. Journal of Mathematical Physics, 57(3):032303, March 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2016:ISA

[628] Luise Adams, Christian Bogner, and Stefan Weinzierl. The iterated structure of the all-order result for the two-loop sunrise integral. Journal of Mathematical Physics, 57(3):032304, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sert:2016:RBH

[629] Özcan Sert. Regular black hole solutions of the non-minimally coupled Y (R)F^{2} gravity. Journal of Mathematical Physics, 57(3):032501, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alaee:2016:RMA

[630] Aghil Alaee and Hari K. Kunduri. Remarks on mass and angular momenta for U(1)^{2}-invariant initial data. Journal of Mathematical Physics, 57(3):032502, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delabays:2016:MPL

[631] Robin Delabays, Tommaso Coletta, and Philippe Jacquod. Multistability of phase-locking and topological winding numbers in locally coupled Kuramoto models on single-loop networks. Journal of Mathematical Physics, 57(3):032701, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:2016:SEM

[632] V. V. Nesterenko and A. V. Nesterenko. Symmetric energy-momentum tensor: The Abraham form and the explicitly covariant formula. Journal of Mathematical Physics, 57(3):032901, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chandre:2016:HRO

[633] C. Chandre and M. Perin. Hamiltonian reductions of the one-dimensional Vlasov equation using phase-space moments. Journal of Mathematical Physics, 57(3):032902, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baez:2016:CFM

[634] John C. Baez, Brendan Fong, and Blake S. Pollard. A compositional framework for Markov processes. Journal of Mathematical Physics, 57 (3):033301, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konenberg:2016:IDE

[635] M. Könenberg and M. Merkli. On the irreversible dynamics emerging from quantum resonances. Journal of Mathematical Physics, 57(3): 033302, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fanelli:2016:TDF

[636] Cristiano Fanelli, Francesco Sisti, and Gabriele V. Stagno. Time dependent friction in a free gas. Journal of Mathematical Physics, 57(3): 033501, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galapon:2016:CPV

[637] Eric A. Galapon. The Cauchy principal value and the Hadamard finite part integral as values of absolutely convergent integrals. Journal of Mathematical Physics, 57(3):033502, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1017].

Chanyal:2016:DOE

[638] B. C. Chanyal. Dual octonion electrodynamics with the massive field of dyons. Journal of Mathematical Physics, 57(3):033503, March 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Francois:2016:BRS

[639] J. François, S. Lazzarini, and T. Masson. Becchi–Rouet–Stora–Tyutin structure for the mixed Weyl-diffeomorphism residual symmetry. Journal of Mathematical Physics, 57(3):033504, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magdziarz:2016:CFP

[640] Marcin Magdziarz and Tomasz Zorawik. Comment on “Fokker–Planck equations for nonlinear dynamical systems driven by non-Gaussian Lévy processes” [J. Math. Phys. 53, 072701 (2012)]. Journal of Mathematical Physics, 57(3):034101, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [10].

Frewer:2016:CAE

[641] Michael Frewer, George Khujadze, and Holger Foysi. Comment on “Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation” [J. Math. Phys. 54, 072901 (2013)]. Journal of Mathematical Physics, 57(3):034102, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [15] and response [642].

Waclawczyk:2016:RCA

[642] Marta Waclawczyk and Martin Oberlack. Response to “Comment on ‘Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation”’ [J. Math. Phys. 57, 034102 (2016)]. Journal of Mathematical Physics, 57(3):034103, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [15, 641].

Burczak:2016:GBL

[643] Jan Burczak, Rafael Granero-Belinchón, and Garving K. Luli. On the generalized Buckley–Leverett equation. Journal of Mathematical Physics, 57(4):041501, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Talati:2016:TCI

[644] Daryoush Talati and Refik Turhan. Two-component integrable generalizations of Burgers equations with nondiagonal linearity. Journal of Mathematical Physics, 57(4):041502, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:2016:EGS

[645] Yun-Sung Chung and Kyungkeun Kang. Existence of global solutions for a chemotaxis-fluid system with nonlinear diffusion. Journal of Mathematical Physics, 57(4):041503, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:STW

[646] Yan Li, Wan-Tong Li, and Yun-Rui Yang. Stability of traveling waves of a diffusive susceptible-infective-removed (SIR) epidemic model. Journal of Mathematical Physics, 57(4):041504, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tang:2016:GSS

[647] Xianhua Tang and Dongdong Qin. Ground state solutions for semilinear time-harmonic Maxwell equations. Journal of Mathematical Physics, 57 (4):041505, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2016:PSQ

[648] Haidong Liu. Positive solution for a quasilinear elliptic equation involving critical or supercritical exponent. Journal of Mathematical Physics, 57 (4):041506, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2016:GIS

[649] Pavel Exner and Jonathan Rohleder. Generalized interactions supported on hypersurfaces. Journal of Mathematical Physics, 57(4):041507, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2016:IPD

[650] N. Aizawa, Z. Kuznetsova, and F. Toppan. Invariant partial differential equations with two-dimensional exotic centrally extended conformal Galilei symmetry. Journal of Mathematical Physics, 57(4):041701, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nigro:2016:VAR

[651] Alessandro Nigro. A q-Virasoro algebra at roots of unity, free fermions, and Temperley–Lieb Hamiltonians. Journal of Mathematical Physics, 57(4):041702, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guay:2016:TYS

[652] Nicolas Guay, Vidas Regelskis, and Curtis Wendlandt. Twisted Yangians of small rank. Journal of Mathematical Physics, 57(4):041703, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frohlich:2016:PSQ

[653] Jürg Fröhlich and Baptiste Schubnel. The preparation of states in quantum mechanics. Journal of Mathematical Physics, 57(4):042101, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samar:2016:ESP

[654] M. I. Samar and V. M. Tkachuk. Exactly solvable problems in the momentum space with a minimum uncertainty in position. Journal of Mathematical Physics, 57(4):042102, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fedoruk:2016:BHB

[655] Sergey Fedoruk and Andrei Smilga. Bi-HKT and bi–Kähler supersymmetric sigma models. Journal of Mathematical Physics, 57(4):042103, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sandoval-Santana:2016:TET

[656] J. C. Sandoval-Santana, V. G. Ibarra-Sierra, J. L. Cardoso, and A. Kunold. Time evolution of two-dimensional quadratic Hamiltonians: A Lie algebraic approach. Journal of Mathematical Physics, 57(4):

042104, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rademacher:2016:ARB

[657] Simone Rademacher and Heinz Siedentop. Accumulation rate of bound states of dipoles in graphene. Journal of Mathematical Physics, 57(4): 042105, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:2016:EFH

[658] Jun Suzuki. Explicit formula for the Holevo bound for two-parameter qubit-state estimation problem. Journal of Mathematical Physics, 57(4): 042201, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lawrence:2016:VMC

[659] Albion Lawrence and Masoud Soroush. N = (4, 4) vector multiplets on curved two-manifolds. Journal of Mathematical Physics, 57(4):042301, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Futakuchi:2016:TOE

[660] Shinichiro Futakuchi and Kouta Usui. Time-ordered exponential on the complex plane and Gell-Mann–Low formula as a mathematical theorem. Journal of Mathematical Physics, 57(4):042302, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haberichter:2016:TDF

[661] Mareike Haberichter, Richard MacKenzie, M. B. Paranjape, and Yvan Ung. Tunneling decay of false domain walls: The silence of the lambs. Journal of Mathematical Physics, 57(4):042303, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Czubak:2016:WPR

[662] Magdalena Czubak and Marcelo M. Disconzi. On the well-posedness of relativistic viscous fluids with non-zero vorticity. Journal of Mathematical Physics, 57(4):042501, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ertem:2016:SOK

[663] Ümit Ertem. Symmetry operators of Killing spinors and superalgebras in AdS_{5}. Journal of Mathematical Physics, 57(4):042502, April 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andrade:2016:DRT

[664] J. Andrade, S. Boatto, and C. Vidal. Dynamics of restricted three and four vortices problem on the plane. Journal of Mathematical Physics, 57(4):042901, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2016:OPG

[665] Alexander Figotin and Aaron Welters. On overdamping phenomena in gyroscopic systems composed of high-loss and lossless components. Journal of Mathematical Physics, 57(4):042902, April 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:2016:SME

[666] Dionisios Margetis and Mitchell Luskin. On solutions of Maxwell’s equations with dipole sources over a thin conducting film. Journal of Mathematical Physics, 57(4):042903, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnik:2016:RPS

[667] W. I. Skrypnik. On regular polygon solutions of Coulomb equation of motion of n + 2 charges n of which are planar. Journal of Mathematical Physics, 57(4):042904, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Webb:2016:HPT

[668] Stephen D. Webb. A Hamiltonian perturbation theory for the nonlinear Vlasov equation. Journal of Mathematical Physics, 57(4):042905, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolte:2016:IBE

[669] Jens Bolte and Joachim Kerner. Instability of Bose–Einstein condensation into the one-particle ground state on quantum graphs under repulsive perturbations. Journal of Mathematical Physics, 57(4):043301, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lei:2016:CSP

[670] Qiang Lei, Xiaochao Su, and Junde Wu. Continuity of the sequential product of sequential quantum effect algebras. Journal of Mathematical Physics, 57(4):043501, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

G:2016:PDR

[671] Ramesh G and Santhosh Kumar P. On the polar decomposition of right linear operators in quaternionic Hilbert spaces. Journal of Mathematical Physics, 57(4):043502, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

G:2016:WNB

[672] Ramesh G. Weyl–von Neumann–Berg theorem for quaternionic operators. Journal of Mathematical Physics, 57(4):043503, April 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2016:APB

[673] F. Bagarello. Appearances of pseudo-bosons from Black–Scholes equation. Journal of Mathematical Physics, 57(4):043504, April 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaudreau:2016:DES

[674] P. Gaudreau, R. Slevinsky, and H. Safouhi. The double exponential sinc collocation method for singular Sturm–Liouville problems. Journal of Mathematical Physics, 57(4):043505, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grammaticos:2016:SAD

[675] B. Grammaticos, A. Ramani, K. M. Tamizhmani, T. Tamizhmani, and J. Satsuma. Strongly asymmetric discrete Painlevé equations: The multiplicative case. Journal of Mathematical Physics, 57(4):043506, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cimmelli:2016:NMT

[676] V. A. Cimmelli, P. Rogolino, and A. Sellitto. A nonlinear model of thermoelectricity with two temperatures: Application to quasicrystalline nanowires. Journal of Mathematical Physics, 57(4):043507, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2016:HOR

[677] Li-Chen Zhao, Boling Guo, and Liming Ling. High-order rogue wave solutions for the coupled nonlinear Schrödinger equations — II. Journal of Mathematical Physics, 57(4):043508, April 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantica:2016:ECP

[678] Carlo Alberto Mantica, Luca Guido Molinari, and Uday Chand De. Erratum: “A condition for a perfect-fluid space–time to be a generalized Robertson–Walker space–time” [j. math. phys. 57, 022508 (2016)]. Journal of Mathematical Physics, 57(4):049901, April 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [571].

Naumkin:2016:SAB

[679] I. P. Naumkin. Sharp asymptotic behavior of solutions for cubic nonlinear Schrödinger equations with a potential. Journal of Mathematical Physics, 57(5):051501, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ambrosio:2016:GSS

[680] Vincenzo Ambrosio. Ground states solutions for a non-linear equation involving a pseudo-relativistic Schrödinger operator. Journal of Mathematical Physics, 57(5):051502, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khmelnytskaya:2016:MEF

[681] Kira V. Khmelnytskaya, Vladislav V. Kravchenko, and Sergii M. Torba. Modulated electromagnetic fields in inhomogeneous media, hyperbolic pseudoanalytic functions, and transmutations. Journal of Mathematical Physics, 57(5):051503, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2016:OTD

[682] Jihong Zhao. The optimal temporal decay estimates for the fractional power dissipative equation in negative Besov spaces. Journal of Mathematical Physics, 57(5):051504, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cantarini:2016:CLC

[683] Nicoletta Cantarini and Victor G. Kac. Classification of linearly compact simple Nambu–Poisson algebras. Journal of Mathematical Physics, 57(5): 051701, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poletaeva:2016:PFA

[684] Elena Poletaeva. On principal finite W -algebras for the Lie superalgebra D(2, 1; α). Journal of Mathematical Physics, 57(5):051702, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:RMR

[685] Chih-Whi Chen. Reduction method for representations of queer Lie superalgebras. Journal of Mathematical Physics, 57(5):051703, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meljanac:2016:WRL

[686] Stjepan Meljanac, Sasa Kresić-Jurić, and Tea Martinić. The Weyl realizations of Lie algebras, and left-right duality. Journal of Mathematical Physics, 57(5):051704, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zheng:2016:SGE

[687] Quan Zheng. Shifted genus expanded W_{∞} algebra and shifted Hurwitz numbers. Journal of Mathematical Physics, 57(5):051705, May 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mudrov:2016:MIS

[688] Andrey Mudrov. R-matrix and inverse Shapovalov form. Journal of Mathematical Physics, 57(5):051706, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kadets:2016:CQG

[689] Boris Kadets, Eugene Karolinsky, Iulia Pop, and Alexander Stolin. Classification of quantum groups and Belavin–Drinfeld cohomologies for orthogonal and symplectic Lie algebras. Journal of Mathematical Physics, 57(5):051707, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2016:SQS

[690] Hoil Kim. Symmetry of quantum supertorus. Journal of Mathematical Physics, 57(5):051708, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marquette:2016:CBQ

[691] Ian Marquette and Christiane Quesne. Connection between quantum systems involving the fourth Painlevé transcendent and k-step rational extensions of the harmonic oscillator related to Hermite exceptional orthogonal polynomial. Journal of Mathematical Physics, 57(5):052101, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gazeau:2016:CAI

[692] Jean Pierre Gazeau and Romain Murenzi. Covariant affine integral quantization(s). Journal of Mathematical Physics, 57(5):052102, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neate:2016:SSM

[693] Andrew Neate and Aubrey Truman. Semiclassical stochastic mechanics for the Coulomb potential with applications to modelling dark matter. Journal of Mathematical Physics, 57(5):052103, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:SCS

[694] Joe P. Chen and Alexander Teplyaev. Singularly continuous spectrum of a self-similar Laplacian on the half-line. Journal of Mathematical Physics, 57(5):052104, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spee:2016:MES

[695] C. Spee, J. I. de Vicente, and B. Kraus. The maximally entangled set of 4-qubit states. Journal of Mathematical Physics, 57(5):052201, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Audenaert:2016:EPI

[696] Koenraad Audenaert, Nilanjana Datta, and Maris Ozols. Entropy power inequalities for qudits. Journal of Mathematical Physics, 57(5):052202, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [724].

Datta:2016:UBS

[697] Nilanjana Datta, Min-Hsiu Hsieh, and Jonathan Oppenheim. An upper bound on the second order asymptotic expansion for the quantum communication cost of state redistribution. Journal of Mathematical Physics, 57(5):052203, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zucchini:2016:LBD

[698] Roberto Zucchini. A Lie based 4-dimensional higher Chern–Simons theory. Journal of Mathematical Physics, 57(5):052301, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz:2016:GDI

[699] Pablo Diaz, Hai Lin, and Alvaro Veliz-Osorio. Graph duality as an instrument of gauge–string correspondence. Journal of Mathematical Physics, 57(5):052302, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyaura:2016:FTS

[700] Ryosuke Miyaura and Atsushi Mukaihira. Factorial-type Schur functions, orthogonal rational functions, and discrete dressing chains. Journal of Mathematical Physics, 57(5):052701, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rashid:2016:ENU

[701] A. A. Rashid and A. Zamani Bahabadi. Ergodicity of non-uniformly expanding transitive group (or semigroup) actions. Journal of Mathematical Physics, 57(5):052702, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2016:BHS

[702] Manuel F. Rañada. Bi-Hamiltonian structure of the bi-dimensional superintegrable nonlinear isotonic oscillator. Journal of Mathematical Physics, 57(5):052703, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rodrigues:2016:STP

[703] Fagner B. Rodrigues and Paulo Varandas. Specification and thermodynamical properties of semigroup actions. Journal of Mathematical Physics, 57(5):052704, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maltsev:2016:CFM

[704] A. Ya. Maltsev. On the canonical forms of the multi-dimensional averaged Poisson brackets. Journal of Mathematical Physics, 57(5):053501, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benini:2016:OSL

[705] Marco Benini. Optimal space of linear classical observables for Maxwell k-forms via spacelike and timelike compact de Rham cohomologies. Journal of Mathematical Physics, 57(5):053502, May 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:SDK

[706] Chun-Xia Li, Stéphane Lafortune, and Shou-Feng Shen. A semi-discrete Kadomtsev–Petviashvili equation and its coupled integrable system. Journal of Mathematical Physics, 57(5):053503, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cresson:2016:MFS

[707] Jacky Cresson and Frédéric Pierret. Multiscale functions, scale dynamics, and applications to partial differential equations. Journal of Mathematical Physics, 57(5):053504, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sechin:2016:AYB

[708] Ivan Sechin and Andrei Zotov. Associative Yang–Baxter equation for quantum (semi-)dynamical R-matrices. Journal of Mathematical Physics, 57(5):053505, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeNittis:2016:DGI

[709] Giuseppe De Nittis and Kiyonori Gomi. Differential geometric invariants for time-reversal symmetric Bloch-bundles: The “real” case. Journal of Mathematical Physics, 57(5):053506, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Troessaert:2016:HSC

[710] Cédric Troessaert. Hamiltonian surface charges using external sources. Journal of Mathematical Physics, 57(5):053507, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2016:LSB

[711] Lamei Yuan and Caixia He. Lie super-bialgebra and quantization of the super Virasoro algebra. Journal of Mathematical Physics, 57(5): 053508, May 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pang:2016:GST

[712] Yicheng Pang, Shaohong Cai, and Yuanying Zhao. Global solutions to the two-dimensional Riemann problem for a system of conservation laws. Journal of Mathematical Physics, 57(6):061501, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mejri:2016:SEA

[713] Youssef Mejri. Stability estimate for the aligned magnetic field in a periodic quantum waveguide from Dirichlet-to-Neumann map. Journal of Mathematical Physics, 57(6):061502, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bakalov:2016:TLM

[714] Bojko Bakalov and McKay Sullivan. Twisted logarithmic modules of free field algebras. Journal of Mathematical Physics, 57(6):061701, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Urichuk:2016:AAF

[715] Andrew Urichuk and Mark A. Walton. Adjoint affine fusion and tadpoles. Journal of Mathematical Physics, 57(6):061702, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferialdi:2016:ENM

[716] L. Ferialdi and D. Dürr. Effective non-Markovian description of a system interacting with a bath. Journal of Mathematical Physics, 57(6):062101, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikolaev:2016:KEQ

[717] Andrey Nikolaev. Kato expansion in quantum canonical perturbation theory. Journal of Mathematical Physics, 57(6):062102, June 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tosiek:2016:WKB

[718] J. Tosiek, R. Cordero, and F. J. Turrubiates. The Wentzel–Kramers– Brillouin approximation method applied to the Wigner function. Journal of Mathematical Physics, 57(6):062103, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ojeda-Guillen:2016:GJC

[719] D. Ojeda-Guillén, R. D. Mota, and V. D. Granados. A generalized Jaynes–Cummings model: The relativistic parametric amplifier and a single trapped ion. Journal of Mathematical Physics, 57(6):062104, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amir:2016:ASS

[720] Naila Amir and Shahid Iqbal. Algebraic solutions of shape-invariant position-dependent effective mass systems. Journal of Mathematical Physics, 57(6):062105, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yadav:2016:PSE

[721] Rajesh Kumar Yadav, Avinash Khare, Bijan Bagchi, Nisha Kumari, and Bhabani Prasad Mandal. Parametric symmetries in exactly solvable real and PT symmetric complex potentials. Journal of Mathematical Physics, 57(6):062106, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2016:FTF

[722] Qing Gao, Daoyi Dong, Ian R. Petersen, and Herschel Rabitz. Fault tolerant filtering and fault detection for quantum systems driven by fields in single photon states. Journal of Mathematical Physics, 57(6):062201, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Somma:2016:TSA

[723] Rolando D. Somma. A Trotter–Suzuki approximation for Lie groups with applications to Hamiltonian simulation. Journal of Mathematical Physics, 57(6):062202, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carlen:2016:QEP

[724] Eric A. Carlen, Elliott H. Lieb, and Michael Loss. On a quantum entropy power inequality of Audenaert, Datta, and Ozols. Journal of Mathematical Physics, 57(6):062203, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosati:2016:AHB

[725] Matteo Rosati and Vittorio Giovannetti. Achieving the Holevo bound via a bisection decoding protocol. Journal of Mathematical Physics, 57 (6):062204, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:2016:LUE

[726] Naihuan Jing, Min Yang, and Hui Zhao. Local unitary equivalence of quantum states and simultaneous orthogonal equivalence. Journal of Mathematical Physics, 57(6):062205, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2016:QMF

[727] Shawn X. Cui, Michael H. Freedman, Or Sattath, Richard Stong, and Greg Minton. Quantum max-flow/min-cut. Journal of Mathematical Physics, 57(6):062206, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Datta:2016:SOA

[728] Nilanjana Datta, Yan Pautrat, and Cambyse Rouzé. Second-order asymptotics for quantum hypothesis testing in settings beyond i.i.d.- quantum lattice systems and more. Journal of Mathematical Physics, 57(6):062207, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benatti:2016:DEQ

[729] F. Benatti, F. Carollo, and R. Floreanini. Dissipative entanglement of quantum spin fluctuations. Journal of Mathematical Physics, 57(6): 062208, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contreras:2016:GMS

[730] I. Contreras, E. Ercolessi, and M. Schiavina. On the geometry of mixed states and the Fisher information tensor. Journal of Mathematical Physics, 57(6):062209, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Challifour:2016:CQL

[731] John L. Challifour and Edward J. Timko. Canonical quantization of lattice Higgs–Yang–Mills fields: Krein essential selfadjointness of the Hamiltonian. Journal of Mathematical Physics, 57(6):062301, June 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Mallick:2016:SPP

[732] Sahanous Mallick, Young Jin Suh, and Uday Chand De. A spacetime with pseudo-projective curvature tensor. Journal of Mathematical Physics, 57(6):062501, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diacu:2016:CTH

[733] Florin Diacu, Slim Ibrahim, and Jedrzej Śniatycki. The continuous transition of Hamiltonian vector fields through manifolds of constant curvature. Journal of Mathematical Physics, 57(6):062701, June 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigorescu:2016:CDR

[734] Marius Grigorescu. Coherent distributions for the rigid rotator. Journal of Mathematical Physics, 57(6):062702, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2016:SNA

[735] Stephen C. Anco, Tyler Meadows, and Vincent Pascuzzi. Some new aspects of first integrals and symmetries for central force dynamics. Journal of Mathematical Physics, 57(6):062901, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beyer:2016:CNW

[736] Horst Reinhard Beyer, Burak Aksoylu, and Fatih Celiker. On a class of nonlocal wave equations from applications. Journal of Mathematical Physics, 57(6):062902, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cruz:2016:HAL

[737] Miguel Cruz, Rosario Gómez-Cortés, Alberto Molgado, and Efraín Rojas. Hamiltonian analysis for linearly acceleration-dependent Lagrangians. Journal of Mathematical Physics, 57(6):062903, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:2016:CDW

[738] Silas L. Carvalho and César R. de Oliveira. Correlation dimension Wonderland theorems. Journal of Mathematical Physics, 57(6):063501, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gungor:2016:VCD

[739] F. Güngör and C. Özemir. Variable coefficient Davey–Stewartson system with a Kac–Moody–Virasoro symmetry algebra. Journal of Mathematical Physics, 57(6):063502, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cirilo-Lombardo:2016:RDG

[740] Diego Julio Cirilo-Lombardo. Relativistic dynamics, Green function and pseudodifferential operators. Journal of Mathematical Physics, 57(6): 063503, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeVecchi:2016:SSD

[741] Francesco C. De Vecchi, Paola Morando, and Stefania Ugolini. Symmetries of stochastic differential equations: A geometric approach. Journal of Mathematical Physics, 57(6):063504, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaillard:2016:FWR

[742] Pierre Gaillard. Fredholm and Wronskian representations of solutions to the KPI equation and multi-rogue waves. Journal of Mathematical Physics, 57(6):063505, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takahashi:2016:QSP

[743] Yuki Takahashi. Quantum and spectral properties of the labyrinth model. Journal of Mathematical Physics, 57(6):063506, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hryniv:2016:EAD

[744] Rostyslav O. Hryniv and Yaroslav V. Mykytyuk. Eigenvalue asymptotics for Dirac–Bessel operators. Journal of Mathematical Physics, 57 (6):063507, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campoamor-Stursberg:2016:FRP

[745] R. Campoamor-Stursberg. A functional realization of sl(3, R) providing minimal Vessiot–Guldberg–Lie algebras of nonlinear second-order ordinary differential equations as proper subalgebras. Journal of Mathematical Physics, 57(6):063508, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stottmeister:2016:CSQa

[746] Alexander Stottmeister and Thomas Thiemann. Coherent states, quantum gravity, and the Born–Oppenheimer approximation. I. General considerations. Journal of Mathematical Physics, 57(6):063509, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goffeng:2016:CEM

[747] Magnus Goffeng, Ayman Kachmar, and Mikael Persson Sundqvist. Clusters of eigenvalues for the magnetic Laplacian with Robin condition. Journal of Mathematical Physics, 57(6):063510, June 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2016:LTB

[748] Young-Pil Choi. Large-time behavior for the Vlasov/compressible Navier–Stokes equations. Journal of Mathematical Physics, 57(7):071501, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ueki:2016:WEL

[749] Naomasa Ueki. Wegner estimates, Lifshitz tails, and Anderson localization for Gaussian random magnetic fields. Journal of Mathematical Physics, 57(7):071502, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2016:IVP

[750] A. Alexandrou Himonas and Dionyssios Mantzavinos. The initial value problem for a Novikov system. Journal of Mathematical Physics, 57(7): 071503, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:SSL

[751] Jianjun Zhang, David G. Costa, and João Marcos do Ó. Semiclassical states of p-Laplacian equations with a general nonlinearity in critical case. Journal of Mathematical Physics, 57(7):071504, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:IBT

[752] Rui Li, Xing Lin, and Zongwei Ma. Improved blowup theorems for the Euler–Poisson equations with attractive forces. Journal of Mathematical Physics, 57(7):071505, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miki:2016:AQT

[753] Kei Miki. Actions of the quantum toroidal algebra of type rmsl_{2} on the space of vertex operators for U_{q}(ĝl2) modules. Journal of Mathematical Physics, 57(7):071701, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patrascu:2016:BHI

[754] Andrei T. Patrascu. Black holes, information, and the universal coefficient theorem. Journal of Mathematical Physics, 57(7):071702, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Budinich:2016:ST

[755] Marco Budinich. On spinors transformations. Journal of Mathematical Physics, 57(7):071703, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Correggi:2016:TCS

[756] M. Correggi and D. Dimonte. On the third critical speed for rotating Bose–Einstein condensates. Journal of Mathematical Physics, 57(7): 071901, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelasCuevas:2016:FLP

[757] G. De las Cuevas, T. S. Cubitt, J. I. Cirac, M. M. Wolf, and D. Pérez-García. Fundamental limitations in the purifications of tensor networks. Journal of Mathematical Physics, 57(7):071902, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tonchev:2016:MRM

[758] N. S. Tonchev. Monotone Riemannian metrics and dynamic structure factor in condensed matter physics. Journal of Mathematical Physics, 57(7):071903, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fishman:2016:SCP

[759] S. Fishman and A. Soffer. Slowly changing potential problems in quantum mechanics: Adiabatic theorems, ergodic theorems, and scattering. Journal of Mathematical Physics, 57(7):072101, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taufer:2016:WEL

[760] Matthias Täufer and Ivan Veselić. Wegner estimate for Landau-breather Hamiltonians. Journal of Mathematical Physics, 57(7):072102, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chhaiba:2016:AGN

[761] Hassan Chhaiba, Nizar Demni, and Zouhair Mouayn. Analysis of generalized negative binomial distributions attached to hyperbolic Landau levels. Journal of Mathematical Physics, 57(7):072103, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moumni:2016:SEN

[762] M. Moumni and M. Falek. Schrödinger equation for non-pure dipole potential in 2D systems. Journal of Mathematical Physics, 57(7):072104, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celeghini:2016:ARH

[763] E. Celeghini, M. Gadella, and M. A. del Olmo. Applications of rigged Hilbert spaces in quantum mechanics and signal processing. Journal of Mathematical Physics, 57(7):072105, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deckert:2016:CMT

[764] Dirk-André Deckert and Lukas Nickel. Consistency of multi-time Dirac equations with general interaction potentials. Journal of Mathematical Physics, 57(7):072301, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zaba:2016:UBS

[765] Mariusz Zaba and Piotr Garbaczewski. Ultrarelativistic bound states in the spherical well. Journal of Mathematical Physics, 57(7):072302, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2016:FPT

[766] Felix Finster, Simone Murro, and Christian Röken. The fermionic projector in a time-dependent external potential: Mass oscillation property and Hadamard states. Journal of Mathematical Physics, 57(7):072303, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chicone:2016:NNC

[767] C. Chicone and B. Mashhoon. Nonlocal Newtonian cosmology. Journal of Mathematical Physics, 57(7):072501, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2016:NIB

[768] F. Calogero. Novel isochronous N-body problems featuring N arbitrary rational coupling constants. Journal of Mathematical Physics, 57(7): 072901, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alberti:2016:FKS

[769] Angelo Alberti and Claudio Vidal. First kind symmetric periodic solutions of the generalized van der Waals Hamiltonian. Journal of Mathematical Physics, 57(7):072902, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lukacs:2016:HPB

[770] Árpád Lukács. “Half a proton” in the Bogomolńyi–Prasad–Sommerfield Skyrme model. Journal of Mathematical Physics, 57(7):072903, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Machida:2016:RTE

[771] Manabu Machida. The radiative transport equation in flatland with separation of variables. Journal of Mathematical Physics, 57(7):073301, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:LDP

[772] Yong Chen, Hao Ge, Jie Xiong, and Lihu Xu. The large deviation principle and steady-state fluctuation theorem for the entropy production rate of a stochastic process in magnetic fields. Journal of Mathematical Physics, 57(7):073302, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjornberg:2016:FEC

[773] J. E. Björnberg. The free energy in a class of quantum spin systems and interchange processes. Journal of Mathematical Physics, 57(7):073303, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stottmeister:2016:CSQb

[774] Alexander Stottmeister and Thomas Thiemann. Coherent states, quantum gravity, and the Born–Oppenheimer approximation. II. Compact Lie groups. Journal of Mathematical Physics, 57(7):073501, July 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Franchetti:2016:MIH

[775] Guido Franchetti and Rafael Maldonado. Monopoles, instantons, and the Helmholtz equation. Journal of Mathematical Physics, 57(7):073502, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Covolo:2016:CS

[776] Tiffany Covolo, Janusz Grabowski, and Norbert Poncin. The category of P Z2n-supermanifolds. Journal of Mathematical Physics, 57(7):073503, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braga:2016:MAA

[777] Gastão A. Braga and William R. P. Conti. A multiscale asymptotic analysis of time evolution equations on the complex plane. Journal of Mathematical Physics, 57(7):073504, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2016:RHA

[778] Zhen Wang and Zhijun Qiao. Riemann–Hilbert approach for the FQXL model: A generalized Camassa–Holm equation with cubic and quadratic nonlinearity. Journal of Mathematical Physics, 57(7):073505, July 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2016:CNF

[779] Claudianor O. Alves and Olimpio H. Miyagaki. A critical nonlinear fractional elliptic equation with saddle-like potential in R^{N}. Journal of Mathematical Physics, 57(8):081501, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2016:NVR

[780] Yanqiu Guo, Irma Hacinliyan, and Edriss S. Titi. Non-viscous regularization of the Davey–Stewartson equations: Analysis and modulation theory. Journal of Mathematical Physics, 57(8):081502, August 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2016:RIF

[781] Tarek Saanouni. Remarks on the inhomogeneous fractional nonlinear Schrödinger equation. Journal of Mathematical Physics, 57(8):081503, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:ESS

[782] Kui Li and Zhitao Zhang. Existence of solutions for a Schrödinger system with linear and nonlinear couplings. Journal of Mathematical Physics, 57(8):081504, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dirr:2016:EGF

[783] Nicolas Dirr, Marios Stamatakis, and Johannes Zimmer. Entropic and gradient flow formulations for nonlinear diffusion. Journal of Mathematical Physics, 57(8):081505, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christodoulou:2016:FSE

[784] Demetrios Christodoulou and Daniel Raoul Perez. On the formation of shocks of electromagnetic plane waves in non-linear crystals. Journal of Mathematical Physics, 57(8):081506, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zanelli:2016:SSE

[785] Lorenzo Zanelli. Schrödinger spectra and the effective Hamiltonian of weak KAM theory on the flat torus. Journal of Mathematical Physics, 57(8):081507, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:GCS

[786] Xiuting Li and Xianwen Zhang. On global classical solutions of the three dimensional relativistic Vlasov–Darwin system. Journal of Mathematical Physics, 57(8):081508, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:DPP

[787] Xiaojun Li. Dynamics of parabolic problems with memory. Subcritical and critical nonlinearities. Journal of Mathematical Physics, 57(8):

081509, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foias:2016:ENS

[788] Ciprian Foias, Jing Tian, and Bingsheng Zhang. On the emergence of the Navier–Stokes–α model for turbulent channel flows. Journal of Mathematical Physics, 57(8):081510, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huh:2016:LGS

[789] Hyungjin Huh and Guanghui Jin. Local and global solutions of Chern– Simons gauged O(3) sigma equations in one space dimension. Journal of Mathematical Physics, 57(8):081511, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:2016:CPE

[790] Sa Jun Park and Seok-Bae Yun. Cauchy problem for the ellipsoidal-BGK model of the Boltzmann equation. Journal of Mathematical Physics, 57 (8):081512, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheung:2016:EUS

[791] Ka Luen Cheung and Anthony Suen. Existence and uniqueness of small energy weak solution to multi-dimensional compressible Navier–Stokes equations with large external potential force. Journal of Mathematical Physics, 57(8):081513, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2016:WPC

[792] Junxiong Jia, Jigen Peng, and Jinghuai Gao. Well-posedness for compressible MHD systems with highly oscillating initial data. Journal of Mathematical Physics, 57(8):081514, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:GSN

[793] Shuai Li, Jianlin Xiang, and Xiaoyu Zeng. Ground states of nonlinear Choquard equations with multi-well potentials. Journal of Mathematical Physics, 57(8):081515, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2016:GEE

[794] Fei Liang, Yucong Chen, and Junbing Li. Global existence and explosion of the stochastic viscoelastic wave equation driven by multiplicative noises. Journal of Mathematical Physics, 57(8):081516, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:DVL

[795] Yinghui Zhang. Decay of the 3D viscous liquid-gas two-phase flow model with damping. Journal of Mathematical Physics, 57(8):081517, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besse:2016:ASA

[796] Nicolas Besse and David Coulette. Asymptotic and spectral analysis of the gyrokinetic-waterbag integro-differential operator in toroidal geometry. Journal of Mathematical Physics, 57(8):081518, August 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsutani:2016:EEM

[797] Shigeki Matsutani and Emma Previato. From Euler’s elastica to the mKdV hierarchy, through the Faber polynomials. Journal of Mathematical Physics, 57(8):081519, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fontelos:2016:SSD

[798] M. A. Fontelos, R. Lecaros, J. C. López-Ríos, and J. H. Ortega. Stationary shapes for 2-d water-waves and hydraulic jumps. Journal of Mathematical Physics, 57(8):081520, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2016:IRS

[799] Guangzhe Fan, Yucai Su, and Chunguang Xia. Infinite rank Schrödinger– Virasoro type Lie conformal algebras. Journal of Mathematical Physics, 57(8):081701, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szehr:2016:CCI

[800] Oleg Szehr and Michael M. Wolf. Connected components of irreducible maps and 1D quantum phases. Journal of Mathematical Physics, 57 (8):081901, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Griffin:2016:PSD

[801] Jory Griffin. On the phase-space distribution of Bloch eigenmodes for periodic point scatterers. Journal of Mathematical Physics, 57(8):082101, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioffe:2016:SSO

[802] M. V. Ioffe, E. V. Kolevatova, and D. N. Nishnianidze. Solution of second order supersymmetrical intertwining relations in Minkowski plane. Journal of Mathematical Physics, 57(8):082102, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujiwara:2016:FTB

[803] Kazumasa Fujiwara and Tohru Ozawa. Finite time blowup of solutions to the nonlinear Schrödinger equation without gauge invariance. Journal of Mathematical Physics, 57(8):082103, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:PEF

[804] Hui Zhang, Xiaoyun Jiang, and Wenping Fan. Parameter estimation for the fractional Schrödinger equation using Bayesian method. Journal of Mathematical Physics, 57(8):082104, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eshghi:2016:EDP

[805] M. Eshghi and H. Mehraban. Effective of the q-deformed pseudoscalar magnetic field on the charge carriers in graphene. Journal of Mathematical Physics, 57(8):082105, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2042].

Bountis:2016:TWS

[806] Tassos Bountis and Fernando D. Nobre. Travelling-wave and separated variable solutions of a nonlinear Schroedinger equation. Journal of Mathematical Physics, 57(8):082106, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golding:2016:QMA

[807] William M. Golding. Quantum modes of atomic waveguides by series techniques. Journal of Mathematical Physics, 57(8):082107, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samar:2016:ODC

[808] M. I. Samar and V. M. Tkachuk. One-dimensional Coulomb-like problem in general case of deformed space with minimal length. Journal of Mathematical Physics, 57(8):082108, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toranzo:2016:HLU

[809] I. V. Toranzo, A. Martínez-Finkelshtein, and J. S. Dehesa. Heisenberglike uncertainty measures for D-dimensional hydrogenic systems at large D. Journal of Mathematical Physics, 57(8):082109, August 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ligabo:2016:TPS

[810] Marilena Ligabò. Torus as phase space: Weyl quantization, dequantization, and Wigner formalism. Journal of Mathematical Physics, 57 (8):082110, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurochkin:2016:CSH

[811] Yu. Kurochkin, I. Rybak, and Dz. Shoukavy. Coherent states on horospheric three-dimensional Lobachevsky space. Journal of Mathematical Physics, 57(8):082111, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2016:DSK

[812] H. Boche and G. Janßen. Distillation of secret-key from a class of compound memoryless quantum sources. Journal of Mathematical Physics, 57(8):082201, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leditzky:2016:SCT

[813] Felix Leditzky, Mark M. Wilde, and Nilanjana Datta. Strong converse theorems using Rényi entropies. Journal of Mathematical Physics, 57 (8):082202, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigi:2016:DQI

[814] Salman Beigi, Nilanjana Datta, and Felix Leditzky. Decoding quantum information via the Petz recovery map. Journal of Mathematical Physics, 57(8):082203, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2016:QNS

[815] Timothy Nguyen. Quantization of the nonlinear sigma model revisited. Journal of Mathematical Physics, 57(8):082301, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Purkart:2016:SPG

[816] Julian Purkart. The signed permutation group on Feynman graphs. Journal of Mathematical Physics, 57(8):082302, August 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hennig:2016:GSC

[817] Jörg Hennig. Gowdy-symmetric cosmological models with Cauchy horizons ruled by non-closed null generators. Journal of Mathematical Physics, 57(8):082501, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohmann:2016:SOS

[818] Manuel Hohmann. Spacetime and observer space symmetries in the language of Cartan geometry. Journal of Mathematical Physics, 57(8): 082502, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zu:2016:NTN

[819] Qi hang Zu and Jian qing Zhu. Noether theorem for nonholonomic nonconservative mechanical systems in phase space on time scales. Journal of Mathematical Physics, 57(8):082701, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bai:2016:PMG

[820] Zhanqiang Bai. Particle motion in generalized Dirac’s monopoles of dimension 2k + 1. Journal of Mathematical Physics, 57(8):082702, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colombo:2016:RPF

[821] Leonardo Colombo and Pedro Daniel Prieto-Martínez. Regularity properties of fiber derivatives associated with higher-order mechanical systems. Journal of Mathematical Physics, 57(8):082901, August 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Feireisl:2016:MVC

[822] Eduard Feireisl, Antonín Novotný, and Yongzhong Sun. On the motion of viscous, compressible, and heat-conducting liquids. Journal of Mathematical Physics, 57(8):083101, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2016:GRM

[823] Haibo Yu. Global regularity to the 3D MHD equations with large initial data in bounded domains. Journal of Mathematical Physics, 57(8): 083102, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rabinowitch:2016:PSN

[824] Alexander S. Rabinowitch. On a particular solution to the 3D Navier– Stokes equations for liquids with cavitation. Journal of Mathematical Physics, 57(8):083103, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lukkarinen:2016:WPT

[825] Jani Lukkarinen and Matteo Marcozzi. Wick polynomials and timeevolution of cumulants. Journal of Mathematical Physics, 57(8):083301, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:LTI

[826] Wei-Kuo Chen. Limit theorems in the imitative monomer-dimer meanfield model via Stein’s method. Journal of Mathematical Physics, 57 (8):083302, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2016:ABB

[827] Ho Lee. Asymptotic behaviour of the Boltzmann equation as a cosmological model. Journal of Mathematical Physics, 57(8):083303, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nicolaescu:2016:CPM

[828] Liviu I. Nicolaescu. Critical points of multidimensional random Fourier series: Variance estimates. Journal of Mathematical Physics, 57(8): 083304, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kwong:2016:NMB

[829] Man Kam Kwong and Manwai Yuen. New method for blowup of the Euler–Poisson system. Journal of Mathematical Physics, 57(8):083501, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ji:2016:IPB

[830] Un Cig Ji and Nobuaki Obata. An implementation problem for boson fields and quantum Girsanov transform. Journal of Mathematical Physics, 57(8):083502, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2016:RBP

[831] H. Inoue and M. Takakura. Regular biorthogonal pairs and pseudobosonic operators. Journal of Mathematical Physics, 57(8):083503, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibragimov:2016:GCC

[832] N. H. Ibragimov, M. L. Gandarias, L. R. Galiakberova, M. S. Bruzon, and E. D. Avdonina. Group classification and conservation laws of anisotropic wave equations with a source. Journal of Mathematical Physics, 57(8): 083504, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2016:NSA

[833] H. Inoue and M. Takakura. Non-self-adjoint Hamiltonians defined by generalized Riesz bases. Journal of Mathematical Physics, 57(8):083505, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xue:2016:GHS

[834] Lingling Xue, Q. P. Liu, and Dengshan Wang. A generalized Hirota– Satsuma coupled KdV system: Darboux transformations and reductions. Journal of Mathematical Physics, 57(8):083506, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2016:NNS

[835] Li-Yuan Ma and Zuo-Nong Zhu. Nonlocal nonlinear Schrödinger equation and its discrete version: Soliton solutions and gauge equivalence. Journal of Mathematical Physics, 57(8):083507, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2016:NDA

[836] Francesco Calogero. Nonlinear differential algorithm to compute all the zeros of a generic polynomial. Journal of Mathematical Physics, 57(8): 083508, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [915].

Stottmeister:2016:CSQc

[837] Alexander Stottmeister and Thomas Thiemann. Coherent states, quantum gravity, and the Born–Oppenheimer approximation. III: Applications to loop quantum gravity. Journal of Mathematical Physics, 57(8): 083509, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prinari:2016:DSF

[838] Barbara Prinari. Discrete solitons of the focusing Ablowitz–Ladik equation with nonzero boundary conditions via inverse scattering. Journal of Mathematical Physics, 57(8):083510, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2016:GTR

[839] H. Inoue. General theory of regular biorthogonal pairs and its physical operators. Journal of Mathematical Physics, 57(8):083511, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hrivnak:2016:WLD

[840] Jirí Hrivnák and Mark A. Walton. Weight-lattice discretization of Weylorbit functions. Journal of Mathematical Physics, 57(8):083512, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2016:CLH

[841] Carl M. Bender, Mariagiovanna Gianfreda, Nima Hassanpour, and Hugh F. Jones. Comment on “On the Lagrangian and Hamiltonian description of the damped linear harmonic oscillator” [j. math. phys. 48, 032701 (2007)]. Journal of Mathematical Physics, 57(8):084101, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1].

Kuramochi:2016:EMS

[842] Yui Kuramochi. Erratum: “Minimal sufficient positive-operator valued measure on a separable Hilbert space” [j. math. phys. 56, 102205 (2015)].

Journal of Mathematical Physics, 57(8):089901, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [402].

Shubina:2016:PPP

[843] Maria Shubina. The 1D parabolic-parabolic Patlak–Keller–Segel model of chemotaxis: The particular integrable case and soliton solution. Journal of Mathematical Physics, 57(9):091501, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2016:EMB

[844] Weiming Liu. Existence of multi-bump solutions for the fractional Schrödinger–Poisson system. Journal of Mathematical Physics, 57(9): 091502, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Primc:2016:CBB

[845] Mirko Primc and Tomislav Sikić. Combinatorial bases of basic modules for affine Lie algebras C_{n}^{(1)}. Journal of Mathematical Physics, 57(9): 091701, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:2016:TPT

[846] Naihuan Jing and Honglian Zhang. Two-parameter twisted quantum affine algebras. Journal of Mathematical Physics, 57(9):091702, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2016:DNQ

[847] Jian-Jian Jiang and S. Hasibul Hassan Chowdhury. Deformation of noncommutative quantum mechanics. Journal of Mathematical Physics, 57 (9):091703, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fakhri:2016:MCS

[848] H. Fakhri and M. Sayyah-Fard. sl(2)-modules by sl(2)-coherent states. Journal of Mathematical Physics, 57(9):091704, September 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gosset:2016:LGT

[849] David Gosset and Evgeny Mozgunov. Local gap threshold for frustrationfree spin systems. Journal of Mathematical Physics, 57(9):091901, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2016:FGR

[850] Minjae Lee and Maciej Zworski. A Fermi golden rule for quantum graphs. Journal of Mathematical Physics, 57(9):092101, September 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Punosevac:2016:DUB

[851] Predrag Punosevac and Sam L. Robinson. Dynamics of uncertainties for bound one-dimensional semiclassical wave packets. Journal of Mathematical Physics, 57(9):092102, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aguiar:2016:EIS

[852] V. Aguiar and I. Guedes. Entropy and information of a spinless charged particle in time-varying magnetic fields. Journal of Mathematical Physics, 57(9):092103, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoque:2016:QAS

[853] Md Fazlul Hoque, Ian Marquette, and Yao-Zhong Zhang. Quadratic algebra for superintegrable monopole system in a Taub–NUT space. Journal of Mathematical Physics, 57(9):092104, September 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lami:2016:BDM

[854] Ludovico Lami and Marcus Huber. Bipartite depolarizing maps. Journal of Mathematical Physics, 57(9):092201, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leung:2016:MPC

[855] Debbie Leung and Nengkun Yu. Maximum privacy without coherence, zero-error. Journal of Mathematical Physics, 57(9):092202, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Westerbaan:2016:UPS

[856] Abraham Westerbaan and Bas Westerbaan. A universal property for sequential measurement. Journal of Mathematical Physics, 57(9):092203, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2016:PAP

[857] Timothy Nguyen. The perturbative approach to path integrals: A succinct mathematical treatment. Journal of Mathematical Physics, 57(9): 092301, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maharaj:2016:REB

[858] S. D. Maharaj, A. K. Tiwari, R. Mohanlal, and R. Narain. Riccati equations for bounded radiating systems. Journal of Mathematical Physics, 57(9):092501, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hsu:2016:MSS

[859] Cheng-Hsiung Hsu and Tetu Makino. Monotone-short solutions of the Tolman–Oppenheimer–Volkoff–de Sitter equation. Journal of Mathematical Physics, 57(9):092502, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:2016:SPP

[860] Robert J. Low. Spaces of paths and the path topology. Journal of Mathematical Physics, 57(9):092503, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyle:2016:HSS

[861] Michael Boyle. How should spin-weighted spherical functions be defined? Journal of Mathematical Physics, 57(9):092504, September 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lasenby:2016:SIG

[862] A. N. Lasenby and M. P. Hobson. Scale-invariant gauge theories of gravity: Theoretical foundations. Journal of Mathematical Physics, 57(9): 092505, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berkeley:2016:DTT

[863] George Berkeley, Alexander V. Mikhailov, and Pavlos Xenitidis. Darboux transformations with tetrahedral reduction group and related integrable systems. Journal of Mathematical Physics, 57(9):092701, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rui:2016:APS

[864] Jie Rui, Bingchen Liu, and Jian Zhang. Almost periodic solutions for a class of linear Schrödinger equations with almost periodic forcing. Journal of Mathematical Physics, 57(9):092702, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fortunati:2016:ISN

[865] Alessandro Fortunati and Stephen Wiggins. Integrability and strong normal forms for non-autonomous systems in a neighbourhood of an equilibrium. Journal of Mathematical Physics, 57(9):092703, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2016:SAP

[866] Jie Xu, Yu Miao, and Jicheng Liu. Strong averaging principle for twotime-scale non-autonomous stochastic FitzHugh–Nagumo system with jumps. Journal of Mathematical Physics, 57(9):092704, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masterov:2016:RHD

[867] Ivan Masterov. Remark on higher-derivative mechanics with l-conformal Galilei symmetry. Journal of Mathematical Physics, 57(9):092901, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:2016:PFE

[868] V. V. Nesterenko and A. V. Nesterenko. Ponderomotive forces in electrodynamics of moving media: The Minkowski and Abraham approaches. Journal of Mathematical Physics, 57(9):092902, September 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostrobij:2016:GDE

[869] P. Kostrobij, B. Markovych, O. Viznovych, and M. Tokarchuk. Generalized diffusion equation with fractional derivatives within Renyi statistics. Journal of Mathematical Physics, 57(9):093301, September 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sako:2016:TFR

[870] Akifumi Sako and Hiroshi Umetsu. Twisted Fock representations of noncommutative Kähler manifolds. Journal of Mathematical Physics, 57(9):

093501, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dohnal:2016:BEN

[871] Tomás Dohnal and Petr Siegl. Bifurcation of eigenvalues in nonlinear problems with antilinear symmetry. Journal of Mathematical Physics, 57(9):093502, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melnikov:2016:NSE

[872] Yuri B. Melnikov. New spectral estimations for a class of integraldifference operators and generalisation to higher dimensions. Journal of Mathematical Physics, 57(9):093503, September 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crampe:2016:SHA

[873] N. Crampé, L. Frappat, E. Ragoucy, and M. Vanicat. 3-state Hamiltonians associated to solvable 33-vertex models. Journal of Mathematical Physics, 57(9):093504, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blower:2016:KPP

[874] Gordon Blower and Yang Chen. Kernels and point processes associated with Whittaker functions. Journal of Mathematical Physics, 57(9): 093505, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chulaevsky:2016:NPA

[875] Victor Chulaevsky. Non-perturbative Anderson localization in heavytailed potentials via large deviations moment analysis. Journal of Mathematical Physics, 57(9):093506, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akhmedova:2016:DPT

[876] V. Akhmedova and A. Zabrodin. Dispersionless pfaff-Toda hierarchy and elliptic Löwner equation. Journal of Mathematical Physics, 57(9):093507, September 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2016:OSS

[877] Chuangye Liu, Nghiem V. Nguyen, and Zhi-Qiang Wang. Orbital stability of spatially synchronized solitary waves of an m-coupled nonlinear Schrödinger system. Journal of Mathematical Physics, 57(10):101501, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Darwich:2016:SRN

[878] Mohamad Darwich and Luc Molinet. Some remarks on the nonlinear Schrödinger equation with fractional dissipation. Journal of Mathematical Physics, 57(10):101502, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyko:2016:SRM

[879] Vaycheslav M. Boyko, Michael Kunzinger, and Roman O. Popovych. Singular reduction modules of differential equations. Journal of Mathematical Physics, 57(10):101503, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singla:2016:IAS

[880] Komal Singla and R. K. Gupta. On invariant analysis of some time fractional nonlinear systems of partial differential equations. I. Journal of Mathematical Physics, 57(10):101504, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2016:EGS

[881] Xiaoli Zhu, Fuyi Li, and Zhanping Liang. Existence of ground state solutions to a generalized quasilinear Schrödinger–Maxwell system. Journal of Mathematical Physics, 57(10):101505, October 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:2016:ABV

[882] Jaywan Chung and Ohsang Kwon. Asymptotic behavior for the viscous Burgers equation with a stationary source. Journal of Mathematical Physics, 57(10):101506, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yin:2016:CIE

[883] Rong Yin and Jihui Zhang. The characters for integral equations system on exterior domain. Journal of Mathematical Physics, 57(10):101507, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:HAR

[884] Tianjie Zhang, Xing Gao, and Li Guo. Hopf algebras of rooted forests, cocyles, and free Rota–Baxter algebras. Journal of Mathematical Physics, 57(10):101701, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2016:PCM

[885] Chengkang Xu. Partial classification of modules for the algebra of skew-derivations over the d-dimensional torus. Journal of Mathematical Physics, 57(10):101702, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2016:GMA

[886] Kang Lu, E. Mukhin, and A. Varchenko. On the Gaudin model associated to Lie algebras of classical types. Journal of Mathematical Physics, 57 (10):101703, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Draper:2016:GRF

[887] Cristina Draper and Valerio Guido. Gradings on the real form e_{6,−26}. Journal of Mathematical Physics, 57(10):101704, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2016:CSW

[888] Haibo Chen, Jianzhi Han, and Yucai Su. A class of simple weight modules over the twisted Heisenberg–Virasoro algebra. Journal of Mathematical Physics, 57(10):101705, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cagliero:2016:CFD

[889] Leandro Cagliero, Luis Gutiérrez Frez, and Fernando Szechtman. Classification of finite dimensional uniserial representations of conformal Galilei algebras. Journal of Mathematical Physics, 57(10):101706, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:2016:QOK

[890] C. Quesne. Quantum oscillator and Kepler–Coulomb problems in curved spaces: Deformed shape invariance, point canonical transformations, and rational extensions. Journal of Mathematical Physics, 57(10):102101, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rokhlenko:2016:ETN

[891] A. Rokhlenko. Exact theory and numeric results for short pulse ionization of simple model atom in one dimension. Journal of Mathematical Physics, 57(10):102102, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2016:GQN

[892] Axel Schulze-Halberg and Barnana Roy. Generalized quantum nonlinear oscillators: Exact solutions and rational extensions. Journal of Mathematical Physics, 57(10):102103, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayashi:2016:CDO

[893] Masahito Hayashi and Marco Tomamichel. Correlation detection and an operational interpretation of the Rényi mutual information. Journal of Mathematical Physics, 57(10):102201, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farenick:2016:FDO

[894] Douglas Farenick, Samuel Jaques, and Mizanur Rahaman. The fidelity of density operators in an operator-algebraic framework. Journal of Mathematical Physics, 57(10):102202, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Junge:2016:RNQ

[895] M. Junge, T. Oikhberg, and C. Palazuelos. Reducing the number of questions in nonlocal games. Journal of Mathematical Physics, 57(10): 102203, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harrison:2016:ZFD

[896] J. M. Harrison, T. Weyand, and K. Kirsten. Zeta functions of the Dirac operator on quantum graphs. Journal of Mathematical Physics, 57(10): 102301, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lowdon:2016:CVC

[897] Peter Lowdon. Conditions on the violation of the cluster decomposition property in QCD. Journal of Mathematical Physics, 57(10):102302, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Devchand:2016:PHP

[898] Chandrashekar Devchand and Andrea Spiro. On pseudo-hyperkähler prepotentials. Journal of Mathematical Physics, 57(10):102501, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantica:2016:WRT

[899] Carlo Alberto Mantica and Luca Guido Molinari. On the Weyl and Ricci tensors of generalized Robertson–Walker space-times. Journal of Mathematical Physics, 57(10):102502, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2016:ESQ

[900] Yunfeng Shi. On the existence of Sobolev quasi-periodic solutions of multidimensional nonlinear beam equation. Journal of Mathematical Physics, 57(10):102701, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goto:2016:CGD

[901] Shin itiro Goto. Contact geometric descriptions of vector fields on dually flat spaces and their applications in electric circuit models and nonequilibrium statistical mechanics. Journal of Mathematical Physics, 57(10): 102702, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krishnaswami:2016:CGI

[902] Govind S. Krishnaswami and Himalaya Senapati. Curvature and geodesic instabilities in a geometrical approach to the planar three-body problem. Journal of Mathematical Physics, 57(10):102901, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosenstock:2016:CEB

[903] Sarita Rosenstock and James Owen Weatherall. A categorical equivalence between generalized holonomy maps on a connected manifold and principal connections on bundles over that manifold. Journal of Mathematical Physics, 57(10):102902, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1489].

Tracy:2016:GSE

[904] Craig A. Tracy and Harold Widom. On the ground state energy of the delta-function Fermi gas. Journal of Mathematical Physics, 57(10): 103301, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katori:2016:EBP

[905] Makoto Katori. Elliptic Bessel processes and elliptic Dyson models realized as temporally inhomogeneous processes. Journal of Mathematical Physics, 57(10):103302, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hong:2016:NDP

[906] Liu Hong, Chen Jia, Yi Zhu, and Wen-An Yong. Novel dissipative properties of the master equation. Journal of Mathematical Physics, 57(10): 103303, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2016:ION

[907] F. Bagarello. Intertwining operators for non-self-adjoint Hamiltonians and bicoherent states. Journal of Mathematical Physics, 57(10):103501, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xia:2016:DTM

[908] Baoqiang Xia, Ruguang Zhou, and Zhijun Qiao. Darboux transformation and multi-soliton solutions of the Camassa–Holm equation and modified Camassa–Holm equation. Journal of Mathematical Physics, 57(10): 103502, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shlivinski:2016:CMA

[909] Amir Shlivinski. Coupled mode analysis of a periodic one-dimensional multimodal fiber bundle. Journal of Mathematical Physics, 57(10): 103503, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:2016:IMN

[910] Yu. G. Smirnov and D. V. Valovik. On the infinitely many nonperturbative solutions in a transmission eigenvalue problem for Maxwell’s equations with cubic nonlinearity. Journal of Mathematical Physics, 57 (10):103504, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirwin:2016:ECS

[911] William D. Kirwin, José Mourão, João P. Nunes, and Tao Qian. Extending coherent state transforms to Clifford analysis. Journal of Mathematical Physics, 57(10):103505, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grebenev:2016:STI

[912] V. N. Grebenev, M. Oberlack, A. G. Megrabov, and A. N. Grishkov. Symmetry transformations of an ideal steady fluid flow determined by a potential function. Journal of Mathematical Physics, 57(10):103506, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:2016:TWS

[913] Metin Gürses and Asli Pekcan. Traveling wave solutions of degenerate coupled multi-KdV equations. Journal of Mathematical Physics, 57(10): 103507, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2016:LTA

[914] Chueh-Hsin Chang, Ching-Hao Yu, and Tony Wen-Hann Sheu. Longtime asymptotic solution structure of Camassa–Holm equation subject to an initial condition with non-zero reflection coefficient of the scattering data. Journal of Mathematical Physics, 57(10):103508, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2016:CND

[915] Francesco Calogero. Comment on “Nonlinear differential algorithm to compute all the zeros of a generic polynomial” [j. math. phys. 57, 083508 (2016)]. Journal of Mathematical Physics, 57(10):104101, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [836].

Chuah:2016:ECS

[916] Meng-Kiat Chuah and Mingjing Zhang. Erratum: “Contragredient supersymmetric pairs” [j. math. phys. 54, 091702 (2013)]. Journal of Mathematical Physics, 57(10):109901, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [16].

Messaoudi:2016:WPE

[917] Salim A. Messaoudi, Abdelfeteh Fareh, and Nadjet Doudi. Well posedness and exponential stability in a wave equation with a strong damping and a strong delay. Journal of Mathematical Physics, 57(11):111501, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2016:MCP

[918] Claudianor O. Alves and Ailton R. da Silva. Multiplicity and concentration of positive solutions for a class of quasilinear problems through Orlicz–Sobolev space. Journal of Mathematical Physics, 57(11):111502, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2016:EVM

[919] Fei Jiang. On effects of viscosity and magnetic fields on the largest growth rate of linear Rayleigh–Taylor instability. Journal of Mathematical Physics, 57(11):111503, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2016:CVP

[920] Hanchun Yang and Jinhuan Wang. Concentration in vanishing pressure limit of solutions to the modified Chaplygin gas equations. Journal of Mathematical Physics, 57(11):111504, November 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2016:EMS

[921] Sihua Liang and Jihui Zhang. Existence and multiplicity of solutions for fourth-order elliptic equations of Kirchhoff type with critical growth in R^{N}. Journal of Mathematical Physics, 57(11):111505, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2016:ODR

[922] Yin Li, Ruiying Wei, and Zheng an Yao. Optimal decay rates for the compressible viscoelastic flows. Journal of Mathematical Physics, 57(11): 111506, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:2016:IMC

[923] Yoshimasa Matsuno. Integrable multi-component generalization of a modified short pulse equation. Journal of Mathematical Physics, 57(11): 111507, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tarlini:2016:BGR

[924] M. Tarlini. Braid group representations from a deformation of the harmonic oscillator algebra. Journal of Mathematical Physics, 57(11): 111701, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boos:2016:OVP

[925] Hermann Boos, Frank Göhmann, Andreas Klümper, Khazret S. Nirov, and Alexander V. Razumov. Oscillator versus prefundamental representations. Journal of Mathematical Physics, 57(11):111702, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brothier:2016:FPP

[926] Arnaud Brothier. On fixed point planar algebras. Journal of Mathematical Physics, 57(11):111703, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cunden:2016:LET

[927] Fabio Deelan Cunden, Francesco Mezzadri, Nick Simm, and Pierpaolo Vivo. Large-N expansion for the time-delay matrix of ballistic chaotic cavities. Journal of Mathematical Physics, 57(11):111901, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weder:2016:TFM

[928] Ricardo Weder. Trace formulas for the matrix Schrödinger operator on the half-line with general boundary conditions. Journal of Mathematical Physics, 57(11):112101, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horvath:2016:IEP

[929] Miklós Horváth and Orsolya Sáfár. Inverse eigenvalue problems. Journal of Mathematical Physics, 57(11):112102, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skowronek:2016:TND

[930] Lukasz Skowronek. There is no direct generalization of positive partial transpose criterion to the three-by-three case. Journal of Mathematical Physics, 57(11):112201, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eckstein:2016:MS

[931] Michal Eckstein, Andrzej Sitarz, and Raimar Wulkenhaar. The Moyal sphere. Journal of Mathematical Physics, 57(11):112301, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koroteev:2016:EAL

[932] Peter Koroteev and Antonio Sciarappa. On elliptic algebras and large-n supersymmetric gauge theories. Journal of Mathematical Physics, 57(11): 112302, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clark:2016:ACB

[933] T. E. Clark and T. ter Veldhuis. AdS–Carroll branes. Journal of Mathematical Physics, 57(11):112303, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guha:2016:RKP

[934] Partha Guha, E. Harikumar, and N. S. Zuhair. Regularization of Kepler problem in κ-spacetime. Journal of Mathematical Physics, 57(11): 112501, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cataldo:2016:CAD

[935] Mauricio Cataldo, Antonella Cid, Pedro Labraña, and Patricio Mella. Cosmic anisotropic doomsday in Bianchi type I universes. Journal of Mathematical Physics, 57(11):112502, November 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gomes:2016:ERG

[936] Henrique Gomes and Vasudev Shyam. Extending the rigidity of general relativity. Journal of Mathematical Physics, 57(11):112503, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diacu:2016:BLO

[937] Florin Diacu. Bifurcations of the Lagrangian orbits from the classical to the curved 3-body problem. Journal of Mathematical Physics, 57(11): 112701, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2016:FDT

[938] Xindong Xu and Dongfeng Zhang. Full dimension tori of Schrödinger equation. Journal of Mathematical Physics, 57(11):112702, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yao:2016:WSU

[939] Lei Yao and Haibo Cui. Weak-strong uniqueness for the isothermal Navier–Stokes equations. Journal of Mathematical Physics, 57(11): 113101, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sogo:2016:MCC

[940] Kiyoshi Sogo. Methods of computing Campbell–Hausdorff formula. Journal of Mathematical Physics, 57(11):113501, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2016:SRB

[941] H. Inoue. Semi-regular biorthogonal pairs and generalized Riesz bases. Journal of Mathematical Physics, 57(11):113502, November 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2016:RRMb

[942] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials. IV. Closure relations and creation/annihilation operators. Journal of Mathematical Physics, 57(11):113503, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vakhnenko:2016:SBC

[943] Oleksiy O. Vakhnenko. Symmetry-broken canonizations of the semidiscrete integrable nonlinear Schrödinger system with backgroundcontrolled inter-site coupling. Journal of Mathematical Physics, 57(11): 113504, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnad:2016:QHN

[944] J. Harnad. Quantum Hurwitz numbers and Macdonald polynomials. Journal of Mathematical Physics, 57(11):113505, November 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:2016:EUM

[945] Paulo R. S. Carvalho and Marcelo M. Leite. Erratum: “Unconventional minimal subtraction and Bogoliubov–Parasyuk–Hepp– Zimmermann method: Massive scalar theory and critical exponents” [J. Math. Phys. 54, 093301 (2013)]. Journal of Mathematical Physics, 57(11):119901, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [17].

Merino-Aceituno:2016:IWT

[946] Sara Merino-Aceituno. Isotropic wave turbulence with simplified kernels: Existence, uniqueness, and mean-field limit for a class of instantaneous coagulation-fragmentation processes. Journal of Mathematical Physics, 57(12):121501, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2016:OSV

[947] Cong He and Yuanjie Lei. One-species Vlasov–Poisson–Landau system for soft potentials in R^{3}. Journal of Mathematical Physics, 57(12):121502, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2016:NMQ

[948] John E. Gough. Non-Markovian quantum feedback networks I: Quantum transmission lines, lossless bounded real property and limit Markovian channels. Journal of Mathematical Physics, 57(12):122101, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2016:EDN

[949] Hong-Biao Zhang and Gangcheng Wang. Exact diagonalization of non-Hermitian so(3, 2) models: Generalized two-mode boson systems. Journal of Mathematical Physics, 57(12):122102, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakayama:2016:TQT

[950] Kunji Nakayama. Topos quantum theory reduced by context-selection functors. Journal of Mathematical Physics, 57(12):122103, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2016:ETQ

[951] Andreas Andersson. Electromagnetism in terms of quantum measurements. Journal of Mathematical Physics, 57(12):122104, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novaes:2016:EDC

[952] Marcel Novaes. Energy-dependent correlations in the S-matrix of chaotic systems. Journal of Mathematical Physics, 57(12):122105, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kenmoe:2016:DKM

[953] M. B. Kenmoe, A. B. Tchapda, and L. C. Fai. Demkov–Kunike models with decay. Journal of Mathematical Physics, 57(12):122106, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buyukasik:2016:ESH

[954] Sirin A. Büyükasik and Zehra Çayiç. Exactly solvable Hermite, Laguerre, and Jacobi type quantum parametric oscillators. Journal of Mathematical Physics, 57(12):122107, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anapolitanos:2016:SPC

[955] Ioannis Anapolitanos and Michael Hott. A simple proof of convergence to the Hartree dynamics in Sobolev trace norms. Journal of Mathematical Physics, 57(12):122108, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetti:2016:BWS

[956] Riccardo Giachetti and Vincenzo Grecchi. Bender–Wu singularities. Journal of Mathematical Physics, 57(12):122109, December 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aerts:2016:EBR

[957] Diederik Aerts and Massimiliano Sassoli de Bianchi. The extended Bloch representation of quantum mechanics: Explaining superposition, interference, and entanglement. Journal of Mathematical Physics, 57(12): 122110, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torres-Vega:2016:STL

[958] Gabino Torres-Vega. A symmetric time-like operator. Journal of Mathematical Physics, 57(12):122111, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kech:2016:DQT

[959] Michael Kech. Dynamical quantum tomography. Journal of Mathematical Physics, 57(12):122201, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Renes:2016:RSU

[960] Joseph M. Renes. Relative submajorization and its use in quantum resource theories. Journal of Mathematical Physics, 57(12):122202, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jencova:2016:COI

[961] Anna Jencová and Martin Plávala. Conditions for optimal input states for discrimination of quantum channels. Journal of Mathematical Physics, 57(12):122203, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frob:2016:AOB

[962] Markus B. Fröb, Jan Holland, and Stefan Hollands. All-order bounds for correlation functions of gauge-invariant operators in Yang–Mills theory. Journal of Mathematical Physics, 57(12):122301, December 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2016:KIA

[963] Luise Adams, Christian Bogner, Armin Schweitzer, and Stefan Weinzierl. The kite integral to all orders in terms of elliptic polylogarithms. Journal of Mathematical Physics, 57(12):122302, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nassar:2016:TDC

[964] Ali Nassar. Two-dimensional conformal field theories with matrix-valued level. Journal of Mathematical Physics, 57(12):122303, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lanery:2016:PLQ

[965] Suzanne Lanéry and Thomas Thiemann. Projective loop quantum gravity. I. State space. Journal of Mathematical Physics, 57(12):122304, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delaParra:2016:HAC

[966] Alejandro Betancourt de la Parra, Andrew S. Dancer, and McKenzie Y. Wang. A Hamiltonian approach to the cohomogeneity one Ricci soliton equations and explicit examples of non–Kähler solitons. Journal of Mathematical Physics, 57(12):122501, December 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pal:2016:AMU

[967] Sridip Pal and Narayan Banerjee. Anisotropic models are unitary: A rejuvenation of standard quantum cosmology. Journal of Mathematical Physics, 57(12):122502, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurianovich:2016:ESF

[968] E. A. Kurianovich. Exact solutions of Friedmann equation. Journal of Mathematical Physics, 57(12):122503, December 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bayin:2016:DRD

[969] Selçuk S. Bayin. Definition of the Riesz derivative and its application to space fractional quantum mechanics. Journal of Mathematical Physics, 57(12):123501, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2016:AID

[970] N. Aizawa and J. Segar. Aspects of infinite dimensional l-super Galilean conformal algebra. Journal of Mathematical Physics, 57(12):123502, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luu:2016:KFQ

[971] Martin T. Luu. KP flows and quantization. Journal of Mathematical Physics, 57(12):123503, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s correction [977].

Skrypnyk:2016:RFD

[972] T. Skrypnyk. Reductions in finite-dimensional integrable systems and special points of classical r-matrices. Journal of Mathematical Physics, 57(12):123504, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2016:SFQ

[973] John E. Gough. The Stratonovich formulation of quantum feedback network rules. Journal of Mathematical Physics, 57(12):123505, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jarosz:2016:FSE

[974] Stefânia Jarosz and Jayme Vaz, Jr. Fractional Schrödinger equation with Riesz–Feller derivative for delta potentials. Journal of Mathematical Physics, 57(12):123506, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hirota:2016:GQD

[975] Yuji Hirota. Geometric quantization of Dirac manifolds. Journal of Mathematical Physics, 57(12):123507, December 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1199].

DeVecchi:2016:RRS

[976] Francesco C. De Vecchi, Paola Morando, and Stefania Ugolini. Reduction and reconstruction of stochastic differential equations via symmetries. Journal of Mathematical Physics, 57(12):123508, December 2016. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luu:2016:PNK

[977] Martin T. Luu. Publisher’s note: “KP flows and quantization” [j. math. phys. 57, 123503 (2016)]. Journal of Mathematical Physics, 57(12): 129901, December 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [971].

Tran:2017:NPS

[978] Chuong V. Tran and Xinwei Yu. Note on Prodi–Serrin–Ladyzhenskaya type regularity criteria for the Navier–Stokes equations. Journal of Mathematical Physics, 58(1):011501, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ou:2017:GCS

[979] Yaobin Ou. Global classical solutions to the 1-D vacuum free boundary problem for full compressible Navier–Stokes equations with large data. Journal of Mathematical Physics, 58(1):011502, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2017:GEA

[980] Lili Fan and Hai-Yang Jin. Global existence and asymptotic behavior to a chemotaxis system with consumption of chemoattractant in higher dimensions. Journal of Mathematical Physics, 58(1):011503, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2017:ISW

[981] Zhipeng Cheng, Zifei Shen, and Minbo Yang. Instability of standing waves for a generalized Choquard equation with potential. Journal of Mathematical Physics, 58(1):011504, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Medina:2017:EOV

[982] Luciano Medina. On the existence of optical vortex solitons propagating in saturable nonlinear media. Journal of Mathematical Physics, 58(1): 011505, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Battaglia:2017:TSC

[983] Luca Battaglia. B_{2} and G_{2} Toda systems on compact surfaces: A variational approach. Journal of Mathematical Physics, 58(1):011506, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2017:RSS

[984] Yong-Chao Zhang and Qinglong Zhou. Radially symmetric sign-definite solutions for a class of nonlocal Schrödinger equations. Journal of Mathematical Physics, 58(1):011507, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2017:DCS

[985] Xing-Bin Pan. Directional curl spaces and applications to the Meissner states of anisotropic superconductors. Journal of Mathematical Physics, 58(1):011508, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bavula:2017:PIE

[986] V. V. Bavula and T. Lu. Prime ideals of the enveloping algebra of the Euclidean algebra and a classification of its simple weight modules. Journal of Mathematical Physics, 58(1):011701, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bringmann:2017:ABP

[987] Kathrin Bringmann, Amanda Folsom, and Antun Milas. Asymptotic behavior of partial and false theta functions arising from Jacobi forms and regularized characters. Journal of Mathematical Physics, 58(1):011702, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carlsson:2017:ASS

[988] Erik Carlsson. AGT and the Segal–Sugawara construction. Journal of Mathematical Physics, 58(1):011703, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kehle:2017:MEC

[989] Christoph Kehle. The maximal excess charge for a family of densitymatrix-functional theories including Hartree–Fock and Müller theories. Journal of Mathematical Physics, 58(1):011901, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atai:2017:DCS

[990] Farrokh Atai and Edwin Langmann. Deformed Calogero–Sutherland model and fractional quantum Hall effect. Journal of Mathematical Physics, 58(1):011902, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovacik:2017:MMN

[991] Samuel Kovácik and Peter Presnajder. Magnetic monopoles in noncommutative quantum mechanics. Journal of Mathematical Physics, 58(1): 012101, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fine:2017:PIS

[992] Dana S. Fine and Stephen Sawin. Path integrals, supersymmetric quantum mechanics, and the Atiyah–Singer index theorem for twisted Dirac. Journal of Mathematical Physics, 58(1):012102, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dykema:2017:KQS

[993] Ken Dykema and Kunal Mukherjee. KMS quantum symmetric states. Journal of Mathematical Physics, 58(1):012103, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Correa:2017:CSE

[994] R. A. C. Correa, A. de Souza Dutra, J. A. de Oliveira, and M. G. Garcia. A complete set of eigenstates for position-dependent massive particles in a Morse-like scenario. Journal of Mathematical Physics, 58(1):012104, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laptev:2017:PEE

[995] Ari Laptev and Sara Maad Sasane. Perturbations of embedded eigenvalues for a magnetic Schrödinger operator on a cylinder. Journal of Mathematical Physics, 58(1):012105, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felice:2017:VGS

[996] Domenico Felice, Minh Hà Quang, and Stefano Mancini. The volume of Gaussian states by information geometry. Journal of Mathematical Physics, 58(1):012201, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cleve:2017:PCO

[997] Richard Cleve, Li Liu, and William Slofstra. Perfect commuting-operator strategies for linear system games. Journal of Mathematical Physics, 58 (1):012202, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levay:2017:CCM

[998] Péter Lévay, Szilvia Nagy, János Pipek, and Gábor Sárosi. The coupled cluster method and entanglement in three fermion systems. Journal of Mathematical Physics, 58(1):012203, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cleve:2017:PEE

[999] Richard Cleve, Li Liu, and Vern I. Paulsen. Perfect embezzlement of entanglement. Journal of Mathematical Physics, 58(1):012204, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Datta:2017:CPQ

[1000] Nilanjana Datta, Yan Pautrat, and Cambyse Rouzé. Contractivity properties of a quantum diffusion semigroup. Journal of Mathematical Physics, 58(1):012205, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huber:2017:GIP

[1001] Stefan Huber, Robert König, and Anna Vershynina. Geometric inequalities from phase space translations. Journal of Mathematical Physics, 58 (1):012206, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heinze:2017:CCA

[1002] Martin Heinze, George Jorjadze, and Luka Megrelidze. Coset construction of AdS particle dynamics. Journal of Mathematical Physics, 58 (1):012301, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maharana:2017:APA

[1003] Jnanadeva Maharana. Analyticity properties and asymptotic behavior of scattering amplitude in higher dimensional theories. Journal of Mathematical Physics, 58(1):012302, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krongos:2017:RCD

[1004] D. S. Krongos and C. G. Torre. Rainich conditions in (2+1)-dimensional gravity. Journal of Mathematical Physics, 58(1):012501, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avalos:2017:EWM

[1005] R. Avalos, F. Dahia, and C. Romero. On the embedding of Weyl manifolds. Journal of Mathematical Physics, 58(1):012502, January 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eremenko:2017:ESP

[1006] Alexandre Eremenko, Andrei Gabrielov, and Aimo Hinkkanen. Exceptional solutions to the Painlevé VI equation. Journal of Mathematical Physics, 58(1):012701, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamiya:2017:CPN

[1007] Ryo Kamiya, Masataka Kanki, Takafumi Mase, and Tetsuji Tokihiro. Coprimeness-preserving non-integrable extension to the two-dimensional discrete Toda lattice equation. Journal of Mathematical Physics, 58(1): 012702, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Imaykin:2017:IPE

[1008] Valeriy Imaykin, Alexander Komech, and Herbert Spohn. On invariants for the Poincaré equations and applications. Journal of Mathematical Physics, 58(1):012901, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gubbiotti:2017:ACS

[1009] G. Gubbiotti and M. C. Nucci. Are all classical superintegrable systems in two-dimensional space linearizable? Journal of Mathematical Physics, 58(1):012902, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bogoyavlenskij:2017:MSV

[1010] Oleg Bogoyavlenskij. Moduli spaces of vortex knots for an exact fluid flow. Journal of Mathematical Physics, 58(1):013101, January 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Machida:2017:TFR

[1011] Manabu Machida. The time-fractional radiative transport equation — continuous-time random walk, diffusion approximation, and Legendrepolynomial expansion. Journal of Mathematical Physics, 58(1):013301, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amour:2017:WCQ

[1012] L. Amour, R. Lascar, and J. Nourrigat. Weyl calculus in QED I. The unitary group. Journal of Mathematical Physics, 58(1):013501, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gerdjikov:2017:CIN

[1013] V. S. Gerdjikov and A. Saxena. Complete integrability of nonlocal nonlinear Schrödinger equation. Journal of Mathematical Physics, 58(1): 013502, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mourao:2017:CST

[1014] José Mourão, João P. Nunes, and Tao Qian. Coherent state transforms and the Weyl equation in Clifford analysis. Journal of Mathematical Physics, 58(1):013503, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delisle:2017:EHB

[1015] Laurent Delisle. A N = 2 extension of the Hirota bilinear formalism and the supersymmetric KdV equation. Journal of Mathematical Physics, 58 (1):013504, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2017:UIS

[1016] Yu-Juan Zhang, Dun Zhao, and Wen-Xiu Ma. A unified inverse scattering transformation for the local and nonlocal nonautonomous Gross– Pitaevskii equations. Journal of Mathematical Physics, 58(1):013505, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galapon:2017:ECP

[1017] Eric A. Galapon. Erratum: “The Cauchy principal value and the Hadamard finite part integral as values of absolutely convergent integrals” [J. Math. Phys. 57, 033502 (2016)]. Journal of Mathematical Physics, 58(1):019901, January 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [637].

Li:2017:SCQ

[1018] Zhouxin Li and Yimin Zhang. Solutions for a class of quasilinear Schrödinger equations with critical Sobolev exponents. Journal of Mathematical Physics, 58(2):021501, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2017:NUD

[1019] Haiquan Wang and Ying Fu. Non-uniform dependence on initial data for the two-component Novikov system. Journal of Mathematical Physics, 58 (2):021502, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Watson:2017:WIP

[1020] Alexander B. Watson, Jianfeng Lu, and Michael I. Weinstein. Wavepackets in inhomogeneous periodic media: Effective particle-field dynamics and Berry curvature. Journal of Mathematical Physics, 58(2):021503, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2017:EHS

[1021] Liangwei Yang and Chengjie Yu. Estimates for higher Steklov eigenvalues. Journal of Mathematical Physics, 58(2):021504, February 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bai:2017:CHL

[1022] Wei Bai and Wende Liu. Cohomology of Heisenberg Lie superalgebras. Journal of Mathematical Physics, 58(2):021701, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2017:WRQ

[1023] A. D. Alhaidari and T. J. Taiwo. Wilson–Racah quantum system. Journal of Mathematical Physics, 58(2):022101, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorsch:2017:ARB

[1024] Florian Dorsch. Accumulation rate of bound states of dipoles generated by point charges in strained graphene. Journal of Mathematical Physics, 58(2):022102, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pandey:2017:QPY

[1025] Mahul Pandey and Sachindeo Vaidya. Quantum phases of Yang–Mills matrix model coupled to fundamental fermions. Journal of Mathematical Physics, 58(2):022103, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holweck:2017:EFQ

[1026] Frédéric Holweck, Jean-Gabriel Luque, and Jean-Yves Thibon. Entanglement of four-qubit systems: A geometric atlas with polynomial compass II (the tame world). Journal of Mathematical Physics, 58(2):022201, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mas:2017:DOS

[1027] Albert Mas. Dirac operators, shell interactions, and discontinuous gauge functions across the boundary. Journal of Mathematical Physics, 58(2): 022301, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2017:NIT

[1028] Ang Wei. On the numbers of images of two stochastic gravitational lensing models. Journal of Mathematical Physics, 58(2):022501, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sanchez:2017:GHS

[1029] Yafet Sanchez Sanchez and James A. Vickers. Generalised hyperbolicity in spacetimes with Lipschitz regularity. Journal of Mathematical Physics, 58(2):022502, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bautista:2017:CBS

[1030] A. Bautista, A. Ibort, J. Lafuente, and R. Low. A conformal boundary for space-times based on light-like geodesics: The 3-dimensional case. Journal of Mathematical Physics, 58(2):022503, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carinena:2017:SSD

[1031] Jose F. Cariñena, Francisco J. Herranz, and Manuel F. Rañada. Superintegrable systems on 3-dimensional curved spaces: Eisenhart formalism and separability. Journal of Mathematical Physics, 58(2):022701, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2017:SPO

[1032] Jaume Llibre and Durval J. Tonon. Symmetric periodic orbits for the collinear charged 3-body problem. Journal of Mathematical Physics, 58 (2):022702, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gentile:2017:RSF

[1033] Guido Gentile and Faenia Vaia. Response solutions for forced systems with large dissipation and arbitrary frequency vectors. Journal of Mathematical Physics, 58(2):022703, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2017:TDC

[1034] Shuqiang Zhu and Suo Zhao. Three-dimensional central configurations in H^{3} and S^{3}. Journal of Mathematical Physics, 58(2):022901, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Efstathiou:2017:RFL

[1035] K. Efstathiou, A. Giacobbe, P. Mardesić, and D. Sugny. Rotation forms and local Hamiltonian monodromy. Journal of Mathematical Physics, 58 (2):022902, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Byun:2017:WEG

[1036] Sun-Sig Byun and Hyoungsuk So. Weighted estimates for generalized steady Stokes systems in nonsmooth domains. Journal of Mathematical Physics, 58(2):023101, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2017:FFM

[1037] Emine Celik, Luan Hoang, Akif Ibragimov, and Thinh Kieu. Fluid flows of mixed regimes in porous media. Journal of Mathematical Physics, 58 (2):023102, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aristoff:2017:LGM

[1038] David Aristoff and Lingjiong Zhu. Lattice gas models with long range interactions. Journal of Mathematical Physics, 58(2):023301, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2017:FDT

[1039] Yong Chen, Chen Jia, and Da-Quan Jiang. Fluctuation-dissipation theorems for inhomogeneous Markov jump processes and a biochemical application. Journal of Mathematical Physics, 58(2):023302, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez:2017:APC

[1040] José C. Hernández, Yevgeniy Kovchegov, and Peter T. Otto. The aggregate path coupling method for the Potts model on bipartite graph. Journal of Mathematical Physics, 58(2):023303, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falek:2017:DKP

[1041] M. Falek, M. Merad, and T. Birkandan. Duffin–Kemmer–Petiau oscillator with Snyder–de Sitter algebra. Journal of Mathematical Physics, 58 (2):023501, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palese:2017:TOL

[1042] Marcella Palese and Ekkehart Winterroth. Topological obstructions in Lagrangian field theories, with an application to 3D Chern–Simons gauge theory. Journal of Mathematical Physics, 58(2):023502, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tadano:2017:UDB

[1043] Homare Tadano. An upper diameter bound for compact Ricci solitons with application to the Hitchin–Thorpe inequality. Journal of Mathematical Physics, 58(2):023503, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dappiaggi:2017:WSP

[1044] Claudio Dappiaggi, Heiko Gimperlein, Simone Murro, and Alexander Schenkel. Wavefront sets and polarizations on supermanifolds. Journal of Mathematical Physics, 58(2):023504, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grabowska:2017:RGL

[1045] Katarzyna Grabowska and Janusz Grabowski. Remarks on generalized Lie algebroids and related concepts. Journal of Mathematical Physics, 58 (2):023505, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurasov:2017:QGP

[1046] P. Kurasov and B. Majidzadeh Garjani. Quantum graphs: PT-symmetry and reflection symmetry of the spectrum. Journal of Mathematical Physics, 58(2):023506, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farnsworth:2017:GPR

[1047] Shane Farnsworth. The graded product of real spectral triples. Journal of Mathematical Physics, 58(2):023507, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2017:CDC

[1048] Salih Celik. Covariant differential calculi on quantum symplectic superspace SP_{q}^{1|2}. Journal of Mathematical Physics, 58(2):023508, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2017:GRG

[1049] Jian Xu and Engui Fan. The GLM representation of the global relation for the two-component nonlinear Schrödinger equation on the interval. Journal of Mathematical Physics, 58(2):023509, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruggieri:2017:CCL

[1050] M. Ruggieri and M. P. Speciale. On the construction of conservation laws: A mixed approach. Journal of Mathematical Physics, 58(2):023510, February 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Denlinger:2017:FSM

[1051] Ryan Denlinger, Zydrunas Gimbutas, Leslie Greengard, and Vladimir Rokhlin. A fast summation method for oscillatory lattice sums. Journal of Mathematical Physics, 58(2):023511, February 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2017:SHI

[1052] Paul W. Y. Lee. Sharp Harnack inequalities for a family of hypoelliptic diffusions. Journal of Mathematical Physics, 58(3):031501, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chirkunov:2017:CIN

[1053] Yu. A. Chirkunov. Conformal invariance and new exact solutions of the elastostatics equations. Journal of Mathematical Physics, 58(3):031502, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Furtado:2017:ESG

[1054] Marcelo F. Furtado, Edcarlos D. Silva, and Maxwell L. Silva. Existence of solution for a generalized quasilinear elliptic problem. Journal of Mathematical Physics, 58(3):031503, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okano:2017:NGT

[1055] Satoshi Okano and Shinichi Deguchi. A no-go theorem for the n-twistor description of a massive particle. Journal of Mathematical Physics, 58 (3):031701, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2017:EES

[1056] Shuiming Xiao and Huabin Chen. Existence and exponential stability for impulsive stochastic partial functional differential equations. Journal of Mathematical Physics, 58(3):032701, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meng:2017:NSC

[1057] Fengjuan Meng and Cuncai Liu. Necessary and sufficient conditions for the existence of time-dependent global attractor and application. Journal of Mathematical Physics, 58(3):032702, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Domingo:2017:BVU

[1058] Dario Domingo, Alberto d’Onofrio, and Franco Flandoli. Boundedness vs unboundedness of a noise linked to Tsallis q-statistics: The role of the overdamped approximation. Journal of Mathematical Physics, 58 (3):033301, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alpay:2017:CQP

[1059] Daniel Alpay, Fabrizio Colombo, and Irene Sabadini. On a class of quaternionic positive definite functions and their derivatives. Journal of Mathematical Physics, 58(3):033501, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2017:TSP

[1060] Kuo-Ming Lee. Transmission scattering problem via a DtN map. Journal of Mathematical Physics, 58(3):033502, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alarie-Vezina:2017:SSS

[1061] L. Alarie-Vézina, L. Lapointe, and P. Mathieu. N ≥ 2 symmetric superpolynomials. Journal of Mathematical Physics, 58(3):033503, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2017:SEE

[1062] Ying ying Sun, Da jun Zhang, and Frank W. Nijhoff. The Sylvester equation and the elliptic Korteweg–de Vries system. Journal of Mathematical Physics, 58(3):033504, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Movassagh:2017:GFH

[1063] Ramis Movassagh, Gilbert Strang, Yuta Tsuji, and Roald Hoffmann. The Green’s function for the Hückel (tight binding) model. Journal of Mathematical Physics, 58(3):033505, March 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2017:EMC

[1064] Quanqing Li and Xian Wu. Existence, multiplicity, and concentration of solutions for generalized quasilinear Schrödinger equations with critical growth. Journal of Mathematical Physics, 58(4):041501, April 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Cao:2017:ESP

[1065] Xiaofei Cao, Junxiang Xu, and Jun Wang. The existence of solutions with prescribed L^{2}-norm for Kirchhoff type system. Journal of Mathematical Physics, 58(4):041502, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dAvenia:2017:VGS

[1066] Pietro d’Avenia, Jarosław Mederski, and Alessio Pomponio. Vortex ground states for Klein–Gordon–Maxwell–Proca type systems. Journal of Mathematical Physics, 58(4):041503, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2017:CFI

[1067] Henan Wu and Lamei Yuan. Classification of finite irreducible conformal modules over some Lie conformal algebras related to the Virasoro conformal algebra. Journal of Mathematical Physics, 58(4):041701, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pandzic:2017:DOC

[1068] Pavle Pandzić and Petr Somberg. Dirac operator and its cohomology for the quantum group U_{q}(∫↕_{∈}). Journal of Mathematical Physics, 58 (4):041702, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kock:2017:PFC

[1069] Joachim Kock. Polynomial functors and combinatorial Dyson–Schwinger equations. Journal of Mathematical Physics, 58(4):041703, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruillard:2017:FMC

[1070] Paul Bruillard, César Galindo, Tobias Hagge, Siu-Hung Ng, Julia Yael Plavnik, Eric C. Rowell, and Zhenghan Wang. Fermionic modular categories and the 16-fold way. Journal of Mathematical Physics, 58(4): 041704, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2017:POA

[1071] Pavel Exner and Jirí Lipovský. Pseudo-orbit approach to trajectories of resonances in quantum graphs with general vertex coupling: Fermi rule and high-energy asymptotics. Journal of Mathematical Physics, 58 (4):042101, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Le:2017:EAS

[1072] Dai-Nam Le, Ngoc-Tram D. Hoang, and Van-Hoang Le. Exact analytical solutions of a two-dimensional hydrogen atom in a constant magnetic field. Journal of Mathematical Physics, 58(4):042102, April 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernardo:2017:EPI

[1073] Reginald Christian S. Bernardo and Jose Perico H. Esguerra. Euclidean path integral formalism in deformed space with minimum measurable length. Journal of Mathematical Physics, 58(4):042103, April 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Olgar:2017:WET

[1074] H. Olğar and O. Sh. Mukhtarov. Weak eigenfunctions of two-interval Sturm–Liouville problems together with interaction conditions. Journal of Mathematical Physics, 58(4):042201, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almeida:2017:QCS

[1075] Carla R. Almeida, Antonio B. Batista, Júlio C. Fabris, and Paulo V. Moniz. Quantum cosmology of scalar-tensor theories and self-adjointness. Journal of Mathematical Physics, 58(4):042301, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okolow:2017:KPQ

[1076] Andrzej Okolów. Kinematic projective quantum states for loop quantum gravity coupled to tensor fields. Journal of Mathematical Physics, 58(4): 042302, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gasparim:2017:TSP

[1077] Elizabeth Gasparim, Bruno Suzuki, Alexander Torres-Gomez, and Carlos A. B. Varea. Topological string partition function on generalised conifolds. Journal of Mathematical Physics, 58(4):042303, April 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Paetz:2017:ACRa

[1078] Tim-Torben Paetz. Algorithmic characterization results for the Kerr– NUT–(A)dS space-time. I. A space-time approach. Journal of Mathematical Physics, 58(4):042501, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paetz:2017:ACRb

[1079] Tim-Torben Paetz. Algorithmic characterization results for the Kerr– NUT–(A)dS space-time. II. KIDs for the Kerr–(A)(de Sitter) family. Journal of Mathematical Physics, 58(4):042502, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xing:2017:ITP

[1080] Zhitao Xing and Ercai Chen. Induced topological pressure and the zero-dimensional extension for BS dimensions. Journal of Mathematical Physics, 58(4):042701, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hau:2017:OSS

[1081] Jan-Niklas Hau, Martin Oberlack, and George Chagelishvili. On the optimal systems of subalgebras for the equations of hydrodynamic stability analysis of smooth shear flows and their group-invariant solutions. Journal of Mathematical Physics, 58(4):043101, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marner:2017:CVF

[1082] F. Marner, P. H. Gaskell, and M. Scholle. A complex-valued first integral of Navier–Stokes equations: Unsteady Couette flow in a corrugated channel system. Journal of Mathematical Physics, 58(4):043102, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dudnikova:2017:CEO

[1083] T. V. Dudnikova. On convergence to equilibrium for one-dimensional chain of harmonic oscillators on the half-line. Journal of Mathematical Physics, 58(4):043301, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pena:2017:BSS

[1084] J. J. Peña, J. Morales, and J. García-Ravelo. Bound state solutions of Dirac equation with radial exponential-type potentials. Journal of Mathematical Physics, 58(4):043501, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramani:2017:MTD

[1085] A. Ramani, B. Grammaticos, and R. Willox. Miura transformations for discrete Painlevé equations coming from the affine E_{8} Weyl group. Journal of Mathematical Physics, 58(4):043502, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ekanga:2017:MPL

[1086] Trésor Ekanga. Multi-particle localization for weakly interacting Anderson tight-binding models. Journal of Mathematical Physics, 58(4): 043503, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Norris:2017:SSF

[1087] L. K. Norris. On the n-symplectic structure of faithful irreducible representations. Journal of Mathematical Physics, 58(4):043504, April 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melnikov:2017:CSL

[1088] Andrey Melnikov. Construction of a Sturm–Liouville vessel using Gelfand–Levitan theory. On solution of the Korteweg–de Vries equation in the first quadrant. Journal of Mathematical Physics, 58(5):051501, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2017:GSO

[1089] Jianlin Zhang, Changming Song, and Hong Li. Global solutions for the one-dimensional compressible Navier–Stokes–Smoluchowski system. Journal of Mathematical Physics, 58(5):051502, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singla:2017:IAS

[1090] Komal Singla and R. K. Gupta. On invariant analysis of space-time fractional nonlinear systems of partial differential equations. II. Journal of Mathematical Physics, 58(5):051503, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAndrade:2017:OSP

[1091] Thiago Pinguello de Andrade, Fabrício Cristófani, and Fábio Natali. Orbital stability of periodic traveling wave solutions for the Kawahara equation. Journal of Mathematical Physics, 58(5):051504, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2017:LES

[1092] Guangze Gu, Yuanyang Yu, and Fukun Zhao. The least energy signchanging solution for a nonlocal problem. Journal of Mathematical Physics, 58(5):051505, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alcock-Zeilinger:2017:SRB

[1093] J. Alcock-Zeilinger and H. Weigert. Simplification rules for birdtrack operators. Journal of Mathematical Physics, 58(5):051701, May 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alcock-Zeilinger:2017:CHY

[1094] J. Alcock-Zeilinger and H. Weigert. Compact Hermitian Young projection operators. Journal of Mathematical Physics, 58(5):051702, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alcock-Zeilinger:2017:TO

[1095] J. Alcock-Zeilinger and H. Weigert. Transition operators. Journal of Mathematical Physics, 58(5):051703, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ito:2017:RES

[1096] Kenichi Ito and Arne Jensen. Resolvent expansions for the Schrödinger operator on the discrete half-line. Journal of Mathematical Physics, 58 (5):052101, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2017:UPK

[1097] Zhengwei Liu and Jinsong Wu. Uncertainty principles for Kac algebras. Journal of Mathematical Physics, 58(5):052102, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agyo:2017:GBT

[1098] S. Agyo, C. Lei, and A. Vourdas. The groupoid of bifractional transformations. Journal of Mathematical Physics, 58(5):052103, May 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Quesne:2017:FQE

[1099] C. Quesne. Families of quasi-exactly solvable extensions of the quantum oscillator in curved spaces. Journal of Mathematical Physics, 58(5): 052104, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

You:2017:EMG

[1100] Yuan You, Chang-Yuan Chen, Farida Tahir, and Shi-Hai Dong. Evaluation of more general integrals involving universal associated Legendre polynomials. Journal of Mathematical Physics, 58(5):052105, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Have:2017:EIB

[1101] Jonas Have, Hynek Kovarík, Thomas G. Pedersen, and Horia D. Cornean. On the existence of impurity bound excitons in one-dimensional systems with zero range interactions. Journal of Mathematical Physics, 58(5):052106, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gamel:2017:QBI

[1102] Omar E. Gamel and Graham R. Fleming. Quasi-Bell inequalities from symmetrized products of noncommuting qubit observables. Journal of Mathematical Physics, 58(5):052107, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guta:2017:IGL

[1103] Madalin Guta and Jukka Kiukas. Information geometry and local asymptotic normality for multi-parameter estimation of quantum Markov dynamics. Journal of Mathematical Physics, 58(5):052201, May 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loubenets:2017:NCU

[1104] Elena R. Loubenets. New concise upper bounds on quantum violation of general multipartite Bell inequalities. Journal of Mathematical Physics, 58(5):052202, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakata:2017:UDR

[1105] Yoshifumi Nakata, Christoph Hirche, Ciara Morgan, and Andreas Winter. Unitary 2-designs from random X- and Z-diagonal unitaries. Journal of Mathematical Physics, 58(5):052203, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mozrzymas:2017:SIC

[1106] Marek Mozrzymas, Michal Studziński, and Nilanjana Datta. Structure of irreducibly covariant quantum channels for finite groups. Journal of Mathematical Physics, 58(5):052204, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonzom:2017:DCS

[1107] Valentin Bonzom, Luca Lionni, and Adrian Tanasa. Diagrammatics of a colored SYK model and of an SYK-like tensor model, leading and next-to-leading orders. Journal of Mathematical Physics, 58(5):052301, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lanery:2017:PLQ

[1108] Suzanne Lanéry and Thomas Thiemann. Projective loop quantum gravity. II. Searching for semi-classical states. Journal of Mathematical Physics, 58(5):052302, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hansraj:2017:ASC

[1109] Sudan Hansraj, Sunil D. Maharaj, Sphakamiso Mlaba, and Nkululeko Qwabe. All static charged dust spheres in general relativity. Journal of Mathematical Physics, 58(5):052501, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bihun:2017:GSD

[1110] Oksana Bihun and Francesco Calogero. Generations of solvable discretetime dynamical systems. Journal of Mathematical Physics, 58(5):052701, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Godo:2017:FIT

[1111] B. Godó and Á. Nagy. Fisher information and topological pressure. Journal of Mathematical Physics, 58(5):052702, May 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boulter:2017:BPS

[1112] Eric Boulter, Florin Diacu, and Shuqiang Zhu. The N-body problem in spaces with uniformly varying curvature. Journal of Mathematical Physics, 58(5):052703, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajeev:2017:CHM

[1113] S. G. Rajeev. Curvature in Hamiltonian mechanics and the Einstein– Maxwell-dilaton action. Journal of Mathematical Physics, 58(5):052901, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeCaer:2017:CSR

[1114] G. Le Caër. Circumspheres of sets of n + 1 random points in the ddimensional Euclidean unit ball (1 ≤ n ≤ d). Journal of Mathematical Physics, 58(5):053301, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjornberg:2017:FOT

[1115] Jakob E. Björnberg. Fermionic observables in the transverse Ising chain. Journal of Mathematical Physics, 58(5):053302, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coletti:2017:CLT

[1116] Cristian F. Coletti, Renato Gava, and Gunter M. Schütz. Central limit theorem and related results for the elephant random walk. Journal of Mathematical Physics, 58(5):053303, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramos:2017:VLA

[1117] A. D. Ramos, F. S. G. Silva, C. S. Sousa, and A. Toom. Variable-length analog of stavskaya process: A new example of misleading simulation. Journal of Mathematical Physics, 58(5):053304, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

G:2017:BFC

[1118] Ramesh G and Santhosh Kumar P. Borel functional calculus for quaternionic normal operators. Journal of Mathematical Physics, 58(5):053501, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2017:QST

[1119] Hoil Kim. Quantum super torus and super mirror symmetry. Journal of Mathematical Physics, 58(5):053502, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2017:RLP

[1120] Giuseppe Gaeta and Francesco Spadaro. Random Lie-point symmetries of stochastic differential equations. Journal of Mathematical Physics, 58(5):053503, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1389].

Dutta:2017:PSD

[1121] Parikshit Dutta and Suvankar Dutta. Phase space distribution of Riemann zeros. Journal of Mathematical Physics, 58(5):053504, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sten:2017:TDE

[1122] J. C.-E. Sten, G. Fragoyiannis, P. Vafeas, P. K. Koivisto, and G. Dassios. Theoretical development of elliptic cross-sectional hyperboloidal harmonics and their application to electrostatics. Journal of Mathematical Physics, 58(5):053505, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grothaus:2017:FSS

[1123] Martin Grothaus and Felix Riemann. A fundamental solution to the Schrödinger equation with Doss potentials and its smoothness. Journal of Mathematical Physics, 58(5):053506, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gubbiotti:2017:NAY

[1124] G. Gubbiotti, C. Scimiterna, and D. Levi. The non-autonomous YdKN equation and generalized symmetries of Boll equations. Journal of Mathematical Physics, 58(5):053507, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Faghfouri:2017:CCG

[1125] Morteza Faghfouri, Ali Haji-Badali, and Fateme Gholami. On cosmological constant of generalized Robertson–Walker space-times. Journal of Mathematical Physics, 58(5):053508, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daoud:2017:GGV

[1126] M. Daoud and L. Gouba. Generalized Grassmann variables for quantum kit (k-level) systems and Barut–Girardello coherent states for su(r + 1) algebras. Journal of Mathematical Physics, 58(5):053509, May 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Chen:2017:LSS

[1127] Jinshu Chen and Caishi Wang. Linear stochastic Schrödinger equations in terms of quantum Bernoulli noises. Journal of Mathematical Physics, 58(5):053510, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liemert:2017:RTE

[1128] André Liemert and Alwin Kienle. Radiative transport equation for the Mittag-Leffler path length distribution. Journal of Mathematical Physics, 58(5):053511, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2017:TIH

[1129] Xiang Gu, Wen-Xiu Ma, and Wen-Ying Zhang. Two integrable Hamiltonian hierarchies in sl(2, R) and so(3, R) with three potentials. Journal of Mathematical Physics, 58(5):053512, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singla:2017:CLS

[1130] Komal Singla and R. K. Gupta. Comment on “Lie symmetries and group classification of a class of time fractional evolution systems” [J. Math. Phys. 56, 123504 (2015)]. Journal of Mathematical Physics, 58 (5):054101, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [464].

Singla:2017:GLS

[1131] Komal Singla and R. K. Gupta. Generalized Lie symmetry approach for fractional order systems of differential equations. III. Journal of Mathematical Physics, 58(6):061501, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kong:2017:LSS

[1132] De-Xing Kong and Chang-Hua Wei. Lifespan of smooth solutions for timelike extremal surface equation in de Sitter spacetime. Journal of Mathematical Physics, 58(6):061502, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2017:SQL

[1133] Shun-Jen Cheng. Supercharacters of queer Lie superalgebras. Journal of Mathematical Physics, 58(6):061701, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chowdhury:2017:PRA

[1134] S. Hasibul Hassan Chowdhury. On the plethora of representations arising in noncommutative quantum mechanics and an explicit construction of noncommutative 4-tori. Journal of Mathematical Physics, 58(6):061702, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2017:ESD

[1135] H. Xu, P. G. Kevrekidis, and T. Kapitula. Existence, stability, and dynamics of harmonically trapped one-dimensional multi-component solitary waves: The near-linear limit. Journal of Mathematical Physics, 58(6):061901, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cunden:2017:DSE

[1136] Fabio Deelan Cunden, Anna Maltsev, and Francesco Mezzadri. Density and spacings for the energy levels of quadratic Fermi operators. Journal of Mathematical Physics, 58(6):061902, June 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [1160] with corrected author affiliations.

Garmon:2017:CDN

[1137] Savannah Garmon and Gonzalo Ordonez. Characteristic dynamics near two coalescing eigenvalues incorporating continuum threshold effects. Journal of Mathematical Physics, 58(6):062101, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Trushechkin:2017:SEQ

[1138] Anton Trushechkin. Semiclassical evolution of quantum wave packets on the torus beyond the Ehrenfest time in terms of Husimi distributions. Journal of Mathematical Physics, 58(6):062102, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maciazek:2017:HGP

[1139] Tomasz Maciazek and Adam Sawicki. Homology groups for particles on one-connected graphs. Journal of Mathematical Physics, 58(6):062103, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braunss:2017:QDP

[1140] G. Braunss. Quantum dynamics in phase space: Moyal trajectories 3. Journal of Mathematical Physics, 58(6):062104, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kermani:2017:QSA

[1141] Maryam Mansoori Kermani and Ali Maghari. Quantum scattering approach for investigation of two interacting atoms trapped in a onedimensional Morse potential via Lippmann–Schwinger equation. Journal of Mathematical Physics, 58(6):062105, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarlette:2017:DSA

[1142] Alain Sarlette and Pierre Rouchon. Deterministic submanifolds and analytic solution of the quantum stochastic differential master equation describing a monitored qubit. Journal of Mathematical Physics, 58(6): 062106, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pereira:2017:AS

[1143] Rajesh Pereira and Connor Paul-Paddock. Anticoherent subspaces. Journal of Mathematical Physics, 58(6):062107, June 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Comelli:2017:HFT

[1144] S. Comelli and C. J. García-Cervera. Hartree–Fock theory with a selfgenerated magnetic field. Journal of Mathematical Physics, 58(6):062108, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2017:TCA

[1145] Xiwang Cao, Jiafu Mi, and Shanding Xu. Two constructions of approximately symmetric informationally complete positive operator-valued measures. Journal of Mathematical Physics, 58(6):062201, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernad:2017:LRQ

[1146] J. Z. Bernád and A. B. Frigyik. On the limit relation for the quantum relative entropy. Journal of Mathematical Physics, 58(6):062202, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuramochi:2017:MSS

[1147] Yui Kuramochi. Minimal sufficient statistical experiments on von Neumann algebras. Journal of Mathematical Physics, 58(6):062203, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zucchini:2017:AFH

[1148] Roberto Zucchini. Algebraic formulation of higher gauge theory. Journal of Mathematical Physics, 58(6):062301, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harris:2017:NHC

[1149] Pietro Artale Harris and Roberto Garra. Nonlinear heat conduction equations with memory: Physical meaning and analytical results. Journal of Mathematical Physics, 58(6):063501, June 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cances:2017:GKF

[1150] Eric Cancès, Paul Cazeaux, and Mitchell Luskin. Generalized Kubo formulas for the transport properties of incommensurate 2D atomic heterostructures. Journal of Mathematical Physics, 58(6):063502, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cherroud:2017:GLP

[1151] Othmane Cherroud, Sid-Ahmed Yahiaoui, and Mustapha Bentaiba. Generalized Laguerre polynomials with position-dependent effective mass visualized via Wigner’s distribution functions. Journal of Mathematical Physics, 58(6):063503, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valtancoli:2017:DOM

[1152] P. Valtancoli. Dirac oscillator and minimal length. Journal of Mathematical Physics, 58(6):063504, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaschke:2017:PCD

[1153] Petr Blaschke. Pedal coordinates, dark Kepler, and other force problems. Journal of Mathematical Physics, 58(6):063505, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciaglia:2017:HJA

[1154] Florio M. Ciaglia, Fabio Di Cosmo, Domenico Felice, Stefano Mancini, Giuseppe Marmo, and Juan M. Pérez-Pardo. Hamilton–Jacobi approach to potential functions in information geometry. Journal of Mathematical Physics, 58(6):063506, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yahiaoui:2017:IHP

[1155] Sid-Ahmed Yahiaoui and Mustapha Bentaiba. Isospectral Hamiltonian for position-dependent mass for an arbitrary quantum system and coherent states. Journal of Mathematical Physics, 58(6):063507, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doubrov:2017:CIS

[1156] B. Doubrov, E. V. Ferapontov, B. Kruglikov, and V. S. Novikov. On a class of integrable systems of Monge–Ampère type. Journal of Mathematical Physics, 58(6):063508, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cockburn:2017:DVM

[1157] Alexander Cockburn, Steffen Krusch, and Abera A. Muhamed. Dynamics of vortices with magnetic impurities. Journal of Mathematical Physics, 58(6):063509, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dattoli:2017:TRH

[1158] G. Dattoli, K. Górska, A. Horzela, K. A. Penson, and E. Sabia. Theory of relativistic heat polynomials and one-sided Lévy distributions. Journal of Mathematical Physics, 58(6):063510, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yafaev:2017:PID

[1159] D. R. Yafaev. A point interaction for the discrete Schrödinger operator and generalized Chebyshev polynomials. Journal of Mathematical Physics, 58(6):063511, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cunden:2017:PND

[1160] Fabio Deelan Cunden, Anna Maltsev, and Francesco Mezzadri. Publisher’s note: “Density and spacings for the energy levels of quadratic Fermi operators” [j. math. phys. 58, 061902 (2017)]. Journal of Mathematical Physics, 58(6):069901, June 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1136].

Gralak:2017:APE

[1161] Boris Gralak. Analytic properties of the electromagnetic Green’s function. Journal of Mathematical Physics, 58(7):071501, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farah:2017:SCN

[1162] Luiz Gustavo Farah and Ademir Pastor. Scattering for a 3D coupled nonlinear Schrödinger system. Journal of Mathematical Physics, 58(7): 071502, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2017:SSS

[1163] Sze-Guang Yang, Zhi-You Chen, and Jann-Long Chern. The solution structure of the O(3) sigma model in a Maxwell–Chern–Simons theory. Journal of Mathematical Physics, 58(7):071503, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassier:2017:BHF

[1164] Maxence Cassier and Graeme W. Milton. Bounds on Herglotz functions and fundamental limits of broadband passive quasistatic cloaking. Journal of Mathematical Physics, 58(7):071504, July 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2017:LWP

[1165] Jianjun Yuan. Local well-posedness of non-abelian Chern–Simons–Higgs system in the Lorenz gauge. Journal of Mathematical Physics, 58(7): 071505, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agafonov:2017:QSL

[1166] S. I. Agafonov, E. V. Ferapontov, and V. S. Novikov. Quasilinear systems with linearizable characteristic webs. Journal of Mathematical Physics, 58(7):071506, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2017:GEW

[1167] Jianwei Yang and Yong Li. Global existence of weak solution for quantum Navier–Stokes–Poisson equations. Journal of Mathematical Physics, 58 (7):071507, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aicha:2017:SEI

[1168] Ibtissem Ben Aïcha. Stability estimate for an inverse problem for the Schrödinger equation in a magnetic field with time-dependent coefficient. Journal of Mathematical Physics, 58(7):071508, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2017:TAI

[1169] Seung-Yeal Ha, Dongnam Ko, Xiongtao Zhang, and Yinglong Zhang. Time-asymptotic interactions of two ensembles of Cucker–Smale flocking particles. Journal of Mathematical Physics, 58(7):071509, July 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meljanac:2017:RBD

[1170] Stjepan Meljanac, Sasa Kresić-Jurić, and Tea Martinić. Realization of bicovariant differential calculus on the Lie algebra type noncommutative spaces. Journal of Mathematical Physics, 58(7):071701, July 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chalykh:2017:KHC

[1171] Oleg Chalykh and Alexey Silantyev. KP hierarchy for the cyclic quiver. Journal of Mathematical Physics, 58(7):071702, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vinogradov:2017:PLS

[1172] A. M. Vinogradov. Particle-like structure of Lie algebras. Journal of Mathematical Physics, 58(7):071703, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Awata:2017:CDV

[1173] Hidetoshi Awata, Hiroki Fujino, and Yusuke Ohkubo. Crystallization of deformed Virasoro algebra, Ding–Iohara–Miki algebra, and 5D AGT correspondence. Journal of Mathematical Physics, 58(7):071704, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melas:2017:RTB

[1174] Evangelos Melas. On the representation theory of the Bondi–Metzner– Sachs group and its variants in three space-time dimensions. Journal of Mathematical Physics, 58(7):071705, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2017:QFR

[1175] Hengyun Yang and Yang Zhang. A quantum fermion realisation of the finite dimensional spinor representation of U_{q}(osp(1|2n)). Journal of Mathematical Physics, 58(7):071706, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anguelova:2017:SBC

[1176] Iana I. Anguelova. The second bosonization of the CKP hierarchy. Journal of Mathematical Physics, 58(7):071707, July 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Girelli:2017:CSL

[1177] Florian Girelli and Giuseppe Sellaroli. SO*(2N) coherent states for loop quantum gravity. Journal of Mathematical Physics, 58(7):071708, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bachmann:2017:LFD

[1178] Sven Bachmann and Andreas Bluhm. Local factorisation of the dynamics of quantum spin systems. Journal of Mathematical Physics, 58(7): 071901, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Betermin:2017:DRT

[1179] Laurent Bétermin and Mircea Petrache. Dimension reduction techniques for the minimization of theta functions on lattices. Journal of Mathematical Physics, 58(7):071902, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pogosyan:2017:QSZ

[1180] George S. Pogosyan, Cristina Salto-Alegre, Kurt Bernardo Wolf, and Alexander Yakhno. Quantum superintegrable Zernike system. Journal of Mathematical Physics, 58(7):072101, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2017:PWD

[1181] Zhiguang Xiao and Zhi-Yong Zhou. Partial wave decomposition in Friedrichs model with self-interacting continua. Journal of Mathematical Physics, 58(7):072102, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bednarik:2017:ESS

[1182] Michal Bednarik and Milan Cervenka. The exact solution of the Schrödinger equation with a polynomially spatially varying mass. Journal of Mathematical Physics, 58(7):072103, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2017:SNW

[1183] A. D. Alhaidari. Solution of the nonrelativistic wave equation using the tridiagonal representation approach. Journal of Mathematical Physics, 58(7):072104, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioffe:2017:MTD

[1184] M. V. Ioffe, D. N. Nishnianidze, and V. V. Vereshagin. Mapping of twodimensional Schrödinger equation under the point transformation. Journal of Mathematical Physics, 58(7):072105, July 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Phan:2017:BPB

[1185] Thanh Viet Phan. Blow-up profile of Bose–Einstein condensate with singular potentials. Journal of Mathematical Physics, 58(7):072301, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morozov:2017:LBM

[1186] Sergey Morozov and David Müller. Lower bounds on the moduli of threedimensional Coulomb–Dirac operators via fractional Laplacians with applications. Journal of Mathematical Physics, 58(7):072302, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Much:2017:CFS

[1187] Albert Much. Curving flat space-time by deformation quantization? Journal of Mathematical Physics, 58(7):072303, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2017:MSR

[1188] Lars Andersson, Siyuan Ma, Claudio Paganini, and Bernard F. Whiting. Mode stability on the real axis. Journal of Mathematical Physics, 58(7): 072501, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leandro:2017:ISS

[1189] Benedito Leandro and Romildo Pina. Invariant solutions for the static vacuum equation. Journal of Mathematical Physics, 58(7):072502, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohanlal:2017:NER

[1190] R. Mohanlal, R. Narain, and S. D. Maharaj. Nonlinear equations in radiating stellar structures. Journal of Mathematical Physics, 58(7):072503, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pogosyan:2017:SCZ

[1191] George S. Pogosyan, Kurt Bernardo Wolf, and Alexander Yakhno. Superintegrable classical Zernike system. Journal of Mathematical Physics, 58(7):072901, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leyvraz:2017:AOI

[1192] F. Leyvraz. An approach for obtaining integrable Hamiltonians from Poisson-commuting polynomial families. Journal of Mathematical Physics, 58(7):072902, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Littin:2017:QAE

[1193] Jorge Littin and Pierre Picco. Quasi-additive estimates on the Hamiltonian for the one-dimensional long range Ising model. Journal of Mathematical Physics, 58(7):073301, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bozejko:2017:FRQ

[1194] Marek Bozejko, Eugene Lytvynov, and Janusz Wysoczański. Fock representations of Q-deformed commutation relations. Journal of Mathematical Physics, 58(7):073501, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teramoto:2017:CHN

[1195] Hiroshi Teramoto, Kenji Kondo, Shyuichi Izumiya, Mikito Toda, and Tamiki Komatsuzaki. Classification of Hamiltonians in neighborhoods of band crossings in terms of the theory of singularities. Journal of Mathematical Physics, 58(7):073502, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2319].

Ruiz:2017:PDF

[1196] Patricia Alonso Ruiz. Power dissipation in fractal Feynman–Sierpinski AC circuits. Journal of Mathematical Physics, 58(7):073503, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ancarani:2017:DHH

[1197] L. U. Ancarani, J. A. Del Punta, and G. Gasaneo. Derivatives of Horn hypergeometric functions with respect to their parameters. Journal of Mathematical Physics, 58(7):073504, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2017:CQS

[1198] John E. Gough and Guofeng Zhang. Classical and quantum stochastic models of resistive and memristive circuits. Journal of Mathematical Physics, 58(7):073505, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hirota:2017:EGQ

[1199] Yuji Hirota. Erratum: “Geometric quantization of Dirac manifolds” [J. Math. Phys. 57, 123507 (2016)]. Journal of Mathematical Physics, 58(7): 079901, July 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [975].

Salmhofer:2017:EHO

[1200] Manfred Salmhofer and Bruno Nachtergaele. Editorial: In honor of Oscar E. Lanford III. Journal of Mathematical Physics, 58(8):080401, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2017:EMC

[1201] Jianfu Yang and Jinge Yang. Existence and mass concentration of pseudo-relativistic Hartree equation. Journal of Mathematical Physics, 58(8):081501, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shang:2017:ECP

[1202] Xudong Shang and Jihui Zhang. Existence and concentration of positive solutions for fractional nonlinear Schrödinger equation with critical growth. Journal of Mathematical Physics, 58(8):081502, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2017:ERC

[1203] Claudianor O. Alves and Olímpio H. Miyagaki. Existence, regularity, and concentration phenomenon of nontrivial solitary waves for a class of generalized variable coefficient Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 58(8):081503, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meng:2017:HOW

[1204] Yiping Meng and Ya-Guang Wang. Highly oscillatory waves in quasilinear hyperbolic-parabolic coupled equations. Journal of Mathematical Physics, 58(8):081504, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibrahim:2017:TWM

[1205] H. Ibrahim and E. Nasreddine. Traveling waves for a model of individual clustering with logistic growth rate. Journal of Mathematical Physics, 58(8):081505, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovalenko:2017:LGC

[1206] S. Kovalenko and O. Patsiuk. Lie group classification of the nonlinear generalized Kompaneets equations with two functional parameters. Journal of Mathematical Physics, 58(8):081506, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basarab-Horwath:2017:LTO

[1207] P. Basarab-Horwath and F. Güngör. Linearizability for third order evolution equations. Journal of Mathematical Physics, 58(8):081507, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2017:GCS

[1208] Xiaolong Wang and Zhenglu Jiang. Global classical solution to the inelastic Boltzmann equation with potential force. Journal of Mathematical Physics, 58(8):081508, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hintz:2017:ALW

[1209] Peter Hintz and András Vasy. Analysis of linear waves near the Cauchy horizon of cosmological black holes. Journal of Mathematical Physics, 58 (8):081509, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2017:NRL

[1210] Min Ding and Yachun Li. Non-relativistic limits of rarefaction wave to the 1-D piston problem for the isentropic relativistic Euler equations. Journal of Mathematical Physics, 58(8):081510, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Opanasenko:2017:GAG

[1211] Stanislav Opanasenko, Alexander Bihlo, and Roman O. Popovych. Group analysis of general Burgers–Korteweg–de Vries equations. Journal of Mathematical Physics, 58(8):081511, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penafiel:2017:IEI

[1212] D. M. Peñafiel and L. Ravera. Infinite S-expansion with ideal subtraction and some applications. Journal of Mathematical Physics, 58(8):081701, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Subag:2017:GAC

[1213] Eyal M. Subag, Ehud Moshe Baruch, Joseph L. Birman, and Ady Mann. Gabor analysis as contraction of wavelets analysis. Journal of Mathematical Physics, 58(8):081702, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deuchert:2017:LBB

[1214] Andreas Deuchert. A lower bound for the BCS functional with boundary conditions at infinity. Journal of Mathematical Physics, 58(8):081901, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kempf:2017:DQS

[1215] Achim Kempf and Angus Prain. Driving quantum systems with superoscillations. Journal of Mathematical Physics, 58(8):082101, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mas:2017:RSS

[1216] Albert Mas and Fabio Pizzichillo. The relativistic spherical δ-shell interaction in R^{3}: Spectrum and approximation. Journal of Mathematical Physics, 58(8):082102, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kryukov:2017:MMB

[1217] A. Kryukov. On the motion of macroscopic bodies in quantum theory. Journal of Mathematical Physics, 58(8):082103, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clivaz:2017:GME

[1218] Fabien Clivaz, Marcus Huber, Ludovico Lami, and Gláucia Murta. Genuine-multipartite entanglement criteria based on positive maps. Journal of Mathematical Physics, 58(8):082201, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chubb:2017:ASH

[1219] Christopher T. Chubb and Steven T. Flammia. Approximate symmetries of Hamiltonians. Journal of Mathematical Physics, 58(8):082202, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kobayshi:2017:CRH

[1220] Masatomo Kobayshi. Coexistence on reflecting hyperplane in generalized probability theories. Journal of Mathematical Physics, 58(8):082203, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parthasarathy:2017:QSD

[1221] K. R. Parthasarathy and A. R. Usha Devi. From quantum stochastic differential equations to Gisin–Percival state diffusion. Journal of Mathematical Physics, 58(8):082204, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avetisyan:2017:ADS

[1222] Zhirayr Avetisyan, Yan-Long Fang, Nikolai Saveliev, and Dmitri Vassiliev. Analytic definition of spin structure. Journal of Mathematical Physics, 58(8):082301, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babaei:2017:QMP

[1223] Hassan Babaei and Ali Mostafazadeh. Quantum mechanics of a photon. Journal of Mathematical Physics, 58(8):082302, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigore:2017:GGI

[1224] Dan-Radu Grigore. A generalization of gauge invariance. Journal of Mathematical Physics, 58(8):082303, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herfray:2017:TFT

[1225] Yannick Herfray and Kirill Krasnov. Topological field theories of 2- and 3forms in six dimensions. Journal of Mathematical Physics, 58(8):082304, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osuga:2017:NWD

[1226] Kento Osuga and Don N. Page. A new way to derive the Taub–NUT metric with positive cosmological constant. Journal of Mathematical Physics, 58(8):082501, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dominy:2017:DCE

[1227] Jason M. Dominy and Daniele Venturi. Duality and conditional expectations in the Nakajima–Mori–Zwanzig formulation. Journal of Mathematical Physics, 58(8):082701, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2017:EUS

[1228] Linfang Liu and Xianlong Fu. Existence and upper semi-continuity of pullback attractors of a p-Laplacian equation with delay. Journal of Mathematical Physics, 58(8):082702, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karpathopoulos:2017:LNP

[1229] Leonidas Karpathopoulos, Andronikos Paliathanasis, and Michael Tsamparlis. Lie and Noether point symmetries for a class of nonautonomous dynamical systems. Journal of Mathematical Physics, 58(8):082901, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertin:2017:HJF

[1230] M. C. Bertin, B. M. Pimentel, C. E. Valcárcel, and G. E. R. Zambrano. Hamilton–Jacobi formalism for Podolsky’s electromagnetic theory on the null-plane. Journal of Mathematical Physics, 58(8):082902, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baranovsky:2017:HDG

[1231] Oleg Baranovsky. Higher-derivative generalization of conformal mechanics. Journal of Mathematical Physics, 58(8):082903, August 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mironov:2017:SEI

[1232] Victor L. Mironov and Sergey V. Mironov. Sedeonic equations of ideal fluid. Journal of Mathematical Physics, 58(8):083101, August 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Al-Buradah:2017:ESG

[1233] Sadig A. Al-Buradah, Hocine Bahlouli, and Abdulaziz D. Alhaidari. Energy spectrum of a generalized Scarf potential using the asymptotic iteration method and the tridiagonal representation approach. Journal of Mathematical Physics, 58(8):083501, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grammaticos:2017:MER

[1234] B. Grammaticos, A. Ramani, R. Willox, and J. Satsuma. Multiplicative equations related to the affine Weyl group E_{8}. Journal of Mathematical Physics, 58(8):083502, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guay-Paquet:2017:GFW

[1235] Mathieu Guay-Paquet and J. Harnad. Generating functions for weighted Hurwitz numbers. Journal of Mathematical Physics, 58(8):083503, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galleas:2017:CRS

[1236] W. Galleas. Continuous representations of scalar products of Bethe vectors. Journal of Mathematical Physics, 58(8):083504, August 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrogowska:2017:TLB

[1237] Alina Dobrogowska and Grzegorz Jakimowicz. Tangent lifts of bi-Hamiltonian structures. Journal of Mathematical Physics, 58(8):083505, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zakharevich:2017:LSP

[1238] Valentin Zakharevich. Localization and stationary phase approximation on supermanifolds. Journal of Mathematical Physics, 58(8):083506, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez-Ramirez:2017:TBP

[1239] M. Alvarez-Ramírez, A. García, J. Meléndez, and J. G. Reyes-Victoria. The three-body problem and equivariant Riemannian geometry. Journal of Mathematical Physics, 58(8):083507, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikitin:2017:KIG

[1240] A. G. Nikitin. Kinematical invariance groups of the 3 d Schrödinger equations with position dependent masses. Journal of Mathematical Physics, 58(8):083508, August 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agaoglou:2017:DBS

[1241] M. Agaoglou, E. G. Charalampidis, T. A. Ioannidou, and P. G. Kevrekidis. Discrete BPS skyrmions. Journal of Mathematical Physics, 58(9):091501, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Przedborski:2017:TWC

[1242] Michelle Przedborski and Stephen C. Anco. Traveling waves and conservation laws for highly nonlinear wave equations modeling Hertz chains. Journal of Mathematical Physics, 58(9):091502, September 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sheng:2017:SCR

[1243] Weimin Sheng and Haobin Yu. Singularity of connection Ricci flow for three-manifolds. Journal of Mathematical Physics, 58(9):091503, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krizka:2017:DIS

[1244] Libor Krizka and Petr Somberg. Differential invariants on symplectic spinors in contact projective geometry. Journal of Mathematical Physics, 58(9):091701, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouarroudj:2017:CDS

[1245] Sofiane Bouarroudj and Guowu Meng. The classical dynamic symmetry for the Sp(1)-Kepler problems. Journal of Mathematical Physics, 58(9): 091702, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motegi:2017:CPS

[1246] Kohei Motegi. Combinatorial properties of symmetric polynomials from integrable vertex models in finite lattice. Journal of Mathematical Physics, 58(9):091703, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Su:2017:RRS

[1247] Dong Su and Shilin Yang. Representation rings of small quantum groups Ū_{q}(sl_{2}). Journal of Mathematical Physics, 58(9):091704, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anantharaman:2017:QEA

[1248] Nalini Anantharaman and Mostafa Sabri. Quantum ergodicity for the Anderson model on regular graphs. Journal of Mathematical Physics, 58 (9):091901, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanki:2017:EDC

[1249] Kazuki Kanki, Savannah Garmon, Satoshi Tanaka, and Tomio Petrosky. Exact description of coalescing eigenstates in open quantum systems in terms of microscopic Hamiltonian dynamics. Journal of Mathematical Physics, 58(9):092101, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerini:2017:OFQ

[1250] Leonardo Guerini, Jessica Bavaresco, Marcelo Terra Cunha, and Antonio Acín. Operational framework for quantum measurement simulability. Journal of Mathematical Physics, 58(9):092102, September 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colombo:2017:ESS

[1251] F. Colombo, J. Gantner, and D. C. Struppa. Evolution of superoscillations for Schrödinger equation in a uniform magnetic field. Journal of Mathematical Physics, 58(9):092103, September 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2017:SSO

[1252] Felix Finster and José M. Isidro. L^{p}-spectrum of the Schrödinger operator with inverted harmonic oscillator potential. Journal of Mathematical Physics, 58(9):092104, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohsawa:2017:GDG

[1253] Tomoki Ohsawa and Cesare Tronci. Geometry and dynamics of Gaussian wave packets and their Wigner transforms. Journal of Mathematical Physics, 58(9):092105, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kribs:2017:OSQ

[1254] David W. Kribs, Comfort Mintah, Michael Nathanson, and Rajesh Pereira. Operator structures and quantum one-way LOCC conditions. Journal of Mathematical Physics, 58(9):092201, September 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2017:EDM

[1255] Hao-Chung Cheng, Min-Hsiu Hsieh, and Marco Tomamichel. Exponential decay of matrix Φ-entropies on Markov semigroups with applications to dynamical evolutions of quantum ensembles. Journal of Mathematical Physics, 58(9):092202, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lancien:2017:FCF

[1256] Cécilia Lancien and Andreas Winter. Flexible constrained de Finetti reductions and applications. Journal of Mathematical Physics, 58(9): 092203, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gour:2017:AAM

[1257] Gilad Gour, Barbara Kraus, and Nolan R. Wallach. Almost all multipartite qubit quantum states have trivial stabilizer. Journal of Mathematical Physics, 58(9):092204, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bizi:2017:DCS

[1258] Nadir Bizi and Fabien Besnard. The disappearance of causality at small scale in almost-commutative manifolds. Journal of Mathematical Physics, 58(9):092301, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2017:BSE

[1259] Ho Lee and Ernesto Nungesser. Bianchi I solutions of the Einstein– Boltzmann system with a positive cosmological constant. Journal of Mathematical Physics, 58(9):092501, September 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurihara:2017:CCG

[1260] Yoshimasa Kurihara. Characteristic classes in general relativity on a modified Poincaré curvature bundle. Journal of Mathematical Physics, 58(9):092502, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ceccato:2017:ASH

[1261] Alessandro Ceccato, Paolo Nicolini, and Diego Frezzato. Attracting subspaces in a hyper-spherical representation of autonomous dynamical systems. Journal of Mathematical Physics, 58(9):092701, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2017:HJT

[1262] Manuel de León, Pedro Daniel Prieto-Martínez, Narciso Román-Roy, and Silvia Vilariño. Hamilton–Jacobi theory in multisymplectic classical field theories. Journal of Mathematical Physics, 58(9):092901, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osano:2017:TAT

[1263] Bob Osano and Patrick W. M. Adams. Toward the analogue of a thermally generated electromagnetic field. Journal of Mathematical Physics, 58(9):093101, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nobili:2017:LBF

[1264] Camilla Nobili and Felix Otto. Limitations of the background field method applied to Rayleigh–Bénard convection. Journal of Mathematical Physics, 58(9):093102, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guarnieri:2017:SFP

[1265] F. Guarnieri, W. Moon, and J. S. Wettlaufer. Solution of the Fokker– Planck equation with a logarithmic potential and mixed eigenvalue spectrum. Journal of Mathematical Physics, 58(9):093301, September 2017.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazzolo:2017:COU

[1266] Alain Mazzolo. Constraint Ornstein–Uhlenbeck bridges. Journal of Mathematical Physics, 58(9):093302, September 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avendano-Camacho:2017:PTA

[1267] M. Avendaño-Camacho, J. A. Vallejo, and Yu. Vorobiev. A perturbation theory approach to the stability of the Pais–Uhlenbeck oscillator. Journal of Mathematical Physics, 58(9):093501, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bourget:2017:RUR

[1268] Olivier Bourget, Víctor H. Cortés, Rafael Del Río, and Claudio Fernández. Resonances under rank-one perturbations. Journal of Mathematical Physics, 58(9):093502, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Efremov:2017:RSY

[1269] Alexander N. Efremov, Riccardo Guida, and Christoph Kopper. Renormalization of SU(2) Yang–Mills theory with flow equations. Journal of Mathematical Physics, 58(9):093503, September 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boos:2017:OVP

[1270] Hermann Boos, Frank Göhmann, Andreas Klümper, Khazret S. Nirov, and Alexander V. Razumov. Oscillator versus prefundamental representations. II. Arbitrary higher ranks. Journal of Mathematical Physics, 58 (9):093504, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bogdanov:2017:LDH

[1271] L. V. Bogdanov and M. V. Pavlov. Linearly degenerate hierarchies of quasiclassical SDYM type. Journal of Mathematical Physics, 58(9):093505, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kong:2017:SWS

[1272] Fanchao Kong, Zhiguo Luo, and Fulai Chen. Solitary wave solutions for singular non-Newtonian filtration equations. Journal of Mathematical Physics, 58(9):093506, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pocheketa:2017:ESA

[1273] Oleksandr A. Pocheketa and Roman O. Popovych. Extended symmetry analysis of generalized Burgers equations. Journal of Mathematical Physics, 58(10):101501, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sekhar:2017:EWI

[1274] T. Raja Sekhar and Minhajul. Elementary wave interactions in blood flow through artery. Journal of Mathematical Physics, 58(10):101502, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2017:VIB

[1275] Xiaoliang Li and Baiyu Liu. Vacuum isolating, blow up threshold, and asymptotic behavior of solutions for a nonlocal parabolic equation. Journal of Mathematical Physics, 58(10):101503, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaneeva:2017:EGC

[1276] Olena Vaneeva and Severin Posta. Equivalence groupoid of a class of variable coefficient Korteweg–de Vries equations. Journal of Mathematical Physics, 58(10):101504, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jamreh:2017:CPR

[1277] Maryam Jamreh and Mehdi Nadjafikhah. Closed pseudo-Riemannian Ricci solitons. Journal of Mathematical Physics, 58(10):101505, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Du:2017:SNI

[1278] Linglong Du and Zhigang Wu. Solving the non-isentropic Navier–Stokes equations in odd space dimensions: The Green function method. Journal of Mathematical Physics, 58(10):101506, October 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2017:VVL

[1279] Ya-Guang Wang, Jierong Yin, and Shiyong Zhu. Vanishing viscosity limit for incompressible Navier–Stokes equations with Navier boundary conditions for small slip length. Journal of Mathematical Physics, 58 (10):101507, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antonic:2017:CFS

[1280] Nenad Antonić, Kresimir Burazin, Ivana Crnjac, and Marko Erceg. Complex Friedrichs systems and applications. Journal of Mathematical Physics, 58(10):101508, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaudreau:2017:DES

[1281] Philippe Gaudreau and Hassan Safouhi. Double exponential sinccollocation method for solving the energy eigenvalues of harmonic oscillators perturbed by a rational function. Journal of Mathematical Physics, 58(10):101509, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2017:PSS

[1282] Changhong Guo and Shaomei Fang. Planar, solitary, and spiral waves of the Burgers–CGL equations for flames governed by a sequential reaction. Journal of Mathematical Physics, 58(10):101510, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2017:CSS

[1283] Yuan Zhou and Wen-Xiu Ma. Complexiton solutions to soliton equations by the Hirota method. Journal of Mathematical Physics, 58(10):101511, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Raposo:2017:HLT

[1284] C. A. Raposo, O. V. Villagrán, J. E. Muñoz Rivera, and M. S. Alves. Hybrid laminated Timoshenko beam. Journal of Mathematical Physics, 58 (10):101512, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2017:DDM

[1285] Mi-Ran Choi, Dirk Hundertmark, and Young-Ran Lee. Discrete diffraction managed solitons: Threshold phenomena and rapid decay for general nonlinearities. Journal of Mathematical Physics, 58(10):101513, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2017:LSA

[1286] Zhonglong Zhao and Bo Han. Lie symmetry analysis, Bäcklund transformations, and exact solutions of a (2 + 1)-dimensional Boiti–Leon– Pempinelli system. Journal of Mathematical Physics, 58(10):101514, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2017:GEL

[1287] Min Li and Zhaoyang Yin. Global existence and local well-posedness of the single-cycle pulse equation. Journal of Mathematical Physics, 58 (10):101515, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hong:2017:RCD

[1288] Jin Hong and Hyeonmi Lee. Rigged configuration descriptions of the crystals B(∞) and B(λ) for special linear Lie algebras. Journal of Mathematical Physics, 58(10):101701, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibort:2017:NAC

[1289] A. Ibort, A. López Yela, and J. Moro. A new algorithm for computing branching rules and Clebsch–Gordan coefficients of unitary representations of compact groups. Journal of Mathematical Physics, 58(10): 101702, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nicoleau:2017:DLT

[1290] F. Nicoleau, D. Parra, and S. Richard. Does Levinson’s theorem count complex eigenvalues? Journal of Mathematical Physics, 58(10):102101, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:2017:ABE

[1291] César R. de Oliveira and Renan G. Romano. Aharonov–Bohm effect without contact with the solenoid. Journal of Mathematical Physics, 58 (10):102102, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fonseca:2017:SAI

[1292] I. C. Fonseca and K. Bakke. Some aspects of the interaction of a magnetic quadrupole moment with an electric field in a rotating frame. Journal of Mathematical Physics, 58(10):102103, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferreira:2017:SRC

[1293] F. J. S. Ferreira and V. B. Bezerra. Some remarks concerning the centrifugal term approximation. Journal of Mathematical Physics, 58(10): 102104, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Love:2017:TPC

[1294] S. T. Love. Transition probability for class of two level system time dependent Hamiltonians. Journal of Mathematical Physics, 58(10):102105, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Przanowski:2017:DWW

[1295] Maciej Przanowski and Jaromir Tosiek. From the discrete Weyl–Wigner formalism for symmetric ordering to a number-phase Wigner function. Journal of Mathematical Physics, 58(10):102106, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weder:2017:NEM

[1296] Ricardo Weder. The number of eigenvalues of the matrix Schrödinger operator on the half line with general boundary conditions. Journal of Mathematical Physics, 58(10):102107, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1947].

Hirota:2017:REN

[1297] K. Hirota. Real eigenvalues of a non-self-adjoint perturbation of the selfadjoint Zakharov–Shabat operator. Journal of Mathematical Physics, 58 (10):102108, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vecsernyes:2017:ETM

[1298] Péter Vecsernyés. An effective toy model in Mn(C) for selective measurements in quantum mechanics. Journal of Mathematical Physics, 58 (10):102109, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karthiga:2017:QSG

[1299] S. Karthiga, V. Chithiika Ruby, M. Senthilvelan, and M. Lakshmanan. Quantum solvability of a general ordered position dependent mass system: Mathews–Lakshmanan oscillator. Journal of Mathematical Physics, 58(10):102110, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2017:RPD

[1300] Kyung Hoon Han and Seung-Hyeok Kye. The role of phases in detecting three-qubit entanglement. Journal of Mathematical Physics, 58(10): 102201, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirokov:2017:TUC

[1301] M. E. Shirokov. Tight uniform continuity bounds for the quantum conditional mutual information, for the Holevo quantity, and for capacities of quantum channels. Journal of Mathematical Physics, 58(10):102202, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2017:CQA

[1302] Holger Boche, Minglai Cai, Christian Deppe, and Janis Nötzel. Classicalquantum arbitrarily varying wiretap channel: Secret message transmission under jamming attacks. Journal of Mathematical Physics, 58(10): 102203, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mathieu:2017:ATV

[1303] Ph. Mathieu and F. Thuillier. Abelian Turaev-Virelizier theorem and U(1) BF surgery formulas. Journal of Mathematical Physics, 58(10): 102301, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okolow:2017:CPQ

[1304] Andrzej Okolów. Constrained projective quantum states for the degenerate Plebański gravity. Journal of Mathematical Physics, 58(10):102302, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ida:2017:UBD

[1305] Daisuke Ida. Upper bound for diameter of cosmological black holes and nonexistence of black strings. Journal of Mathematical Physics, 58(10): 102501, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avalos:2017:PTS

[1306] R. Avalos, F. Dahia, C. Romero, and J. H. Lira. On the proof of the thin sandwich conjecture in arbitrary dimensions. Journal of Mathematical Physics, 58(10):102502, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2017:QPS

[1307] Jing Li. Quasi-periodic solutions of a fractional nonlinear Schrödinger equation. Journal of Mathematical Physics, 58(10):102701, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2017:LTL

[1308] Guangming Pan, Shaochen Wang, and Wang Zhou. Limit theorems for linear spectrum statistics of orthogonal polynomial ensembles and their applications in random matrix theory. Journal of Mathematical Physics, 58(10):103301, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Puertas-Centeno:2017:EUM

[1309] D. Puertas-Centeno, N. M. Temme, I. V. Toranzo, and J. S. Dehesa. Entropic uncertainty measures for large dimensional hydrogenic systems. Journal of Mathematical Physics, 58(10):103302, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frederiksen:2017:QDI

[1310] Jorgen S. Frederiksen. Quasi-diagonal inhomogeneous closure for classical and quantum statistical dynamics. Journal of Mathematical Physics, 58(10):103303, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2017:SSG

[1311] Li-Yuan Ma, Shou-Feng Shen, and Zuo-Nong Zhu. Soliton solution and gauge equivalence for an integrable nonlocal complex modified Korteweg– de Vries equation. Journal of Mathematical Physics, 58(10):103501, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghosh:2017:SMP

[1312] Amalendu Ghosh and Ramesh Sharma. Sasakian manifolds with purely transversal Bach tensor. Journal of Mathematical Physics, 58(10): 103502, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alama:2017:GSL

[1313] Stan Alama, Lia Bronsard, Rustum Choksi, and Ihsan Topaloglu. Ground-states for the liquid drop and TFDW models with long-range attraction. Journal of Mathematical Physics, 58(10):103503, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hrivnak:2017:DTC

[1314] Jirí Hrivnák and Michal Juránek. On E-discretization of tori of compact simple Lie groups. II. Journal of Mathematical Physics, 58(10):103504, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atakishiyev:2017:IEZ

[1315] Natig M. Atakishiyev, George S. Pogosyan, Kurt Bernardo Wolf, and Alexander Yakhno. Interbasis expansions in the Zernike system. Journal of Mathematical Physics, 58(10):103505, October 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popowicz:2017:LRM

[1316] Z. Popowicz. Lax representations for matrix short pulse equations. Journal of Mathematical Physics, 58(10):103506, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pena:2017:SBF

[1317] Dixan Peña Peña, Irene Sabadini, and Franciscus Sommen. Segal– Bargmann–Fock modules of monogenic functions. Journal of Mathematical Physics, 58(10):103507, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conte:2017:GBS

[1318] Robert Conte. Generalized Bonnet surfaces and Lax pairs of P_{VI}. Journal of Mathematical Physics, 58(10):103508, October 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiao:2017:CEC

[1319] Li Guang Jiao and Li Rong Zan. Comment on “Entropy and complexity analysis of hydrogenic Rydberg atoms” [j. math. phys. 54, 052109 (2013)]. Journal of Mathematical Physics, 58(10):104101, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [12] and response [1320].

Lopez-Rosa:2017:RCE

[1320] S. López-Rosa, I. V. Toranzo, P. Sánchez-Moreno, and J. S. Dehesa. Response to “Comment on ‘Entropy and complexity analysis of hydrogenic Rydberg atoms”’ [j. math. phys. 58, 104101 (2017)]. Journal of Mathematical Physics, 58(10):104102, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [12, 1319].

Zheng:2017:PPT

[1321] Pan Zheng, Chunlai Mu, and Xuegang Hu. Persistence property in a two-species chemotaxis system with two signals. Journal of Mathematical Physics, 58(11):111501, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2017:EAP

[1322] Youjun Wang and Qing Li. Existence and asymptotic profiles of positive solutions of quasilinear Schrödinger equations in R^{3}. Journal of Mathematical Physics, 58(11):111502, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez-Caudevilla:2017:HOS

[1323] P. Álvarez-Caudevilla, Eduardo Colorado, and Rasiel Fabelo. A higher order system of some coupled nonlinear Schrödinger and Korteweg– de Vries equations. Journal of Mathematical Physics, 58(11):111503, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2017:GEA

[1324] Qingshan Zhang, Xiaopan Liu, and Xiaofei Yang. Global existence and asymptotic behavior of solutions to a two-species chemotaxis system with two chemicals. Journal of Mathematical Physics, 58(11):111504, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ali:2017:MMI

[1325] K. Ben Ali, A. Ghanmi, and K. Kefi. Minimax method involving singular p(x)-Kirchhoff equation. Journal of Mathematical Physics, 58(11): 111505, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramsey:2017:SGD

[1326] Scott D. Ramsey and Roy S. Baty. Symmetries of the gas dynamics equations using the differential form method. Journal of Mathematical Physics, 58(11):111506, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2017:EMS

[1327] Claudianor O. Alves and César E. Torres Ledesma. Existence and multiplicity of solutions for a non-linear Schrödinger equation with nonlocal regional diffusion. Journal of Mathematical Physics, 58(11):111507, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2017:BMT

[1328] Youngae Lee. Bubbling mixed type solutions of the SU(3) models on a torus. Journal of Mathematical Physics, 58(11):111508, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ran:2017:KGS

[1329] Yanping Ran and Qihong Shi. Klein–Gordon–Schrödinger system: Dinucleon field. Journal of Mathematical Physics, 58(11):111509, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavlov:2017:TDR

[1330] Maxim V. Pavlov and Nikola M. Stoilov. Three dimensional reductions of four-dimensional quasilinear systems. Journal of Mathematical Physics, 58(11):111510, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2017:ETM

[1331] Jongmin Han and Juhee Sohn. Existence of topological multi-string solutions in Abelian gauge field theories. Journal of Mathematical Physics, 58(11):111511, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pandzic:2017:BCS

[1332] Pavle Pandžić and Vladimír Souček. BGG complexes in singular infinitesimal character for type A. Journal of Mathematical Physics, 58(11): 111512, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joudioux:2017:GCH

[1333] Jérémie Joudioux. Gluing for the constraints for higher spin fields. Journal of Mathematical Physics, 58(11):111513, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poletaeva:2017:FWA

[1334] Elena Poletaeva and Vera Serganova. On the finite W -algebra for the Lie superalgebra Q(N) in the non-regular case. Journal of Mathematical Physics, 58(11):111701, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2017:DGD

[1335] Jun Zhao, Liangyun Chen, and Lamei Yuan. Deformations and generalized derivations of Lie conformal superalgebras. Journal of Mathematical Physics, 58(11):111702, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lauret:2017:MFF

[1336] Emilio A. Lauret and Fiorela Rossi Bertone. Multiplicity formulas for fundamental strings of representations of classical Lie algebras. Journal of Mathematical Physics, 58(11):111703, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bausch:2017:CTI

[1337] Johannes Bausch and Stephen Piddock. The complexity of translationally invariant low-dimensional spin lattices in 3D. Journal of Mathematical Physics, 58(11):111901, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Appleby:2017:DTS

[1338] Marcus Appleby, Ingemar Bengtsson, Irina Dumitru, and Steven Flammia. Dimension towers of SICs. I. Aligned SICs and embedded tight frames. Journal of Mathematical Physics, 58(11):112201, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aguillon:2017:NRG

[1339] Cesar A. Aguillón, Albert Much, Marcos Rosenbaum, and J. David Vergara. Noncommutative Riemannian geometry from quantum spacetime generated by twisted Poincaré group. Journal of Mathematical Physics, 58(11):112301, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mihaylov:2017:TET

[1340] Deyan P. Mihaylov and Jonathan R. Gair. Transition of EMRIs through resonance: Corrections to higher order in the on-resonance flux modification. Journal of Mathematical Physics, 58(11):112501, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chudecki:2017:GCN

[1341] Adam Chudecki. On geometry of congruences of null strings in 4dimensional complex and real pseudo-Riemannian spaces. Journal of Mathematical Physics, 58(11):112502, November 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samann:2017:GNI

[1342] Clemens Sämann and Roland Steinbauer. Geodesics in nonexpanding impulsive gravitational waves with Λ. II. Journal of Mathematical Physics, 58(11):112503, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osetrin:2017:PRS

[1343] Evgeny Osetrin and Konstantin Osetrin. Pure radiation in space-time models that admit integration of the eikonal equation by the separation of variables method. Journal of Mathematical Physics, 58(11):112504, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herfray:2017:PCF

[1344] Yannick Herfray. Pure connection formulation, twistors, and the chase for a twistor action for general relativity. Journal of Mathematical Physics, 58(11):112505, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bialy:2017:SPS

[1345] Misha Bialy and Andrey E. Mironov. In search of periodic solutions for a reduction of the Benney chain. Journal of Mathematical Physics, 58(11): 112701, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ban:2017:MPT

[1346] Jung-Chao Ban and Chih-Hung Chang. Mixing properties of tree-shifts. Journal of Mathematical Physics, 58(11):112702, November 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadha:2017:SAN

[1347] Alka Chadha and Swaroop Nandan Bora. Stability analysis for neutral stochastic differential equation of second order driven by Poisson jumps. Journal of Mathematical Physics, 58(11):112703, November 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khamehchi:2017:ESG

[1348] Masoumeh Khamehchi and Farhad H. Jafarpour. Effective stochastic generator with site-dependent interactions. Journal of Mathematical Physics, 58(11):113301, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Slyusarenko:2017:KTW

[1349] Yurii V. Slyusarenko and Oleksii Yu. Sliusarenko. Kinetic theory of weakly ionized dilute gas of hydrogen-like atoms of the first principles of quantum statistics and dispersion laws of eigenwaves. Journal of Mathematical Physics, 58(11):113302, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2017:LEL

[1350] Peter J. Forrester, Anas A. Rahman, and Nicholas S. Witte. Large N expansions for the Laguerre and Jacobi β-ensembles from the loop equations. Journal of Mathematical Physics, 58(11):113303, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2017:GSS

[1351] N. Aizawa and J. Segar. Z_{2} × Z_{2} generalizations of N = 2 super Schrödinger algebras and their representations. Journal of Mathematical Physics, 58(11):113501, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2017:EEF

[1352] Wu-Sheng Dai and Mi Xie. The explicit expression of the fugacity for weakly interacting Bose and Fermi gases. Journal of Mathematical Physics, 58(11):113502, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zapata:2017:FLT

[1353] J. P. Arias Zapata, A. Belokogne, E. Huguet, J. Queva, J. Renaud, et al. Friedmann–Lemaître–Robertson–Walker spaces as submanifolds of R^{6}: Restriction to the Klein–Gordon operator. Journal of Mathematical Physics, 58(11):113503, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manibo:2017:LEB

[1354] Neil Mañibo. Lyapunov exponents for binary substitutions of constant length. Journal of Mathematical Physics, 58(11):113504, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2017:QTS

[1355] Qiufang Liu and Chuanzhong Li. Quantum torus symmetries of the CKP and multi-component CKP hierarchies. Journal of Mathematical Physics, 58(11):113505, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kallinikos:2017:GCC

[1356] N. Kallinikos. Group classification of charged particle motion in stationary electromagnetic fields. Journal of Mathematical Physics, 58(11): 113506, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2017:BST

[1357] Axel Schulze-Halberg and Pinaki Roy. Bound states of the twodimensional Dirac equation for an energy-dependent hyperbolic Scarf potential. Journal of Mathematical Physics, 58(11):113507, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kundu:2017:CCL

[1358] Prasun K. Kundu. Comment on “Comments on ‘The Euclidean gravitational action as black hole entropy, singularities and space-time voids’ ” [J. Math. Phys. 50, 042502 (2009)] — Schwarzschild black hole lives to fight another day. Journal of Mathematical Physics, 58(11):114101, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4].

Huang:2017:EAL

[1359] Qing Huang and Shoufeng Shen. Erratum and addendum: “Lie symmetries and group classification of a class of time fractional evolution systems” [J. Math. Phys. 56, 123504 (2015)]. Journal of Mathematical Physics, 58(11):119901, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [464].

Derezinski:2017:BQH

[1360] Jan Dereziński. Bosonic quadratic Hamiltonians. Journal of Mathematical Physics, 58(12):121101, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Denlinger:2017:SCB

[1361] Ryan Denlinger. Structure of correlations for the Boltzmann–Grad limit of hard spheres. Journal of Mathematical Physics, 58(12):121501, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borisov:2017:SFQ

[1362] Denis Borisov, Martin Tautenhahn, and Ivan Veselić. Scale-free quantitative unique continuation and equidistribution estimates for solutions of elliptic differential equations. Journal of Mathematical Physics, 58(12): 121502, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Idelon-Riton:2017:EFM

[1363] Guillaume Idelon-Riton. Explicit formula and meromorphic extension of the resolvent for the massive Dirac operator in the Schwarzschild-anti-de Sitter spacetime. Journal of Mathematical Physics, 58(12):121503, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsenko:2017:VGF

[1364] A. A. Bytsenko, M. Chaichian, and R. Luna. Vector generation functions, q-spectral functions of hyperbolic geometry, and vertex operators for quantum affine algebras. Journal of Mathematical Physics, 58(12): 121701, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelasCuevas:2017:IFM

[1365] Gemma De las Cuevas, J. Ignacio Cirac, Norbert Schuch, and David Perez-Garcia. Irreducible forms of matrix product states: Theory and applications. Journal of Mathematical Physics, 58(12):121901, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nathanson:2017:TLF

[1366] Ekaterina S. Nathanson and Palle E. T. Jørgensen. Trotter’s limit formula for the Schrödinger equation with singular potential. Journal of Mathematical Physics, 58(12):122101, December 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aleixo:2017:IGA

[1367] A. N. F. Aleixo and A. B. Balantekin. Inverse generators in the algebraic description of quantum confined systems: Properties and connections with nonlinear quantum deformation process and phase operators. Journal of Mathematical Physics, 58(12):122102, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kermani:2017:GSA

[1368] Maryam Mansoori Kermani. Ground state atoms confined in a real Rydberg and complex Rydberg–Scarf II potential. Journal of Mathematical Physics, 58(12):122103, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haapasalo:2017:OQO

[1369] Erkka Haapasalo and Juha-Pekka Pellonpää. Optimal quantum observables. Journal of Mathematical Physics, 58(12):122104, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2017:LES

[1370] Yu Cao and Jianfeng Lu. Lindblad equation and its semiclassical limit of the Anderson–Holstein model. Journal of Mathematical Physics, 58(12): 122105, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolotin:2017:QS

[1371] Arkady Bolotin. Quantum supervaluationism. Journal of Mathematical Physics, 58(12):122106, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kravchenko:2017:ARS

[1372] Vladislav V. Kravchenko and Sergii M. Torba. Asymptotics with respect to the spectral parameter and Neumann series of Bessel functions for solutions of the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 58(12):122107, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samar:2017:EST

[1373] M. I. Samar and V. M. Tkachuk. Exact solutions for two-body problems in 1D deformed space with minimal length. Journal of Mathematical Physics, 58(12):122108, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grassl:2017:FLS

[1374] Markus Grassl and Andrew J. Scott. Fibonacci–Lucas SIC–POVMs. Journal of Mathematical Physics, 58(12):122201, December 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Volkoff:2017:MTD

[1375] T. J. Volkoff. Maximal trace distance between isoenergetic bosonic Gaussian states. Journal of Mathematical Physics, 58(12):122202, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harris:2017:CEP

[1376] Samuel J. Harris. Connes’ embedding problem and winning strategies for quantum XOR games. Journal of Mathematical Physics, 58(12):122203, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krumnow:2017:FFT

[1377] Christian Krumnow, Zoltán Zimborás, and Jens Eisert. A fermionic de Finetti theorem. Journal of Mathematical Physics, 58(12):122204, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandenHoogen:2017:TCS

[1378] R. J. van den Hoogen. Towards a covariant smoothing procedure for gravitational theories. Journal of Mathematical Physics, 58(12):122501, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Searight:2017:DMD

[1379] Trevor P. Searight. On degenerate metrics, dark matter and unification. Journal of Mathematical Physics, 58(12):122502, December 2017. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Jiang:2017:WZA

[1380] Tao Jiang, Xianming Liu, and Jinqiao Duan. A Wong–Zakai approximation for random invariant manifolds. Journal of Mathematical Physics, 58(12):122701, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2017:SCP

[1381] Eunghyun Lee. Some conditional probabilities in the TASEP with second class particles. Journal of Mathematical Physics, 58(12):123301, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tracy:2017:BAS

[1382] Craig A. Tracy and Harold Widom. Blocks in the asymmetric simple exclusion process. Journal of Mathematical Physics, 58(12):123302, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez:2017:PMC

[1383] J. Cerda Hernández. Potts model coupled to random causal triangulations. Journal of Mathematical Physics, 58(12):123303, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lunardi:2017:NDK

[1384] José T. Lunardi. A note on the Duffin–Kemmer–Petiau equation in (1+1) space-time dimensions. Journal of Mathematical Physics, 58(12):123501, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2017:SUM

[1385] Hongmei Chen, Min Chen, Gordon Blower, and Yang Chen. Single-user MIMO system, painlevé transcendents, and double scaling. Journal of Mathematical Physics, 58(12):123502, December 2017. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2017:SIH

[1386] Baiying He and Liangyun Chen. Super integrable hierarchies associated with spl(2, 1) and Darboux transformations. Journal of Mathematical Physics, 58(12):123503, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willox:2017:SMC

[1387] R. Willox, A. Ramani, and B. Grammaticos. A systematic method for constructing discrete Painlevé equations in the degeneration cascade of the E_{8} group. Journal of Mathematical Physics, 58(12):123504, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hartung:2017:RFP

[1388] Tobias Hartung. Regularizing Feynman path integrals using the generalized Kontsevich–Vishik trace. Journal of Mathematical Physics, 58(12): 123505, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2017:ERL

[1389] Giuseppe Gaeta and Francesco Spadaro. Erratum: “Random Lie-point symmetries of stochastic differential equations” [J. Math. Phys. 58, 053503 (2017)]. Journal of Mathematical Physics, 58(12):129901, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1120].

Ablowitz:2018:IST

[1390] Mark J. Ablowitz, Xu-Dan Luo, and Ziad H. Musslimani. Inverse scattering transform for the nonlocal nonlinear Schrödinger equation with nonzero boundary conditions. Journal of Mathematical Physics, 59(1): 011501, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pasquali:2018:AGE

[1391] S. Pasquali. Almost global existence for the nonlinear Klein–Gordon equation in the nonrelativistic limit. Journal of Mathematical Physics, 59(1):011502, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deng:2018:TBU

[1392] Yinbin Deng, Yujin Guo, and Lu Lu. Threshold behavior and uniqueness of ground states for mass critical inhomogeneous Schrödinger equations. Journal of Mathematical Physics, 59(1):011503, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2018:PSS

[1393] Yanheng Ding and Xiaoying Liu. Periodic solutions of superlinear Dirac equations with perturbations from symmetry. Journal of Mathematical Physics, 59(1):011504, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2018:GEF

[1394] Hui Wu and Xiaoping Yang. Global existence and finite time blow-up for a parabolic system on hyperbolic space. Journal of Mathematical Physics, 59(1):011505, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iturriaga:2018:NCC

[1395] Leonelo Iturriaga and Eugenio Massa. On necessary conditions for the comparison principle and the sub- and supersolution method for the stationary Kirchhoff equation. Journal of Mathematical Physics, 59(1): 011506, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:EIW

[1396] Xin Liu. On the energy inequality for weak solutions to the threedimensional compressible magnetohydrodynamic equations. Journal of Mathematical Physics, 59(1):011507, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peng:2018:SSW

[1397] Congming Peng and Qihong Shi. Stability of standing wave for the fractional nonlinear Schrödinger equation. Journal of Mathematical Physics, 59(1):011508, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arbunich:2018:RDN

[1398] Jack Arbunich and Christof Sparber. Rigorous derivation of nonlinear Dirac equations for wave propagation in honeycomb structures. Journal of Mathematical Physics, 59(1):011509, January 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2018:APS

[1399] Yuhua Sun. Absence of positive solutions to the system of differential inequalities on manifolds. Journal of Mathematical Physics, 59(1):011510, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2018:ASF

[1400] Zaihong Jiang, Jinbo Geng, Li Li, and Ning Zhong. Asymptotic stability of a fluid-particles interaction model with external forces. Journal of Mathematical Physics, 59(1):011511, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amara:2018:NBC

[1401] Jamel Ben Amara and Hedi Bouzidi. Null boundary controllability of a one-dimensional heat equation with an internal point mass and variable coefficients. Journal of Mathematical Physics, 59(1):011512, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2018:TWD

[1402] Shou-Peng Zhang, Yun-Rui Yang, and Yong-Hui Zhou. Traveling waves in a delayed SIR model with nonlocal dispersal and nonlinear incidence. Journal of Mathematical Physics, 59(1):011513, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Afra:2018:RWR

[1403] M. Afra, H. Fakhri, and M. Sayyah-Fard. A reducible Weil representation of sp(4) realized by differential operators in the space of smooth functions on H^{2} × S^{1}. Journal of Mathematical Physics, 59(1):011701, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruillard:2018:CRP

[1404] Paul Bruillard and Carlos M. Ortiz-Marrero. Classification of rank 5 premodular categories. Journal of Mathematical Physics, 59(1):011702, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashida:2018:STM

[1405] Sohei Ashida. Scattering theory for multistate Schrödinger operators. Journal of Mathematical Physics, 59(1):012101, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brewster:2018:GDF

[1406] R. A. Brewster and J. D. Franson. Generalized delta functions and their use in quantum optics. Journal of Mathematical Physics, 59(1):012102, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DiMartino:2018:FGL

[1407] Sara Di Martino, Paolo Facchi, and Giuseppe Florio. Feynman graphs and the large dimensional limit of multipartite entanglement. Journal of Mathematical Physics, 59(1):012201, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iten:2018:SMS

[1408] Raban Iten and Roger Colbeck. Smooth manifold structure for extreme channels. Journal of Mathematical Physics, 59(1):012202, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zanardi:2018:QCG

[1409] Paolo Zanardi and Lorenzo Campos Venuti. Quantum coherence generating power, maximally abelian subalgebras, and Grassmannian geometry. Journal of Mathematical Physics, 59(1):012203, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2018:ABS

[1410] Pavel Exner, Sylwia Kondej, and Vladimir Lotoreichik. Asymptotics of the bound state induced by δ-interaction supported on a weakly deformed plane. Journal of Mathematical Physics, 59(1):013501, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vinogradov:2018:PLS

[1411] A. M. Vinogradov. Particle-like structure of coaxial Lie algebras. Journal of Mathematical Physics, 59(1):011703, January 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baseilhac:2018:BHB

[1412] Pascal Baseilhac and Xavier Martin. A bispectral q-hypergeometric basis for a class of quantum integrable models. Journal of Mathematical Physics, 59(1):011704, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samajdar:2018:EEA

[1413] Rhine Samajdar and Sudhir R. Jain. Exact eigenfunction amplitude distributions of integrable quantum billiards. Journal of Mathematical Physics, 59(1):012103, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sandev:2018:GTD

[1414] Trifce Sandev, Irina Petreska, and Ervin K. Lenzi. Generalized timedependent Schrödinger equation in two dimensions under constraints.

Journal of Mathematical Physics, 59(1):012104, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andrade:2018:NMS

[1415] M. A. De Andrade and C. Neves. Noncommutative mapping from the symplectic formalism. Journal of Mathematical Physics, 59(1):012105, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:ISP

[1416] Yongxia Guo and Guangsheng Wei. Inverse scattering problems where the potential is not absolutely continuous on the known interior subinterval. Journal of Mathematical Physics, 59(1):012106, January 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akbas:2018:BOA

[1417] Haci Akbas and O. Teoman Turgut. Born–Oppenheimer approximation for a singular system. Journal of Mathematical Physics, 59(1):012107, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dias:2018:WFN

[1418] Nuno Costa Dias and João Nuno Prata. Wigner functions on nonstandard symplectic vector spaces. Journal of Mathematical Physics, 59 (1):012108, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beaud:2018:BEE

[1419] V. Beaud and S. Warzel. Bounds on the entanglement entropy of droplet states in the XXZ spin chain. Journal of Mathematical Physics, 59(1): 012109, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lemm:2018:MTI

[1420] Marius Lemm. On multivariate trace inequalities of Sutter, Berta, and Tomamichel. Journal of Mathematical Physics, 59(1):012204, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:2018:FLQ

[1421] Siddhartha Das, Sumeet Khatri, George Siopsis, and Mark M. Wilde. Fundamental limits on quantum dynamics based on entropy change.

Journal of Mathematical Physics, 59(1):012205, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:2018:POP

[1422] Christopher King. Potential output purity of completely positive maps. Journal of Mathematical Physics, 59(1):012206, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimock:2018:URQ

[1423] J. Dimock. Ultraviolet regularity for QED in d = 3. Journal of Mathematical Physics, 59(1):012301, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gay-Balmaz:2018:DSN

[1424] François Gay-Balmaz and Hiroaki Yoshimura. Dirac structures in nonequilibrium thermodynamics. Journal of Mathematical Physics, 59 (1):012701, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amore:2018:HOA

[1425] Paolo Amore, John P. Boyd, and Francisco M. Fernández. High order analysis of the limit cycle of the van der Pol oscillator. Journal of Mathematical Physics, 59(1):012702, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bilogliadov:2018:MRE

[1426] Mykhailo Bilogliadov. Minimum Riesz energy problem on the hyperdisk. Journal of Mathematical Physics, 59(1):013301, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peng:2018:MPA

[1427] Liangrong Peng, Yi Zhu, and Liu Hong. The Markov process admits a consistent steady-state thermodynamic formalism. Journal of Mathematical Physics, 59(1):013302, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mavi:2018:GSE

[1428] Rajinder Mavi and Mei Yin. Ground states for exponential random graphs. Journal of Mathematical Physics, 59(1):013303, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Annaby:2018:SIE

[1429] M. H. Annaby and R. M. Asharabi. Sinc-interpolants in the energy plane for regular solution, Jost function, and its zeros of quantum scattering. Journal of Mathematical Physics, 59(1):013502, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2018:OPD

[1430] A. D. Alhaidari. Orthogonal polynomials derived from the tridiagonal representation approach. Journal of Mathematical Physics, 59(1):013503, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2018:OPH

[1431] Satoru Odake and Ryu Sasaki. Orthogonal polynomials from Hermitian matrices. II. Journal of Mathematical Physics, 59(1):013504, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarrico:2018:NSL

[1432] C. O. R. Sarrico and A. Paiva. Newton’s second law and the multiplication of distributions. Journal of Mathematical Physics, 59(1):013505, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giusti:2018:DRT

[1433] Andrea Giusti. Dispersion relations for the time-fractional Cattaneo– Maxwell heat equation. Journal of Mathematical Physics, 59(1):013506, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2328].

Conache:2018:GSC

[1434] Diana Conache, Alexei Daletskii, Yuri Kondratiev, and Tanja Pasurek. Gibbs states of continuum particle systems with unbounded spins: Existence and uniqueness. Journal of Mathematical Physics, 59(1):013507, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Graham:2018:FDL

[1435] Curtis Graham, Klaus Kirsten, Pedro Morales-Almazan, and Brian Quantz Streit. Functional determinants for Laplacians on annuli and elliptical regions. Journal of Mathematical Physics, 59(1):013508, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chulaevsky:2018:KAL

[1436] Victor Chulaevsky. The KAM approach to the localization in “haarsch” quasi-periodic media. Journal of Mathematical Physics, 59(1):013509, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ucgun:2018:RTM

[1437] Filiz Çagatay Uçgun, Ogul Esen, and Hasan Gümral. Reductions of topologically massive gravity I: Hamiltonian analysis of second order degenerate Lagrangians. Journal of Mathematical Physics, 59(1):013510, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordon:2018:EWL

[1438] James Gordon and Gordon W. Semenoff. Erratum: “World-line instantons and the Schwinger effect as a Wentzel–Kramers–Brillouin exact path integral” [j. math. phys. 56, 022111 (2015)]. Journal of Mathematical Physics, 59(1):019901, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [85].

Kirsten:2018:EBF

[1439] Klaus Kirsten and Yoonweon Lee. Erratum: “The Burghelea– Friedlander–Kappeler-gluing formula for zeta-determinants on a warped product manifold and a product manifold” [J. Math. Phys. 56, 123501 (2015)]. Journal of Mathematical Physics, 59(1):019902, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [498].

Wei:2018:BCP

[1440] L. Wei and Y. Wang. Blowup criterion and persistent decay for a modified Camassa–Holm system. Journal of Mathematical Physics, 59(2):021501, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cappiello:2018:CDS

[1441] Marco Cappiello, René Schulz, and Patrik Wahlberg. Conormal distributions in the Shubin calculus of pseudodifferential operators. Journal of Mathematical Physics, 59(2):021502, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fang:2018:PSF

[1442] Daoyuan Fang, Zheng Han, and Weimin Wang. Point spectrum of the Floquet Hamiltonian for Klein–Gordon equation under quasi-periodic perturbations. Journal of Mathematical Physics, 59(2):021503, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bracken:2018:CFA

[1443] Paul Bracken. Cartan frames and algebras with links to integrable systems differential equations and surfaces. Journal of Mathematical Physics, 59(2):021504, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:ENP

[1444] Jianhua Chen, Xianhua Tang, and Bitao Cheng. Existence and nonexistence of positive solutions for a class of generalized quasilinear Schrödinger equations involving a Kirchhoff-type perturbation with critical Sobolev exponent. Journal of Mathematical Physics, 59(2):021505, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naumkin:2018:HOD

[1445] Pavel I. Naumkin and Jhon J. Perez. Higher-order derivative nonlinear Schrödinger equation in the critical case. Journal of Mathematical Physics, 59(2):021506, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2018:OSO

[1446] Chunyu Shen. Optimal solution and optimality condition of the Hunter– Saxton equation. Journal of Mathematical Physics, 59(2):021507, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:DBS

[1447] C. X. Guo, C. Y. Li, and Y. Q. Han. Dynamical behaviors of stochastic Hasegawa–Mima equation in torus. Journal of Mathematical Physics, 59 (2):021508, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:CPG

[1448] Suhong Li, Jiongyue Li, and Yu Chen. On concentration of positive ground states for the asymptotically linear fractional Schrödinger equation. Journal of Mathematical Physics, 59(2):021509, February 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meljanac:2018:GWR

[1449] Stjepan Meljanac, Sasa Kresić-Jurić, and Danijel Pikutić. Generalization of Weyl realization to a class of Lie superalgebras. Journal of Mathematical Physics, 59(2):021701, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havlicek:2018:CRP

[1450] Miloslav Havlícek, Jan Kotrbatý, Patrick Moylan, and Severin Posta. Construction of representations of Poincaré group using Lie fields. Journal of Mathematical Physics, 59(2):021702, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bols:2018:ALB

[1451] Alex Bols and Wojciech De Roeck. Asymptotic localization in the Bose– Hubbard model. Journal of Mathematical Physics, 59(2):021901, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:2018:GIC

[1452] Andras Molnar, Yimin Ge, Norbert Schuch, and J. Ignacio Cirac. A generalization of the injectivity condition for projected entangled pair states. Journal of Mathematical Physics, 59(2):021902, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2018:QNE

[1453] N. Aizawa, Z. Kuznetsova, and F. Toppan. The quasi-nonassociative exceptional F (4) deformed quantum oscillator. Journal of Mathematical Physics, 59(2):022101, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoffmann:2018:UAA

[1454] Scott E. Hoffmann. Uniform analytic approximation of Wigner rotation matrices. Journal of Mathematical Physics, 59(2):022102, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miller:2018:KMA

[1455] Carl A. Miller, Roger Colbeck, and Yaoyun Shi. Keyring models: An approach to steerability. Journal of Mathematical Physics, 59(2):022103, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abouamal:2018:FOS

[1456] Ismail Abouamal and Pavel Winternitz. Fifth-order superintegrable quantum systems separating in Cartesian coordinates: Doubly exotic potentials. Journal of Mathematical Physics, 59(2):022104, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Combot:2018:IOD

[1457] Thierry Combot. Integrability of the one dimensional Schrödinger equation. Journal of Mathematical Physics, 59(2):022105, February 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laissaoui:2018:PSD

[1458] Aïda Laissaoui and Lyazid Chetouani. Propagator for symmetric double delta potential using path decomposition method. Journal of Mathematical Physics, 59(2):022106, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2018:DBQ

[1459] Andreas Andersson. Detailed balance as a quantum-group symmetry of Kraus operators. Journal of Mathematical Physics, 59(2):022107, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turbiner:2018:TBP

[1460] Alexander V. Turbiner, Willard Miller, Jr., and M. A. Escobar-Ruiz. Three-body problem in d-dimensional space: Ground state, (quasi)- exact-solvability. Journal of Mathematical Physics, 59(2):022108, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bach:2018:BDS

[1461] V. Bach, W. de Siqueira Pedra, and S. N. Lakaev. Bounds on the discrete spectrum of lattice Schrödinger operators. Journal of Mathematical Physics, 59(2):022109, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bildstein:2018:HTF

[1462] Steve Bildstein. Half theory fractional angular momentum and the application of fractional derivatives to quantum mechanics. Journal of Mathematical Physics, 59(2):022110, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:2018:TIF

[1463] Tapas Das, Uttam Ghosh, Susmita Sarkar, and Shantanu Das. Time independent fractional Schrödinger equation for generalized Mie-type potential in higher dimension framed with Jumarie type fractional derivative. Journal of Mathematical Physics, 59(2):022111, February 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiao:2018:ACE

[1464] Yong Jiao, Eyuri Wakakuwa, and Tomohiro Ogawa. Asymptotic convertibility of entanglement: An information-spectrum approach to entanglement concentration and dilution. Journal of Mathematical Physics, 59 (2):022201, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

OCarroll:2018:SLF

[1465] Michael O’Carroll and Paulo A. Faria da Veiga. Scaled lattice fermion fields, stability bounds, and regularity. Journal of Mathematical Physics, 59(2):022301, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beyer:2018:ABL

[1466] Florian Beyer and Tim-Torben Paetz. Analysis of a Bianchi-like equation satisfied by the Mars–Simon tensor. Journal of Mathematical Physics, 59 (2):022501, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2018:PSS

[1467] Jia-Fang Zhang and Shaoli Wang. Positive steady states for a nonlinear diffusion Beddington–DeAngelis model. Journal of Mathematical Physics, 59(2):022701, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakazono:2018:RLE

[1468] Nobutaka Nakazono. Reduction of lattice equations to the Painlevé equations: P_{IV} and P_{V}. Journal of Mathematical Physics, 59(2):022702, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2018:DCG

[1469] Feng Zhou, Chunyou Sun, and Jiaqi Cheng. Dynamics for the complex Ginzburg–Landau equation on non-cylindrical domains II: The monotone case. Journal of Mathematical Physics, 59(2):022703, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:CBO

[1470] Zhijing Chen, Tingwen Huang, Yu Huang, and Xin Liu. Chaotic behaviors of one dimensional wave equations with van der Pol nonlinear boundary conditions. Journal of Mathematical Physics, 59(2):022704, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tibboel:2018:PRC

[1471] Pieter Tibboel. Polygonal rotopulsators of the curved n-body problem. Journal of Mathematical Physics, 59(2):022901, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Esen:2018:HJT

[1472] Ogul Esen, Manuel de León, and Cristina Sardón. A Hamilton–Jacobi theory for implicit differential systems. Journal of Mathematical Physics, 59(2):022902, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delaFuente:2018:GGR

[1473] Daniel de la Fuente and Rafael M. Rubio. Galilean generalized Robertson–Walker spacetimes: A new family of Galilean geometrical models. Journal of Mathematical Physics, 59(2):022903, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ward:2018:HSC

[1474] R. S. Ward. Hopf solitons on compact manifolds. Journal of Mathematical Physics, 59(2):022904, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yin:2018:SSS

[1475] Haiyan Yin. Stability of stationary solutions for inflow problem on the planar magnetohydrodynamics. Journal of Mathematical Physics, 59(2):

023101, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2018:DCM

[1476] Peixin Zhang. Decay of the compressible magneto-micropolar fluids. Journal of Mathematical Physics, 59(2):023102, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borodin:2018:SHS

[1477] Alexei Borodin. Stochastic higher spin six vertex model and Macdonald measures. Journal of Mathematical Physics, 59(2):023301, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Acik:2018:FEK

[1478] Özgür Açik. Field equations from Killing spinors. Journal of Mathematical Physics, 59(2):023501, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luef:2018:CBQ

[1479] Franz Luef and Eirik Skrettingland. Convolutions for Berezin quantization and Berezin–Lieb inequalities. Journal of Mathematical Physics, 59 (2):023502, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castro:2018:EBA

[1480] E. Castro. Equivalence between the Arquès–Walsh sequence formula and the number of connected Feynman diagrams for every perturbation order in the fermionic many-body problem. Journal of Mathematical Physics, 59(2):023503, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakubczyk:2018:CAR

[1481] Dorota Jakubczyk and Paweł Jakubczyk. Combinatorial approach to the representation of the Schur–Weyl duality in one-dimensional spin systems. Journal of Mathematical Physics, 59(2):023504, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dwivedi:2018:HMF

[1482] Ravi Dwivedi and Vivek Sahai. On the hypergeometric matrix functions of several variables. Journal of Mathematical Physics, 59(2):023505, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baleanu:2018:TWS

[1483] Dumitru Baleanu, Mustafa Inc, Abdullahi Yusuf, and Aliyu Isa Aliyu. Traveling wave solutions and conservation laws for nonlinear evolution equation. Journal of Mathematical Physics, 59(2):023506, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contatto:2018:MPE

[1484] Felipe Contatto and Maciej Dunajski. Metrisability of Painlevé equations. Journal of Mathematical Physics, 59(2):023507, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brouder:2018:PFF

[1485] Christian Brouder, Nguyen Viet Dang, Camille Laurent-Gengoux, and Kasia Rejzner. Properties of field functionals and characterization of local functionals. Journal of Mathematical Physics, 59(2):023508, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tica:2018:FPI

[1486] Christian D. Tica and Eric A. Galapon. Finite-part integration of the generalized Stieltjes transform and its dominant asymptotic behavior for small values of the parameter. I. Integer orders. Journal of Mathematical Physics, 59(2):023509, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yahiaoui:2018:NFT

[1487] Sid-Ahmed Yahiaoui and Mustapha Bentaiba. New Fourier transform containing a pair of complex Euler Γ-functions with a monomial: Mathematical and physical applications. Journal of Mathematical Physics, 59 (2):023510, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tutiya:2018:GTS

[1488] Y. Tutiya. On genus-two solutions for the ILW equation. Journal of Mathematical Physics, 59(2):023511, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosenstock:2018:ECE

[1489] Sarita Rosenstock and James Owen Weatherall. Erratum: “A categorical equivalence between generalized holonomy maps on a connected manifold and principal connections on bundles over that manifold” [J. Math. Phys. 57, 102902 (2016)]. Journal of Mathematical Physics, 59 (2):029901, February 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [903].

Si:2018:BDN

[1490] Xin Si, Huashui Zhan, and Hongjun Yuan. The boundary degeneracy of a nonlinear equation related to electrorheological fluid. Journal of Mathematical Physics, 59(3):031501, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2018:NUD

[1491] Shouming Zhou, Bin Wang, and Rong Chen. Non-uniform dependence on initial data for the periodic Constantin–Lannes equation. Journal of Mathematical Physics, 59(3):031502, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2018:LWP

[1492] Jishan Fan, Fucai Li, and Gen Nakamura. Local well-posedness for a compressible non-isothermal model for nematic liquid crystals. Journal of Mathematical Physics, 59(3):031503, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luo:2018:GSS

[1493] Huxiao Luo, Xianhua Tang, and Zu Gao. Ground state sign-changing solutions for fractional Kirchhoff equations in bounded domains. Journal of Mathematical Physics, 59(3):031504, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:MSC

[1494] Caisheng Chen, Jinfeng Bao, and Hongxue Song. Multiple solutions for a class of fractional (p, q)-Laplacian system in RN. Journal of Mathematical Physics, 59(3):031505, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ridout:2018:SVW

[1495] David Ridout, Steve Siu, and Simon Wood. Singular vectors for the W_{N} algebras. Journal of Mathematical Physics, 59(3):031701, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peano:2018:TES

[1496] Vittorio Peano and Hermann Schulz-Baldes. Topological edge states for disordered bosonic systems. Journal of Mathematical Physics, 59(3): 031901, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:RCT

[1497] Li Chen, Ji Oon Lee, and Jinyeop Lee. Rate of convergence toward Hartree dynamics with singular interaction potential. Journal of Mathematical Physics, 59(3):031902, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katsura:2018:NIT

[1498] Hosho Katsura and Tohru Koma. The noncommutative index theorem and the periodic table for disordered topological insulators and superconductors. Journal of Mathematical Physics, 59(3):031903, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdul-Rahman:2018:ECN

[1499] Houssam Abdul-Rahman. Entanglement of a class of non-Gaussian states in disordered harmonic oscillator systems. Journal of Mathematical Physics, 59(3):031904, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tada:2018:DSC

[1500] Yasuhiro Tada and Tohru Koma. Decay of superconducting correlations for gauged electrons in dimensions D ≤ 4. Journal of Mathematical Physics, 59(3):031905, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Le:2018:EAS

[1501] Dai-Nam Le, Ngoc-Tram D. Hoang, and Van-Hoang Le. Exact analytical solutions of the Schrödinger equation for a two dimensional purely sextic double-well potential. Journal of Mathematical Physics, 59(3):032101, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Phan:2018:VSS

[1502] Ngoc-Hung Phan, Dai-Nam Le, Tuan-Quoc N. Thoi, and Van-Hoang Le. Variables separation and superintegrability of the nine-dimensional MICZ–Kepler problem. Journal of Mathematical Physics, 59(3):032102, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takahashi:2018:QMBa

[1503] K. Takahashi. On the quantization of the massless Bateman system. Journal of Mathematical Physics, 59(3):032103, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2018:SGB

[1504] Se-Jin Kim, Vern Paulsen, and Christopher Schafhauser. A synchronous game for binary constraint systems. Journal of Mathematical Physics, 59(3):032201, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tylec:2018:IBM

[1505] Tomasz I. Tylec and Marek Kuś. Ignorance is a bliss: Mathematical structure of many-box models. Journal of Mathematical Physics, 59(3): 032202, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arenz:2018:DDU

[1506] Christian Arenz, Daniel Burgarth, Paolo Facchi, and Robin Hillier. Dynamical decoupling of unbounded Hamiltonians. Journal of Mathematical Physics, 59(3):032203, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2018:DTC

[1507] Hai Lin and Keyou Zeng. Detecting topology change via correlations and entanglement from gauge/gravity correspondence. Journal of Mathematical Physics, 59(3):032301, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:TPS

[1508] Xiao-Kan Guo. Transition probability spaces in loop quantum gravity. Journal of Mathematical Physics, 59(3):032302, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakel:2018:SRQ

[1509] Christian D. Jäkel and Walter F. Wreszinski. Stability of relativistic quantum electrodynamics in the Coulomb gauge. Journal of Mathematical Physics, 59(3):032303, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Perlov:2018:UPL

[1510] Leonid Perlov. Uncertainty principle in loop quantum cosmology by Moyal formalism. Journal of Mathematical Physics, 59(3):032304, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2018:DAL

[1511] Chenmei Xu and Yisong Yang. Determination of angle of light deflection in higher-derivative gravity theories. Journal of Mathematical Physics, 59(3):032501, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaset:2018:MUF

[1512] Jordi Gaset and Narciso Román-Roy. Multisymplectic unified formalism for Einstein–Hilbert gravity. Journal of Mathematical Physics, 59(3): 032502, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goto:2018:MSM

[1513] Shin itiro Goto and Ken Umeno. Maps on statistical manifolds exactly reduced from the Perron–Frobenius equations for solvable chaotic maps. Journal of Mathematical Physics, 59(3):032701, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2018:RSR

[1514] Seung-Yeal Ha and Jinwook Jung. Remarks on the slow relaxation for the fractional Kuramoto model for synchronization. Journal of Mathematical Physics, 59(3):032702, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:USA

[1515] Dingshi Li and Lin Shi. Upper semicontinuity of attractors of stochastic delay reaction-diffusion equations in the delay. Journal of Mathematical Physics, 59(3):032703, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [1569].

Brzykcy:2018:NHG

[1516] Przemysław Brzykcy. (1 + 1) Newton–Hooke group for the simple and damped harmonic oscillator. Journal of Mathematical Physics, 59(3): 032901, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lazo:2018:APA

[1517] Matheus J. Lazo, Juilson Paiva, João T. S. Amaral, and Gastão S. F. Frederico. An action principle for action-dependent Lagrangians: Toward an action principle to non-conservative systems. Journal of Mathematical Physics, 59(3):032902, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krupka:2018:VSE

[1518] D. Krupka, Z. Urban, and J. Volná. Variational submanifolds of Euclidean spaces. Journal of Mathematical Physics, 59(3):032903, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-Perez:2018:LDD

[1519] N. E. Martínez-Pérez and C. Ramírez. On the Lagrangian description of dissipative systems. Journal of Mathematical Physics, 59(3):032904, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakobsen:2018:LME

[1520] P. Jakobsen, M. Mansuripur, and M. Kolesik. Leaky-mode expansion of the electromagnetic field inside dispersive spherical cavity. Journal of Mathematical Physics, 59(3):033501, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campoamor-Stursberg:2018:RIP

[1521] R. Campoamor-Stursberg. Reduction by invariants and projection of linear representations of Lie algebras applied to the construction of nonlinear realizations. Journal of Mathematical Physics, 59(3):033502, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fucci:2018:SNR

[1522] Guglielmo Fucci and Klaus Kirsten. Some new results for the one-loop mass correction to the compactified λφ^{4} theory. Journal of Mathematical Physics, 59(3):033503, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:GSF

[1523] Quanqing Li, Kaimin Teng, and Xian Wu. Ground states for fractional Schrödinger equations with critical growth. Journal of Mathematical Physics, 59(3):033504, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:2018:CDA

[1524] Shahn Majid and Anna Pachol. Classification of digital affine noncommutative geometries. Journal of Mathematical Physics, 59(3):033505, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2018:BVS

[1525] F. Bagarello, H. Inoue, and C. Trapani. Biorthogonal vectors, sesquilinear forms, and some physical operators. Journal of Mathematical Physics, 59(3):033506, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shestopalov:2018:STC

[1526] Yury Shestopalov. Singularities of the transmission coefficient and anomalous scattering by a dielectric slab. Journal of Mathematical Physics, 59(3):033507, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Siudzinska:2018:QCI

[1527] Katarzyna Siudzińska and Dariusz Chruściński. Quantum channels irreducibly covariant with respect to the finite group generated by the Weyl operators. Journal of Mathematical Physics, 59(3):033508, March 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz:2018:FBS

[1528] Victor Alfonzo Diaz and Andrea Giusti. Fractional bosonic strings. Journal of Mathematical Physics, 59(3):033509, March 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almog:2018:SSC

[1529] Y. Almog, D. S. Grebenkov, and B. Helffer. Spectral semi-classical analysis of a complex Schrödinger operator in exterior domains. Journal of Mathematical Physics, 59(4):041501, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheung:2018:BPI

[1530] Ka Luen Cheung, Sen Wong, and Manwai Yuen. Blowup phenomenon for the initial-boundary value problem of the non-isentropic compressible Euler equations. Journal of Mathematical Physics, 59(4):041502, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borrelli:2018:MSS

[1531] William Borrelli. Multiple solutions for a self-consistent Dirac equation in two dimensions. Journal of Mathematical Physics, 59(4):041503, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lewin:2018:GMB

[1532] Mathieu Lewin, Phan Thành Nam, and Nicolas Rougerie. Gibbs measures based on 1 d (an)harmonic oscillators as mean-field limits. Journal of Mathematical Physics, 59(4):041901, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daCosta:2018:PDM

[1533] Bruno G. da Costa and Ernesto P. Borges. A position-dependent mass harmonic oscillator and deformed space. Journal of Mathematical Physics, 59(4):042101, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Busch:2018:SUR

[1534] Paul Busch, Jukka Kiukas, and R. F. Werner. Sharp uncertainty relations for number and angle. Journal of Mathematical Physics, 59(4): 042102, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marzlin:2018:SEA

[1535] Karl-Peter Marzlin and Bryce Fitzgerald. Spontaneous emission and atomic line shift in causal perturbation theory. Journal of Mathematical Physics, 59(4):042103, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:2018:DSI

[1536] C. Quesne. Deformed shape invariance symmetry and potentials in curved space with two known eigenstates. Journal of Mathematical Physics, 59(4):042104, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2018:GIT

[1537] Pavel Exner, Tomás Kalvoda, and Matej Tusek. A geometric Iwatsuka type effect in quantum layers. Journal of Mathematical Physics, 59(4): 042105, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerini:2018:UJM

[1538] Leonardo Guerini and Marcelo Terra Cunha. Uniqueness of the joint measurement and the structure of the set of compatible quantum measurements. Journal of Mathematical Physics, 59(4):042106, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2018:VPS

[1539] Le Bin Ho and Nobuyuki Imoto. Various pointer states approaches to polar modular values. Journal of Mathematical Physics, 59(4):042107, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khodja:2018:PID

[1540] A. Khodja, F. Benamira, and L. Guechi. Path integral discussion of the improved Tietz potential. Journal of Mathematical Physics, 59(4): 042108, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lawson:2018:LPT

[1541] Latévi M. Lawson and Gabriel Y. H. Avossevou. Landau problem with time dependent mass in time dependent electric and harmonic background fields. Journal of Mathematical Physics, 59(4):042109, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lammert:2018:SHK

[1542] Paul E. Lammert. In search of the Hohenberg–Kohn theorem. Journal of Mathematical Physics, 59(4):042110, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Griesemer:2018:DNH

[1543] M. Griesemer and A. Wünsch. On the domain of the Nelson Hamiltonian. Journal of Mathematical Physics, 59(4):042111, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alicata:2018:QMS

[1544] G. Alicata, F. Bagarello, F. Gargano, and S. Spagnolo. Quantum mechanical settings inspired by RLC circuits. Journal of Mathematical Physics, 59(4):042112, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holzapfel:2018:PRV

[1545] Milan Holzäpfel, Marcus Cramer, Nilanjana Datta, and Martin B. Plenio. Petz recovery versus matrix reconstruction. Journal of Mathematical Physics, 59(4):042201, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leverrier:2018:CSG

[1546] Anthony Leverrier. SU(p, q) coherent states and a Gaussian de Finetti theorem. Journal of Mathematical Physics, 59(4):042202, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuramochi:2018:QIC

[1547] Yui Kuramochi. Quantum incompatibility of channels with general outcome operator algebras. Journal of Mathematical Physics, 59(4):042203, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hanson:2018:MME

[1548] Eric P. Hanson and Nilanjana Datta. Maximum and minimum entropy states yielding local continuity bounds. Journal of Mathematical Physics, 59(4):042204, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albeverio:2018:AES

[1549] Sergio Albeverio and Hiroshi Tamura. Asymptotics of the evolution semigroup associated with a scalar field in the presence of a non-linear electromagnetic field. Journal of Mathematical Physics, 59(4):042301, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:2018:CCC

[1550] J. Schmidt and J. Bicák. Covariant conserved currents for scalar–tensor Horndeski theory. Journal of Mathematical Physics, 59(4):042501, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Starko:2018:WEF

[1551] Darij Starko and Walter Craig. The wave equation in Friedmann– Robertson–Walker space-times and asymptotics of the intensity and distance relationship of a localised source. Journal of Mathematical Physics, 59(4):042502, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurtug:2018:QPH

[1552] O. Gurtug and M. Mangut. Quantum probe of Horava–Lifshitz gravity. Journal of Mathematical Physics, 59(4):042503, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chudecki:2018:TTR

[1553] Adam Chudecki and Maciej Przanowski. On twisting type [N] ⊗ [N] Ricci flat complex spacetimes with two homothetic symmetries. Journal of Mathematical Physics, 59(4):042504, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vishwakarma:2018:NWL

[1554] Ram Gopal Vishwakarma. A new Weyl-like tensor of geometric origin. Journal of Mathematical Physics, 59(4):042505, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castellana:2018:SRT

[1555] Michele Castellana. Symmetry reduction of the three-body problem based on Euler angles. Journal of Mathematical Physics, 59(4):042901, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scholle:2018:EIU

[1556] M. Scholle, P. H. Gaskell, and F. Marner. Exact integration of the unsteady incompressible Navier–Stokes equations, gauge criteria, and applications. Journal of Mathematical Physics, 59(4):043101, April 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Pennisi:2018:RER

[1557] Sebastiano Pennisi and Tommaso Ruggeri. Relativistic Eulerian rarefied gas with internal structure. Journal of Mathematical Physics, 59(4): 043102, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heydenreich:2018:CAE

[1558] Markus Heydenreich and Leonid Kolesnikov. The critical 1-arm exponent for the ferromagnetic Ising model on the Bethe lattice. Journal of Mathematical Physics, 59(4):043301, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carillo:2018:NNK

[1559] Sandra Carillo, Mauro Lo Schiavo, Egmont Porten, and Cornelia Schiebold. A novel noncommutative KdV-type equation, its recursion operator, and solitons. Journal of Mathematical Physics, 59(4):043501, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samann:2018:TMB

[1560] Christian Sämann and Lennart Schmidt. Towards an M5-brane model I: A 6 d superconformal field theory. Journal of Mathematical Physics, 59(4):043502, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biswas:2018:GME

[1561] Indranil Biswas and Varun Thakre. Generalised monopole equations on Kähler surfaces. Journal of Mathematical Physics, 59(4):043503, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odzijewicz:2018:ICC

[1562] A. Odzijewicz and E. Wawreniuk. Integrability and correspondence of classical and quantum non-linear three-mode systems. Journal of Mathematical Physics, 59(4):043504, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2018:SDM

[1563] Jianqing Sun, Xingbiao Hu, and Yingnan Zhang. A semi-discrete modified KdV equation. Journal of Mathematical Physics, 59(4):043505, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krishnan:2018:CG

[1564] Chethan Krishnan and Avinash Raju. Contractions from grading. Journal of Mathematical Physics, 59(4):043506, April 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tadano:2018:UDB

[1565] Homare Tadano. An upper diameter bound for compact Ricci solitons with application to the Hitchin–Thorpe inequality. II. Journal of Mathematical Physics, 59(4):043507, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2018:GCS

[1566] Axel Schulze-Halberg and Özlem YesIltas. The generalized confluent supersymmetry algorithm: Representations and integral formulas. Journal of Mathematical Physics, 59(4):043508, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:2018:GPS

[1567] John E. Gough. The Gisin–Percival stochastic Schrödinger equation from standard quantum filtering theory. Journal of Mathematical Physics, 59 (4):043509, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huh:2018:CPS

[1568] Hyungjin Huh and Jihyun Yim. The Cauchy problem for space-time monopole equations in Sobolev spaces. Journal of Mathematical Physics, 59(4):043510, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:PNU

[1569] Dingshi Li and Lin Shi. Publisher’s note: “Upper semicontinuity of attractors of stochastic delay reaction-diffusion equations in the delay” [J. Math. Phys. 59, 032703 (2018)]. Journal of Mathematical Physics, 59(4):049901, April 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1515].

Gurses:2018:NNS

[1570] Metin Gürses and Asli Pekcan. Nonlocal nonlinear Schrödinger equations and their soliton solutions. Journal of Mathematical Physics, 59 (5):051501, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghosh:2018:AIC

[1571] Tuhin Ghosh and Ashwin Tarikere. Approximate isotropic cloak for the Maxwell equations. Journal of Mathematical Physics, 59(5):051502, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:NED

[1572] Zhengguang Liu and Xiaoli Li. A novel equivalent definition of Caputo fractional derivative without singular kernel and superconvergent analysis. Journal of Mathematical Physics, 59(5):051503, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolbeault:2018:MR

[1573] Jean Dolbeault, Maria J. Esteban, Ari Laptev, and Michael Loss. Magnetic rings. Journal of Mathematical Physics, 59(5):051504, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2018:CPD

[1574] Shengqi Yu. Continuous properties of the data-to-solution map for a generalized μ-Camassa–Holm integrable equation. Journal of Mathematical Physics, 59(5):051505, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rogers:2018:MSN

[1575] Colin Rogers and Boris Malomed. On Madelung systems in nonlinear optics: A reciprocal invariance. Journal of Mathematical Physics, 59(5): 051506, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gungor:2018:ESVa

[1576] F. Güngör. Equivalence and symmetries for variable coefficient linear heat type equations. I. Journal of Mathematical Physics, 59(5):051507, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mustafa:2018:BCL

[1577] Muhammad I. Mustafa. Boundary control of laminated beams with interfacial slip. Journal of Mathematical Physics, 59(5):051508, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohamad:2018:HCC

[1578] Haidar Mohamad and Marcel Oliver. H^{s}-class construction of an almost invariant slow subspace for the Klein–Gordon equation in the nonrelativistic limit. Journal of Mathematical Physics, 59(5):051509, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fischbacher:2018:DSQ

[1579] C. Fischbacher and G. Stolz. Droplet states in quantum XXZ spin systems on general graphs. Journal of Mathematical Physics, 59(5):051901, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:MNI

[1580] Li Chen and Heinz Siedentop. The maximal negative ion of the timedependent Thomas–Fermi and the Vlasov atom. Journal of Mathematical Physics, 59(5):051902, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoque:2018:RAH

[1581] Md Fazlul Hoque, Ian Marquette, and Yao-Zhong Zhang. Recurrence approach and higher order polynomial algebras for superintegrable monopole systems. Journal of Mathematical Physics, 59(5):052101, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albertini:2018:CSS

[1582] Francesca Albertini and Domenico D’Alessandro. Controllability of symmetric spin networks. Journal of Mathematical Physics, 59(5):052102, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kryukov:2018:OPQ

[1583] A. Kryukov. On observation of position in quantum theory. Journal of Mathematical Physics, 59(5):052103, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gholizadehkalkhoran:2018:EDH

[1584] Hossein Gholizadehkalkhoran, Ilkka Ruokosenmäki, and Tapio T. Rantala. Eigenstates and dynamics of Hooke’s atom: Exact results and path integral simulations. Journal of Mathematical Physics, 59(5):

052104, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fresneda:2018:QLC

[1585] Rodrigo Fresneda, Jean Pierre Gazeau, and Diego Noguera. Quantum localisation on the circle. Journal of Mathematical Physics, 59(5):052105, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGosson:2018:RPS

[1586] Maurice A. de Gosson and Fernando Nicacio. Relative phase shifts for metaplectic isotopies acting on mixed Gaussian states. Journal of Mathematical Physics, 59(5):052106, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nechita:2018:AAQ

[1587] Ion Nechita, Zbigniew Puchała, Łukasz Pawela, and Karol Życzkowski. Almost all quantum channels are equidistant. Journal of Mathematical Physics, 59(5):052201, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drumond:2018:SVB

[1588] Raphael C. Drumond, Cristhiano Duarte, and Roberto I. Oliveira. Small violations of Bell inequalities for multipartite pure random states. Journal of Mathematical Physics, 59(5):052202, May 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Junker:2018:PIS

[1589] G. Junker and A. Inomata. Path integral and spectral representations for supersymmetric Dirac–Hamiltonians. Journal of Mathematical Physics, 59(5):052301, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barnich:2018:NLB

[1590] Glenn Barnich and Nicolas Boulanger. A note on local BRST cohomology of Yang–Mills type theories with free Abelian factors. Journal of Mathematical Physics, 59(5):052302, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agrawal:2018:FTL

[1591] Gunjan Agrawal and Soami Pyari Sinha. Fine topology and locally Minkowskian manifolds. Journal of Mathematical Physics, 59(5):052501, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baxter:2018:ETD

[1592] J. Erik Baxter. On the existence of topological dyons and dyonic black holes in anti-de Sitter Einstein–Yang–Mills theories with compact semisimple gauge groups. Journal of Mathematical Physics, 59(5): 052502, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avalos:2018:CEL

[1593] R. Avalos, F. Dahia, and C. Romero. On the C^{k}-embedding of Lorentzian manifolds in Ricci-flat spaces. Journal of Mathematical Physics, 59(5): 052503, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pinzari:2018:CEM

[1594] Gabriella Pinzari. On the co-existence of maximal and whiskered tori in the planetary three-body problem. Journal of Mathematical Physics, 59(5):052701, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz:2018:GLA

[1595] Bogar Díaz and Merced Montesinos. Geometric Lagrangian approach to the physical degree of freedom count in field theory. Journal of Mathematical Physics, 59(5):052901, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davis:2018:IPC

[1596] Candice Davis, Scott Geyer, William Johnson, and Zhifu Xie. Inverse problem of central configurations in the collinear 5-body problem. Journal of Mathematical Physics, 59(5):052902, May 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyberg:2018:PSS

[1597] I. Lyberg, V. Korepin, G. A. P. Ribeiro, and J. Viti. Phase separation in the six-vertex model with a variety of boundary conditions. Journal of Mathematical Physics, 59(5):053301, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2018:FFL

[1598] Peter J. Forrester and Allan K. Trinh. Functional form for the leading correction to the distribution of the largest eigenvalue in the GUE and LUE. Journal of Mathematical Physics, 59(5):053302, May 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bianucci:2018:USR

[1599] Marco Bianucci. Using some results about the Lie evolution of differential operators to obtain the Fokker–Planck equation for non-Hamiltonian dynamical systems of interest. Journal of Mathematical Physics, 59(5): 053303, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karayer:2018:SSE

[1600] H. Karayer, D. Demirhan, and F. Büyükkiliç. Solution of Schrödinger equation for two different potentials using extended Nikiforov–Uvarov method and polynomial solutions of biconfluent Heun equation. Journal of Mathematical Physics, 59(5):053501, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2327].

Celeghini:2018:SHR

[1601] E. Celeghini, M. Gadella, and M. A. del Olmo. Spherical harmonics and rigged Hilbert spaces. Journal of Mathematical Physics, 59(5):053502, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eremenko:2018:PSE

[1602] Alexandre Eremenko and Andrei Gabrielov. PT-symmetric eigenvalues for homogeneous potentials. Journal of Mathematical Physics, 59(5): 053503, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vakhnenko:2018:INS

[1603] Oleksiy O. Vakhnenko. Integrable nonlinear Schrödinger system on a lattice with three structural elements in the unit cell. Journal of Mathematical Physics, 59(5):053504, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motegi:2018:SFW

[1604] Kohei Motegi. Symmetric functions and wavefunctions of XXZ-type sixvertex models and elliptic Felderhof models by Izergin–Korepin analysis. Journal of Mathematical Physics, 59(5):053505, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arik:2018:QMP

[1605] Metin Arik and Medine Ildes. Quantum mechanics on periodic and non-periodic lattices and almost unitary Schwinger operators. Journal of Mathematical Physics, 59(5):053506, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bassi:2018:STJ

[1606] Jacopo Bassi and Ludwik Dabrowski. Spectral triples on the Jiang–Su algebra. Journal of Mathematical Physics, 59(5):053507, May 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boccagna:2018:FLN

[1607] Roberto Boccagna. Fick’s law in non-local evolution equations. Journal of Mathematical Physics, 59(5):053508, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Norris:2018:GTS

[1608] L. K. Norris. Gromov’s theorem in n-symplectic geometry on LR^{n}. Journal of Mathematical Physics, 59(5):053509, May 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2018:ESG

[1609] Ruifeng Zhang. Existence of solutions for a generalized Yang–Mills–Higgs theory. Journal of Mathematical Physics, 59(6):061501, June 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hovhannisyan:2018:SDE

[1610] Gro Hovhannisyan. Schwarzian derivative and Ermakov equation on a time scale. Journal of Mathematical Physics, 59(6):061502, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2018:GSF

[1611] Runzhang Xu, Xingchang Wang, Yanbing Yang, and Shaohua Chen. Global solutions and finite time blow-up for fourth order nonlinear damped wave equation. Journal of Mathematical Physics, 59(6):061503, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:LBM

[1612] Yujin Guo, Yong Luo, and Zhi-Qiang Wang. Limit behavior of mass critical Hartree minimization problems with steep potential wells. Journal of Mathematical Physics, 59(6):061504, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:LTA

[1613] Boling Guo and Nan Liu. Long-time asymptotics for the Kundu–Eckhaus equation on the half-line. Journal of Mathematical Physics, 59(6):061505, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaikina:2018:IIB

[1614] Elena I. Kaikina. Inhomogeneous initial-boundary value problem for the 2D nonlinear Schrödinger equation. Journal of Mathematical Physics, 59(6):061506, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gungor:2018:ESVb

[1615] F. Güngör. Equivalence and symmetries for variable coefficient linear heat type equations. II. Fundamental solutions. Journal of Mathematical Physics, 59(6):061507, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milton:2018:NRF

[1616] Graeme W. Milton. A new route to finding bounds on the generalized spectrum of many physical operators. Journal of Mathematical Physics, 59(6):061508, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Janda:2018:ESS

[1617] Artur Z. Janda. Exact solutions and singularities of an X-point collapse in Hall magnetohydrodynamics. Journal of Mathematical Physics, 59(6): 061509, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1980] and response [1981].

Figueroa-OFarrill:2018:KLA

[1618] José M. Figueroa-O’Farrill. Kinematical Lie algebras via deformation theory. Journal of Mathematical Physics, 59(6):061701, June 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Figueroa-OFarrill:2018:HDK

[1619] José M. Figueroa-O’Farrill. Higher-dimensional kinematical Lie algebras via deformation theory. Journal of Mathematical Physics, 59(6):061702, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andrzejewski:2018:KLA

[1620] Tomasz Andrzejewski and José Miguel Figueroa-O’Farrill. Kinematical lie algebras in 2 + 1 dimensions. Journal of Mathematical Physics, 59(6): 061703, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ikeda:2018:HBL

[1621] Kazuki Ikeda. Hofstadter’s butterfly and Langlands duality. Journal of Mathematical Physics, 59(6):061704, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bley:2018:LBR

[1622] Gonzalo A. Bley. A lower bound on the renormalized Nelson model. Journal of Mathematical Physics, 59(6):061901, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolotin:2018:HPA

[1623] Arkady Bolotin. Hardy’s paradox according to non-classical semantics. Journal of Mathematical Physics, 59(6):062101, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koppe:2018:QHE

[1624] J. Köppe, M. Patzold, W. Grecksch, and W. Paul. Quantum Hamilton equations of motion for bound states of one-dimensional quantum systems. Journal of Mathematical Physics, 59(6):062102, June 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kumari:2018:CES

[1625] Nisha Kumari, Rajesh Kumar Yadav, Avinash Khare, and Bhabani Prasad Mandal. A class of exactly solvable rationally extended non-central potentials in two and three dimensions. Journal of Mathematical Physics, 59(6):062103, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Unal:2018:QCS

[1626] Nuri Ünal. Quasi-coherent states for the Hermite oscillator. Journal of Mathematical Physics, 59(6):062104, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2087].

H:2018:NAG

[1627] Bharath H. M. Non-Abelian geometric phases carried by the spin fluctuation tensor. Journal of Mathematical Physics, 59(6):062105, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyazawa:2018:SCS

[1628] Toru Miyazawa. SL(3, C) structure of one-dimensional Schrödinger equation. Journal of Mathematical Physics, 59(6):062106, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lebl:2018:UMF

[1629] Jirí Lebl, Asif Shakeel, and Nolan Wallach. Unentangled measurements and frame functions. Journal of Mathematical Physics, 59(6):062107, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gallone:2018:DSC

[1630] M. Gallone and A. Michelangeli. Discrete spectra for critical Dirac– Coulomb Hamiltonians. Journal of Mathematical Physics, 59(6):062108, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rahaman:2018:EEB

[1631] Mizanur Rahaman, Samuel Jaques, and Vern I. Paulsen. Eventually entanglement breaking maps. Journal of Mathematical Physics, 59(6): 062201, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duarte:2018:RTC

[1632] Cristhiano Duarte and Barbara Amaral. Resource theory of contextuality for arbitrary prepare-and-measure experiments. Journal of Mathematical Physics, 59(6):062202, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2018:HCH

[1633] L. Gao, M. Junge, and N. Laracuente. Heralded channel Holevo superadditivity bounds from entanglement monogamy. Journal of Mathematical Physics, 59(6):062203, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hollands:2018:REE

[1634] Stefan Hollands, Onirban Islam, and Ko Sanders. Relative entanglement entropy for widely separated regions in curved spacetime. Journal of Mathematical Physics, 59(6):062301, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valenzuela:2018:PSM

[1635] Mauricio Valenzuela. From phase space to multivector matrix models. Journal of Mathematical Physics, 59(6):062302, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bizi:2018:STD

[1636] Nadir Bizi, Christian Brouder, and Fabien Besnard. Space and time dimensions of algebras with application to Lorentzian noncommutative geometry and quantum electrodynamics. Journal of Mathematical Physics, 59(6):062303, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sati:2018:FMB

[1637] Hisham Sati. Framed m-branes, corners, and topological invariants. Journal of Mathematical Physics, 59(6):062304, June 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablinger:2018:IEH

[1638] J. Ablinger, J. Blümlein, A. De Freitas, M. van Hoeij, E. Imamoglu, C. G. Raab, C.-S. Radu, and C. Schneider. Iterated elliptic and hypergeometric integrals for Feynman diagrams. Journal of Mathematical Physics, 59(6): 062305, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barboza:2018:ISE

[1639] Marcelo Barboza, Benedito Leandro, and Romildo Pina. Invariant solutions for the Einstein field equation. Journal of Mathematical Physics, 59(6):062501, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avalos:2018:CPW

[1640] R. Avalos, I. P. Lobo, T. Sanomiya, and C. Romero. On the Cauchy problem for Weyl-geometric scalar-tensor theories of gravity. Journal of Mathematical Physics, 59(6):062502, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chudecki:2018:CCR

[1641] Adam Chudecki. Classification of complex and real vacuum spaces of the type [N] ⊗ [N]. Journal of Mathematical Physics, 59(6):062503, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2018:EHI

[1642] F. Calogero and F. Leyvraz. Examples of Hamiltonians isochronous in configuration space only and their quantization. Journal of Mathematical Physics, 59(6):062701, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez-Bermejo:2018:CMG

[1643] Benito Hernández-Bermejo. Congruence method for global Darboux reduction of finite-dimensional Poisson systems. Journal of Mathematical Physics, 59(6):062901, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cork:2018:SCR

[1644] Josh Cork. Symmetric calorons and the rotation map. Journal of Mathematical Physics, 59(6):062902, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1719].

Freistuhler:2018:CDR

[1645] Heinrich Freistühler and Blake Temple. Causal dissipation in the relativistic dynamics of barotropic fluids. Journal of Mathematical Physics, 59(6):063101, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Starr:2018:PUM

[1646] Shannon Starr and Meg Walters. Phase uniqueness for the Mallows measure on permutations. Journal of Mathematical Physics, 59(6):063301, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Redig:2018:NCC

[1647] Frank Redig, Wioletta M. Ruszel, and Ellen Saada. Non-criticality criteria for Abelian sandpile models with sources and sinks. Journal of Mathematical Physics, 59(6):063302, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:CQS

[1648] Wencai Liu. Continuous quasiperiodic Schrödinger operators with Gordon type potentials. Journal of Mathematical Physics, 59(6):063501, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conforti:2018:EFW

[1649] Giovanni Conforti and Michele Pavon. Extremal flows in Wasserstein space. Journal of Mathematical Physics, 59(6):063502, June 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hrivnak:2018:DCS

[1650] Jirí Hrivnák and Lenka Motlochová. Discrete cosine and sine transforms generalized to honeycomb lattice. Journal of Mathematical Physics, 59 (6):063503, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerner:2018:PIE

[1651] Joachim Kerner. On pairs of interacting electrons in a quantum wire. Journal of Mathematical Physics, 59(6):063504, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khalkhali:2018:GBM

[1652] Masoud Khalkhali and Andrzej Sitarz. Gauss–Bonnet for matrix conformally rescaled Dirac. Journal of Mathematical Physics, 59(6):063505, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konstantinou-Rizos:2018:NDP

[1653] S. Konstantinou-Rizos and T. E. Kouloukas. A noncommutative discrete potential KdV lift. Journal of Mathematical Physics, 59(6):063506, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandenDungen:2018:FST

[1654] Koen van den Dungen. Families of spectral triples and foliations of space(time). Journal of Mathematical Physics, 59(6):063507, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2018:SSL

[1655] A. D. Alhaidari. Series solutions of Laguerre- and Jacobi-type differential equations in terms of orthogonal polynomials and physical applications. Journal of Mathematical Physics, 59(6):063508, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alpay:2018:GFS

[1656] Daniel Alpay and Motke Porat. Generalized Fock spaces and the Stirling numbers. Journal of Mathematical Physics, 59(6):063509, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kandirmaz:2018:CSK

[1657] Nalan Kandirmaz. Coherent states for Kratzer-type potentials. Journal of Mathematical Physics, 59(6):063510, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joziak:2018:QBT

[1658] Paweł Józiak and Kamil Szpojankowski. Quantum Bernstein’s theorem and the hyperoctahedral quantum group. Journal of Mathematical Physics, 59(6):063511, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braun:2018:MMQ

[1659] Christopher Braun and James Maunder. Minimal models of quantum homotopy Lie algebras via the BV-formalism. Journal of Mathematical Physics, 59(6):063512, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2018:IPM

[1660] Byoung Jin Choi, Un Cig Ji, and Yongdo Lim. Inequalities for positive module operators on von Neumann algebras. Journal of Mathematical Physics, 59(6):063513, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fu:2018:LIE

[1661] Wei Fu and Frank W. Nijhoff. Linear integral equations, infinite matrices, and soliton hierarchies. Journal of Mathematical Physics, 59(7): 071101, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khorbatly:2018:DWP

[1662] Bashar Khorbatly, Ibtissam Zaiter, and Samer Isrwai. Derivation and well-posedness of the extended Green–Naghdi equations for flat bottoms with surface tension. Journal of Mathematical Physics, 59(7):071501, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgiev:2018:ZRL

[1663] V. Georgiev and J. Wirth. Zero resonances for localised potentials. Journal of Mathematical Physics, 59(7):071502, July 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2018:GSS

[1664] Chongsheng Cao, Yanqiu Guo, and Edriss S. Titi. Global strong solutions for the three-dimensional Hasegawa–Mima model with partial dissipation. Journal of Mathematical Physics, 59(7):071503, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:DEP

[1665] Huyuan Chen. The Dirichlet elliptic problem involving regional fractional Laplacian. Journal of Mathematical Physics, 59(7):071504, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2018:PRD

[1666] Boling Guo and Guoquan Qin. On the propagation of regularity and decay of solutions to the Benjamin equation. Journal of Mathematical Physics, 59(7):071505, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dinh:2018:SBS

[1667] Van Duong Dinh. A study on blowup solutions to the focusing L^{2}- supercritical nonlinear fractional Schrödinger equation. Journal of Mathematical Physics, 59(7):071506, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:FNS

[1668] Guoyuan Chen and Youquan Zheng. Fractional nonlinear Schrödinger equations with singular potential in R^{n}. Journal of Mathematical Physics, 59(7):071507, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gou:2018:EOS

[1669] Tianxiang Gou. Existence and orbital stability of standing waves to nonlinear Schrödinger system with partial confinement. Journal of Mathematical Physics, 59(7):071508, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ortner:2018:CPS

[1670] Norbert Ortner and Peter Wagner. Calculation of the propagator of Schrödinger’s equation on (0, ∞) with the potential kx^{−2} + ω^{2}x^{2} by Laplace’s method. Journal of Mathematical Physics, 59(7):071509, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuehn:2018:VAE

[1671] Christian Kuehn and Sebastian Throm. Validity of amplitude equations for nonlocal nonlinearities. Journal of Mathematical Physics, 59(7): 071510, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2018:ACO

[1672] T. Saanouni. Asymptotics for a class of oscillating fourth-order nonlinear Schrödinger equations. Journal of Mathematical Physics, 59(7):071511, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2018:MRS

[1673] Yueqiang Song and Shaoyun Shi. Multiplicity results for stationary Kirchhoff problems involving fractional elliptic operator and critical nonlinearity in RN. Journal of Mathematical Physics, 59(7):071512, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1839].

Mihailescu:2018:ITC

[1674] Mihai Mihailescu and Mayte Pérez-Llanos. Inhomogeneous torsional creep problems in anisotropic Orlicz Sobolev spaces. Journal of Mathematical Physics, 59(7):071513, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rassokha:2018:GCC

[1675] I. Rassokha, M. Serov, S. Spichak, and V. Stogniy. Group classification of a class of generalized nonlinear Kolmogorov equations and exact solutions. Journal of Mathematical Physics, 59(7):071514, July 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barbosa:2018:CPN

[1676] Isnaldo Isaac Barbosa. The Cauchy problem for nonlinear quadratic interactions of the Schrödinger type in one dimensional space. Journal of Mathematical Physics, 59(7):071515, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flandoli:2018:LCF

[1677] Ilaria Flandoli and Simon Lentner. Logarithmic conformal field theories of type B_{n}, l = 4 and symplectic fermions. Journal of Mathematical Physics, 59(7):071701, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Subag:2018:SHA

[1678] Eyal M. Subag. Symmetries of the hydrogen atom and algebraic families. Journal of Mathematical Physics, 59(7):071702, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haldane:2018:MIM

[1679] F. D. M. Haldane. A modular-invariant modified Weierstrass sigmafunction as a building block for lowest-Landau-level wavefunctions on the torus. Journal of Mathematical Physics, 59(7):071901, July 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pinna:2018:ACJ

[1680] Lorenzo Pinna and Gianluca Panati. Approximate controllability of the Jaynes–Cummings dynamics. Journal of Mathematical Physics, 59(7): 072101, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2018:RHT

[1681] Hyungrok Kim and Ingmar Saberi. Real homotopy theory and supersymmetric quantum mechanics. Journal of Mathematical Physics, 59(7):

072102, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koussa:2018:PHI

[1682] Walid Koussa, Naima Mana, Oum Kaltoum Djeghiour, and Mustapha Maamache. The pseudo Hermitian invariant operator and timedependent non-Hermitian Hamiltonian exhibiting a SU(1, 1) and SU(2) dynamical symmetry. Journal of Mathematical Physics, 59(7):072103, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basti:2018:TBP

[1683] Giulia Basti, Claudio Cacciapuoti, Domenico Finco, and Alessandro Teta. The three-body problem in dimension one: From short-range to contact interactions. Journal of Mathematical Physics, 59(7):072104, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langhoff:2018:APT

[1684] P. W. Langhoff, J. D. Mills, and J. A. Boatz. Atomic-pair theorem for universal matrix representatives of molecules and atomic clusters in nonrelativistic Born–Oppenheimer approximation. Journal of Mathematical Physics, 59(7):072105, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Michelangeli:2018:FPS

[1685] A. Michelangeli, A. Ottolini, and R. Scandone. Fractional powers and singular perturbations of quantum differential Hamiltonians. Journal of Mathematical Physics, 59(7):072106, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giller:2018:SSR

[1686] Stefan Giller. Superscar states in rational polygon billiards — a reality or an illusion? Journal of Mathematical Physics, 59(7):072107, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takahashi:2018:QMBb

[1687] K. Takahashi. On the quantization of the massive Bateman system. Journal of Mathematical Physics, 59(7):072108, July 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mahdifar:2018:PAP

[1688] A. Mahdifar, E. Amooghorban, and M. Jafari. Photon-added and photon-subtracted coherent states on a sphere. Journal of Mathematical Physics, 59(7):072109, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helsen:2018:RMQ

[1689] Jonas Helsen, Joel J. Wallman, and Stephanie Wehner. Representations of the multi-qubit Clifford group. Journal of Mathematical Physics, 59 (7):072201, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amaral:2018:NCE

[1690] Barbara Amaral, Cristhiano Duarte, and Roberto I. Oliveira. Necessary conditions for extended noncontextuality in general sets of random variables. Journal of Mathematical Physics, 59(7):072202, July 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seveso:2018:CGE

[1691] Luigi Seveso, Dardo Goyeneche, and Karol Zyczkowski. Coarse-grained entanglement classification through orthogonal arrays. Journal of Mathematical Physics, 59(7):072203, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seshadreesan:2018:RRE

[1692] Kaushik P. Seshadreesan, Ludovico Lami, and Mark M. Wilde. Rényi relative entropies of quantum Gaussian states. Journal of Mathematical Physics, 59(7):072204, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:FTB

[1693] Jiaojiao Li and Li Ma. Finite time blowup and global solutions of Euler type equations in matrix geometry. Journal of Mathematical Physics, 59(7):072205, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takaesu:2018:GSQ

[1694] Toshimitsu Takaesu. Ground states of quantum electrodynamics with cutoffs. Journal of Mathematical Physics, 59(7):072301, July 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Lachaume:2018:CEL

[1695] Xavier Lachaume. The constraint equations of Lovelock gravity theories: A new σ_{k}-Yamabe problem. Journal of Mathematical Physics, 59(7): 072501, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eskin:2018:HRA

[1696] G. Eskin. Hawking radiation from acoustic black holes in two space dimensions. Journal of Mathematical Physics, 59(7):072502, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bekaert:2018:CDT

[1697] Xavier Bekaert and Kevin Morand. Connections and dynamical trajectories in generalised Newton–Cartan gravity. II. An ambient perspective. Journal of Mathematical Physics, 59(7):072503, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2018:ZRF

[1698] Francesco Calogero. Zeros of rational functions and solvable nonlinear evolution equations. Journal of Mathematical Physics, 59(7):072701, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geng:2018:IDK

[1699] Jiansheng Geng and Shidi Zhou. An infinite dimensional KAM theorem with application to two dimensional completely resonant beam equation. Journal of Mathematical Physics, 59(7):072702, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:2018:EDC

[1700] A. V. Tsiganov. On exact discretization of cubic-quintic Duffing oscillator. Journal of Mathematical Physics, 59(7):072703, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contreras:2018:PPS

[1701] Ivan Contreras and Nicolas Martinez Alba. Poly-Poisson sigma models and their relational poly-symplectic groupoids. Journal of Mathematical Physics, 59(7):072901, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:ERP

[1702] Parry Y. Chen, Michael J. A. Smith, and Ross C. McPhedran. Evaluation and regularization of phase-modulated Eisenstein series and application to double Schlömilch-type sums. Journal of Mathematical Physics, 59(7): 072902, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramenko:2018:ARG

[1703] Andriy A. Avramenko, Igor V. Shevchuk, Alexander V. Kravchuk, Andrii I. Tyrinov, and Vladimir I. Shevchuk. Application of renormalization group analysis to two-phase turbulent flows with solid dust particles. Journal of Mathematical Physics, 59(7):073101, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2018:GSS

[1704] Haibo Yu. Global strong solutions to the 3D incompressible Navier– Stokes equations with large external force. Journal of Mathematical Physics, 59(7):073102, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McComb:2018:DDR

[1705] W. D. McComb and R. B. Fairhurst. The dimensionless dissipation rate and the Kolmogorov (1941) hypothesis of local stationarity in freely decaying isotropic turbulence. Journal of Mathematical Physics, 59(7): 073103, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jolly:2018:ESP

[1706] M. S. Jolly and D. Wirosoetisno. Energy spectra and passive tracer cascades in turbulent flows. Journal of Mathematical Physics, 59(7): 073104, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:2018:SNC

[1707] Yue-Hong Feng, Xin Li, and Shu Wang. Stability of non-constant equilibrium solutions for two-fluid non-isentropic Euler–Maxwell systems arising in plasmas. Journal of Mathematical Physics, 59(7):073105, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Youn:2018:EUR

[1708] Sang-Gyun Youn. Entropic uncertainty relations under localizations on discrete quantum groups. Journal of Mathematical Physics, 59(7): 073501, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2018:DCS

[1709] Xiaopeng Zhao and Mingxuan Zhu. Decay characterization of solutions to generalized Hall–MHD system in R^{3}. Journal of Mathematical Physics, 59(7):073502, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atakishiyev:2018:EBZ

[1710] Natig M. Atakishiyev, George S. Pogosyan, Kurt Bernardo Wolf, and Alexander Yakhno. Elliptic basis for the Zernike system: Heun function solutions. Journal of Mathematical Physics, 59(7):073503, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ouvry:2018:HPS

[1711] Stéphane Ouvry, Stephan Wagner, and Shuang Wu. Hofstadter point spectrum trace and the almost Mathieu operator. Journal of Mathematical Physics, 59(7):073504, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2018:QPW

[1712] Xiu-Bin Wang, Shou-Fu Tian, Lian-Li Feng, and Tian-Tian Zhang. On quasi-periodic waves and rogue waves to the (4 + 1)-dimensional nonlinear Fokas equation. Journal of Mathematical Physics, 59(7):073505, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choreno:2018:AAT

[1713] E. Choreño, D. Ojeda-Guillén, and V. D. Granados. Algebraic approach to the Tavis–Cummings model with three modes of oscillation. Journal of Mathematical Physics, 59(7):073506, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adan:2018:DIS

[1714] Alí Guzmán Adán and Franciscus Sommen. Distributions and integration in superspace. Journal of Mathematical Physics, 59(7):073507, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2018:IBV

[1715] Jian Xu, Qiaozhen Zhu, and Engui Fan. The initial-boundary value problem for the Sasa–Satsuma equation on a finite interval via the Fokas method. Journal of Mathematical Physics, 59(7):073508, July 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shah:2018:FWP

[1716] Firdous A. Shah, O. Ahmad, and P. E. Jorgensen. Fractional wave packet systems in L^{2}(R). Journal of Mathematical Physics, 59(7):073509, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ling:2018:NIL

[1717] Si-Tao Ling, Ming-Hui Wang, and Xue-Han Cheng. A new implementation of LSMR algorithm for the quaternionic least squares problem. Journal of Mathematical Physics, 59(7):073510, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wen:2018:MID

[1718] Xiao-Yong Wen and Zhenya Yan. Modulational instability and dynamics of multi-rogue wave solutions for the discrete Ablowitz–Ladik equation. Journal of Mathematical Physics, 59(7):073511, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cork:2018:PNS

[1719] Josh Cork. Publisher’s note: “Symmetric calorons and the rotation map” [J. Math. Phys. 59, 062902 (2018)]. Journal of Mathematical Physics, 59(7):079901, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1644].

Arrizabalaga:2018:ESA

[1720] Naiara Arrizabalaga, Javier Duoandikoetxea, and Luis Vega. Erratum: “Self-adjoint extensions of Dirac operators with Coulomb type singularity” [J. Math. Phys. 54, 041504 (2013)]. Journal of Mathematical Physics, 59(7):079902, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11].

DePalma:2018:GOE

[1721] Giacomo De Palma, Dario Trevisan, Vittorio Giovannetti, and Luigi Ambrosio. Gaussian optimizers for entropic inequalities in quantum information. Journal of Mathematical Physics, 59(8):081101, August 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alexandrov:2018:WHN

[1722] A. Alexandrov, G. Chapuy, B. Eynard, and J. Harnad. Weighted Hurwitz numbers and topological recursion: An overview. Journal of Mathematical Physics, 59(8):081102, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2018:LES

[1723] Liejun Shen and Xiaohua Yao. Least energy solutions for a class of fractional Schrödinger–Poisson systems. Journal of Mathematical Physics, 59(8):081501, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiang:2018:SLS

[1724] Tian Xiang. Sub-logistic source can prevent blow-up in the 2D minimal Keller–Segel chemotaxis system. Journal of Mathematical Physics, 59 (8):081502, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2018:EMS

[1725] Xianyong Yang, Wei Zhang, and Fukun Zhao. Existence and multiplicity of solutions for a quasilinear Choquard equation via perturbation method. Journal of Mathematical Physics, 59(8):081503, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorjgotov:2018:ESC

[1726] Khongorzul Dorjgotov, Hiroyuki Ochiai, and Uuganbayar Zunderiya. Exact solutions to a class of time fractional evolution systems with variable coefficients. Journal of Mathematical Physics, 59(8):081504, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2018:GCS

[1727] Yi Zhou and Yi Zhu. Global classical solutions of 2D MHD system with only magnetic diffusion on periodic domain. Journal of Mathematical Physics, 59(8):081505, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luo:2018:UGS

[1728] Yong Luo. Uniqueness of ground states for nonlinear Hartree equations. Journal of Mathematical Physics, 59(8):081506, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2018:GCS

[1729] Shijin Ding, Bingyuan Huang, and Xiaoling Liu. Global classical solutions to the 2D compressible Navier–Stokes equations with vacuum. Journal of Mathematical Physics, 59(8):081507, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:GSS

[1730] Sitong Chen and Xianhua Tang. Ground state solutions for generalized quasilinear Schrödinger equations with variable potentials and Berestycki–Lions nonlinearities. Journal of Mathematical Physics, 59 (8):081508, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2018:SIS

[1731] Tarek Saanouni. Strong instability of standing waves for the fractional Choquard equation. Journal of Mathematical Physics, 59(8):081509, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ou:2018:LTB

[1732] Yaobin Ou, Pan Shi, and Peter Wittwer. Large time behaviors of strong solutions to magnetohydrodynamic equations with free boundary and degenerate viscosity. Journal of Mathematical Physics, 59(8):081510, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tan:2018:MCM

[1733] Zhong Tan and Qiuju Xu. On the motion of the 3D compressible micropolar fluids with time periodic external forces. Journal of Mathematical Physics, 59(8):081511, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Satapathy:2018:NSC

[1734] Purnima Satapathy and T. Raja Sekhar. Nonlocal symmetries classifications and exact solution of Chaplygin gas equations. Journal of Mathematical Physics, 59(8):081512, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:TQA

[1735] Fulin Chen, Naihuan Jing, Fei Kong, and Shaobin Tan. Twisted quantum affinizations and their vertex representations. Journal of Mathematical Physics, 59(8):081701, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2018:KDN

[1736] Yuwang Hu, Jiefeng Liu, and Yunhe Sheng. Kupershmidt–(dual- )Nijenhuis structures on a Lie algebra with a representation. Journal of Mathematical Physics, 59(8):081702, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:IFB

[1737] Zhongming Chen, Jinjie Liu, Liqun Qi, Quanshui Zheng, and Wennan Zou. An irreducible function basis of isotropic invariants of a third order three-dimensional symmetric tensor. Journal of Mathematical Physics, 59(8):081703, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hezenci:2018:NEG

[1738] Fatih Hezenci and Yasar Sozen. A note on exceptional groups and Reidemeister torsion. Journal of Mathematical Physics, 59(8):081704, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lauret:2018:WMF

[1739] Emilio A. Lauret and Fiorela Rossi Bertone. Weight multiplicity formulas for bivariate representations of classical Lie algebras. Journal of Mathematical Physics, 59(8):081705, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Musto:2018:CAQ

[1740] Benjamin Musto, David Reutter, and Dominic Verdon. A compositional approach to quantum functions. Journal of Mathematical Physics, 59 (8):081706, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haldane:2018:OHS

[1741] F. D. M. Haldane. The origin of holomorphic states in Landau levels from non-commutative geometry and a new formula for their overlaps on the torus. Journal of Mathematical Physics, 59(8):081901, August 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bildstein:2018:HTI

[1742] Steve Bildstein. Half theory. II. The application of fractional spherical harmonics to chemical bonding. Journal of Mathematical Physics, 59 (8):082101, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lasser:2018:NHP

[1743] Caroline Lasser, Roman Schubert, and Stephanie Troppmann. Non-Hermitian propagation of Hagedorn wavepackets. Journal of Mathematical Physics, 59(8):082102, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeFloch:2018:BFS

[1744] Yohann Le Floch. Bounds for fidelity of semiclassical Lagrangian states in Kähler quantization. Journal of Mathematical Physics, 59(8):082103, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buyukasik:2018:SRG

[1745] Sirin A. Büyükasik. Squeezing and resonance in a generalized Caldirola– Kanai type quantum parametric oscillator. Journal of Mathematical Physics, 59(8):082104, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Serhan:2018:QDH

[1746] M. Serhan, M. Abusini, Ahmed Al-Jamel, H. El-Nasser, and Eqab M. Rabei. Quantization of the damped harmonic oscillator. Journal of Mathematical Physics, 59(8):082105, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haapasalo:2018:UIF

[1747] Erkka Haapasalo, Teiko Heinosaari, and Takayuki Miyadera. The unavoidable information flow to environment in quantum measurements. Journal of Mathematical Physics, 59(8):082106, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovacik:2018:MMN

[1748] Samuel Kovácik and Peter Presnajder. Magnetic monopoles in noncommutative quantum mechanics 2. Journal of Mathematical Physics, 59 (8):082107, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuda:2018:LOD

[1749] Toru Fuda, Daiju Funakawa, and Akito Suzuki. Localization for a onedimensional split-step quantum walk with bound states robust against perturbations. Journal of Mathematical Physics, 59(8):082201, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeWetering:2018:TCS

[1750] John van de Wetering. Three characterisations of the sequential product. Journal of Mathematical Physics, 59(8):082202, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torres:2018:MMS

[1751] Marcus A. C. Torres. Metric in the moduli of SU(2) monopoles from spectral curves and Gauss–Manin connection in disguise. Journal of Mathematical Physics, 59(8):082301, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arici:2018:RPH

[1752] Francesca Arici, Daniel Becker, Chris Ripken, Frank Saueressig, and Walter D. van Suijlekom. Reflection positivity in higher derivative scalar theories. Journal of Mathematical Physics, 59(8):082302, August 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Remy:2018:LQG

[1753] Guillaume Remy. Liouville quantum gravity on the annulus. Journal of Mathematical Physics, 59(8):082303, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chargui:2018:GUP

[1754] Y. Chargui and A. Dhahbi. Generalized uncertainty principle and the asymmetrical spinless Salpeter Coulomb problem. Journal of Mathematical Physics, 59(8):082304, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allen:2018:SPR

[1755] Brian Allen. Stability of the PMT and RPI for asymptotically hyperbolic manifolds foliated by IMCF. Journal of Mathematical Physics, 59(8):

082501, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melas:2018:LSP

[1756] Evangelos Melas. On the Liouvillian solutions to the perturbation equations of the Schwarzschild black hole. Journal of Mathematical Physics, 59(8):082502, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2018:IUS

[1757] Sun-Ho Choi and Seung-Yeal Ha. Interplay of the unit-speed constraint and time-delay in Cucker–Smale flocking. Journal of Mathematical Physics, 59(8):082701, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Possanner:2018:GVA

[1758] Stefan Possanner. Gyrokinetics from variational averaging: Existence and error bounds. Journal of Mathematical Physics, 59(8):082702, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khlevniuk:2018:CTP

[1759] A. Khlevniuk and V. Tymchyshyn. Classical treatment of particle with position-dependent mass m(r) = 1/(1 + r^{4}) in 1D and 2D subjected to harmonic potential. Journal of Mathematical Physics, 59(8):082901, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Papageorgiou:2018:LSI

[1760] Ioannis Papageorgiou. The log-Sobolev inequality with quadratic interactions. Journal of Mathematical Physics, 59(8):083301, August 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fedoruk:2018:GHG

[1761] Sergey Fedoruk, Evgeny Ivanov, and Andrei Smilga. Generic HKT geometries in the harmonic superspace approach. Journal of Mathematical Physics, 59(8):083501, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meklachi:2018:AAR

[1762] Taoufik Meklachi, Shari Moskow, and John C. Schotland. Asymptotic analysis of resonances of small volume high contrast linear and nonlinear scatterers. Journal of Mathematical Physics, 59(8):083502, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beasley:2018:RLS

[1763] Chris Beasley, Brendan McLellan, and Ruoran Zhang. Remarks on Legendrian self-linking. Journal of Mathematical Physics, 59(8):083503, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ozer:2018:EAW

[1764] H. T. Özer and Aytül Filiz. On the explicit asymptotic W_{5} symmetry of 3D Chern–Simons higher spin AdS_{3} gravity. Journal of Mathematical Physics, 59(8):083504, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuchs:2018:HSC

[1765] E. Fuchs, P. D. Jarvis, G. Rudolph, and M. Schmidt. The Hilbert space costratification for the orbit type strata of SU(2)-lattice gauge theory. Journal of Mathematical Physics, 59(8):083505, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaz:2018:DAD

[1766] Jayme Vaz, Jr. and Stephen Mann. DKP algebra, DKP equation, and differential forms. Journal of Mathematical Physics, 59(8):083506, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:2018:ABS

[1767] S. Y. Lou. Alice–Bob systems, P̂–T̂–Ĉ symmetry invariant and symmetry breaking soliton solutions. Journal of Mathematical Physics, 59 (8):083507, August 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Its:2018:PIS

[1768] Alexander Its and Nicolai Reshetikhin. Preface: Introduction to special issue: In memory of Ludwig Faddeev. Journal of Mathematical Physics, 59(9):091201, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tracy:2018:BGA

[1769] Craig A. Tracy and Harold Widom. Blocks and gaps in the asymmetric simple exclusion process: Asymptotics. Journal of Mathematical Physics, 59(9):091401, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spohn:2018:INI

[1770] Herbert Spohn. Interacting and noninteracting integrable systems. Journal of Mathematical Physics, 59(9):091402, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biswas:2018:GCM

[1771] Indranil Biswas and Jacques Hurtubise. Geometry of Calogero–Moser systems. Journal of Mathematical Physics, 59(9):091403, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basor:2018:SMI

[1772] Estelle Basor, Fan Ge, and Michael O. Rubinstein. Some multidimensional integrals in number theory and connections with the Painlevé V equation. Journal of Mathematical Physics, 59(9):091404, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubrovin:2018:SVL

[1773] B. Dubrovin and T. Skrypnyk. Separation of variables for linear Lax algebras and classical r-matrices. Journal of Mathematical Physics, 59 (9):091405, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablowitz:2018:SKV

[1774] M. J. Ablowitz, X.-D. Luo, and J. T. Cole. Solitons, the Kortewegde Vries equation with step boundary values, and pseudo-embedded eigenvalues. Journal of Mathematical Physics, 59(9):091406, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bleher:2018:DMF

[1775] Pavel Bleher, Brad Elwood, and Drazen Petrović. Dimer model: Full asymptotic expansion of the partition function. Journal of Mathematical Physics, 59(9):091407, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlowski:2018:TLF

[1776] Karol K. Kozlowski. On the thermodynamic limit of form factor expansions of dynamical correlation functions in the massless regime of the XXZ spin 1/2 chain. Journal of Mathematical Physics, 59(9):091408, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lisovyy:2018:ICB

[1777] O. Lisovyy, H. Nagoya, and J. Roussillon. Irregular conformal blocks and connection formulae for Painlevé V functions. Journal of Mathematical Physics, 59(9):091409, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bobenko:2018:CLP

[1778] Alexander I. Bobenko and Sergey Tsarev. The curvature line parametrization from circular nets on a surface. Journal of Mathematical Physics, 59(9):091410, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnad:2018:SPG

[1779] J. Harnad and Eunghyun Lee. Symmetric polynomials, generalized Jacobi–Trudi identities and τ-functions. Journal of Mathematical Physics, 59(9):091411, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korotkin:2018:TFH

[1780] Dmitry Korotkin and Peter Zograf. Tau functions, Hodge classes, and discriminant loci on moduli spaces of Hitchin’s spectral covers. Journal of Mathematical Physics, 59(9):091412, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fokas:2018:CIM

[1781] A. S. Fokas and M. C. van der Weele. Complexification and integrability in multidimensions. Journal of Mathematical Physics, 59(9):091413, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bothner:2018:SPU

[1782] Thomas Bothner, Percy Deift, Alexander Its, and Igor Krasovsky. The sine process under the influence of a varying potential. Journal of Mathematical Physics, 59(9):091414, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moon:2018:SGG

[1783] Alvin Moon and Bruno Nachtergaele. Stability of gapped ground state phases of spins and fermions in one dimension. Journal of Mathematical Physics, 59(9):091415, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babich:2018:AII

[1784] Mikhail Babich and Sergey Slavyanov. Antiquantization, isomonodromy, and integrability. Journal of Mathematical Physics, 59(9): 091416, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maillet:2018:QSV

[1785] J. M. Maillet and G. Niccoli. On quantum separation of variables. Journal of Mathematical Physics, 59(9):091417, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:2018:PDR

[1786] Mark Adler and Pierre van Moerbeke. Probability distributions related to tilings of non-convex polygons. Journal of Mathematical Physics, 59 (9):091418, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matveev:2018:ANH

[1787] Vladimir B. Matveev and Aleksandr O. Smirnov. AKNS and NLS hierarchies, MRW solutions, P_{n} breathers, and beyond. Journal of Mathematical Physics, 59(9):091419, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keating:2018:RTG

[1788] David Keating and Ananth Sridhar. Random tilings with the GPU. Journal of Mathematical Physics, 59(9):091420, September 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:NTL

[1789] Yong Liu and Juncheng Wei. Nondegeneracy of the traveling lump solution to the 2 + 1 Toda lattice. Journal of Mathematical Physics, 59 (10):101501, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stanislavova:2018:ASS

[1790] Milena Stanislavova and Atanas G. Stefanov. Asymptotic stability for spectrally stable Lugiato–Lefever solitons in periodic waveguides. Journal of Mathematical Physics, 59(10):101502, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pecher:2018:LWP

[1791] Hartmut Pecher. Local well-posedness for low regularity data for the higher dimensional Maxwell–Klein–Gordon system in Lorenz gauge. Journal of Mathematical Physics, 59(10):101503, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elong:2018:SFI

[1792] Ouissam Elong, Didier Robert, and Abderrahmane Senoussaoui. On semiclassical Fourier integral operators, Schrödinger propagators and coherent states. Journal of Mathematical Physics, 59(10):101504, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bensouilah:2018:SIS

[1793] Abdelwahab Bensouilah, Van Duong Dinh, and Shihui Zhu. On stability and instability of standing waves for the nonlinear Schrödinger equation with an inverse-square potential. Journal of Mathematical Physics, 59 (10):101505, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drouot:2018:RRH

[1794] Alexis Drouot. Resonances for random highly oscillatory potentials. Journal of Mathematical Physics, 59(10):101506, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abhinav:2018:SQI

[1795] Kumar Abhinav, Partha Guha, and Indranil Mukherjee. Study of quasiintegrable and non-holonomic deformation of equations in the NLS and DNLS hierarchy. Journal of Mathematical Physics, 59(10):101507, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:FAA

[1796] Jinjing Liu and Wanyi Xiao. Flux approximation to the Aw–Rascle model of traffic flow. Journal of Mathematical Physics, 59(10):101508, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:EVN

[1797] Xiao Chen, Shouxin Chen, and Shu Wang. Existence of vortices in nonlinear optics. Journal of Mathematical Physics, 59(10):101509, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See retraction [1898].

Aouadi:2018:LTD

[1798] Moncef Aouadi. Long-time dynamics for nonlinear porous thermoelasticity with second sound and delay. Journal of Mathematical Physics, 59 (10):101510, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kojima:2018:CRV

[1799] Takeo Kojima. Commutation relations of vertex operators for U_{q}(sl^{(M|N)})U_{q}(sl^{(M|N)})U_{q}(sl^{(M|N)}). Journal of Mathematical Physics, 59(10):101701, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Draper:2018:GRF

[1800] Cristina Draper and Valerio Guido. Gradings on the real form e_{6,−14}. Journal of Mathematical Physics, 59(10):101702, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koshida:2018:LMA

[1801] Shinji Koshida. Local martingales associated with Schramm–Loewner evolutions with internal symmetry. Journal of Mathematical Physics, 59 (10):101703, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnal:2018:RCS

[1802] Didier Arnal, Mabrouk Ben Ammar, Wafa Mtaouaa, and Zeineb Selmi. Relative cohomology spaces for some osp(n|2)-modules. Journal of Mathematical Physics, 59(10):101704, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grabowiecka:2018:ROF

[1803] Zofia Grabowiecka, Jirí Patera, and Marzena Szajewska. Reduction of orbits of finite Coxeter groups of non-crystallographic type. Journal of Mathematical Physics, 59(10):101705, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:2018:MBC

[1804] Alexandre G. M. Schmidt and Anderson L. de Jesus. Mapping between charge-monopole and position-dependent mass systems. Journal of Mathematical Physics, 59(10):102101, October 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dePrunele:2018:TDQ

[1805] E. de Prunelé. Two-dimensional quantum scattering by non-isotropic interactions localized on a circle, applications to open billiards. Journal of Mathematical Physics, 59(10):102102, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delValle:2018:TTY

[1806] J. C. del Valle and D. J. Nader. Toward the theory of the Yukawa potential. Journal of Mathematical Physics, 59(10):102103, October 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurasov:2018:SOG

[1807] P. Kurasov and R. Suhr. Schrödinger operators on graphs and geometry. III. General vertex conditions and counterexamples. Journal of Mathematical Physics, 59(10):102104, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2018:DDS

[1808] Richard L. Hall, Nasser Saad, and Kyle R. Bryenton. The d-dimensional softcore Coulomb potential and the generalized confluent Heun equation. Journal of Mathematical Physics, 59(10):102105, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivetic:2018:SOD

[1809] Boris Ivetić. On the symmetry of a one-dimensional hydrogen atom. Journal of Mathematical Physics, 59(10):102106, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanda:2018:MJJ

[1810] Tomohiro Kanda. A model of Josephson junctions on Boson systems — currents and entropy production rate. Journal of Mathematical Physics, 59(10):102107, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2018:RSP

[1811] Siendong Huang. Remote state preparation using Einstein–Podolsky– Rosen representations. Journal of Mathematical Physics, 59(10):102108, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeWetering:2018:OQS

[1812] John van de Wetering. Ordering quantum states and channels based on positive Bayesian evidence. Journal of Mathematical Physics, 59(10): 102201, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiai:2018:QDN

[1813] Fumio Hiai. Quantum f-divergences in von Neumann algebras. I. Standard f-divergences. Journal of Mathematical Physics, 59(10):102202, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Hermes:2018:DLM

[1814] Alexander Müller-Hermes. Decomposability of linear maps under tensor powers. Journal of Mathematical Physics, 59(10):102203, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bialonczyk:2018:AST

[1815] Michal Bialończyk, Andrzej Jamiolkowski, and Karol Zyczkowski. Application of Shemesh theorem to quantum channels. Journal of Mathematical Physics, 59(10):102204, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gesmundo:2018:MPS

[1816] Fulvio Gesmundo, J. M. Landsberg, and Michael Walter. Matrix product states and the quantum max-flow/min-cut conjectures. Journal of Mathematical Physics, 59(10):102205, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuramochi:2018:EBC

[1817] Yui Kuramochi. Entanglement-breaking channels with general outcome operator algebras. Journal of Mathematical Physics, 59(10):102206, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardona:2018:MDK

[1818] S. A. H. Cardona, H. García-Compeán, and A. Martínez-Merino. On the Moyal deformation of Kapustin–Witten systems. Journal of Mathematical Physics, 59(10):102301, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:2018:FFS

[1819] Stephen L. Adler. Free field structure of the model with a spin-32 Rarita– Schwinger field directly coupled to a spin-12 field. Journal of Mathematical Physics, 59(10):102302, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paetz:2018:SCS

[1820] Tim-Torben Paetz. On the smoothness of the critical sets of the cylinder at spatial infinity in vacuum spacetimes. Journal of Mathematical Physics, 59(10):102501, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Makino:2018:SRA

[1821] Tetu Makino. On slowly rotating axisymmetric solutions of the Einstein– Euler equations. Journal of Mathematical Physics, 59(10):102502, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huh:2018:ABS

[1822] Hyungjin Huh, Seung-Yeal Ha, and Dohyun Kim. Asymptotic behavior and stability for the Schrödinger–Lohe model. Journal of Mathematical Physics, 59(10):102701, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Escobar-Ruiz:2018:TCP

[1823] M. A. Escobar-Ruiz and C. A. Escobar. Three charges on a plane in a magnetic field: Special trajectories. Journal of Mathematical Physics, 59 (10):102901, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schoberl:2018:EBC

[1824] Markus Schöberl and Kurt Schlacher. On the extraction of the boundary conditions and the boundary ports in second-order field theories. Journal of Mathematical Physics, 59(10):102902, October 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basu:2018:RIR

[1825] B. Basu and C. I. Martin. Resonant interactions of rotational water waves in the equatorial f-plane approximation. Journal of Mathematical Physics, 59(10):103101, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhan:2018:CMD

[1826] Longjun Zhan, Gordon Blower, Yang Chen, and Mengkun Zhu. Center of mass distribution of the Jacobi unitary ensembles: Painlevé V, asymptotic expansions. Journal of Mathematical Physics, 59(10):103301, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2018:EMZ

[1827] Yuanran Zhu, Jason M. Dominy, and Daniele Venturi. On the estimation of the Mori–Zwanzig memory integral. Journal of Mathematical Physics, 59(10):103501, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Igarashi:2018:TPS

[1828] Hikaru Igarashi and Kouichi Takemura. On two-parameter solutions of simultaneous ultradiscrete Painlevé II equation with parity variables. Journal of Mathematical Physics, 59(10):103502, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2018:CAK

[1829] Shoufeng Shen, Chunxia Li, Yongyang Jin, and Wen-Xiu Ma. Completion of the Ablowitz–Kaup–Newell–Segur integrable coupling. Journal of Mathematical Physics, 59(10):103503, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2018:DWA

[1830] Rui Wang, Shao-Kui Yao, Min-Li Li, Ke Wu, and Wei-Zhong Zhao. On deformations of the Witt n-algebra. Journal of Mathematical Physics, 59 (10):103504, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2018:MEM

[1831] Congming Jin and Jiu Ding. A maximum entropy method for solving the boundary value problem of second order ordinary differential equations. Journal of Mathematical Physics, 59(10):103505, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muraleetharan:2018:FOE

[1832] B. Muraleetharan and K. Thirulogasanthar. Fredholm operators and essential S-spectrum in the quaternionic setting. Journal of Mathematical Physics, 59(10):103506, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chamazkoti:2018:GTP

[1833] Rohollah Bakhshandeh Chamazkoti and Martin Oberlack. Geometry of two-point velocity correlation tensor of homogeneous isotropic turbulence with helicity in sol-manifold coordinates. Journal of Mathematical Physics, 59(10):103507, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mansour:2018:CED

[1834] Toufik Mansour and Ranya Rayan. On Cauchy–Euler’s differential equation involving a para-Grassmann variable. Journal of Mathematical Physics, 59(10):103508, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kharchev:2018:QCH

[1835] S. Kharchev, A. Levin, M. Olshanetsky, and A. Zotov. Quasi-compact Higgs bundles and Calogero–Sutherland systems with two types of spins. Journal of Mathematical Physics, 59(10):103509, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guin:2018:SAN

[1836] Satyajit Guin. Subadditivity and additivity of the noncommutative Yang–Mills action functional. Journal of Mathematical Physics, 59(10): 103510, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hilfer:2018:MLP

[1837] R. Hilfer. Multiscale local porosity theory, weak limits, and dielectric response in composite and porous media. Journal of Mathematical Physics, 59(10):103511, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penz:2018:REE

[1838] Markus Penz. Regularity for evolution equations with non-autonomous perturbations in Banach spaces. Journal of Mathematical Physics, 59 (10):103512, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2018:EMR

[1839] Yueqiang Song and Shaoyun Shi. Erratum: “Multiplicity results for stationary Kirchhoff problems involving fractional elliptic operator and critical nonlinearity in RN” [j. math. phys. 59, 071512 (2018)]. Journal of Mathematical Physics, 59(10):109901, October 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1673].

Nachtergaele:2018:ENS

[1840] Bruno Nachtergaele. Editor’s note: Spontaneous SU_{2}(C) symmetry breaking in the ground states of quantum spin chain [J. Math. Phys. 59, 111701 (2018)]. Journal of Mathematical Physics, 59(11):110401, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1845].

Conde:2018:ASG

[1841] J. M. Conde and F. Güngör. Analysis of the symmetry group and exact solutions of the dispersionless KP equation in n + 1 dimensions. Journal of Mathematical Physics, 59(11):111501, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zheng:2018:FNI

[1842] Jiqiang Zheng. Focusing NLS with inverse square potential. Journal of Mathematical Physics, 59(11):111502, November 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2018:TPS

[1843] Yu-Zhu Wang and Yanshuo Li. Time periodic solutions to the beam equation with weak damping. Journal of Mathematical Physics, 59(11): 111503, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2018:SHT

[1844] Dian Hu and Wancheng Sheng. Stability of e-h type regular shock refraction. Journal of Mathematical Physics, 59(11):111504, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohari:2018:SSC

[1845] Anilesh Mohari. Spontaneous SU_{2}(C) symmetry breaking in the ground states of quantum spin chain. Journal of Mathematical Physics, 59(11): 111701, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See editor’s note [1840].

Lawson:2018:LRQ

[1846] Latévi M. Lawson, Gabriel Y. H. Avossevou, and Laure Gouba. Lewis– Riesenfeld quantization and SU(1, 1) coherent states for 2D damped harmonic oscillator. Journal of Mathematical Physics, 59(11):112101, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feintzeig:2018:CLS

[1847] Benjamin H. Feintzeig. The classical limit of a state on the Weyl algebra. Journal of Mathematical Physics, 59(11):112102, November 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2018:LPI

[1848] Jianping Dong and Han Geng. Lévy path integrals of particle on circle and some applications. Journal of Mathematical Physics, 59(11):112103, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaspard:2018:CFB

[1849] David Gaspard. Connection formulas between Coulomb wave functions. Journal of Mathematical Physics, 59(11):112104, November 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harris:2018:SMM

[1850] Samuel J. Harris, Rupert H. Levene, Vern I. Paulsen, Sarah Plosker, and Mizanur Rahaman. Schur multipliers and mixed unitary maps. Journal of Mathematical Physics, 59(11):112201, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bluhm:2018:JMQ

[1851] Andreas Bluhm and Ion Nechita. Joint measurability of quantum effects and the matrix diamond. Journal of Mathematical Physics, 59(11): 112202, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ertem:2018:HST

[1852] Ümit Ertem. Harmonic spinors from twistors and potential forms. Journal of Mathematical Physics, 59(11):112301, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiessling:2018:QMS

[1853] Michael K.-H. Kiessling and A. Shadi Tahvildar-Zadeh. On the quantummechanics of a single photon. Journal of Mathematical Physics, 59(11):

112302, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geloun:2018:RET

[1854] Joseph Ben Geloun and Reiko Toriumi. Renormalizable enhanced tensor field theory: The quartic melonic case. Journal of Mathematical Physics, 59(11):112303, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fletcher:2018:GSS

[1855] Samuel C. Fletcher. Global spacetime similarity. Journal of Mathematical Physics, 59(11):112501, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:QPS

[1856] Hepeng Li and Xiaoping Yuan. Quasi-periodic solution for the complex Ginzburg–Landau equation with continuous spectrum. Journal of Mathematical Physics, 59(11):112701, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2018:RAS

[1857] Seung-Yeal Ha and Doheon Kim. Robustness and asymptotic stability for the winfree model on a general network under the effect of time-delay. Journal of Mathematical Physics, 59(11):112702, November 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2018:QPS

[1858] Yanling Shi and Xindong Xu. Quasi-periodic solutions for generalized Kaup system with quasi-periodic forcing. Journal of Mathematical Physics, 59(11):112703, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonzalez-Navarrete:2018:NMR

[1859] Manuel González-Navarrete and Rodrigo Lambert. Non-Markovian random walks with memory lapses. Journal of Mathematical Physics, 59(11): 113301, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Betz:2018:PTL

[1860] Volker Betz, Johannes Ehlert, and Benjamin Lees. Phase transition for loop representations of quantum spin systems on trees. Journal of Mathematical Physics, 59(11):113302, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juhasz:2018:CCS

[1861] David Juhasz, Miro Kolesik, and Per Kristen Jakobsen. Convergence and completeness for square-well Stark resonant state expansions. Journal of Mathematical Physics, 59(11):113501, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adem:2018:TET

[1862] Alejandro Adem, José Cantarero, and José Manuel Gómez. Twisted equivariant K-theory of compact Lie group actions with maximal rank isotropy. Journal of Mathematical Physics, 59(11):113502, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2018:GSE

[1863] Axel Schulze-Halberg and Özlem Yesiltas. Generalized Schrödinger equations with energy-dependent potentials: Formalism and applications. Journal of Mathematical Physics, 59(11):113503, November 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koutvitsky:2018:PAH

[1864] Vladimir A. Koutvitsky and Eugene M. Maslov. Perturbation approach to the Hill equation with a slowly varying parameter with application to the inflationary cosmology. Journal of Mathematical Physics, 59(11): 113504, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ge:2018:NCE

[1865] Yanyan Ge and Dafeng Zuo. A new class of Euler equation on the dual of the n = 1 extended Neveu–Schwarz algebra. Journal of Mathematical Physics, 59(11):113505, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramani:2018:IRB

[1866] A. Ramani, B. Grammaticos, and T. Tamizhmani. Investigating relations between discrete Painlevé equations: The multistep approach. Journal of Mathematical Physics, 59(11):113506, November 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2018:SSH

[1867] A. D. Alhaidari. Series solutions of Heun-type equation in terms of orthogonal polynomials. Journal of Mathematical Physics, 59(11):113507, November 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2043].

Nachtergaele:2018:GW

[1868] Bruno Nachtergaele. Goodbye and welcome! Journal of Mathematical Physics, 59(12):120401, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:SSC

[1869] Jianli Liu and Yaoli Fang. Singularities of solutions to the compressible Euler equations and Euler–Poisson equations with damping. Journal of Mathematical Physics, 59(12):121501, December 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2018:CFP

[1870] Wenjing Chen. Critical fractional p-Kirchhoff type problem with a generalized Choquard nonlinearity. Journal of Mathematical Physics, 59(12): 121502, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:TGS

[1871] Wulong Liu and Guowei Dai. Three ground state solutions for double phase problem. Journal of Mathematical Physics, 59(12):121503, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:SSN

[1872] Bingchen Liu, Mengzhen Dong, and Fengjie Li. Singular solutions in nonlinear parabolic equations with anisotropic nonstandard growth conditions. Journal of Mathematical Physics, 59(12):121504, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2018:SES

[1873] Xingxing Liu. Stability in the energy space of the sum of n peakons for a modified Camassa–Holm equation with higher-order nonlinearity. Journal of Mathematical Physics, 59(12):121505, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:MPS

[1874] Fuyi Li, Ting Rong, and Zhanping Liang. Multiple positive solutions for a class of (2, p)-Laplacian equation. Journal of Mathematical Physics, 59(12):121506, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jenssen:2018:SFC

[1875] Helge Kristian Jenssen and Charis Tsikkou. On similarity flows for the compressible Euler system. Journal of Mathematical Physics, 59(12): 121507, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2018:SPT

[1876] Jing Wang and Shixiang Ma. On the steady Prandtl type equations with magnetic effects arising from 2D incompressible MHD equations in a half plane. Journal of Mathematical Physics, 59(12):121508, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qian:2018:GSN

[1877] Aixia Qian, Jingmei Liu, and Anmin Mao. Ground state and nodal solutions for a Schrödinger–Poisson equation with critical growth. Journal of Mathematical Physics, 59(12):121509, December 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bryenton:2018:CTC

[1878] Nicolas Bryenton, Cameron Davies, Andrew Douglas, Saminul Haque, Kieran Hauser, Thaddeus Janisse, Matthew Koster, Jessica (Chengjin) Liu, Thomas Ma, Yuveshen Mooroogen, Arthur Rabinovich, Joe Repka, Saied Sorkhou, and Matt Wu. A comparison of two classifications of solvable Lie algebras. Journal of Mathematical Physics, 59(12):121701, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2018:SPA

[1879] Sangjib Kim, Soo Teck Lee, and Yi Wang. Skew Pieri algebras of the general linear group. Journal of Mathematical Physics, 59(12):121702, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lemay:2018:BBI

[1880] Jean-Michel Lemay and Luc Vinet. Bivariate bannai-Ito polynomials. Journal of Mathematical Physics, 59(12):121703, December 2018. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spitkovsky:2018:SQP

[1881] Ilya M. Spitkovsky and Stephan Weis. Signatures of quantum phase transitions from the boundary of the numerical range. Journal of Mathematical Physics, 59(12):121901, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heim:2018:RFM

[1882] D. M. Heim. Recursive formulation of Madelung continuity equation leads to propagation equation. Journal of Mathematical Physics, 59(12): 122101, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maioli:2018:ESL

[1883] A. C. Maioli and Alexandre G. M. Schmidt. Exact solution to Lippmann– Schwinger equation for a circular billiard. Journal of Mathematical Physics, 59(12):122102, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2018:ENE

[1884] Richard L. Hall, Nasser Saad, and K. D. Sen. Exact normalized eigenfunctions for general deformed Hulthén potentials. Journal of Mathematical Physics, 59(12):122103, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See correction [1982].

Palma:2018:CEP

[1885] Giacomo De Palma and Stefan Huber. The conditional entropy power inequality for quantum additive noise channels. Journal of Mathematical Physics, 59(12):122201, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2018:CBO

[1886] Li Gao, Marius Junge, and Nicholas LaRacuente. Capacity bounds via operator space methods. Journal of Mathematical Physics, 59(12): 122202, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adamo:2018:MYM

[1887] Tim Adamo, David Skinner, and Jack Williams. Minitwistors and 3d Yang–Mills–Higgs theory. Journal of Mathematical Physics, 59(12): 122301, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Svoboda:2018:ASP

[1888] David Svoboda. Algebroid structures on para-Hermitian manifolds. Journal of Mathematical Physics, 59(12):122302, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zarmi:2018:SGF

[1889] Yair Zarmi. Soliton-generating τ-functions revisited. Journal of Mathematical Physics, 59(12):122701, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bambusi:2018:RDS

[1890] D. Bambusi and R. Montalto. Reducibility of 1-d Schrödinger equation with unbounded time quasiperiodic perturbations. III. Journal of Mathematical Physics, 59(12):122702, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kwon:2018:DQC

[1891] Young-Sam Kwon. From the degenerate quantum compressible Navier– Stokes–Poisson system to incompressible Euler equations. Journal of Mathematical Physics, 59(12):123101, December 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kluczek:2018:PFP

[1892] Mateusz Kluczek. Physical flow properties for pollard-like internal water waves. Journal of Mathematical Physics, 59(12):123102, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:RDB

[1893] Yachun Li, Yue-Jun Peng, and Shuai Xi. Rigorous derivation of a Boltzmann relation from isothermal Euler–Poisson systems. Journal of Mathematical Physics, 59(12):123501, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2018:KWL

[1894] Gehao Wang. From Kontsevich–Witten to linear Hodge integrals via Virasoro operators. Journal of Mathematical Physics, 59(12):123502, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2018:SRN

[1895] Chuanzhong Li. Symmetries and reductions on the noncommutative Kadomtsev–Petviashvili and Gelfand–Dickey hierarchies. Journal of Mathematical Physics, 59(12):123503, December 2018. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dehesa:2018:LKL

[1896] J. S. Dehesa, J. J. Moreno-Balcázar, and I. V. Toranzo. Linearization and Krein-like functionals of hypergeometric orthogonal polynomials. Journal of Mathematical Physics, 59(12):123504, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blumenhagen:2018:UBQ

[1897] Ralph Blumenhagen, Max Brinkmann, Vladislav Kupriyanov, and Matthias Traube. On the uniqueness of L_{∞} bootstrap: Quasiisomorphisms are Seiberg–Witten maps. Journal of Mathematical Physics, 59(12):123505, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nachtergaele:2018:REV

[1898] Bruno Nachtergaele. Retraction: “Existence of vortices in nonlinear optics” [j. math. phys. 59, 101509 (2018)]. Journal of Mathematical Physics, 59(12):129901, December 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1797].

Solovej:2019:LJF

[1899] Jan Philip Solovej. Leading JMP into the future. Journal of Mathematical Physics, 60(1):010401, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samoilenko:2019:ASL

[1900] Valerii Samoilenko and Yuliia Samoilenko. Asymptotic soliton-like solutions to the singularly perturbed Benjamin–Bona–Mahony equation with variable coefficients. Journal of Mathematical Physics, 60(1):011501, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:PPS

[1901] Hengling Wang and Yuxiang Li. On a parabolic-parabolic system with gradient dependent chemotactic coefficient and consumption. Journal of Mathematical Physics, 60(1):011502, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2019:EPS

[1902] Claudianor O. Alves and Giovany M. Figueiredo. Existence of positive solution for a planar Schrödinger–Poisson system with exponential growth. Journal of Mathematical Physics, 60(1):011503, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2019:LTA

[1903] Nan Liu and Boling Guo. Long-time asymptotics for the Sasa–Satsuma equation via nonlinear steepest descent method. Journal of Mathematical Physics, 60(1):011504, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chemetov:2019:WSU

[1904] Nikolai V. Chemetov, Sárka Necasová, and Boris Muha. Weak-strong uniqueness for fluid-rigid body interaction problem with slip boundary condition. Journal of Mathematical Physics, 60(1):011505, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:IMS

[1905] Wang Li, Vicentiu D. Radulescu, and Binlin Zhang. Infinitely many solutions for fractional Kirchhoff–Schrödinger–Poisson systems. Journal of Mathematical Physics, 60(1):011506, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:BHD

[1906] Jianping Wang and Mingxin Wang. Boundedness in the higherdimensional Keller–Segel model with signal-dependent motility and logistic growth. Journal of Mathematical Physics, 60(1):011507, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:DSV

[1907] Yu Zhang and Yanyan Zhang. Delta-shocks and vacuums in the relativistic Euler equations for isothermal fluids with the flux approximation. Journal of Mathematical Physics, 60(1):011508, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaibhav:2019:NBO

[1908] Vishal Vaibhav. On the nonreflecting boundary operators for the general two dimensional Schrödinger equation. Journal of Mathematical Physics, 60(1):011509, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benhassine:2019:NDE

[1909] Abderrazek Benhassine. On nonlinear Dirac equations. Journal of Mathematical Physics, 60(1):011510, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:WPC

[1910] Fei Chen, Boling Guo, and Lan Zeng. The well-posedness for the Cauchy problem of the double-diffusive convection system. Journal of Mathematical Physics, 60(1):011511, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2019:MPS

[1911] Xiaomei Sun and Anqiang Zhu. Multi-peak solutions for nonlinear Choquard equation in the plane. Journal of Mathematical Physics, 60 (1):011512, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bie:2019:BBG

[1912] Hendrik De Bie, Plamen Iliev, and Luc Vinet. Bargmann and Barut– Girardello models for the Racah algebra. Journal of Mathematical Physics, 60(1):011701, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Asai:2019:PTO

[1913] Nobuhiro Asai and Hiroaki Yoshida. Poisson type operators on the Fock space of type B. Journal of Mathematical Physics, 60(1):011702, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marinho:2019:HPF

[1914] Andre A. Marinho and Francisco A. Brito. Hermite polynomials and Fibonacci oscillators. Journal of Mathematical Physics, 60(1):012101, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loran:2019:PIS

[1915] Farhang Loran and Ali Mostafazadeh. Potentials with identical scattering properties below a critical energy. Journal of Mathematical Physics, 60 (1):012102, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maiz:2019:DPT

[1916] F. Maiz. Development of the perturbation theory using polynomial solutions. Journal of Mathematical Physics, 60(1):012103, January 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bebiano:2019:NSA

[1917] N. Bebiano and J. da Providência. Non-self-adjoint operators with real spectra and extensions of quantum mechanics. Journal of Mathematical Physics, 60(1):012104, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aleixo:2019:CBC

[1918] A. N. F. Aleixo and A. B. Balantekin. A consistent bridge for the construction of generalized nonclassical states for confined systems through the inclusion of inverse generators in the algebraic formalism. Journal of Mathematical Physics, 60(1):012105, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramirez:2019:DFD

[1919] Romina Ramírez and Marta Reboiro. Dynamics of finite dimensional non-Hermitian systems with indefinite metric. Journal of Mathematical Physics, 60(1):012106, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cedzich:2019:QWE

[1920] C. Cedzich, T. Geib, A. H. Werner, and R. F. Werner. Quantum walks in external gauge fields. Journal of Mathematical Physics, 60(1):012107, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reuvers:2019:LBE

[1921] Robin Reuvers. Lower bound on entanglement in subspaces defined by Young diagrams. Journal of Mathematical Physics, 60(1):012201, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rouze:2019:CQS

[1922] Cambyse Rouzé and Nilanjana Datta. Concentration of quantum states from quantum functional and transportation cost inequalities. Journal of Mathematical Physics, 60(1):012202, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiai:2019:QFD

[1923] Fumio Hiai. Quantum f-divergences in von Neumann algebras. II. Maximal f-divergences. Journal of Mathematical Physics, 60(1):012203, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anshu:2019:NOO

[1924] Anurag Anshu, Rahul Jain, and Naqueeb Ahmad Warsi. On the nearoptimality of one-shot classical communication over quantum channels. Journal of Mathematical Physics, 60(1):012204, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Livine:2019:HFL

[1925] Etera R. Livine. Hamiltonian flows of Lorentzian polyhedra: Kapovich– Millson phase space and SU(1, 1) intertwiners. Journal of Mathematical Physics, 60(1):012301, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsenko:2019:DHF

[1926] A. A. Bytsenko, M. Chaichian, and A. E. Gonçalves. The Duistermaat– Heckman formula with application to circle actions and Poincaré qpolynomials in twisted equivariant K-theory. Journal of Mathematical Physics, 60(1):012302, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zucchini:2019:ERG

[1927] Roberto Zucchini. Exact renormalization group and effective action: A Batalin–Vilkovisky algebraic formulation. Journal of Mathematical Physics, 60(1):012303, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tehrani:2019:QBC

[1928] Mojtaba Taslimi Tehrani. Quantum BRST charge in gauge theories in curved space-time. Journal of Mathematical Physics, 60(1):012304, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kremer:2019:EGI

[1929] Attay Kremer. An electro-gravitational induction effect. Journal of Mathematical Physics, 60(1):012501, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vatandoost:2019:NCP

[1930] Mehdi Vatandoost, Rahimeh Pourkhandani, and Neda Ebrahimi. On null and causal pseudoconvex space-times. Journal of Mathematical Physics, 60(1):012502, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bialecki:2019:CNS

[1931] Mariusz Bialecki. Catalan numbers out of a stochastic cellular automaton. Journal of Mathematical Physics, 60(1):012701, January 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grochowski:2019:CTS

[1932] Marek Grochowski. Causality and topology in sub-Lorentzian geometry. Journal of Mathematical Physics, 60(1):012702, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McKinley:2019:NRE

[1933] Damaris McKinley, Daniel Pasca, and Cristina Stoica. Notes on relative equilibria of isosceles molecules in classical approximation. Journal of Mathematical Physics, 60(1):012901, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katori:2019:MDA

[1934] Makoto Katori. Macdonald denominators for affine root systems, orthogonal theta functions, and elliptic determinantal point processes. Journal of Mathematical Physics, 60(1):013301, January 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durand:2019:ABF

[1935] Loyal Durand. Asymptotic Bessel-function expansions for Legendre and Jacobi functions. Journal of Mathematical Physics, 60(1):013501, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tica:2019:FPI

[1936] Christian D. Tica and Eric A. Galapon. Finite-part integration of the generalized Stieltjes transform and its dominant asymptotic behavior for small values of the parameter. II. non-integer orders. Journal of Mathematical Physics, 60(1):013502, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ustinov:2019:IGS

[1937] N. V. Ustinov. Integrable generalizations of the sine-Gordon, short pulse, and reduced Maxwell–Bloch equations. Journal of Mathematical Physics, 60(1):013503, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banerjee:2019:GRL

[1938] R. Banerjee and M. Niedermaier. Graph rules for the linked cluster expansion of the Legendre effective action. Journal of Mathematical Physics, 60(1):013504, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falek:2019:BOU

[1939] M. Falek, M. Merad, and M. Moumni. Bosonic oscillator under a uniform magnetic field with Snyder–de Sitter algebra. Journal of Mathematical Physics, 60(1):013505, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2324].

Sulaiman:2019:VPP

[1940] Samira Alamin Sulaiman. Vortex patch problem for Euler–Boussinesq systems. Journal of Mathematical Physics, 60(1):013506, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2321].

Braga:2019:ANI

[1941] Gastão A. Braga, Jussara M. Moreira, and Camila F. Souza. Asymptotics for nonlinear integral equations with a generalized heat kernel using renormalization group technique. Journal of Mathematical Physics, 60 (1):013507, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alpay:2019:DSN

[1942] Daniel Alpay, Ismael L. Paiva, and Daniele C. Struppa. Distribution spaces and a new construction of stochastic processes associated with the Grassmann algebra. Journal of Mathematical Physics, 60(1):013508, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alshammari:2019:COC

[1943] Fahad Alshammari, Phillip S. Isaac, and Ian Marquette. On Casimir operators of conformal Galilei algebras. Journal of Mathematical Physics, 60(1):013509, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:NAB

[1944] Tao Zhang, Shuanhong Wang, and Dingguo Wang. A new approach to braided monoidal categories. Journal of Mathematical Physics, 60 (1):013510, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calvino-Louzao:2019:CBF

[1945] E. Calviño-Louzao, E. García-Río, P. Gilkey, I. Gutiérrez-Rodríguez, and R. Vázquez-Lorenzo. Constructing Bach flat manifolds of signature (2,

2) using the modified Riemannian extension. Journal of Mathematical Physics, 60(1):013511, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weder:2019:EST

[1946] Ricardo Weder. Erratum: Scattering theory for the matrix Schrödinger operator on the half line with general boundary conditions [J. Math. Phys. 56, 092103 (2015)]. Journal of Mathematical Physics, 60(1): 019901, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [362].

Weder:2019:ENE

[1947] Ricardo Weder. Erratum: The number of eigenvalues of the matrix Schrödinger operator on the half line with general boundary conditions [J. Math. Phys. 58, 102107 (2017)]. Journal of Mathematical Physics, 60 (1):019902, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1296].

Lu:2019:RA

[1948] Jan Philip Solovej Lu. Referee acknowledgment for 2018. Journal of Mathematical Physics, 60(2):020201, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

An:2019:SCS

[1949] Xiaoming An, Shuangjie Peng, and Chaodong Xie. Semi-classical solutions for fractional Schrödinger equations with potential vanishing at infinity. Journal of Mathematical Physics, 60(2):021501, February 2019.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2019:SCQ

[1950] Fengshuang Gao and Yuxia Guo. Solutions for critical quasilinear elliptic equations in RN. Journal of Mathematical Physics, 60(2):021502, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valli:2019:VIB

[1951] Alberto Valli. A variational interpretation of the Biot–Savart operator and the helicity of a bounded domain. Journal of Mathematical Physics, 60(2):021503, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huo:2019:CSR

[1952] Saisai Huo and Changhua Wei. Classical solution to relativistic Burger’s equation in SdS and SAdS space-times. Journal of Mathematical Physics, 60(2):021504, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhattarai:2019:ESS

[1953] Santosh Bhattarai. Existence and stability of standing waves for coupled nonlinear Hartree type equations. Journal of Mathematical Physics, 60 (2):021505, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyamoto:2019:EEE

[1954] Yasuhito Miyamoto and Tohru Wakasa. Exact eigenvalues and eigenfunctions for a one-dimensional Gel’fand problem. Journal of Mathematical Physics, 60(2):021506, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2019:ESC

[1955] Lan Huang and Ivan Drazic. Exponential stability for the compressible micropolar fluid with cylinder symmetry in R^{3}. Journal of Mathematical Physics, 60(2):021507, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2019:RCS

[1956] Guanghui Jin and Hyungjin Huh. Remarks on Chern–Simons gauged o (3) sigma model in one space dimension. Journal of Mathematical Physics, 60(2):021508, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krishnan:2019:ASW

[1957] Chethan Krishnan, K. V. Pavan Kumar, and P. N. Bala Subramanian. (anti-)symmetrizing wave functions. Journal of Mathematical Physics, 60 (2):021701, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueroa-OFarrill:2019:CLA

[1958] José M. Figueroa-O’Farrill. Conformal Lie algebras via deformation theory. Journal of Mathematical Physics, 60(2):021702, February 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milas:2019:POH

[1959] Antun Milas, Michael Penn, and Hanbo Shao. Permutation orbifolds of the Heisenberg vertex algebra H(3). Journal of Mathematical Physics, 60 (2):021703, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:2019:EHS

[1960] César R. de Oliveira, Luiz Hartmann, and Alessandra A. Verri. Effective Hamiltonians in surfaces of thin quantum waveguides. Journal of Mathematical Physics, 60(2):022101, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vershynina:2019:ERR

[1961] Anna Vershynina. Entanglement rates for Rényi, Tsallis, and other entropies. Journal of Mathematical Physics, 60(2):022201, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ge:2019:FGS

[1962] Yimin Ge, Jordi Tura, and J. Ignacio Cirac. Faster ground state preparation and high-precision ground energy estimation with fewer qubits. Journal of Mathematical Physics, 60(2):022202, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luczak:2019:CCO

[1963] Andrzej Luczak. Comparison of channels in operator algebras. Journal of Mathematical Physics, 60(2):022203, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2019:STC

[1964] Holger Boche, Gisbert Janßen, and Sajad Saeedinaeeni. Simultaneous transmission of classical and quantum information under channel uncertainty and jamming attacks. Journal of Mathematical Physics, 60(2): 022204, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wrochna:2019:CEB

[1965] Michal Wrochna. Conformal extension of the Bunch–Davies state across the de Sitter boundary. Journal of Mathematical Physics, 60(2):022301, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abouagwa:2019:SFD

[1966] Mahmoud Abouagwa and Ji Li. Stochastic fractional differential equations driven by Lévy noise under Carathéodory conditions. Journal of Mathematical Physics, 60(2):022701, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qin:2019:NSC

[1967] Hong Qin. A necessary and sufficient condition for the stability of linear Hamiltonian systems with periodic coefficients. Journal of Mathematical Physics, 60(2):022901, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dumitriu:2019:SGW

[1968] Ioana Dumitriu and Yizhe Zhu. Sparse general Wigner-type matrices: Local law and eigenvector delocalization. Journal of Mathematical Physics, 60(2):023301, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leinster:2019:SCR

[1969] Tom Leinster. A short characterization of relative entropy. Journal of Mathematical Physics, 60(2):023302, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:2019:DCV

[1970] Mahouton Norbert Hounkonnou and Fridolin Melong. R(p, q)-deformed conformal Virasoro algebra. Journal of Mathematical Physics, 60(2): 023501, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teixeira:2019:SFP

[1971] Renato Teixeira, Eduardo S. G. Leandro, Luiz C. B. da Silva, and Fernando Moraes. Schrödinger formalism for a particle constrained to a surface in R_{1}^{3}. Journal of Mathematical Physics, 60(2):023502, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galleas:2019:ESS

[1972] W. Galleas. On the elliptic }↕_{∈} solid-on-solid model: Functional relations and determinants. Journal of Mathematical Physics, 60(2):023503, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xia:2019:FMI

[1973] Baoqiang Xia. The Fokas method for integrable evolution equations on a time-dependent interval. Journal of Mathematical Physics, 60(2):023504, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Todino:2019:QCL

[1974] Anna Paola Todino. A quantitative central limit theorem for the excursion area of random spherical harmonics over subdomains of S^{2}. Journal of Mathematical Physics, 60(2):023505, February 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2019:PSI

[1975] Jimmy He. A power series identity and Bessel-type integrals over unitary groups. Journal of Mathematical Physics, 60(2):023506, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2019:ZGN

[1976] N. Aizawa, P. S. Isaac, and J. Segar. Z2×Z2 generalizations of n=1 superconformal Galilei algebras and their representations. Journal of Mathematical Physics, 60(2):023507, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2019:RRM

[1977] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials V: Racah and q-Racah types. Journal of Mathematical Physics, 60(2):023508, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaz:2019:RSU

[1978] Jayme Vaz, Jr. Representation of surfaces using spinor operators. Journal of Mathematical Physics, 60(2):023509, February 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kou:2019:SLQ

[1979] Kit Ian Kou, Wan-Kai Liu, and Yong-Hui Xia. Solve the linear quaternion-valued differential equations having multiple eigenvalues. Journal of Mathematical Physics, 60(2):023510, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brizard:2019:CES

[1980] Alain J. Brizard. Comment on “Exact solutions and singularities of an X-point collapse in Hall magnetohydrodynamics” [j. math. phys. 59, 061509 (2018)]. Journal of Mathematical Physics, 60(2):024101, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1617] and response [1981].

Janda:2019:RCE

[1981] Artur Z. Janda. Response to “Comment on ‘Exact solutions and singularities of an X-point collapse in Hall magnetohydrodynamics”’ [j. math. phys. 60, 024101 (2019)]. Journal of Mathematical Physics, 60(2):024102, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1617, 1980].

Hall:2019:PNE

[1982] Richard L. Hall, Nasser Saad, and K. D. Sen. Publisher’s note: “Exact normalized eigenfunctions for general deformed Hulthén potentials” [j. math. phys. 59, 122103 (2018)]. Journal of Mathematical Physics, 60 (2):029901, February 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1884].

Bar:2019:COO

[1983] Christian Bär. The curl operator on odd-dimensional manifolds. Journal of Mathematical Physics, 60(3):031501, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2019:GWP

[1984] Zhenhua Guo and Wenjing Song. Global well-posedness and large-time behavior of classical solutions to the 3D Navier–Stokes system with changed viscosities. Journal of Mathematical Physics, 60(3):031502, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:CLS

[1985] Jinlu Li, Minghua Yang, and Yanghai Yu. A class large solution of the 2D MHD equations with velocity and magnetic damping. Journal of Mathematical Physics, 60(3):031503, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rybalko:2019:LTA

[1986] Yan Rybalko and Dmitry Shepelsky. Long-time asymptotics for the integrable nonlocal nonlinear Schrödinger equation. Journal of Mathematical Physics, 60(3):031504, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2019:SFC

[1987] Rongfeng Yu. Singularity formation of the compressible non-barotropic flows with zero heat conductivity. Journal of Mathematical Physics, 60 (3):031505, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mourrat:2019:IIQ

[1988] J.-C. Mourrat. An informal introduction to quantitative stochastic homogenization. Journal of Mathematical Physics, 60(3):031506, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:NGW

[1989] Nengqiu Zhang. A new global well-posedness for the damped Boussinesq equations. Journal of Mathematical Physics, 60(3):031507, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qi:2019:MMO

[1990] Fei Qi. On modules for meromorphic open-string vertex algebras. Journal of Mathematical Physics, 60(3):031701, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gontier:2019:NCW

[1991] D. Gontier, A. Levitt, and S. Siraj-dine. Numerical construction of Wannier functions through homotopy. Journal of Mathematical Physics, 60 (3):031901, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bossmann:2019:DNS

[1992] Lea Boßmann. Derivation of the 1 d nonlinear Schrödinger equation from the 3 d quantum many-body dynamics of strongly confined bosons. Journal of Mathematical Physics, 60(3):031902, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petreska:2019:TDS

[1993] Irina Petreska, Antonio S. M. de Castro, Trifce Sandev, and Ervin K. Lenzi. The time-dependent Schrödinger equation in three dimensions under geometric constraints. Journal of Mathematical Physics, 60(3): 032101, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliveira:2019:ESS

[1994] M. D. Oliveira and Alexandre G. M. Schmidt. Exact solutions of Schrödinger and Pauli equations for a charged particle on a sphere and interacting with non-central potentials. Journal of Mathematical Physics, 60(3):032102, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiscop:2019:LFI

[1995] Irina Chiscop, Holger R. Dullin, Konstantinos Efstathiou, and Holger Waalkens. A Lagrangian fibration of the isotropic 3-dimensional harmonic oscillator with monodromy. Journal of Mathematical Physics, 60 (3):032103, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2019:MPS

[1996] Hao-Chung Cheng and Min-Hsiu Hsieh. Matrix Poincaré, Φ-Sobolev inequalities, and quantum ensembles. Journal of Mathematical Physics, 60(3):032201, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kribs:2019:QEC

[1997] David W. Kribs, Comfort Mintah, Michael Nathanson, and Rajesh Pereira. Quantum error correction and one-way LOCC state distinguishability. Journal of Mathematical Physics, 60(3):032202, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bravyi:2019:AAQ

[1998] Sergey Bravyi, David Gosset, Robert König, and Kristan Temme. Approximation algorithms for quantum many-body problems. Journal of Mathematical Physics, 60(3):032203, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anastopoulos:2019:TAL

[1999] Charis Anastopoulos and Ntina Savvidou. Time of arrival and localization of relativistic particles. Journal of Mathematical Physics, 60(3): 032301, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McNutt:2019:PDL

[2000] David Duncan McNutt and Matthew Terje Aadne. I-preserving diffeomorphisms of Lorentzian manifolds. Journal of Mathematical Physics, 60(3):032501, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ryan:2019:SVW

[2001] Paul Ryan and Dmytro Volin. Separated variables and wave functions for rational gl(N) spin chains in the companion twist frame. Journal of Mathematical Physics, 60(3):032701, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2019:PAN

[2002] Kaixuan Zhu, Yongqin Xie, Feng Zhou, and Xin Li. Pullback attractors for the non-autonomous reaction-diffusion equations in RN. Journal of Mathematical Physics, 60(3):032702, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alshal:2019:GHS

[2003] H Alshal and T Curtright. Grounded hyperspheres as squashed wormholes. Journal of Mathematical Physics, 60(3):032901, March 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Owczarek:2019:ESP

[2004] A. L. Owczarek and T. Prellberg. Exact solution of pulled, directed vesicles with sticky walls in two dimensions. Journal of Mathematical Physics, 60(3):033301, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agliari:2019:RHN

[2005] Elena Agliari, Adriano Barra, and Matteo Notarnicola. The relativistic Hopfield network: Rigorous results. Journal of Mathematical Physics, 60(3):033302, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valero:2019:COS

[2006] Carlos Valero and Raymond G. McLenaghan. Classification of the orthogonal separable webs for the Hamilton–Jacobi and Laplace–Beltrami equations on 3-dimensional hyperbolic and de Sitter spaces. Journal of Mathematical Physics, 60(3):033501, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zabrodin:2019:TDS

[2007] A. Zabrodin. Time discretization of the spin Calogero–Moser model and the semi-discrete matrix KP hierarchy. Journal of Mathematical Physics, 60(3):033502, March 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:GSS

[2008] Yuan Li, Dun Zhao, and Qingxuan Wang. Ground state solution and nodal solution for fractional nonlinear Schrödinger equation with indefinite potential. Journal of Mathematical Physics, 60(4):041501, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassano:2019:BTD

[2009] Biagio Cassano and Fabio Pizzichillo. Boundary triples for the Dirac operator with Coulomb-type spherically symmetric perturbations. Journal of Mathematical Physics, 60(4):041502, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manna:2019:HOR

[2010] Utpal Manna and Akash A. Panda. Higher order regularity and blow-up criterion for semi-dissipative and ideal Boussinesq equations. Journal of Mathematical Physics, 60(4):041503, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:EPS

[2011] Qi Zhang. Existence of positive solution to Kirchhoff–Schrödinger– Poisson system with strong singular term. Journal of Mathematical Physics, 60(4):041504, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2019:LWP

[2012] Yingying Guo and Zhaoyang Yin. Local well-posedness and blow-up phenomena of the generalized short pulse equation. Journal of Mathematical Physics, 60(4):041505, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouzettouta:2019:ESF

[2013] Lamine Bouzettouta and Djebabla Abdelhak. Exponential stabilization of the full von Kármán beam by a thermal effect and a frictional damping and distributed delay. Journal of Mathematical Physics, 60(4):041506, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2019:MHW

[2014] Chengkang Xu, Wenqi Luo, and Fen Zhang. Modules of highest weight type for the algebra of derivations over a rational quantum torus. Journal of Mathematical Physics, 60(4):041701, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goto:2019:BAM

[2015] Yukimi Goto. Binding of atoms in Müller theory. Journal of Mathematical Physics, 60(4):041901, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dattoli:2019:GIM

[2016] G. Dattoli and H. Fares. The Gaussian integration method of the Schrödinger equation and quantum 1-D theory of low gain free electron laser. Journal of Mathematical Physics, 60(4):042101, April 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2019:SGS

[2017] N. Aizawa, I. E. Cunha, Z. Kuznetsova, and F. Toppan. On the spectrumgenerating superalgebras of the deformed one-dimensional quantum oscillators. Journal of Mathematical Physics, 60(4):042102, April 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolotin:2019:PCF

[2018] Arkady Bolotin. Propositional counter-factual definiteness and the EPR paradox. Journal of Mathematical Physics, 60(4):042103, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wong:2019:NAT

[2019] Chieh-Lei Wong. Non-adiabatic transitions in a tilted conical intersection. Journal of Mathematical Physics, 60(4):042104, April 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hastings:2019:SFF

[2020] M. B. Hastings. The stability of free Fermi Hamiltonians. Journal of Mathematical Physics, 60(4):042201, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Filippov:2019:QIP

[2021] Sergey N. Filippov and Ksenia V. Kuzhamuratova. Quantum informational properties of the Landau–Streater channel. Journal of Mathematical Physics, 60(4):042202, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soltan:2019:QSS

[2022] P. M. Sołtan. Quantum semigroups from synchronous games. Journal of Mathematical Physics, 60(4):042203, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beghetto:2019:RIM

[2023] D. Beghetto, R. J. Bueno Rogerio, and C. H. Coronado Villalobos. The restricted Inomata–McKinley spinor-plane, homotopic deformations and the lounesto classification. Journal of Mathematical Physics, 60(4): 042301, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lienert:2019:IRQ

[2024] Matthias Lienert and Roderich Tumulka. Interacting relativistic quantum dynamics of two particles on spacetimes with a big bang singularity. Journal of Mathematical Physics, 60(4):042302, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Searight:2019:CDM

[2025] Trevor P. Searight. Completing the dark matter solutions in degenerate Kaluza–Klein theory. Journal of Mathematical Physics, 60(4):042501, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:2019:TSM

[2026] A. V. Tsiganov. Transformation of the Stäckel matrices preserving superintegrability. Journal of Mathematical Physics, 60(4):042701, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:EUS

[2027] Jian Zhang and Ji Shu. Existence and upper semicontinuity of random attractors for non-autonomous fractional stochastic Ginzburg–Landau equations. Journal of Mathematical Physics, 60(4):042702, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2019:POP

[2028] Jaume Llibre and Claudia Valls. Periodic orbits of the planar anisotropic generalized Kepler problem. Journal of Mathematical Physics, 60(4): 042901, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Massa:2019:SON

[2029] Enrico Massa and Stefano Vignolo. Small oscillations of non-dissipative Lagrangian systems. Journal of Mathematical Physics, 60(4):042902, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fukuda:2019:SEN

[2030] Ikuo Fukuda. Symmetric, explicit numerical integrator for molecular dynamics equations of motion with a generalized friction. Journal of Mathematical Physics, 60(4):042903, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Addario-Berry:2019:LOS

[2031] Louigi Addario-Berry, Luc Devroye, Gábor Lugosi, and Roberto I. Oliveira. Local optima of the Sherrington–Kirkpatrick Hamiltonian. Journal of Mathematical Physics, 60(4):043301, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2019:DEN

[2032] Xiaolei Yang and Yanbin Tang. Decay estimates of nonlocal diffusion equations in some particle systems. Journal of Mathematical Physics, 60(4):043302, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouten:2019:AEQ

[2033] Luc Bouten and John E. Gough. Asymptotic equivalence of quantum stochastic models. Journal of Mathematical Physics, 60(4):043501, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buschle:2019:NWI

[2034] L. R. Buschle, F. T. Kurz, H.-P. Schlemmer, and C. H. Ziener. Neumann– Weber integral transform for complex indices. Journal of Mathematical Physics, 60(4):043502, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:PTI

[2035] Daguang Chen, Haizhong Li, and Tailong Zhou. A Penrose type inequality for graphs over Reissner–Nordström-anti-deSitter manifold. Journal of Mathematical Physics, 60(4):043503, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xue:2019:AFC

[2036] Panpan Xue, Guang-Liang Li, Junpeng Cao, Jiao Li, Kun Hao, Tao Yang, and Wen-Li Yang. An anisotropic four-component spin chain with integrable boundary terms. Journal of Mathematical Physics, 60 (4):043504, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thuillier:2019:TMH

[2037] F. Thuillier. 3D topological models and Heegaard splitting. I. Partition function. Journal of Mathematical Physics, 60(4):043505, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parzygnat:2019:TDA

[2038] Arthur J. Parzygnat. Two-dimensional algebra in lattice gauge theory. Journal of Mathematical Physics, 60(4):043506, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poumai:2019:WHF

[2039] K. T. Poumai, S. K. Kaushik, and P. Mantry. Weyl–Heisenberg frames and Balian–Low theorem in ℓ^{2}(Z). Journal of Mathematical Physics, 60(4):043507, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalantarov:2019:BSH

[2040] Varga Kalantarov, Ali Mostafazadeh, and Neslihan Oflaz. Blow-up solutions of Helmholtz equation for a Kerr slab with a complex linear and nonlinear permittivity. Journal of Mathematical Physics, 60(4):043508, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boccagna:2019:CWS

[2041] Roberto Boccagna. Current with “wrong” sign and phase transitions. Journal of Mathematical Physics, 60(4):043509, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Obispo:2019:CEQ

[2042] Angel E. Obispo, Gisele B. Freitas, and Luis B. Castro. Comment on “Effective of the q-deformed pseudoscalar magnetic field on the charge carriers in graphene” [j. math. phys. 57, 082105 (2016)]. Journal of Mathematical Physics, 60(4):044101, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [805].

Alhaidari:2019:ESS

[2043] A. D. Alhaidari. Erratum: “Series solutions of Heun-type equation in terms of orthogonal polynomials” [j. math. phys. 59, 113507 (2018)]. Journal of Mathematical Physics, 60(4):049901, April 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1867].

Zhang:2019:EMC

[2044] Wei Zhang and Xian Wu. Existence, multiplicity, and concentration of positive solutions for a quasilinear Choquard equation with critical exponent. Journal of Mathematical Physics, 60(5):051501, May 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:VST

[2045] Yan-Hong Chen and Wenming Zou. Vector solutions for two coupled Schrödinger equations on Riemannian manifolds. Journal of Mathematical Physics, 60(5):051502, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yagdjian:2019:GES

[2046] Karen Yagdjian. Global existence of the self-interacting scalar field in the de Sitter universe. Journal of Mathematical Physics, 60(5):051503, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paranamana:2019:FMR

[2047] Pushpi Paranamana, Eugenio Aulisa, Akif Ibragimov, and Magdalena Toda. Fracture model reduction and optimization for Forchheimer flows in reservoirs. Journal of Mathematical Physics, 60(5):051504, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhong:2019:SSC

[2048] Xin Zhong. Strong solutions to the 2D Cauchy problem of densitydependent viscous Boussinesq equations with vacuum. Journal of Mathematical Physics, 60(5):051505, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2019:GSI

[2049] Jianjun Huang and Zhenglu Jiang. Global solution to the inelastic Boltzmann equation with hard potentials. Journal of Mathematical Physics, 60(5):051506, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2019:KVE

[2050] A. Alexandrou Himonas, Dionyssios Mantzavinos, and Fangchi Yan. The Korteweg-de Vries equation on an interval. Journal of Mathematical Physics, 60(5):051507, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2019:UTT

[2051] Seung-Yeal Ha, Dohyun Kim, Jeongho Kim, and Xiongtao Zhang. Uniform-in-time transition from discrete to continuous dynamics in the Kuramoto synchronization. Journal of Mathematical Physics, 60(5): 051508, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thompson:2019:CPD

[2052] Ryan C. Thompson. The Cauchy problem for the 1-D gurevich-zybin system. Journal of Mathematical Physics, 60(5):051509, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sen:2019:DSW

[2053] Anupam Sen and T. Raja Sekhar. Delta shock wave as self-similar viscosity limit for a strictly hyperbolic system of conservation laws. Journal of Mathematical Physics, 60(5):051510, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:2019:GSS

[2054] Xiaojing Feng. Ground state solutions for Schrödinger–Poisson systems involving the fractional Laplacian with critical exponent. Journal of Mathematical Physics, 60(5):051511, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:2019:ESS

[2055] Binhua Feng, Ruipeng Chen, and Jiajia Ren. Existence of stable standing waves for the fractional Schrödinger equations with combined power-type and Choquard-type nonlinearities. Journal of Mathematical Physics, 60 (5):051512, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mao:2019:SSC

[2056] Anmin Mao and Wenqing Wang. Signed and sign-changing solutions of bi-nonlocal fourth order elliptic problem. Journal of Mathematical Physics, 60(5):051513, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:2019:SCB

[2057] G. Bergeron, E. Koelink, and L. Vinet. SU_{q} (3) corepresentations and bivariate q-Krawtchouk polynomials. Journal of Mathematical Physics, 60(5):051701, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lemm:2019:SGF

[2058] Marius Lemm and Evgeny Mozgunov. Spectral gaps of frustration-free spin systems with boundary. Journal of Mathematical Physics, 60(5): 051901, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helland:2019:SSC

[2059] Inge S. Helland. Symmetry in a space of conceptual variables. Journal of Mathematical Physics, 60(5):052101, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2320].

Aleixo:2019:CCB

[2060] A. N. F. Aleixo and A. B. Balantekin. The close connection between the definition and construction of raised and decreased coherent states with the inverse generators in the algebraic description of quantum confined systems. Journal of Mathematical Physics, 60(5):052102, May 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mavi:2019:RTS

[2061] Rajinder Mavi and Jeffrey Schenker. Resonant tunneling in a system with correlated pure point spectrum. Journal of Mathematical Physics, 60(5):052103, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hawton:2019:PPE

[2062] Margaret Hawton and Vincent Debierre. Photon position eigenvectors, Wigner’s little group, and Berry’s phase. Journal of Mathematical Physics, 60(5):052104, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoffmann:2019:LOC

[2063] Scott E. Hoffmann, Véronique Hussin, Ian Marquette, and Yao-Zhong Zhang. Ladder operators and coherent states for multi-step supersymmetric rational extensions of the truncated oscillator. Journal of Mathematical Physics, 60(5):052105, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Punosevac:2019:DUO

[2064] Predrag Punosevac and Sam L. Robinson. Dynamics of uncertainties for one-dimensional semiclassical wave packets: Isochronicity, scattering, and capture. Journal of Mathematical Physics, 60(5):052106, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tadano:2019:LRS

[2065] Yukihide Tadano. Long-range scattering theory for discrete Schrödinger operators on graphene. Journal of Mathematical Physics, 60(5):052107, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mobus:2019:QZE

[2066] Tim Möbus and Michael M. Wolf. Quantum Zeno effect generalized. Journal of Mathematical Physics, 60(5):052201, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2019:GFS

[2067] Yu Cao, Jianfeng Lu, and Yulong Lu. Gradient flow structure and exponential decay of the sandwiched Rényi divergence for primitive Lindblad equations with GNS-detailed balance. Journal of Mathematical Physics, 60(5):052202, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:NSA

[2068] Juven Wang, Xiao-Gang Wen, and Edward Witten. A new SU(2) anomaly. Journal of Mathematical Physics, 60(5):052301, May 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fullwood:2019:ECR

[2069] James Fullwood and Martin Helmer. On the Euler characteristic of a relative hypersurface. Journal of Mathematical Physics, 60(5):052302, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sachdev:2019:ULT

[2070] Subir Sachdev. Universal low temperature theory of charged black holes with AdS_{2} horizons. Journal of Mathematical Physics, 60(5):052303, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barman:2019:HES

[2071] Subhajit Barman, Golam Mortuza Hossain, and Chiranjeeb Singha. Is Hawking effect short-lived in polymer quantization? Journal of Mathematical Physics, 60(5):052304, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bryden:2019:PMT

[2072] Edward T. Bryden, Marcus A. Khuri, and Benjamin D. Sokolowsky. The positive mass theorem with angular momentum and charge for manifolds with boundary. Journal of Mathematical Physics, 60(5):052501, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon:2019:SDG

[2073] Erick Chacón and Hugo García-Compeán. Self-dual gravity via Hitchin’s equations. Journal of Mathematical Physics, 60(5):052502, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singh:2019:ECS

[2074] Sayuri Singh, Rituparno Goswami, and Sunil D. Maharaj. Existence of conformal symmetries in locally rotationally symmetric spacetimes: Some covariant results. Journal of Mathematical Physics, 60(5):052503, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barker:2019:SEG

[2075] W. E. V. Barker, A. N. Lasenby, M. P. Hobson, and W. J. Handley. Static energetics in gravity. Journal of Mathematical Physics, 60(5): 052504, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jancewicz:2019:PEW

[2076] Bernard Jancewicz. Plane electromagnetic wave in spatially flat Friedman universe. Journal of Mathematical Physics, 60(5):052505, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantica:2019:PFG

[2077] Carlo Alberto Mantica, Luca Guido Molinari, Young Jin Suh, and Sameh Shenawy. Perfect-fluid, generalized Robertson–Walker space-times, and Gray’s decomposition. Journal of Mathematical Physics, 60(5):052506, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Young:2019:CCR

[2078] Lai-Sang Young. Comparing chaotic and random dynamical systems. Journal of Mathematical Physics, 60(5):052701, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mitchell:2019:WCE

[2079] Chad Mitchell. Weak convergence to equilibrium of statistical ensembles in integrable Hamiltonian systems. Journal of Mathematical Physics, 60(5):052702, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geng:2019:ITT

[2080] Jiansheng Geng and Shuaishuai Xue. Invariant tori for two-dimensional nonlinear Schrödinger equations with large forcing terms. Journal of Mathematical Physics, 60(5):052703, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiron:2019:SED

[2081] D. Chiron and B. Marcos. Series expansions of the deflection angle in the scattering problem for power-law potentials. Journal of Mathematical Physics, 60(5):052901, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casas:2019:NFP

[2082] Gabriela A. Casas and Fernando D. Nobre. Nonlinear Fokker–Planck equations in super-diffusive and sub-diffusive regimes. Journal of Mathematical Physics, 60(5):053301, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2019:SIS

[2083] G. Gaeta. W -symmetries of Itô stochastic differential equations. Journal of Mathematical Physics, 60(5):053501, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2019:MCS

[2084] Min Liu and Zhongwei Tang. Multiplicity and concentration of solutions for a fractional Schrödinger equation via Nehari method and pseudo-index theory. Journal of Mathematical Physics, 60(5):053502, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2019:TEF

[2085] Qinglong Zhou. Trace estimation of a family of periodic Sturm–Liouville operators with application to Robe’s restricted three-body problem. Journal of Mathematical Physics, 60(5):053503, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Androulakis:2019:NQD

[2086] George Androulakis and Duncan Wright. On the nonlinearity of quantum dynamical entropy. Journal of Mathematical Physics, 60(5):053504, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zelaya:2019:CQC

[2087] Kevin Zelaya and Oscar Rosas-Ortiz. Comment on “Quasi-coherent states for the Hermite oscillator” [j. math. phys. 59, 062104 (2018)]. Journal of Mathematical Physics, 60(5):054101, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1626].

Nachtergaele:2019:QLB

[2088] Bruno Nachtergaele, Robert Sims, and Amanda Young. Quasi-locality bounds for quantum lattice systems. I. Lieb–Robinson bounds, quasilocal maps, and spectral flow automorphisms. Journal of Mathematical Physics, 60(6):061101, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2019:LMN

[2089] Bingyuan Huang, Jinrui Huang, and Huanyao Wen. Low Mach number limit of the compressible Navier–Stokes–Smoluchowski equations in multi-dimensions. Journal of Mathematical Physics, 60(6):061501, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferreira:2019:GWP

[2090] Lucas C. F. Ferreira and Monisse Postigo. Global well-posedness and asymptotic behavior in Besov–Morrey spaces for chemotaxis-Navier– Stokes fluids. Journal of Mathematical Physics, 60(6):061502, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:GWP

[2091] Yuhui Chen, Wei Luo, and Xiaoping Zhai. Global well-posedness for the phan-thein-Tanner model in critical Besov spaces without damping. Journal of Mathematical Physics, 60(6):061503, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shao:2019:SPS

[2092] Mengqiu Shao and Anmin Mao. Schrödinger–Poisson system with concave–convex nonlinearities. Journal of Mathematical Physics, 60(6): 061504, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:ABS

[2093] Fuzhi Li and Yangrong Li. Asymptotic behavior of stochastic g-Navier– Stokes equations on a sequence of expanding domains. Journal of Mathematical Physics, 60(6):061505, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huynh:2019:LTK

[2094] Nhat Vy Huynh, Phuong Le, and Dinh Phu Nguyen. Liouville theorems for Kirchhoff equations in RN. Journal of Mathematical Physics, 60(6):

061506, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anishetty:2019:ASG

[2095] Ramesh Anishetty and T. P. Sreeraj. Addition of SU(3) generators and its singlet Hilbert space. Journal of Mathematical Physics, 60(6):061701, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lewin:2019:DRG

[2096] Mathieu Lewin, Phan Thành Nam, and Nicolas Rougerie. Derivation of renormalized Gibbs measures from equilibrium many-body quantum Bose gases. Journal of Mathematical Physics, 60(6):061901, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Escobar-Ruiz:2019:FBP

[2097] M. A. Escobar-Ruiz, Willard Miller, Jr., and Alexander V. Turbiner. Four-body problem in d-dimensional space: Ground state, (quasi)-exact-solvability. IV. Journal of Mathematical Physics, 60(6):062101, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Le:2019:PPS

[2098] Dai-Nam Le, Ngoc-Hung Phan, Tuan-Quoc N. Thoi, and Van-Hoang Le. Parabolic, prolate spheroidal bases and relation between bases of the nine-dimensional MICZ–Kepler problem. Journal of Mathematical Physics, 60(6):062102, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gibara:2019:PET

[2099] Ryan Gibara and Richard L. Hall. Potential envelope theory and the local energy theorem. Journal of Mathematical Physics, 60(6):062103, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buyukasik:2019:TES

[2100] Sirin A. Büyükasik and Zehra Çayiç. Time-evolution of squeezed coherent states of a generalized quantum parametric oscillator. Journal of Mathematical Physics, 60(6):062104, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruschi:2019:TEC

[2101] David Edward Bruschi. Time evolution of coupled multimode and multiresonator optomechanical systems. Journal of Mathematical Physics, 60(6):062105, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hofer-Szabo:2019:QMN

[2102] Gábor Hofer-Szabó. Quantum mechanics as a noncommutative representation of classical conditional probabilities. Journal of Mathematical Physics, 60(6):062106, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeWetering:2019:SPS

[2103] John van de Wetering. Sequential product spaces are Jordan algebras. Journal of Mathematical Physics, 60(6):062201, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2019:SMT

[2104] Holger Boche, Minglai Cai, Janis Nötzel, and Christian Deppe. Secret message transmission over quantum channels under adversarial quantum noise: Secrecy capacity and super-activation. Journal of Mathematical Physics, 60(6):062202, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Appleby:2019:SES

[2105] Marcus Appleby and Ingemar Bengtsson. Simplified exact SICS. Journal of Mathematical Physics, 60(6):062203, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:IBS

[2106] Pin Wang and Guanggan Chen. Invariant behavior of stochastic atmosphere-ocean model with degenerate noise. Journal of Mathematical Physics, 60(6):062701, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chatar:2019:HAV

[2107] W. Chatar, M. Benkhali, I. El Fakkousy, J. Kharbach, A. Rezzouk, and M. Ouazzani-Jamil. The hydrogen atom in the van der Waals potential combined by magnetic and electric fields, Painlevé analysis, and integrability. Journal of Mathematical Physics, 60(6):062702, June 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Gallavotti:2019:QHM

[2108] Giovanni Gallavotti. Quasiperiodic Hamiltonian motions, scale invariance, harmonic oscillators. Journal of Mathematical Physics, 60(6): 062901, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majic:2019:ITS

[2109] Matt Majić. Image theory for a sphere with negative permittivity. Journal of Mathematical Physics, 60(6):062902, June 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guliyev:2019:SOD

[2110] Namig J. Guliyev. Schrödinger operators with distributional potentials and boundary conditions dependent on the eigenvalue parameter. Journal of Mathematical Physics, 60(6):063501, June 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grammaticos:2019:RDP

[2111] B. Grammaticos, A. Ramani, and R. Willox. Restoring discrete Painlevé equations from an E8(1)-associated one. Journal of Mathematical Physics, 60(6):063502, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besnard:2019:DSS

[2112] Fabien Besnard and Nadir Bizi. Doppler shift in semi-Riemannian signature and the non-uniqueness of the Krein space of spinors. Journal of Mathematical Physics, 60(6):063503, June 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jian:2019:ALO

[2113] Wenwen Jian, Yunfeng Shi, and Xiaoping Yuan. Anderson localization for one-frequency quasi-periodic block operators with long-range interactions. Journal of Mathematical Physics, 60(6):063504, June 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pava:2019:OSS

[2114] Jaime Angulo Pava, César A. Hernández Melo, and Ramón G. Plaza. Orbital stability of standing waves for the nonlinear Schrödinger equation with attractive delta potential and double power repulsive nonlinearity. Journal of Mathematical Physics, 60(7):071501, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2019:PFA

[2115] Hanchun Yang and Yu Zhang. Pressure and flux-approximation to the isentropic relativistic Euler equations for modified Chaplygin gas. Journal of Mathematical Physics, 60(7):071502, July 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amour:2019:IDS

[2116] L. Amour, L. Jager, and J. Nourrigat. Infinite dimensional semiclassical analysis and applications to a model in nuclear magnetic resonance. Journal of Mathematical Physics, 60(7):071503, July 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarrico:2019:NCL

[2117] C. O. R. Sarrico. A nonlinear conservation law with a step function involving the flux: A distributional approach. Journal of Mathematical Physics, 60(7):071504, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Remizov:2019:SGF

[2118] Ivan D. Remizov. Solution-giving formula to Cauchy problem for multidimensional parabolic equation with variable coefficients. Journal of Mathematical Physics, 60(7):071505, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frank:2019:NSE

[2119] Rupert L. Frank. Non-spherical equilibrium shapes in the liquid drop model. Journal of Mathematical Physics, 60(7):071506, July 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:2019:EBM

[2120] Byung-Hoon Hwang and Seok-Bae Yun. Ellipsoidal BGK model near a global Maxwellian in the whole space. Journal of Mathematical Physics, 60(7):071507, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2019:BPN

[2121] Dinh-Thi Nguyen. Blow-up profile of neutron stars in the Chandrasekhar theory. Journal of Mathematical Physics, 60(7):071508, July 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Ye:2019:LTB

[2122] Yulin Ye and Zilai Li. The large time behavior of the free boundary for one dimensional compressible Navier–Stokes equations. Journal of Mathematical Physics, 60(7):071509, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benalili:2019:ESF

[2123] Mohammed Benalili and Ali Zouaoui. On the Einstein-scalar field Lichnerowicz equations on compact Riemannian manifolds. Journal of Mathematical Physics, 60(7):071510, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:GWS

[2124] Yang Li and Yongzhong Sun. Global weak solutions and long time behavior for 1D compressible MHD equations without resistivity. Journal of Mathematical Physics, 60(7):071511, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

De:2019:KAS

[2125] Sandipan De. From a Kac algebra subfactor to Drinfeld double. Journal of Mathematical Physics, 60(7):071701, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kac:2019:PTF

[2126] Victor Kac and Johan van de Leur. Polynomial tau-functions of BKP and DKP hierarchies. Journal of Mathematical Physics, 60(7):071702, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adem:2019:MSC

[2127] Alejandro Adem and Man Chuen Cheng. On the moduli spaces of commuting elements in the projective unitary groups. Journal of Mathematical Physics, 60(7):071703, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Terwilliger:2019:APB

[2128] Paul Terwilliger. The alternating PBW basis for the positive part of U_{q}(ˆ∫↕_{∈}). Journal of Mathematical Physics, 60(7):071704, July 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Ito:2019:DFA

[2129] Masahiko Ito and Masatoshi Noumi. A determinant formula associated with the elliptic hypergeometric integrals of type BC_{n}. Journal of Mathematical Physics, 60(7):071705, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khoroshkin:2019:FLC

[2130] S. M. Khoroshkin and M. G. Matushko. Fermionic limit of the Calogero– Sutherland system. Journal of Mathematical Physics, 60(7):071706, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:NWM

[2131] Qiufan Chen and Jianzhi Han. Non-weight modules over the affine-Virasoro algebra of type a_{1}. Journal of Mathematical Physics, 60(7): 071707, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ouyang:2019:CSB

[2132] Yingkai Ouyang. Computing spectral bounds of the Heisenberg ferromagnet from geometric considerations. Journal of Mathematical Physics, 60(7):071901, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rademacher:2019:CLT

[2133] Simone Rademacher and Benjamin Schlein. Central limit theorem for Bose–Einstein condensates. Journal of Mathematical Physics, 60(7): 071902, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fournais:2019:GSE

[2134] Søren Fournais, Marcin Napiórkowski, Robin Reuvers, and Jan Philip Solovej. Ground state energy of a dilute two-dimensional Bose gas from the Bogoliubov free energy functional. Journal of Mathematical Physics, 60(7):071903, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Viesca:2019:CNH

[2135] Edgar Lozano Viesca, Jonas Schober, and Hermann Schulz-Baldes. Chern numbers as half-signature of the spectral localizer. Journal of Mathematical Physics, 60(7):072101, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elgressy:2019:GQR

[2136] Gil Elgressy and Lawrence Horwitz. Geometry of quantum Riemannian Hamiltonian evolution. Journal of Mathematical Physics, 60(7):072102, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bauml:2019:EES

[2137] S. Bäuml, A. Winter, and D. Yang. Every entangled state provides an advantage in classical communication. Journal of Mathematical Physics, 60(7):072201, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Swingle:2019:RMF

[2138] Brian G. Swingle and Yixu Wang. Recovery map for fermionic Gaussian channels. Journal of Mathematical Physics, 60(7):072202, July 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deckert:2019:MTD

[2139] Dirk-André Deckert and Lukas Nickel. Multi-time dynamics of the Dirac– Fock–Podolsky model of QED. Journal of Mathematical Physics, 60(7): 072301, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonzom:2019:DQI

[2140] Valentin Bonzom, Victor Nador, and Adrian Tanasa. Diagrammatics of the quartic O(N)^{3}-invariant Sachdev–Ye–Kitaev-like tensor model. Journal of Mathematical Physics, 60(7):072302, July 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paetz:2019:CCG

[2141] Tim-Torben Paetz. On the choice of a conformal Gauss gauge near the cylinder representing spatial infinity. Journal of Mathematical Physics, 60(7):072501, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lohe:2019:SMR

[2142] M. A. Lohe. Systems of matrix Riccati equations, linear fractional transformations, partial integrability and synchronization. Journal of Mathematical Physics, 60(7):072701, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nobe:2019:GRC

[2143] Atsushi Nobe. Generators of rank 2 cluster algebras of affine types via linearization of seed mutations. Journal of Mathematical Physics, 60(7): 072702, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fagnola:2019:RAD

[2144] Franco Fagnola, Emanuela Sasso, and Veronica Umanità. The role of the atomic decoherence-free subalgebra in the study of quantum Markov semigroups. Journal of Mathematical Physics, 60(7):072703, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:RDF

[2145] Dingshi Li, Bixiang Wang, and Xiaohu Wang. Random dynamics of fractional stochastic reaction-diffusion equations on Rn without uniqueness. Journal of Mathematical Physics, 60(7):072704, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2019:DCS

[2146] Jiu-Gang Dong, Seung-Yeal Ha, and Doheon Kim. From discrete Cucker– Smale model to continuous Cucker–Smale model in a temperature field. Journal of Mathematical Physics, 60(7):072705, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueiredo:2019:DPF

[2147] A. D. Figueiredo, T. M. da Rocha Filho, and M. A. Amato. Distribution probability of force for a physical system of n random particles. Journal of Mathematical Physics, 60(7):073301, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doyon:2019:GHC

[2148] Benjamin Doyon. Generalized hydrodynamics of the classical Toda system. Journal of Mathematical Physics, 60(7):073302, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlowski:2019:LDL

[2149] Karol K. Kozlowski. Long-distance and large-time asymptotic behaviour of dynamic correlation functions in the massless regime of the XXZ spin-1/2 chain. Journal of Mathematical Physics, 60(7):073303, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ye:2019:MGT

[2150] Zichun Ye. Models of gradient type with sub-quadratic actions. Journal of Mathematical Physics, 60(7):073304, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:VVW

[2151] Yun-Zhang Li and Jing Zhao. Vector-valued weak Gabor dual frames on discrete periodic sets. Journal of Mathematical Physics, 60(7):073501, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pascalie:2019:LLS

[2152] R. Pascalie, C. I. Pérez-Sánchez, A. Tanasa, and R. Wulkenhaar. On the large N limit of Schwinger–Dyson equations of a rank-3 tensor field theory. Journal of Mathematical Physics, 60(7):073502, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigin:2019:IOA

[2153] Misha Feigin and Martin Vrabec. Intertwining operator for AG_{2} Calogero–Moser–Sutherland system. Journal of Mathematical Physics, 60(7):073503, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2019:SWT

[2154] Inkang Kim. Seiberg–Witten theory with SpinH structure. Journal of Mathematical Physics, 60(7):073504, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2019:HOD

[2155] Axel Schulze-Halberg. Higher-order Darboux transformations for the massless Dirac equation at zero energy. Journal of Mathematical Physics, 60(7):073505, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Raymond:2019:AEE

[2156] Nicolas Raymond and Julien Royer. Absence of embedded eigenvalues for translationally invariant magnetic Laplacians. Journal of Mathematical Physics, 60(7):073506, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigin:2019:TTD

[2157] B. Feigin, M. Jimbo, and E. Mukhin. Towards trigonometric deformation of ˆ∫↕_{∈} coset VOA. Journal of Mathematical Physics, 60(7):073507, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakabekov:2019:ABV

[2158] A. Sakabekov and Y. Auzhani. Approximation of the boundary value problem for 3-dimensional Boltzmann’s equation with Maxwell’s microscopic conditions. Journal of Mathematical Physics, 60(7):073508, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2019:DAS

[2159] Lei Cao and Yisong Yang. Determination of anti-de Sitter monopole wall via minimization. Journal of Mathematical Physics, 60(7):073509, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biondini:2019:IST

[2160] Gino Biondini, Ildar Gabitov, Gregor Kovacic, and Sitai Li. Inverse scattering transform for two-level systems with nonzero background. Journal of Mathematical Physics, 60(7):073510, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Angelelli:2019:CRS

[2161] Mario Angelelli. Complexity reduction for sign configurations through the KP II equation and its information-theoretic aspects. Journal of Mathematical Physics, 60(7):073511, July 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dePrunele:2019:TDH

[2162] E. de Prunelé. Two-dimensional Helmholtz equation with zero Dirichlet boundary condition on a circle: Analytic results for boundary deformation, the transition disk-lens. Journal of Mathematical Physics, 60(8): 081501, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hakkaev:2019:SSP

[2163] Sevdzhan Hakkaev. Stability of semitrivial periodic waves of a Schrödinger system. Journal of Mathematical Physics, 60(8):081502, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aouadi:2019:PGE

[2164] Moncef Aouadi and Alberto Castejón. Properties of global and exponential attractors for nonlinear thermo-diffusion Timoshenko system. Journal of Mathematical Physics, 60(8):081503, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silva:2019:QSE

[2165] Edcarlos D. Silva and J. S. Silva. Quasilinear Schrödinger equations with nonlinearities interacting with high eigenvalues. Journal of Mathematical Physics, 60(8):081504, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dyatlov:2019:IFU

[2166] Semyon Dyatlov. An introduction to fractal uncertainty principle. Journal of Mathematical Physics, 60(8):081505, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bal:2019:CBI

[2167] Guillaume Bal. Continuous bulk and interface description of topological insulators. Journal of Mathematical Physics, 60(8):081506, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2019:RPI

[2168] Hongjun Cheng. Riemann problem for the isentropic Chaplygin gas Cargo-LeRoux model. Journal of Mathematical Physics, 60(8):081507, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cen:2019:INH

[2169] Julia Cen, Francisco Correa, and Andreas Fring. Integrable nonlocal Hirota equations. Journal of Mathematical Physics, 60(8):081508, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2019:IBV

[2170] A. Alexandrou Himonas, Dionyssios Mantzavinos, and Fangchi Yan. Initial-boundary value problems for a reaction-diffusion equation. Journal of Mathematical Physics, 60(8):081509, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2019:GWP

[2171] Xiaofeng Liu. Global well-posedness to some modified 2-D nonhomogeneous Navier–Stokes equations. Journal of Mathematical Physics, 60 (8):081510, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Che:2019:EMP

[2172] Guofeng Che, Haibo Chen, and Tsung fang Wu. Existence and multiplicity of positive solutions for fractional Laplacian systems with nonlinear coupling. Journal of Mathematical Physics, 60(8):081511, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2019:HNE

[2173] Jie Zhao and Juan Wang. Homogenization of nonlinear equations with mixed boundary conditions. Journal of Mathematical Physics, 60(8): 081512, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cho:2019:WPS

[2174] Yonggeun Cho. Well-posedness and scattering of inhomogeneous cubicquintic NLS. Journal of Mathematical Physics, 60(8):081513, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alharbi:2019:STG

[2175] Majed Ghazi Alharbi and Tarek Saanouni. Sharp threshold of global wellposedness vs finite time blow-up for a class of inhomogeneous Choquard equations. Journal of Mathematical Physics, 60(8):081514, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:GED

[2176] Jingjing Zhang and Ting Zhang. Global existence of discretely self-similar solutions to the generalized MHD system in Besov space. Journal of Mathematical Physics, 60(8):081515, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adamovic:2019:FRI

[2177] Drazen Adamović and Veronika Pedić. On fusion rules and intertwining operators for the Weyl vertex algebra. Journal of Mathematical Physics, 60(8):081701, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luu:2019:SCD

[2178] Martin T. Luu. Spectral curve duality beyond the two-matrix model. Journal of Mathematical Physics, 60(8):081702, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baseilhac:2019:HRC

[2179] Pascal Baseilhac, Nicolas Crampé, and Rodrigo A. Pimenta. Higher rank classical analogs of the Askey–Wilson algebra from the sl_{N} Onsager algebra. Journal of Mathematical Physics, 60(8):081703, August 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohamed:2019:PLA

[2180] Mohamed Belhaj Mohamed. On the pre-Lie algebra of specified Feynman graphs. Journal of Mathematical Physics, 60(8):081704, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neergaard:2019:PCO

[2181] K. Neergård. Proof by characters of the orthogonal-orthogonal duality and relations of Casimir invariants. Journal of Mathematical Physics, 60 (8):081705, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nishiyama:2019:MFT

[2182] Seiya Nishiyama and João da Providência. Mean-field theory based on the jacobi hsp semidirect sum ⟨_{N} [rtimes]∫√(∈N , R)_{C} algebra of boson

operators. Journal of Mathematical Physics, 60(8):081706, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

CorreadaSilva:2019:PKS

[2183] R. Correa da Silva. Perturbations of KMS states and noncommutative L_{p}-spaces. Journal of Mathematical Physics, 60(8):081707, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cenatiempo:2019:BTD

[2184] Serena Cenatiempo. Bogoliubov theory for dilute Bose gases: The Gross– Pitaevskii regime. Journal of Mathematical Physics, 60(8):081901, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gontier:2019:RHF

[2185] David Gontier and Salma Lahbabi. The reduced Hartree–Fock model with self-generated magnetic fields. Journal of Mathematical Physics, 60 (8):081902, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loring:2019:BSK

[2186] Terry A. Loring. Bulk spectrum and K-theory for infinite-area topological quasicrystals. Journal of Mathematical Physics, 60(8):081903, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rangamani:2019:SUR

[2187] Nishant Rangamani. Singular-unbounded random Jacobi matrices. Journal of Mathematical Physics, 60(8):081904, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Viver:2019:GCM

[2188] M. Pons Viver and A. Puente. Generation of Coulomb matrix elements for the 2D quantum harmonic oscillator. Journal of Mathematical Physics, 60(8):081905, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anzaldo-Meneses:2019:SEM

[2189] A. Anzaldo-Meneses. Superlattices with entangled modes and the Hopf bundle. Journal of Mathematical Physics, 60(8):081906, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cuesta:2019:ASA

[2190] Javier Cuesta and Michael M. Wolf. Are almost-symmetries almost linear? Journal of Mathematical Physics, 60(8):082101, August 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2019:MEU

[2191] Jinsong Wu. Minimizers of the entropic uncertainty principle for locally compact quantum groups. Journal of Mathematical Physics, 60 (8):082102, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamamura:2019:RBS

[2192] Ikko Hamamura and Takayuki Miyadera. Relation between the state distinction power and disturbance in quantum measurements. Journal of Mathematical Physics, 60(8):082103, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller:2019:NRQ

[2193] Olaf Müller. On the von Neumann rule in quantization. Journal of Mathematical Physics, 60(8):082104, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golovaty:2019:SOC

[2194] Yuriy Golovaty. 1D Schrödinger operators with Coulomb-like potentials. Journal of Mathematical Physics, 60(8):082105, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mahdifar:2019:FPI

[2195] A. Mahdifar and E. Amooghorban. Free particle and isotropic harmonic oscillator on a spheroidal surface: The Higgs-like approach. Journal of Mathematical Physics, 60(8):082106, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2019:DCQ

[2196] Siendong Huang. Dense coding and quantum teleportation using type II_{1} von Neumann algebras. Journal of Mathematical Physics, 60(8):082107, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castro-Alvaredo:2019:ECQ

[2197] Olalla A. Castro-Alvaredo, Cecilia De Fazio, Benjamin Doyon, and István M. Szécsényi. Entanglement content of quantum particle excitations. III. Graph partition functions. Journal of Mathematical Physics, 60(8):082301, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deMontigny:2019:DKP

[2198] M. de Montigny and E. S. Santos. On the Duffin–Kemmer–Petiau equation in arbitrary dimensions. Journal of Mathematical Physics, 60(8): 082302, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aksteiner:2019:SLG

[2199] Steffen Aksteiner and Thomas Bäckdahl. Symmetries of linearized gravity from adjoint operators. Journal of Mathematical Physics, 60(8): 082501, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2019:PMZ

[2200] Francesco Calogero and Farrin Payandeh. Polynomials with multiple zeros and solvable dynamical systems including models in the plane with polynomial interactions. Journal of Mathematical Physics, 60(8):082701, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2019:LBN

[2201] Lin Shi and Xiliang Li. Limiting behavior of non-autonomous stochastic reaction-diffusion equations on unbounded thin domains. Journal of Mathematical Physics, 60(8):082702, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duston:2019:TSW

[2202] Christopher Levi Duston. Torus solutions to the Weierstrass–Enneper representation of surfaces. Journal of Mathematical Physics, 60(8): 082703, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fortunati:2019:LTA

[2203] Alessandro Fortunati and Stephen Wiggins. A Lie transform approach to the construction of Lyapunov functions in autonomous and nonautonomous systems. Journal of Mathematical Physics, 60(8):082704, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jeffrey:2019:IPF

[2204] Lisa C. Jeffrey and Jia Ji. Intersection pairings in the N-fold reduced product of adjoint orbits. Journal of Mathematical Physics, 60(8):082901, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krishnaswami:2019:ITA

[2205] Govind S. Krishnaswami and T. R. Vishnu. Invariant tori, action-angle variables, and phase space structure of the Rajeev–Ranken model. Journal of Mathematical Physics, 60(8):082902, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diez:2019:CLV

[2206] Tobias Diez and Gerd Rudolph. Clebsch–Lagrange variational principle and geometric constraint analysis of relativistic field theories. Journal of Mathematical Physics, 60(8):082903, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Croydon:2019:IMB

[2207] David A. Croydon and Makiko Sasada. Invariant measures for the boxball system based on stationary Markov chains and periodic Gibbs measures. Journal of Mathematical Physics, 60(8):083301, August 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2019:DFE

[2208] Yi Jiang, Ilya Dumer, Alexey A. Kovalev, and Leonid P. Pryadko. Duality and free energy analyticity bounds for few-body Ising models with extensive homology rank. Journal of Mathematical Physics, 60(8):083302, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hislop:2019:TQP

[2209] Peter D. Hislop, Kay Kirkpatrick, Stefano Olla, and Jeffrey Schenker. Transport of a quantum particle in a time-dependent white-noise potential. Journal of Mathematical Physics, 60(8):083303, August 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itoi:2019:NRS

[2210] C. Itoi and Y. Utsunomiya. No replica symmetry breaking phase in the random field Ginzburg–Landau model. Journal of Mathematical Physics, 60(8):083304, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jagannath:2019:DMF

[2211] Aukosh Jagannath. Dynamics of mean field spin glasses on short and long timescales. Journal of Mathematical Physics, 60(8):083305, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dehghani:2019:EDC

[2212] A. Dehghani, B. Mojaveri, and A. A. Alenabi. Excitation and depression of coherent state of the simple harmonic oscillator. Journal of Mathematical Physics, 60(8):083501, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhoi:2019:SJF

[2213] J. Bhoi, A. K. Behera, and U. Laha. Off-shell Jost function for the Hulthén potential in all partial waves. Journal of Mathematical Physics, 60(8):083502, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2019:IIS

[2214] Yu Liu, Guoliang Shi, and Jun Yan. Incomplete inverse spectral problems for Dirac–Bessel operators. Journal of Mathematical Physics, 60 (8):083503, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deJong:2019:NEP

[2215] J. de Jong and R. Wulkenhaar. Nonperturbative evaluation of the partition function for the real scalar quartic QFT on the Moyal plane at weak coupling. Journal of Mathematical Physics, 60(8):083504, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juhasz:2019:CPT

[2216] David Juhasz and Per Kristen Jakobsen. Constructing a partially transparent computational boundary for UPPE using leaky modes. Journal of Mathematical Physics, 60(8):083505, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diki:2019:BFF

[2217] Kamal Diki, Rolf Sören Krausshar, and Irene Sabadini. On the Bargmann–Fock–Fueter and Bergman–Fueter integral transforms. Journal of Mathematical Physics, 60(8):083506, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leopoldino:2019:DEA

[2218] Andressa P. C. Leopoldino, José Weberszpil, Cresus F. L. Godinho, and José Abdalla Helayël-Neto. Discussing the extension and applications of a variational approach with deformed derivatives. Journal of Mathematical Physics, 60(8):083507, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celeghini:2019:ZFR

[2219] E. Celeghini, M. Gadella, and M. A. del Olmo. Zernike functions, rigged Hilbert spaces, and potential applications. Journal of Mathematical Physics, 60(8):083508, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bailey:2019:MMC

[2220] E. C. Bailey, S. Bettin, G. Blower, J. B. Conrey, A. Prokhorov, M. O. Rubinstein, and N. C. Snaith. Mixed moments of characteristic polynomials of random unitary matrices. Journal of Mathematical Physics, 60 (8):083509, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gamayun:2019:SSS

[2221] O. Gamayun and O. Lisovyy. On self-similar solutions of the vortex filament equation. Journal of Mathematical Physics, 60(8):083510, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yan:2019:IBV

[2222] Zhenya Yan. Initial-boundary value problem for the spin-1 Gross– Pitaevskii system with a 4 × 4 Lax pair on a finite interval. Journal of Mathematical Physics, 60(8):083511, August 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cresson:2019:SSL

[2223] Jacky Cresson, Yasmina Kheloufi, Khadra Nachi, and Frederic Pierret. Selection of a stochastic Landau–Lifshitz equation and the stochastic persistence problem. Journal of Mathematical Physics, 60(8):083512, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lv:2019:WPB

[2224] Wujun Lv, Ping He, and Qinghua Wang. Well-posedness and blow-up solution for the stochastic Dullin–Gottwald–Holm equation. Journal of Mathematical Physics, 60(8):083513, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gustavsson:2019:AMB

[2225] Andreas Gustavsson. Abelian M5-brane on Sq6. Journal of Mathematical Physics, 60(8):083514, August 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zulfiqar:2019:SMN

[2226] Hina Zulfiqar, Shenglan Yuan, Ziying He, and Jinqiao Duan. Slow manifolds for a nonlocal fast-slow stochastic system with stable Lévy noise. Journal of Mathematical Physics, 60(9):091501, September 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:IPT

[2227] Zhiyuan Li, Xing Cheng, and Gongsheng Li. An inverse problem in timefractional diffusion equations with nonlinear boundary condition. Journal of Mathematical Physics, 60(9):091502, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koelink:2019:CID

[2228] Erik Koelink, Jean-Michel Lemay, and Luc Vinet. Convolution identities for Dunkl orthogonal polynomials from the osp(1—2) Lie superalgebra. Journal of Mathematical Physics, 60(9):091701, September 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mozrzymas:2019:BOD

[2229] Marek Mozrzymas, Dariusz Chruściński, and Gniewomir Sarbicki. On bipartite operators defined by sets of completely different permutations. Journal of Mathematical Physics, 60(9):091702, September 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nies:2019:MLR

[2230] André Nies and Volkher B. Scholz. Martin-Löf random quantum states. Journal of Mathematical Physics, 60(9):092201, September 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pitelli:2019:PPB

[2231] Joao Paulo M. Pitelli and Vitor S. Barroso. Particle production between isometric frames on a Poincaré patch of AdS_{2}. Journal of Mathematical Physics, 60(9):092301, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hara:2019:HEM

[2232] Kentaro Hara, Akifumi Sako, and Hyun Seok Yang. Hermitian–Einstein metrics from noncommutative u (1) instantons. Journal of Mathematical Physics, 60(9):092501, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casimiro:2019:DFD

[2233] Ana Casimiro and César Rodrigo. Discrete formulation for the dynamics of rods deforming in space. Journal of Mathematical Physics, 60(9): 092901, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendes:2019:NCP

[2234] R. Vilela Mendes. Non-commutative probability and non-commutative processes: Beyond the Heisenberg algebra. Journal of Mathematical Physics, 60(9):093501, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuchment:2019:ASD

[2235] Peter Kuchment and Jia Zhao. Analyticity of the spectrum and Dirichletto-Neumann operator technique for quantum graphs. Journal of Mathematical Physics, 60(9):093502, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carstea:2019:SQM

[2236] A. S. Carstea and T. Takenawa. Super-QRT and 4D-mappings reduced from the lattice super-KdV equation. Journal of Mathematical Physics, 60(9):093503, September 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rhouma:2019:LTR

[2237] N. Belhaj Rhouma and M. Seddik. Liouville type results for a class of quasilinear parabolic problems. Journal of Mathematical Physics, 60 (10):101501, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2019:PGS

[2238] Zhenyu Guo and Wenming Zou. Positive ground state solutions for an elliptic system with Hardy–Sobolev critical exponent growth. Journal of Mathematical Physics, 60(10):101502, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alimohammady:2019:EMP

[2239] Mohsen Alimohammady, Claudianor O. Alves, and Hassan Kaffash Amiri. Existence and multiplicity of positive solutions for a class of Kirchhoff Laplacian type problems. Journal of Mathematical Physics, 60 (10):101503, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sheng:2019:VAE

[2240] Shouqiong Sheng and Zhiqiang Shao. The vanishing adiabatic exponent limits of Riemann solutions to the isentropic Euler equations for power law with a Coulomb-like friction term. Journal of Mathematical Physics, 60(10):101504, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atakishiyev:2019:SGZ

[2241] Natig M. Atakishiyev, George S. Pogosyan, Kurt Bernardo Wolf, and Alexander Yakhno. Spherical geometry, Zernike’s separability, and interbasis expansion coefficients. Journal of Mathematical Physics, 60(10): 101701, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malkoun:2019:RSA

[2242] Joseph Malkoun. Root systems and the Atiyah–Sutcliffe problem. Journal of Mathematical Physics, 60(10):101702, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naber:2019:TDH

[2243] M. G. Naber. Two-dimensional hydrogen-like atom in a constant magnetic field. Journal of Mathematical Physics, 60(10):102101, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kazakov:2019:NES

[2244] K. V. Kazakov. A novel examination of the similarity between a perturbed hydrogen atom and an anharmonic oscillator. Journal of Mathematical Physics, 60(10):102102, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nicola:2019:AFP

[2245] Fabio Nicola and S. Ivan Trapasso. Approximation of Feynman path integrals with non-smooth potentials. Journal of Mathematical Physics, 60 (10):102103, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vieira:2019:MDS

[2246] S. L. R. Vieira and K. Bakke. Maxwell duality and semiclassical analysis of the interaction of the magnetic quadrupole moment of a neutral particle with external fields. Journal of Mathematical Physics, 60(10):102104, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vieira:2019:SER

[2247] H. S. Vieira, V. B. Bezerra, C. R. Muniz, and M. S. Cunha. Some exact results on quantum Newtonian cosmology. Journal of Mathematical Physics, 60(10):102301, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kocic:2019:APE

[2248] Mikica Kocic, Marcus Högås, Francesco Torsello, and Edvard Mörtsell. Algebraic properties of Einstein solutions in ghost-free bimetric theory. Journal of Mathematical Physics, 60(10):102501, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fox:2019:MHS

[2249] Matthew S. Fox. Multipole hair of Schwarzschild–Tangherlini black holes. Journal of Mathematical Physics, 60(10):102502, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See retraction [2326].

Betermin:2019:MLS

[2250] Laurent Bétermin. Minimizing lattice structures for Morse potential energy in two and three dimensions. Journal of Mathematical Physics, 60 (10):102901, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giuliani:2019:NID

[2251] Alessandro Giuliani and Fabio Lucio Toninelli. Non-integrable dimer models: Universality and scaling relations. Journal of Mathematical Physics, 60(10):103301, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tertychniy:2019:SSS

[2252] Sergey I. Tertychniy. Symmetries of the space of solutions to special double confluent Heun equations of integer order. Journal of Mathematical Physics, 60(10):103501, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hovhannisyan:2019:SSS

[2253] Gro Hovhannisyan. 3 soliton solution to sine–Gordon equation on a space scale. Journal of Mathematical Physics, 60(10):103502, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bihun:2019:TDP

[2254] Oksana Bihun. Time-dependent polynomials with one multiple root and new solvable dynamical systems. Journal of Mathematical Physics, 60 (10):103503, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:MFL

[2255] Lei Li, Jian-Guo Liu, and Pu Yu. On the mean field limit for Brownian particles with Coulomb interaction in 3D. Journal of Mathematical Physics, 60(11):111501, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2019:BCC

[2256] Yimin Sun. A blow-up criterion for classical solutions to the Prandtl equations. Journal of Mathematical Physics, 60(11):111502, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bechtluft-Sachs:2019:GFB

[2257] Stefan Bechtluft-Sachs and Evangelia Samiou. Green’s functions, Biot– Savart operators, and linking numbers on negatively curved symmetric spaces. Journal of Mathematical Physics, 60(11):111503, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoshino:2019:ABS

[2258] Gaku Hoshino. Asymptotic behavior for solutions to the dissipative nonlinear Schrödinger equations with the fractional Sobolev space. Journal of Mathematical Physics, 60(11):111504, November 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Truong:2019:NMC

[2259] Le Xuan Truong. The Nehari manifold for a class of Schrödinger equation involving fractional p-Laplacian and sign-changing logarithmic nonlinearity. Journal of Mathematical Physics, 60(11):111505, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bellassoued:2019:RCL

[2260] Mourad Bellassoued and Yosra Boughanja. Recovery of coefficients in the linear Boltzmann equation. Journal of Mathematical Physics, 60(11): 111506, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Asakura:2019:SIG

[2261] Fumioki Asakura and Andrea Corli. A system of ionized gas dynamics. Journal of Mathematical Physics, 60(11):111507, November 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Serov:2019:IMP

[2262] V. Serov and T. Tyni. Inverse medium problem for a singular contrast. Journal of Mathematical Physics, 60(11):111508, November 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Niu:2019:CRH

[2263] Weisheng Niu and Yue Yuan. Convergence rate in homogenization of elliptic systems with singular perturbations. Journal of Mathematical Physics, 60(11):111509, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:NUC

[2264] Jinlu Li, Li Dai, and Weipeng Zhu. Non-uniform continuous dependence on initial data of solutions to the Euler–Poincaré system. Journal of Mathematical Physics, 60(11):111510, November 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2019:EQM

[2265] Haisheng Li and Qiang Mu. Equivariant quasi modules at infinity for vertex Γ-algebras. Journal of Mathematical Physics, 60(11):111701, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazorchuk:2019:CTM

[2266] Volodymyr Mazorchuk and Christoffer Söderberg. Category O for Takiff sl_{2}. Journal of Mathematical Physics, 60(11):111702, November 2019.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2019:SBR

[2267] Jeong-Ah Kim and Dong-Uy Shin. Strip bundle realization of the crystals over U_{q}(G_{2}^{(1)}). Journal of Mathematical Physics, 60(11):111703, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choreno:2019:BPT

[2268] E. Choreño, D. Ojeda-Guillén, R. Valencia, and V. D. Granados. Berry phase of the Tavis–Cummings model with three modes of oscillation. Journal of Mathematical Physics, 60(11):111704, November 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amorim:2019:CMS

[2269] Érik Amorim. On a center-of-mass system of coordinates for symmetric classical and quantum many-body problems. Journal of Mathematical Physics, 60(11):111901, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DePalma:2019:SEN

[2270] Giacomo De Palma. The squashed entanglement of the noiseless quantum Gaussian attenuator and amplifier. Journal of Mathematical Physics, 60(11):112201, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Somma:2019:UCS

[2271] Rolando D. Somma. Unitary circuit synthesis for tomography of generalized coherent states. Journal of Mathematical Physics, 60(11):112202, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fukuda:2019:TEG

[2272] Motohisa Fukuda and Robert Koenig. Typical entanglement for Gaussian states. Journal of Mathematical Physics, 60(11):112203, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yue:2019:ZEE

[2273] Qiuling Yue and Eric Chitambar. The zero-error entanglement cost is highly non-additive. Journal of Mathematical Physics, 60(11):112204, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pereniguez:2019:PBN

[2274] David Pereñiguez. p-brane Newton–Cartan geometry. Journal of Mathematical Physics, 60(11):112501, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maraner:2019:JMG

[2275] Paolo Maraner. On the Jacobi metric for a general Lagrangian system. Journal of Mathematical Physics, 60(11):112901, November 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeMartino:2019:ENS

[2276] Antonino De Martino and Arianna Passerini. Existence and nonlinear stability of convective solutions for almost compressible fluids in Bénard problem. Journal of Mathematical Physics, 60(11):113101, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonzalez-Navarrete:2019:DOO

[2277] Manuel González-Navarrete and Rodrigo Lambert. The diffusion of opposite opinions in a randomly biased environment. Journal of Mathematical Physics, 60(11):113301, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:OPA

[2278] Yang Chen, Galina Filipuk, and Longjun Zhan. Orthogonal polynomials, asymptotics, and Heun equations. Journal of Mathematical Physics, 60(11):113501, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gresnigt:2019:CAO

[2279] N. Gresnigt. A combing algorithm for orientable braided 3-belts. Journal of Mathematical Physics, 60(11):113502, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciaglia:2019:EEC

[2280] F. M. Ciaglia, F. Di Cosmo, G. Marmo, and L. Schiavone. Evolutionary equations and constraints: Maxwell equations. Journal of Mathematical Physics, 60(11):113503, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barth:2019:VLT

[2281] Simon Barth and Andreas Bitter. On the virtual level of two-body interactions and applications to three-body systems in higher dimensions. Journal of Mathematical Physics, 60(11):113504, November 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Desiraju:2019:FAS

[2282] Harini Desiraju. The τ-function of the Ablowitz–Segur family of solutions to Painlevé II as a Widom constant. Journal of Mathematical Physics, 60(11):113505, November 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:WNO

[2283] Caishi Wang, Yuling Tang, and Suling Ren. Weighted number operators on Bernoulli functionals and quantum exclusion semigroups. Journal of Mathematical Physics, 60(11):113506, November 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2019:BTD

[2284] Yanfang Gao and Zhiyong Wang. Blow-up for trapped dipolar quantum gases with large energy. Journal of Mathematical Physics, 60(12):121501, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2019:GEW

[2285] Boqiang Lü, Rong Zhang, and Xin Zhong. Global existence of weak solutions to the compressible quantum Navier–Stokes equations with degenerate viscosity. Journal of Mathematical Physics, 60(12):121502, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2019:RPS

[2286] Yujin Liu and Wancheng Sheng. The Riemann problem of selfsimilar ZND model for the pressure gradient system in gas dynamical combustion. Journal of Mathematical Physics, 60(12):121503, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taujanskas:2019:LDD

[2287] Grigalius Taujanskas. Large data decay of Yang–Mills–Higgs fields on Minkowski and de Sitter spacetimes. Journal of Mathematical Physics, 60(12):121504, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2019:IBV

[2288] Der-Chen Chang, Tao Luo, and Hua Zhong. On an initial boundary value problem for gas dynamics in thermal nonequilibrium. Journal of Mathematical Physics, 60(12):121505, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naumkin:2019:DCA

[2289] Pavel I. Naumkin. Dissipative character of asymptotics for the nonlinear fractional Schrödinger equation. Journal of Mathematical Physics, 60(12):121506, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laugesen:2019:RLS

[2290] Richard S. Laugesen. The Robin Laplacian-Spectral conjectures, rectangular theorems. Journal of Mathematical Physics, 60(12):121507, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2019:SQN

[2291] Hua Wang and Qixiang Yang. Scattering for the 5D quadratic NLS system without mass-resonance. Journal of Mathematical Physics, 60(12): 121508, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2019:BLC

[2292] Sitong Chen and Xianhua Tang. Berestycki-Lions conditions on ground state solutions for Kirchhoff-type problems with variable potentials. Journal of Mathematical Physics, 60(12):121509, December 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2019:PNS

[2293] Xianyong Yang, Xianhua Tang, and Youpei Zhang. Positive, negative, and sign-changing solutions to a quasilinear Schrödinger equation with a parameter. Journal of Mathematical Physics, 60(12):121510, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bridgeman:2019:FBI

[2294] Jacob C. Bridgeman, Daniel Barter, and Corey Jones. Fusing binary interface defects in topological phases: The Z/pZ case. Journal of Mathematical Physics, 60(12):121701, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aasen:2019:FCS

[2295] David Aasen, Ethan Lake, and Kevin Walker. Fermion condensation and super pivotal categories. Journal of Mathematical Physics, 60(12): 121901, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2019:SAL

[2296] Pavel Exner and Jirí Lipovský. Spectral asymptotics of the Laplacian on Platonic solids graphs. Journal of Mathematical Physics, 60(12):122101, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deJesus:2019:EMB

[2297] Anderson L. de Jesus and Alexandre G. M. Schmidt. Exact mapping between charge-monopole and position-dependent effective mass systems via Pauli equation. Journal of Mathematical Physics, 60(12):122102, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anshu:2019:MAO

[2298] Anurag Anshu, Mario Berta, Rahul Jain, and Marco Tomamichel. A minimax approach to one-shot entropy inequalities. Journal of Mathematical Physics, 60(12):122201, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carmeli:2019:WIQ

[2299] Claudio Carmeli, Teiko Heinosaari, Takayuki Miyadera, and Alessandro Toigo. Witnessing incompatibility of quantum channels. Journal of Mathematical Physics, 60(12):122202, December 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rejzner:2019:LCP

[2300] Kasia Rejzner. Locality and causality in perturbative algebraic quantum field theory. Journal of Mathematical Physics, 60(12):122301, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2019:IAL

[2301] Fang-Jie Zhao. Inductive approach to loop vertex expansion. Journal of Mathematical Physics, 60(12):122302, December 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chanda:2019:JMR

[2302] Sumanto Chanda, G. W. Gibbons, Partha Guha, Paolo Maraner, and Marcus C. Werner. Jacobi–Maupertuis Randers–Finsler metric for curved spaces and the gravitational magnetoelectric effect. Journal of Mathematical Physics, 60(12):122501, December 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia-Parrado:2019:CKI

[2303] Alfonso García-Parrado and Igor Khavkine. Conformal Killing initial data. Journal of Mathematical Physics, 60(12):122502, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rahmoune:2019:EAS

[2304] Abita Rahmoune. Existence and asymptotic stability for the semilinear wave equation with variable-exponent nonlinearities. Journal of Mathematical Physics, 60(12):122701, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vomEnde:2019:UDD

[2305] Frederik vom Ende and Gunther Dirr. Unitary dilations of discretetime quantum-dynamical semigroups. Journal of Mathematical Physics, 60(12):122702, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2019:MCR

[2306] Erchuan Zhang. Magnetic cubics in Riemannian manifolds associated with different magnetic fields. Journal of Mathematical Physics, 60(12): 122703, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2019:IPS

[2307] Seung-Yeal Ha, Jeongho Kim, Peter Kuchling, and Oleksandr Kutoviy. Infinite particle systems with collective behaviour and related mesoscopic equations. Journal of Mathematical Physics, 60(12):122704, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Le:2019:KHI

[2308] Thi Thai Le. Kelvin–Helmholtz instability in a shallow-water flow with a finite width. Journal of Mathematical Physics, 60(12):123101, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Correales:2019:IVS

[2309] Álvaro Correales and Carlos Escudero. Itô vs Stratonovich in the presence of absorbing states. Journal of Mathematical Physics, 60(12): 123301, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lebensztayn:2019:LLN

[2310] Elcio Lebensztayn and Mario Andrés Estrada. Laws of large numbers for the frog model on the complete graph. Journal of Mathematical Physics, 60(12):123302, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kotlyarov:2019:DSW

[2311] Vladimir Kotlyarov and Alexander Minakov. Dispersive shock wave, generalized Laguerre polynomials, and asymptotic solitons of the focusing nonlinear Schrödinger equation. Journal of Mathematical Physics, 60(12):123501, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gutierrez-Reyes:2019:RTL

[2312] E. Gutiérrez-Reyes. Reflection and transmission of light by a half-space with a nonlocal response: A complex variable approach. Journal of Mathematical Physics, 60(12):123502, December 2019. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kassotakis:2019:TMS

[2313] P. Kassotakis, M. Nieszporski, V. Papageorgiou, and A. Tongas. Tetrahedron maps and symmetries of three dimensional integrable discrete equations. Journal of Mathematical Physics, 60(12):123503, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaksic:2019:ISC

[2314] Vojkan Jaksic and Robert Seiringer. Introduction to the special collection: International Congress on Mathematical Physics (ICMP) 2018.

Journal of Mathematical Physics, 60(12):123504, December 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zuo:2019:SEE

[2315] Dafeng Zuo. Supersymmetric Euler equations associated to the N ≤ 3 Neveu–Schwarz algebra. Journal of Mathematical Physics, 60(12): 123505, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besnard:2019:ABNa

[2316] Fabien Besnard. Algebraic backgrounds for noncommutative Kaluza– Klein theory. I. Motivations and general framework. Journal of Mathematical Physics, 60(12):123506, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besnard:2019:ABNb

[2317] Fabien Besnard. Algebraic backgrounds for noncommutative Kaluza– Klein theory. II. The almost-commutative case and the standard model. Journal of Mathematical Physics, 60(12):123507, December 2019. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliveira:2019:RSM

[2318] Lucas H. Oliveira and Marcel Novaes. Random stochastic matrices from classical compact Lie groups and symmetric spaces. Journal of Mathematical Physics, 60(12):123508, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teramoto:2019:ECH

[2319] Hiroshi Teramoto, Kenji Kondo, Shyuichi Izumiya, Mikito Toda, and Tamiki Komatsuzaki. Erratum: “Classification of Hamiltonians in neighborhoods of band crossings in terms of the theory of singularities” [J. Math. Phys. 58, 073502 (2017)]. Journal of Mathematical Physics, 60(12):129901, December 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1195].

Helland:2020:ESS

[2320] Inge S. Helland. Erratum: “Symmetry in a space of conceptual variables” [J. Math. Phys. 60(5), 052101 (2019)]. Journal of Mathematical Physics, 61(1):019901, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2059].

Sulaiman:2020:EVP

[2321] Samira Alamin Sulaiman. Erratum: “Vortex patch problem for Euler– Boussinesq system” [J. Math. Phys. 60, 013506 (2019)]. Journal of Mathematical Physics, 61(2):029901, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1940].

Iacono:2020:CCL

[2322] Roberto Iacono. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 61(4):044101, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13, 2323].

Chandrasekar:2020:RCC

[2323] V. K. Chandrasekar, A. K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Response to “Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]” [j. math. phys. 61, 044101 (2020)]. Journal of Mathematical Physics, 61(4):044102, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13, 2322].

Falek:2020:EBO

[2324] M. Falek, M. Merad, and M. Moumni. Erratum: “Bosonic oscillator under a uniform magnetic field with Snyder–de Sitter algebra” [j. math. phys. 60, 013505 (2019)]. Journal of Mathematical Physics, 61(6):069901, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1939].

Fernandez:2021:CQA

[2325] Francisco M. Fernández. Comment on: “Quantum aspects of a moving magnetic quadrupole moment interacting with an electric field” [j. math. phys. 56, 062107 (2015)]. Journal of Mathematical Physics, 62(1):014101, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [241].

Fox:2021:RMH

[2326] Matthew S. Fox. Retraction: “Multipole hair of Schwarzschild– Tangherlini black holes” [J. Math. Phys. 60, 102502 (2019)]. Journal of Mathematical Physics, 62(5):059901, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2249].

Fernandez:2021:CSS

[2327] Francisco M. Fernández. Comment on “Solution of Schrödinger equation for two different potentials using extended Nikiforov–Uvarov method and polynomial solutions of biconfluent Heun equation” [J. Math. Phys. 59, 053501 (2018)]. Journal of Mathematical Physics, 62(10):104101, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1600].

Giusti:2021:EDR

[2328] Andrea Giusti. Erratum: “Dispersion relations for the time-fractional Cattaneo–Maxwell heat equation” [J. Math. Phys. 59, 013506 (2018)]. Journal of Mathematical Physics, 62(11):119901, November 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1433].

Meyers:2015:NDP

[2329] Catherine Meyers. New derivation of pi links quantum physics and pure math: Researchers stumbled upon a famous pre-Newtonian formula for pi while computing the energy levels of a hydrogen atom,

November 10, 2015. URL https://publishing.aip.org/publishing/
journal-highlights/new-derivation-pi-links-quantum-physics-
and-pure-math. See [442].